A Complete Bibliography of Publications in the Proceedings of the American Philosophical

Society (1950–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 June 2023 Version 1.01

Title word cross-reference

14 [Kam94].

10 [TNN71]. 13 [Kai70, Shi70]. 1398 [Kam71]. 1772 [Rau73]. 1777 [Sio51]. 1786 [CR52]. 1790s [Dur87]. 1875 [Ros75]. 1916 [Bro85]. 1920s [GS86]. 1930s [GS86]. 1940s [Bir93a]. 1956 [Kro57, Sel56]. 1959 [Ano60m]. 1980s [Gar80]. 1988 [Hea88]. 1991 [Gom95]. 1993 [McK94].

2000-Year-Old [Nor73]. 25 [Hea88, McK94]. 27 [Kam71]. 2nd [vH93].

3.7.12-14 [Dum63b]. 3.7.7-10 [Dum63b].

406 [Mer88]. 440 [Mer84].

546 [Gre92].

600 [Ost95].

A. [Pel95]. A.D. [Con58]. Aaron [Woo99]. Abbé [Bei51, Chi50, Per53, Per58]. Abdallah [RT99]. Abdication [Hor65]. Abdus [Dys99]. Abilities [Thu50]. Abode [Men69a]. Abolitionist [Sch71]. Aboriginal [HK77]. Abroad [Wri56]. Abrogation [Ega71]. ABSCAM [Gri82]. Absentee [Mor74a]. Abstract [dT58b]. Academic [Car57a, Gid50, Ing57, Tay57]. Academies [Adr56, Frä99]. Academy [Dup57, DM65, Rai92, Pen50]. Acadia [Olm60]. Acceleration [Dic81]. Accelerators [Sim87]. Acceptance [Lew56b]. Accessibility [Ano50a, Ano50b, Ano50c, Ano50d, Ano50e, Ano50f, Ano51a, Ano51b, Ano51c, Ano51d, Ano51e, Ano51f, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano52f, Ano53a, Ano53b, Ano53c, Ano53d, Ano53e, Ano53f, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano55a, Ano55b, Ano55c, Ano55d, Ano55e, Ano55f, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano58a, Ano58b, Ano58c, Ano58d, Ano58e, Ano58f, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59f, Ano60a, Ano60b, Ano60c, Ano60d, Ano60e, Ano60f, Ano61a, Ano61b, Ano61c, Ano61d, Ano61e, Ano61f, Ano62a, Ano62b, Ano62c, Ano62d, Ano62e, Ano62f, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano65a, Ano65b]. Accessibility [Ano65c, Ano65d, Ano65e, Ano65f, Ano65g, Ano66a, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Ano67f, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c]. Accessibility [Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d].

Accord [Cla78]. According [Che54, Dub60, Dum51, Haa50b, May63a, Eas82]. Account [Gra73, Sch99b, Smi83]. Acculturation [Bar52b, Cha80, VH53]. Acheulean [RGSS72]. Acheulian [Pad77]. Achievements [Lud90]. Acquaintance [Rob83a]. Acquired [Chi57b, Joh65]. Act [Bro85, Mae80, Wol55]. Action [Boy74, Fre91, Str64]. Active [Wes78]. Activities [Pic54]. Activity [EB79, Luc70, Tho50b]. Actor [Ben91]. Actual [Coo76]. Acupuncture [Bow73]. Adam [Wil94b, CC64]. Adams [Joh96, Ano63m, But74, Han76, Hom81, Kai70, Kno63, Mor74b, Mus69, Rob53]. Adamson [Mil95]. Adaptability [Gra75]. Adaptation [Nic64]. Adaptations [Det55, Moc68]. Adapted [DK54, VD53]. Adaptive [Fle75]. Addendum [Kah56]. Address [Ati93a]. Addressed [Ada87b]. Adèle [Haa50b]. Adenohypophysis [Li72]. Adjustments [Fol58]. Administration [Cop54, Lin68, RD73, vD86]. Administrator [Tay57]. Admiral [Con70]. Admirer [Rob53]. Adrenal [Ken53]. Adrian [Hop61]. Adrift [Fri96]. Advance [Bre61, Lok78]. Advancement [McG79, Rho98]. Adventures [Gra96]. Advertising [Viv74]. Advice [Gol86]. Advisers [Zuc80]. Advising [Gol91]. Advisory [Pur65]. Advocacy [Dub79, Ede75, Tei92]. Aegean [Fra83, VW78]. Aemilius [Bat86]. Aero [Hou66, Kar60]. Aero-Biophysical [Hou66]. Aero-Field [Kar60]. Aeronautical [Dry53, Dry54, Hun54]. Aeronautics [Pom54]. Aeschylus [Lon58]. Aesculapius [Cor65]. Aesop [Wol91]. Aesopian [Pye78a]. Aesthetics [Ree86]. Affair [Fos89, Hei94, BL57]. Affecting [Bar54]. Affectional [HZ58]. Affinities [Dav52]. Affinity [Ant63]. Africa [Lam80]. African [FGWP92, Vog73]. Africans [Was63]. Afro [Fra79]. after [Goo58, War77b]. Again [Rai85]. against [Bla71, Bur73, GS86, Mae77, Tal99, War63]. Age [Bad68, Fry88, Jac75, Lin55b, Luk81, McK73, Mur55, Rho98, Sei96, Wil51a, dB79, Car58, PB74, Shu58]. Agencies [Par57, Rus57]. Agency [And57]. Agenda [Str95]. Agents [Was58]. Ages [Byn98, Neu63]. Aggradation [LB79]. Aggregation [Kar99]. Aging [McG88, Not54, Ril88b, Tib54]. Agnicayana [Dum51]. Agnihotra [Dum64]. Agrarian [LBWD86]. Agricultural [Ott86, Ott89, Tol51a]. Agriculture [Ada78, Man65]. Ahead [Ams91, Pie80]. Aid [MS98]. Aids [Mar53, Tat54, Thi93]. Ainsworth [Moo87]. Air [Amd70, Mot75, WM70]. Aires [Pye78b]. Airplane [Hit52]. Akhaeans [Güt84]. Al [Daw60]. Al-Husayn [Daw60]. Alaska [Han50, Her54, Mil64b, Wol82]. Alaskan [Mil64a]. Alatamaha [Plu77]. Albani [Ver58]. Albert [Abr91, Bai89, Pey58]. Albertini [Jar72]. Albums [Sch99a]. Aleksei [Bus97]. Alexander [Per99, Wil98, Cor99, Cra58, Eds91, Hab96, Kro58, Mit58, Whi60, dT58a, dT58b, dT59, dT60]. Alexandre [Far82]. Alexandrian [Sch99a]. Alexis [Gol64a]. Alfonsine [GCM94]. Alfred [Coa91, Haa50b, Mau99]. Algal [Bur84]. Algonquian [Wol88]. ’Ali [Daw60]. Allegany [Sny82, Sny83]. Allegory [Law92, She83]. Alliance [Dai95a, Mer77]. Allocating [Ful68]. Allusion [Lav75]. Almanacs [Het75].

Almost [Mer83]. along [Han50, Ker82]. Alonso [MV94]. Also [DSR50, Ekv80]. Altar [Dum51, Gil82]. Alternative [McG97]. Altitude [Mor77]. Alva [Bry88]. Alvin [RB96]. Amaldi [Sal92]. Amateur [Dri52, Has80, Ros90]. Amateurs [Hol70]. Amazon [Ald76]. Ambiguities [Bai93, Con60]. Ambiguity [Tei92]. Ambivalence [Tei92]. Ambrosian [LL74]. Amendment [Sto87]. Amendments [Mar65]. America [And76, Ano92e, Bok95, Bri61, Bro78, Bur80, Car74, Car58, Chi57a, Cla63, DM53a, Fry89, GS86, Gum55, Haa96, Han53a, Han63, Hea88, LVF98, Lon57, May81, Mil59, Mil82b, Pat97, Pri98, Ran70, Rei68, Rei75, Ric50a, Rob62, RK82, Sch57c, Ser91, Sim78, Smi79b, Smi88, Sta91, Whi58b, Whi60, Whi70a, Wil57, Wil80, Wri50a]. American [Ada87b, Ano60m, Dil67, Ega71, Fre62, GS94, Hea88, Hei51, Kim93, LH85, Mea54, Shr60a, Swa52, VH53, Ada93, And61, App97, Ast82, Bar60, Bel50, Bel51, Bel59, Bel64a, Bel72, Bel75, Bla73a, Bro89, Bun75, Car93, Car57b, Chi54, Cle54, Cle64, Con92a, Con50, Cor65, Cre85, CS52b, Cur61, D’E66, DL57, Dav91, Dew56, Dod63, Dur62, DM50, DM65, Dye76, Dye89, EWW78, Egg65, EJ56, Fra79, Gey62, Gib92, Gil86, Gre60b, Gri72b, Gru67, Har53, HC88, Her61, Hij83, Hol70, Hys60, Jak68, Joh74, Jon63a, Jur66, Kam87, Kee68, Kle71, Kli73, Klu59, Kro57, Lar93, Las91, Law92, Leo53, Lin51, Lin52a, Lin52b, Lin53, Lin55a, Lin56, McP99, Mel75, Mik58, Mor74a, Oak77, Oli50, Pen50, Plo78, Rob72, Rol56, Sch55, Sch53]. American [She85, Shi51, Shr61, Shr64, Shu53, Smi84, Sny53, Sny58, Sta91, Sta50c, Ste93c, Sti77, Sto87, Tho50a, Tho90, Ver64, Viv74, Voe50, Voe52, Was63, WW70, Wil51a, WD96, Zig57, dBdS54, dT58b]. Americanism [Lin52a]. Americanist [Swi98]. Americans [Bak94, Dav91, Gil92, Mil76]. Ames [Dou59]. Amir [Daw60]. Amoeboid [Rap60]. among [Ama59, Bes79, EJ56, Fen56, Gou66, Gra96, Gru89, Hei51, Kin54, Lan89, Mag56, Min73, MS67, Rob83a, Sal67, VH51]. Amount [Dub60]. Amsterdam [Faj74]. Analog [Leo94]. Analysis [BB53, Bla52, McG67, Nic64, Nou50, Per69, Pol77, Sou52, Ste86]. Analytical [Ewa59]. Anastos [Vry99]. Anatolia [Goe53, Mel66]. Anatomical [Win67]. Anatomy [Mil82a, Pur65, Sim59, Win67]. Ancestor [Man58]. Ancestry [Hoe90]. Ancient [Arr99, BB53, Fry98, Glu56, Hab85, Han73, Kam78, Las75, Spe51a, Wil90, Bor87, Bro72, Chr84, Jac63, Kil71, Lan90, Mei66, Por96, Pro84, Wil56]. Ancients [Pay78]. Anderson [FC92]. Andes [Jul87]. André [Chi57a, Sch56b, Chi57a, RH58, Sch57c, Ste97]. Andrei [Dre94]. Andrew [Pat70b]. Anesthesia [Dri64]. Anfinsen [Sel96]. Angevins [May89]. Anglo [Ada87b, Con92a, Dav91, Mor74a]. Animadversiones [Jar72]. Animal [Cha83, Dum62, Dum69, Has79]. Animals [Urt81]. Ann [Boo78]. Annals [McC85, Tay50]. Annie [Ing65]. Anniversary [Bar64a, Coh52b, Sli98]. Announcement [Ano92k]. Antarctic [Hoi75]. Ante [Shi70, Hil69]. Ante-Bellum [Shi70, Hil69]. Antebellum [Bon74]. Antecedents [Pan63]. Anthony [McI75]. Anthracite [Wal83]. Anthropological [Goo81, Kro57, Spo81]. Anthropologists [Mur61b, Pan72]. Anthropology [Egg69, Goo69, Gru67, Her79, Tud69]. Anti [Boy74, Old94, Pey58]. Anti-Christian [Pey58]. Anti-Liberalism [Boy74]. Anti-Semitism [Old94]. Antibiotics [Fle67]. Anticipations [Mum54]. Antigenic [Fra60b]. Antigens [HH76]. Antioch [Dow59, Dow58]. Antique [Car81]. Antiquities [Ver58]. Antiquity [Dav53, Rei70, Neu63]. Antivivisectionists [Vis72]. Antoine [DK54]. Ants [Höl79, Sch57b]. Anuvākas [Dum61, Dum63b]. Anxiety [Bou84, Boy74, Sid64]. Any [Gre98a]. Aphrodisias [Cra58]. Aplanatic [Abr72a]. Apollo [Pen57, Sti92]. Apostasy [Moo87]. Apotheosis [Chi55]. Appalachian [Ott89]. Apparitions [Bai62]. Appeals [Sto85]. Appearances [Dav52]. Appendix [Ano55g, Moy92, Saw92, Ano92e]. Appetite [Ste92, Tei64]. Appleton [Gil82]. Application [Rob89]. Applied [Foo65]. Appointment [Abr87]. Appraisal [Ram69]. Appreciation [Det87]. Apprenticed [Cor65]. Approach [Ber67, BC69, Ram77]. Approaches [HPSH50, Rho76]. April [Bir93b, Bog99, Con92b, Dai95b, Gal91, Gel95b, Gil94, Gli94, Jef99, Lhe93, Loc97, McC93, McK94, Mil98, Ost99, Pel95, Roc93b, Sel56, Shu98, Ste99a, Str97, Str91, Tob97, Vry99, Woo98, Woo92, Zan92, vdH94]. Aquatic [Wal51a]. Arab [Daw60]. Arabic [Lew80, Ros66]. Arachnids [Reh54]. Aragon [Rus62]. Aran [Kea76]. Arbitration [Gar74, Gil74, Gom74]. Arboretum [Sch56b]. Arcadia [Wil57]. Arcanum [Han76]. Arch [Tay50]. Archaeological [CSL50, Kid61]. Archaeology [Ada66, Arr99, Bre61, Car58, Cza76, Dal66, Dau56, Lin52b, Rai66, Ver58, Wer53, Wil66, Wil67]. Archeology [Wit50, Wit51]. Archimedes [Cla82]. Architectural [Sch99a]. Architecture [Hux90, Sim62, Smi71]. Archival [But54]. Archive [Eks99]. Archives [Bel59, Rei90]. Arctic [Gid52, Ric50a]. Area [Car58, McG99, Shr64, Sny53, Spi93, Coc64]. Areas [Abr62, Dil68, Myr51, Sel53, SW50, Spe51b]. Arian [Pat76b]. Arid [Kle82]. Aristocracy [Sym60]. Aristotle [Fre93, Oeh74, Sou52]. Armed [Wal80]. Army [Jac80]. Aron [Led94]. Art [Ano72m, Bar87, Bar90b, Cha71, Ett79, Fel87, Fle86, Fon98, Fri66, Gal74, Gri72a, Han73, Kra58, Lei90, Lin52a, Neu89, Ree86, Smi72, Spi56, Zio87, Mel75, Nor73]. Arthur [Bau94, Her79]. Article [Fos71, Hil70, Mac65b, Shi70]. Articles [For96]. Artifacts [Dup52]. Artificial [Kol65, Whi58a]. Artillery [dLL70]. Artist [Ehr90, Hor65]. Artists [Chi56]. Arts [DM65, Fry76]. Asclepiad [Pen57]. Ascorbate [CP79]. Asher [Fen56]. Asia [Bra72a, Dal66, Jo70, Mov53, She85]. Asian [Luk81]. Asklepios [Bie57]. Aspect [Pat74]. Aspects [Anc74, Aul83, CBL61, Cun57, Fle66, Fru77, Gil89, Har62, Hoo55b, Hyd67, Kat70, Neu72, Not54, Tho50a, Ver58, Wil55]. Aspirations [Pet69]. Assassination [For76]. Assay [Cha83]. Assembly [Bei51]. Assessment [Kae64]. Assignment [Bor65]. Assimilated [Moc68]. Assistants [Mor64a]. Assizes [Smi86]. Associates [Ewa52]. Association [Can77, McG79, Lon57]. Associations [HHDA66]. Associative [AE62]. Assumptions [Gol75]. Assyria [Por96]. Assyrian [CWMM79, PB74]. Assyriology [Neu89].

Astrological [NV64]. Astrology [BR79, Gra98]. Astronomer [Pin99]. Astronomers [Plo78]. Astronomical [Bel64a, CR52, Neu62a, NP67, Neu86, Tho50b]. Astronomy [BG91, Bra85, Cha64, Dys96, Het75, Neu72, Pin78, Pou97, Urt81, vH96]. Astrophysics [Sim87]. Asymmetry [Roc97]. Athanasius [O’M77]. Athenian [Con93, Mer75]. Athens [Fra75, Hab96, Lan90, Mer77, Mer84, Ost95, Tho53]. Atlantic [Coc64, Kay68, Shr64]. Atlas [MMK^{+}86]. Atmosphere [Pat70a]. Atmospheric [Stu70]. Atom [CC64]. Atomic [Mum54]. Attainable [Pan89]. Attempt [Dub60, Lok78]. Attendance [Dav85]. Attention [Coa88]. Attitude [Ros57]. Attitudes [Col72, Hij83, Kli73]. Attributed [Mar72]. Aubrey [Got91]. Auction [Gru89]. August [Ahr99, Bel92, Bog99, Bra92, Coa99, Dun97, Eds91, Fir91, Fis94, Fre96, Gre98b, Jef95, Lec94, LB95, Lip91, Mel98, Nat91, Ost98, Ram99, Sal98, Sch94, Sch99c, Soz92, Sto99b, Tie98, Tob91, Tra98, Wat99, Wes98a, Bor65]. Augustin [Zig57]. Augustine [Fre96, Mor80a]. Augustus [Gol94, Tay50]. Aumale [Tur71]. Aurora [Leh60]. Aurorae [Leh60]. Austin [Ano92k, RR92a]. Australia [Cop54, Dav52, Dav53]. Australian [Sim77]. Australopithecus [Tob96]. Austria [Boy74]. Authenticity [Mar53]. Author [Ech54]. Authoritarianism [Sta50b]. Authority [Haa80, Rog51]. Authors [Mos87]. Autobiographical [Hil78]. Autographs [Bel59]. Autonomic [Pic54]. Avant [Nem68b]. Avant-Garde [Nem68b]. Avian [Sib58]. Awakening [Gel68]. Away [Kri84, Qui60]. Axe [Beg97]. Axes [Jet61a]. Axiomatization [Mou59]. Ayom [Gus61]. Azad [Jet61a]. Azuzul [GB94].

B [Gol94, Loc97, Moy92, Mur61b, Oak98, Wol95b]. B.C. [Ant63, Gre92, Mer84, Mer88, NPZ81]. B.C.E. [Ost95]. Babel [Pep50]. Baby [Cha85, Coh87]. Babylonia [Por96]. Babylonian [Neu51, Neu54, Neu63]. Bach [Sch97]. Back [Ano50g, Ano50h, Ano50i, Ano50j, Ano50k, Ano50l, Ano51g, Ano51h, Ano51i, Ano51j, Ano51k, Ano51l, Ano52g, Ano52h, Ano52i, Ano52j, Ano52k, Ano52l, Ano53g, Ano53h, Ano53i, Ano53j, Ano53k, Ano53l, Ano54g, Ano54h, Ano54i, Ano54j, Ano54k, Ano54l, Ano55h, Ano55i, Ano55j, Ano55k, Ano55l, Ano55m, Ano56g, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano57g, Ano57h, Ano57i, Ano57j, Ano57k, Ano57l, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano59g, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano60g, Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano61g, Ano61h, Ano61i, Ano61j, Ano61k, Ano61l, Ano62g, Ano62h, Ano62i, Ano62j, Ano62k, Ano62l, Ano63g, Ano63h, Ano63i, Ano63j, Ano63k, Ano63l, Ano64g, Ano64h, Ano64i, Ano64j]. Back [Ano64k, Ano64l, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79g]. Back [Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano87h, Ano88e, Ano88f, Ano88g, Ano88h, Ano89e, Ano89f, Ano91e, Ano92f, Ano93e, Ano95e, Ano96e, Ano96f, Ano96g, Ano96h, Ano97e, Ano97f, Ano97g, Ano97h, Ano98e, Ano98f, Ano98g, Ano98h, Ano99e, Ano99f, Ano99g, Ano99h, CH81]. Backbench [Lub78]. Background [Ham84, Lin53, Miz79, Per53, Set56]. Bacon [Mar61]. Baconian [Bev67]. Bacteria [Rob86, SESB65]. Baedeker [Hab85]. Baer [Opp90]. Bagatelles [Chi59]. Baird [Her54]. Balance [Mal58, Mul62, Roo90, Sey57]. Baldi [Big54]. Balkans [Ske75]. Balzac [Ger78]. Bang [AH75, Ree93, Ber95]. Bank [For98]. Banks [Rau73]. Baptiste [Hoi75, Wri97]. Barbara [Fed94, Wil75a]. Barbarorum [Lad75]. Barber [Din55]. Barbeu [Ald51]. Barbeu-Dubourg [Ald51]. Barghoorn [Kno91]. Baring [Kap98]. Bark [For76]. Barkhouse [Spr98]. Barnum [Pow95]. Baron [Tur71]. Baroque [Smi71]. Barrel [Wil75b]. Barrier [Woo66a]. Bart. [Con51]. Bartholomew [Fre97]. Bartlett [Pel95, Wes98a]. Bartolomé [Han53a]. Barton [Ewa83]. Bartram [Har53, Plu77, Por74]. Bartrams [Cap74]. Basalt [BT60]. Base [LB79]. Bases [Rat76]. Basic [Bod63, Foo65, Hum51, HvW53, HvW61, Sam99]. Basis [Abr62, Bon79, Ful51, Mil67, vM80]. Baskerville [Got91]. Bastards [Sym60]. Batak [Bar52b]. Be [Ega78, HC84, Hor58, Kut99, McG96, Por92, Wie71]. Beach [Gli94]. Beaches [Sar78]. Beagle [Key79, Ano79m]. Beauvois [Gil92]. Bebel [Mae77]. Beccaria [Pac52]. Becomes [Coo76]. Bed [Car89a, Car89b]. Been [Ver64]. Bees [vF56]. before [Rai86, Rob72, Som76, Zir59]. Beginnings [Cha86, Wes56, Wol88]. Behalf [Lew56b]. Behavior [Cla66, Cot71, Det81, Dev69, HHDA66, Lil62, Spi69]. Behaviorism [Bla65]. Being [Lur72, Spi56]. Belated [Mil87]. Beleaguered [Far82]. Belgian [Sol71]. Belief [Bor52]. Believe [Wol91]. Bellum [Shi70, Hil69]. ben [Neu62b]. Bendix [Sme93]. Benedict [Kam71]. Benefactor [Bel75]. Benjamin [D’E66, Ald51, AF54, Ald62, Ano60m, Bel55, Ber93, Chi50, Chi53, Chi55, Coh50, Coh52b, Coh56, Cop51, Cor53, Cor58, D’E70, DM52, DM55, Ech54, Han53b, Jam54, Kah58, LB57, Lab64, Lin55a, Lin56, Lop64, Lop77, Lop85, Mar72, Mar51, McB56, Pac50, Pac52, Sel50, Sel56, SC84, Ste87, Tho91, Wol55, dM56]. Bennett [And99]. Benthic [Pat76a]. Bentley [Fry96]. Benyowski [DM55]. Bequests [Ben91]. Berensons [Gil76].

Bergson [Cai60]. Bering [Gid62]. Berlin [KB82]. Bernard [Sti59]. Bernhard [Roa98]. Bernini [Scr91]. Bers [BK96]. Bert [Sax96]. Besant [Ing65]. Bessborough [Bel95]. Best [SC89]. Beta [Esa67]. Better [Reu97, Spi93]. between [Ant63, Can77, Chi57a, Con60, Eri67, Gla81, Gri73, Ken58, Kuz75, Mer77, Ser91, Ste63a, Wei93]. Bevis [Ash81]. Beyond [Bré70, Ada91, Lok78, MS98]. Bible [Bar99, Met91]. Biblical [Ste87]. Bibliographic [TZ66]. Bibliographical [Neu62a, NV64]. Bibliography [McC91]. Bid [Bes79]. Biennis [Cle64]. Big [AH75, Ree93, Ber95]. Bilingual [Rav85]. Bilingualism [Bar52b, Eme62]. Binary [Cam91, Tay93]. Bind [For98]. Biochemical [Cha83, Fru77, Gla65, Hyd67]. Biochemist [Eds85]. Biochemistry [EB79]. Biographer [But74, Gun96]. Biographical [Bei60, Cra78, Rob50]. Biography [Bat86, Fra79, Kir66, Pot62, Scr80]. Biological [Bon79, BP67, Ell83, Hug69, Key92, Mil67, Sch56a, Ste92, Ste86]. Biologically [Wes78]. Biology [Bil71, Eds67, EB79, Per69]. Bioluminescent [CT69]. Biophysical [Hou66]. Biota [Yoc96]. Birch [Ahr99]. Bird [Gri73]. Birds [Ama59, Gre60c, Roo94, Ros90, Mur62]. Birkhoff [Lan98a]. Birth [Bar64a, Llo89, Par93, Wol73]. Bischoff [Roa98]. Bison [Jak68]. Bit [Cho98]. Bite [Gay91]. Bittel [Güt92]. Black [AH75, Cor76, Ree93, Whe81, Wil92, Sod89]. Black-Body [AH75]. Blackall [Abr91]. Blacks [Ano92e]. Bladelets [HK62]. Blalock [Tau76]. Blanckenhagen [Har93]. Blind [GH58]. Bliss [Hor95]. Bloemmark [Faj74]. Blood [Leh59, McP99, Mur55]. Bloodletting [SS57]. Blum [Gil94]. Blyth [EG59]. Board [Pur65]. Boardroom [Bow95]. Boas [Voe52]. Bodily [Pic54]. Bodin [For72]. Bodleian [Fah71]. Body [AH75, Bun91a, Men61, Ste98]. Boehmer [Sel96]. Boekhandelaar [Faj74]. Bohr [Mar88]. Boltwood [Bad68]. Bon [Cha85]. Bonald [Ree86]. Bonaparte [Bro78]. Bondaroy [Ewa59]. Bonham [Gra72]. Bonner [Cle98]. Book [Cor50]. Books [Col72, Dav63, Mor74b, Rob83a, Roc93a]. Booming [Rob67a]. Borisovich [Fri91b]. Boron [O’G78]. Borrowing [Eme62]. Bose [Mos50b]. Bosworth [Nat91]. Botanical [Ano62m, Chi57a, Chi57b, Ewa52, Pen50, Smi62]. Botanists [Rob57a]. Botany [SC84]. Bottlenose [Lil62]. Boulding [Lat95]. Bound [Lon58]. Boundaries [Jul87]. Boundary [CR54, Cop55, Mil64b]. Bourbon [McB56]. Bourne [Par96]. Boussingault [Aul70]. Bow [Eme53]. boyli [Leh59, Leh59]. Brāhmana [Dum57, Dum59, Dum61, Dum63a, Dum63b, Dum64, Dum51, Dum54, Dum60, Dum62, Dum65, Dum69]. Bradbury [Ric50b]. Braga [Smi71]. Brain [KS91, Pen54, Ste74]. Brand [Rob53]. Brant [Boy73]. Breeders [Wei86]. Brennan [LB95]. Brenton [For99]. Bretonne [Chi58]. Brian [Bre92]. Bribes [Ros64]. Bridge [Woo66a]. Bridging [Ser91]. Brief [Gol77, Hij83]. Bright [GHKW95]. Bring [RR92b]. Bristol [Her82, dLL70]. Britain [Joh54, Rau73]. British [Adr88, Bro62a, Con92a, For62, For71, Fra60a, Hal55, Han50, Jon63a, Kra80, Lab64, Lan89, McG79, New90, Rob83b, Sch53]. Britton [Gle60]. Broader [Hol90]. Broadsides [DK54]. Brogan [Nic62]. Broken [Mah94]. Bronze [Bie57, Jet61a, Sca76, Ver64]. Brook [Lik99]. Brooks [Lit96]. Brotherly [Hof83]. Brothers [Kap98]. Broughton [Ken95]. Brunelleschi [Pra68]. Bryan [Hor61]. Bryn [McP98a]. Buchan [Eis50]. Buenos [Pye78b]. Builders [Luk81, Pra68]. Building [Dum51, Mah94, McP98a, Rob67b, Sch95a]. Buildings [Gre60a]. Bulgarian [Bar64b]. Bulkeley [Jod87]. Bullets [Wit99]. Bundle [MS67]. Bundles [TZ66]. Bundy [Kay98]. Bünning [Bon94]. Bunyan [Kin76]. Buonaccorso [Sca76]. Burbank [GS80]. Bureau [Viv74]. Bureaucracy [Clo80]. Buren [Ray80]. Burger [O’C95]. Burgess [Yoc96]. Burins [HK62]. Burke [Cop50b, GB56, Hof50]. Burr [Woo99]. Burrhus [Bra92]. Business [Fab58, Joh74, Rie84, Roc63]. Businessman [Wyl59]. Businessmen [Rie84]. Bustamante [Ben79]. Busts [Lav75]. Butler [Lan76]. Byte [McC91]. Byte-Sized [McC91]. Byzantine [May63a, Set56].

C [Bel92, Her79, Kam94, LL74, Mau99, Sno58, Ano92e]. C. [Boe58, Fis94]. ca [Ash81, Chr84, Gre92, O’M77, vH81]. Cabot [Fos71, Hew70]. Cacao [Ald76]. Caddoan [Tay63]. Cadwallader [Lok78]. Cahen [Lew97]. Cairo [GP81]. Calculating [Gre93a]. Calcutta [Ewa83, Gor88]. Calendar [Fen53]. California [Gou66, HK62, Leh59]. Called [Fry83, Jon63b]. Calvin [Bou84]. Cambrai [Fin82]. Cambrian [Yoc96]. Cambridge [Sok72]. Cambyses [Lan72]. Camden [DeM84, Fus57]. Came [Llo89]. Campaign [Sto85, Wil67]. Campaigns [Por96]. Campbell [Sta98]. Campsite [GBN77]. Camus [Pey58]. Can [Coo84, Cus50, For98, Hol90, Hor58, Kut99, Mat99, Por92, Wie71]. Canada [Mil64b, Edw61, Hut79]. Canal [Sio51]. Cancer [Bay69, CP79, Dul97, Now95, Par76, Spi76, Sta57, Tal99, Zub76]. Candid [Ste89]. Candide [Wad57]. Canoes [Dra60]. Canons [Lew94]. Can’t [Reu97]. Capital [Bec92, Leo53]. Capitalism [Sil95]. Capito [Lew54]. Captain [Cla54]. Captivities [Bar50]. Caradon [Ano97m]. Carbohydrates [Sta51]. Carbon [Kam94, Kee70]. Cardano [Gra98]. Cardiac [C.65, Jar72]. Cardinals [Hir53]. Care [Bun91a, Kee63, Lan99b, Sho93]. Career [Cor50, Fre97, Kam86, Ken71]. Carl [Dun97, FC92, Gin99, Wil94b]. Carnegie [Dub79, Pat70b]. Carroll [KWB97, Wea54]. Cartoon [Wol55]. Casas [Han53a]. Case [Bau93, Ben79, Big54, BG91, Bra85, Bur73, Bus97, Cha64, Fen71, Fra79, Gor80, Gra72, Hou90, Kim62, Kin76, Moo87, Nas68, Opp90, Pat74, Pet63, Rav85, Roo94, Sym72, vH96]. Castañega [Dar79]. Caste [Mor60, Wil79]. Castellio [For72]. Castiglioni [Li55]. Castile [Llo89]. Castilian [GCM94]. Castle [Nat91]. Cat [Tri80]. Catalog [Mor74b]. Catalogue [Ano56m, Jon63a, Sel56]. Catalysis [Fru77, Sin99]. Catawba [Mer83]. Catcher [Ric68]. Catechist [vH81]. Category [Jul87]. Cathedrals [Bun91b]. Catherine [Wei66, Ati52]. Catherinian [Sch99a]. Catholic [Boy74]. Cattle [Ott86]. Cattle-Ranching [Ott86]. Caucus [Mor69]. Causality [Gra75]. Cause [But50, Jar72, Lop85]. Causes [Dew56]. Cave [Ang52, Coo52b, Dup52, Mur74, Ste77]. Cavendish [McC98]. Caxton [Wol91]. Celebation [Ano56t]. Celebration [Sca87]. Célèbre [Lop85]. Celestial [Wee54]. Cell [Col65, Cri88, ES56b, HH76]. Cell-Mediated [HH76]. Cells [Par76]. Cellular [Rap60]. Cemeteries [Myl55]. Censorship [Som76]. Centennial [Ler59]. Center [Lan55]. Central [CSL50, Jac63, Mov53, Fre93]. Centrifuge [Mac53]. Centuries [Alb60, Ant63, Con58, Han53a, Nem80]. Century [Boy74, Cor57, Dry54, Fra93, Gla65, Hun54, Kam94, Ler77, McP98a, NPZ81, Tur71, Woo53, Zei88, Bis79, Bla73b, Boa54, Chi58, Coh95, Col72, Com56, Dav91, DS56, DM53a, Fen51, For76, Fos82, Frä99, Gla60, Gol88c, Gou89, Haa80, Han53b, Han54, Het75, Jen68, Kan89, Kli73, Lav75, Lok78, Mae80, Nas68, Nic56, Pac50, Pat74, Per53, Per58, Rei75, Sch95a, Sch71, SC89, Ske75, Smi71, Smi86, Smi88, Smi89, Spo81, Sti92, Str61, Ted87, Ver58, Wal83, War85, WW70, Wit51]. Certain [BR79, CBL61, Con60, Kni79]. Cessac [HK62]. Chōsen [Con56]. Chaim [Gil95]. Chain [Hug69, Jen71, She85]. Chains [Smi79a]. Chair [Pal78]. Challenge [CWH^{+}71, Dea67, Fuc98, Kli59, Zwo61]. Challenges [Lud90, Roo94]. Chamber [Laf75]. Champollion [All60]. Chance [Gra75, Gre93a]. Chandrasekhar [Ost98]. Change [Coc70, Dea67, Led92, LS55, Rob76, Roo94, Sim55, Tei88, vD77]. Changed [CR54]. Changes [LVF98, Mil64a, Ric70, Roo94]. Changing [Cla63, EWW78, Gid50, Hum89, Kee70, Kin78, Roc63, Roo72, Ste52, Tru59, Whi70a]. Channing [Puk57, Ric53]. Chant [Bar52b]. Chaos [Dil68]. Chapter [Gla81, Mer61, Spo81]. Character [Hoe68, Met91]. Characteristics [Gou89, Ros89]. Characters [DK54, Lan76]. Charcot [Hoi75]. Charge [Smi86]. Charity [Bel55]. Charles [Bor65, Cas92, Con51, Eag98, McG99, Bai97, Bar52a, Bro78, Can61, Con51, CR52, Dar59b, Eis57, Eis63, EG59, For99, Kau95, Ken58, Law92, Len65, RC67, Tre74, Wil80, Yoc96]. Charms [For71]. Charter [Wol52]. Chatham [Kni79]. Chaucer [Mal50]. Chemical [Bry88, Eis94, Sta50a, DK54]. Chemicals [RD73]. Chemist [Eds85]. Chemistry [Mer53, Mul62, Sin99, SC84]. Chemists [Coh82]. Chesapeake [Dav61]. Chest [Kol65]. Chicago [HC88]. Chickahominy [Ste52]. Chickens [Rob67a]. Child [Ede75]. Childhood [Boo75, Bri75]. Children [Ede97, Hof75, Kie80]. Chile [Heu66, TNN71]. Chimpanzee [HH51]. China [Bod63, Bro89, Cho89, Coa88, CB96, CS52b, Dav86, Hum89, Pet73]. Chinese [Bod63, Bod85, Cam91, Fon71, Fon98, Hum51, Pye78a]. Chivalry [Sal50]. Chlorella [Spo51]. Choice [Ful68, Joh54]. Choices [Dub63, Gun96]. Cholera [Van70]. Choose [Hea88]. Chrétien [Abr72a]. Christ [Mar77]. Christendom [Moe61]. Christian [Fra75, Fry83, Law72, Pey58, Ric80, Sel96]. Christmas [Lop77]. Christoph [TG91]. Christopher [Lar93]. Chromosome [DSR50]. Chromosomes [For96]. Chronology [Bar99, Mer71, Mer78, Sar78]. Chumash [MS67]. Church [Bri61, Nem68a, Bré68]. Churingas [Dav53]. Cicadas [Fle75]. Cincinnati [Bie57]. Circles [Per58]. Circulation [Dri64]. Cirrhosis [Bes52]. Cirripedes [Tre74]. Cities [Pri98, Sch70]. Citizenship [Bes60, Gol64a]. Citoyen [Chi54]. City [Led92, Nas68, Sul68, Wil92, Wol52]. Civic [Con93, McP99]. Civil [Dur62, Gey62, Jon63a, Jon67, McP99, War66, Bur80, Nic62]. Civilization [Bod85, Fen56, Klu58, Kro51, PB74, Spe63]. Civilizations [Dia94, Eis84, Mat99]. Clactonian [Ohe78]. Claim [Dow58, Ros54]. Claims [Raw78]. Clans [Goo76]. Clark [DeV55, Ehr90, Mou83]. Clash [Mat99]. Classical [Gra98, Rei68, Tho66, Ver58, Wes56]. Classics [Gum55]. Classification [Coo52a, Mil64b, Sim59, Tay63]. Classify [Dub60]. Claude [Lew97]. Claudel [Pey70]. Clavicula [Car81]. Cleanth [Lit96]. Clear [Bel56]. Clerical [Don76, Nem68b]. Clerks [Pie65]. Clews [Rei50]. Climate [Frä99, Mur55, Rob76, Rob79]. Climatic [Roo94]. Climbing [Bau93]. Clinic [Bol62]. Clinton [Hop61]. Clio [Hof83, Sei90]. Clock [Bed64, Cop50a, Llo50, Smi84]. Closed [Kin54]. Closing [Ste93a]. Cloud [Gri73, Rod92]. Cnidus [Pen57]. Co. [Hou90, Kap98]. Coach [Whi70b]. Coalition [Cor76]. Coast [Rau73]. Coastal [Mil64a, Sar74, Sar78]. Cobbett [Dur87]. Coca [Par83]. Code [Row91]. Codification [Mil59]. Cohen [Din55]. Cohort [Wol73]. Colbert [Olm60]. Colden [Lok78]. Coldspring [CS53]. Coleman [Moy97]. Coleridge [Abr72b]. Collaboration [Mor58]. Collaborators [Spo81]. Collapses [Dia94]. Collected [Ste63b, Bai89]. Collecting [Dor57]. Collection [Dav63, Yoc96, Eis54a, Fen71, GS94, HK62, Ran53, Reh54, Rei50, Sel51, Wal51b]. Collections [Pen50, Sny58, Tay54, Ver64, Zig57]. College [Bar60, Con60, Bar90a, Gor88, Tho71]. Colliders [Pan89]. Collision [Her93, PG52]. Colloid [Eds85]. Colombian [Sib58]. Colonial [Ald76, Ben79, Gor88, Gum55, Har89, Jur66, Kni79, Mel57, Mur56, Smi79b, Wer53, Wok89, vH81]. Colonies [Men89, Hin64]. Colonists [Dys78]. Colophons [Büh70]. Color [App84, GH58, Rat76]. Color-Blind [GH58]. Colors [GBF60]. Columbia [Mos55b, Han50]. Columbus [Lar93]. Combination [Sta50a]. Combray [Bré68]. Comes [McK73, Zie82, dB79]. Comet [Whi76]. Coming [Cor72, Red78]. Comment [Fos71, Mac65b, Shi70]. Commentariolus [Swe73]. Commentary [Bun79, Jon63a, Kam86, Ric80, Swe73]. Comments [Bac77, Din55, End57, Jef65, Ken55, Mea54, Mos55a]. Commerce [Kap98, Laf75]. Commercial [Viv74]. Commission [Cai60, Lum78]. Commissioner [Sec65]. Commissioners [Cop55]. Commissions [Kim77]. Committee [Ano65v, LH85]. Common [Gre92, Mal98, Ric68, Mil59]. Commoner [Kni79]. Commons [Lub78, Ros75]. Commonwealth [Hal55]. Commune [Bla88]. Communism [Hoo55a]. Communist [Nem68a, Rob55]. Communities [Ham92, Has79, Mah94, Pat76a, RR92b]. Community [Gib63, Sat82, Sch98, Ste52, Tin70]. Compact [Hen87]. Company [Pet51, dLL70]. Comparative [Aar74, Ald76, Dev69, Gre60b, Har89, LS55, Ott86, Ott89, Spe69, SW77, Swi98, Wil76]. Comparison [Cro81, Dys78, Kle82, Kuz74, Gre93a]. Comparisons [But54]. Competition [Tay72, WE79]. Complex [MS67, Ree93, Sei90, Ste86]. Complexity [Sim62].

Components [Kuz75]. Composer [And76]. Compositae [Mag56]. Composite [Eme53]. Computation [Neu51]. Computer [App84, Cho77, Cla87, Gol77, Rob67b]. Computer-Generated [Cla87]. Computers [BG77, Coo84, Gol64b, Sho93, War77a, Gol92b]. Computing [Mor64a]. Concept [Böw97, For98, Ger62, Gol64a, Ler59, Par63]. Conceptions [Gun93]. Concepts [Bod63, Car57a, Fel87, Gla60, Mur61b]. Concern [Goo90]. Concerning [Can60, Pat84, Fus57, Gre93a, Kah56]. Concerns [Jac63]. Conditions [Mor77, Sim58]. Condorcet [Gil72]. Conduct [Dri53]. Conference [Gre54, Shr55, Ada87b, HC88]. Conflict [Ohl74, Wil70, vD77]. Confrontation [Mar61]. Confucius [Hau67]. Confusion [Gor80]. Congress [Kro57, Ano56m]. Congressional [Mor69, Tru59]. Conklin [Bar64a, BB84]. Connecticut [Jod87]. Conquest [BC69]. Conscience [Han63]. Consciousness [Ong76]. Consensus [Ful61]. Consequences [Mum54, Pay88, Reu97, Roc97, dT60]. Conservation [DG73]. Conservative [Hau67, Mid85, Wes98b]. Conservative-Liberal [Hau67]. Consideration [Sch57b]. Considerations [Bur60b]. Considered [Mom50]. Conspectus [Bor52]. Constant [MD64, Wri50b]. Constitution [Blu88, Bla73a, But51, Chr84, Cor56, Gre87, Hen87, Hom81, Kam87, Tru59, Wes56]. Constitutional [Jen71, Mar65, Sca87]. Constitutions [Pet99, Kim93]. Constraints [Gib81, dD98]. Construct [ER90]. Construction [Hos89]. Consulship [Tay50]. Consults [Cor53]. Consumption [Kap98, Sch95a]. Contact [Bec89, Tol63]. Contacts [Cop51, Dod63, Elk63, Swa52, Was63]. Contamination [Pat70a]. Contemporary [Klu58]. Content [Fre93]. Context [Boo75, Jon99b]. Contexts [Nis75]. Contextual [Pat76b]. Continent [Dav52]. Continental [Bur93a, Els68, Kay68, Pie68]. Continuations [Roa66]. Continue [Wie71]. Continued [Kra50]. Continuing [Det81, Sin99]. Continuity [Ger62]. Contour [Rat71]. Contractility [Nic50]. Contractions [Fab58]. Contradictions [Ast82]. Contrast [Rat71]. Contrasts [But54, Fru85]. Contribution [Bie59, Gib92, Per69, Rat68]. Contributions [Len65, New90, New61, OC81]. Control [Ber54, Coa67, Mea54, Zim74]. Controlled [Har58]. Controlling [Sec65]. Controversial [Lin55b]. Controversialist [Eds85]. Controversy [Aul83, Hol70, Mag85, Men61, Pat76b, Sim78, Wol59, Fos89, Mor74a]. Conversion [Mil87]. Conybeare [RC67]. Cook [Cla54]. Cooke [Bar52a, McI75]. Cooper [Wal52]. Cooperation [Ber57, Cai60, Cor76, LC62, Pai54, Sch73]. Cope [Maa81]. Copernican [Eas82, Gra73, Tho50b]. Copernicus [Har73, Gin73, Ros57, Ros60, Swe73]. Copies [Bie62]. Coping [Whi94]. Copying [Bie59]. Coral [Bur84]. Corn [Man58, Man83, Wal56]. Cornpone [Hil69, Shi70]. Corporation [Cha86, Moo60]. Correction [Kel57]. Correctional [Ohl74]. Corrections [Wil74]. Correspondence [CBA97, Eis57, Lew63, Pac55, Ril51, dT59]. Correspondent [Bel51, Pac52]. Corruption [Dur87, Lew54]. Corsica [Pal61]. Cortex [Dav68, Ken53]. Cortical [KHO52]. Cos [Pen57]. Cosmic [AH75, Mer53]. Cosmological [Hub51]. Cosmology [AH75, Boa54, Gra85, RC99a]. Cosmos [McG95, Pel61, Ree93]. Costs [Bau93]. Coulee [Dau56]. Counselled [Cop55]. Count [DM55, Tur71]. Countess [Hei94]. Countries [Cop56b]. Country [SC89, Sum60]. Counts [DSR50]. County [McG99]. Courageous [Bow57]. Court [Abr87, Fra54, Kat97, McK94]. Courtier [DL57]. Courts [Pol85, Wol94]. Covenant [Cha51, Jen71]. Covert [Cot71]. Covington [Zie82]. Cowper [QC53]. Craft [Bar90b]. Craig [Sid99]. Crawford [Jef95]. Creation [Aul83, Cro70, Fry83, Mel57, New73, Pel61, Tho98, Wel89]. Creation-Science [Fry83]. Creative [Sas80]. Creativity [Kam86]. Creel [Roy96]. Cremin [Rav91]. Crime [Wol73, Zim74]. Criminal [Row91, Wol94]. Criminality [Sel53]. Crises [Cas69]. Crisis [Adr88, Alb81, CWH^{+}71, Con92a, For98, Pat94b]. Critical [Bla65, VD53, Yod82]. Criticism [Bar90b, Hux90, Wel75]. Critics [Hab85]. Critique [Mur55, RC67]. Croce [Dav63]. Crown [Fus57]. Crowned [Mor77]. Crucial [Cla59]. Crusades [May81]. Crust [Bur93a, TAT53]. Cruz [HK62]. Cryptozoology [Sim84]. Crystal [SE98]. Crystals [Sui53]. Cuban [Riv58]. Cullen [Bel50]. Cultivation [Bar52b]. Cults [Sim55]. Cultural [Coc58, Egg69, Fen71, Goo69, LS55, Miz79, Moc68]. Culture [Bon79, Boy73, Car58, Col72, Elk63, Hum51, Jon99b, Kap98, Mur61a, Mur61b, Nic61a, Pad77, Rie84, Roc93a, Sim55, Smi60, Ste52, Sum60, Wit52, vH81]. Cultures [Arr99, Goe53]. Cuneiform [Kra63, Spe63]. Cuno [Cor72]. Curator [Tat54]. Cure [Bay69, Zub76]. Curious [Bau93]. Current [Car90a, Wil68]. Currents [Gid52, Key92, KHO52]. Curti [Boy98]. Curves [See51]. Cuss [Rei86]. Custodies [Gri82]. Custodiet [Gri82]. Cycle [Aul70, Lew56a, Mor77]. Cyclic [Sch57b]. Cynical [Rho98]. Cyprus [Dow58, VW78]. Cyrano [Nie86]. Cyriacotheriidae [RK82]. Cyril [Lec94]. Cytogenetics [OC81]. Cytopathogenicity [End57]. Cytotaxonomy [Men51]. Czech [Nem68b, Nem80]. Czechoslovakia [Nem68a].

D [RC67, Fre97, For80]. Daily [Hol90]. Dal [Ver58]. D’Alverny [Con92b]. Dámaso [MV94]. Dame [EoL99]. Damião [Hir51, Hir53]. Daniel [Bai62, She85]. Dante [Maz55]. D’Aranda [Pet63]. Dardistan [Jet61b]. d’Arimathie [O’G78]. Dark [Ree93, Rub88a, Rub88b]. Darwin [Key79, Ano79m, Bar74, Bur86, Con51, Dar59a, Dar59b, Eis54a, EG59, Gre59, Irv59, Kot80, Leo94, May91, May95, Mün59, Sta60, Tre74, Zir59]. Darwinism [Dob65, Her61, Lin51, Wyl59]. Dashed [Lan76]. Dashkova [WD96]. Data [HvW53, HvW61, Oli50]. Date [Ste77]. Dates [Coh56]. Dating [Bad68, Swa52, Tre53]. Dauphiné [vD77, vD86]. Daux [Jam95]. David [FC92, Bow93, D’E70, Lai98, Rob60, Sio51]. Davis [Coa99, Sch94]. Dawn [Han63]. Day [Cor50, CSL50, Jon55, Mün59, Pot72]. Days [Cor50, McG97]. Daytime [BG91]. Dead [Ben91, Gol80, Tre53]. Deafferentation [Mor64b]. Dean [Tho91]. Deane [Bor93]. Dear [Ber93]. Death [Byn98, Hij83, Jen68, Mae80, Bev89]. Debate [Hau67]. Debates [Bla73a]. Deborah [Ril51]. Decade [Pie80, Shi51]. Decades [CB96]. Decapitated [Win79]. Decays [Hal63]. December [Bea96, Bec94, Bel95, Con99, Dem94, Dre94, Dul95, Eag98, Eis98, Fri91b, Gin99, Gla93, GHKW95, Gre98c, Hen97, Him99, Jam95, Jef94, KWB97, dK95, Kit96, Lei98, Mac97, Mel97, Oak98, Par96, Pow95, Roa98, Sal92, Tra98, Wil94a, Wil94b, Wol95b, vH93]. Deciduous [HvW53]. Decisions [Koo90]. Declaration [Sta57, But51, Chi54]. Decline [Bur93b, Rai92, Rot77]. Decrease [Hof75]. Decreasing [MD64]. Dedication [Ano60m]. Deed [Bur80]. Deep [Ste63b]. Defects [Sui53]. Defense [Lew99, Mae80, Sec65, VK70, War66]. Defenses [Fle86]. Deficiencies [Kel57]. Defoe [Bai62]. Degradation [Whi94]. Dehumanize [Sho93]. Dehydrogenases [Smi79a]. Deistical [Ald50b]. Delaware [Bec89, Cro73, Gru89]. Deletion [Smi79a]. Deliberate [Hen82]. Delight [McG99]. Delivery [Sho93]. Delmarva [Cop56a]. Delmas [Set92]. Delta [Ric50a]. Demand [Fri61]. Demanded [Dub60]. Demise [Red78]. Democracies [Son54]. Democracy [Bur52, For98, Gre98a]. Democratic [Jam93]. Demographic [BC69, Gou89, Har89, Kin78, Kle89b, Men89, Not54, Ros89, Tei88]. Demography [Kle89a, Tho50a, Wel89]. Demotic [NPZ81]. Dense [Her93]. Densely [Myr51, Spe51b]. Density [CS52a]. Dental [Luk81]. Department [And57, Ega71, Viv74]. Dependence [Sta57]. Depositories [Fen53]. Depression [Lub78]. Deprived [HHDA66]. Depth [Sch69]. Derivation [Swe73]. Derived [Lei90]. Descent [Hoe90, Kra50]. Describing [Goo81]. Description [HR55, Llo50, Yoc96]. Desert [May63a, Dil67]. Desertion [Bri55]. Desertions [But50]. Deserts [DW65]. Design [Cha71, Cha87, Wil55]. Designer [Bed64, Kar99]. Designing [BG77, Ste69]. Destiny [Bro70]. Destroyed [Gre92, Pen94]. Deteriorated [Shi51]. Determinants [BP67]. Determination [All66, Ker82]. Determining [VH51]. Determinism [Hoo55a]. Dethier [HE96]. Deuteroscopy [Pat76b]. Developed [Cop56b, Kuz75]. Developing [Tay99]. Development [Aus81, Bai89, Bev67, Bri75, Bro75a, Cop54, Gre69, Gru67, HZ58, Har89, HH51, Hou90, Hun54, Jor61, Kel57, Kim93, Kno93, Lon57, Mac53, Mag99, Mel79, Mor69, Mor64b, Nic50, Pic54, RD73, Spe51b, Sul68, Tau76, Too89]. Developmental [Mag99]. Developments [Dua66]. Devil [Lat62]. Devils [CS52b]. Devise [Kat97]. Devotio [Nem80]. Dharmaśāstras [Roc93a]. Diabetes [Bes52]. Diagrams [Heu66]. Dialects [VH51]. Dialogue [Tay53]. Diamonds [Sui64]. Diaries [Haa50b]. Diary [Jod87, Kid59, Sno58]. Diatom [Pat76a]. Dickinson [Ano63m, Kno63, Law92]. Dictators [Son54]. Dictatorships [Bun75]. Dictionaries [VH53]. Dictionary [Cra78, Scr80]. Did [For76, Lan72, Ros54]. Difference [Ant63, Lan84]. Differences [Ama59, Kuz74]. Different [Coa88]. Differentials [Kuz75]. Difficult [Haa96]. Diffraction [Har58]. Digest [McC91]. Digit [LG78, Rai85]. Dilemma [Goo81]. Diller [Rai86]. Dimensions [For72, KS84]. Dimorphism [Kot80]. Dinner [Ano60m, Ati93a]. Dioxide [Kee70].

Diplomacy [Con70, Lin51, Nic67]. Diplomat [Sha58, dM56]. Diplomatic [RT99, Woo53]. Direction [Chi56]. Directions [CWH^{+}71, Wil74]. Disadvantaged [RR92b]. Disadvantageous [Gre77]. Disagree [Mac65c]. Disarmament [Cla53, Cla78]. Disarray [Gem89]. Disciple [Ald51]. Discipleship [Bar83]. Discipline [New90, Roo90]. Disciplines [Gre80]. Discourse [Jac89, Jam93, Fus57]. Discover [Coo84]. Discoveries [Gid62, Tob96]. Discovery [Bad78, BHT86, Kil71, Mar51, Mer61, Mul72, Yoc96, Hux86]. Discretionary [McK94]. Discriminations [GH58]. Discussion [BCGR81, CWH^{+}71, Dav52, Jac68, Mal58, Nic62, Par57, Sny82, Was63]. Disease [Joh65, Ken53, Mal55, Pau52, Rad93, Rat68, Sin94]. Diseases [Jar72, Sch63]. Disenchantment [Bok95]. Disintegrating [Kro51]. Dismal [Bau93]. Disobedience [Jon67]. Dispersal [Riv58]. Dispersals [Ohe78]. Display [Eml62, Spi93]. Dispute [Eri67]. Disputes [Aar74, Gom74]. Dissection [Fle67]. Dissent [Bar83, Fry89]. Dissidence [Fry89]. Distance [Men69b]. Distinctions [Sch50]. Distinguished [Haa96]. Distorting [Nic61c]. Distribution [Mag56, Riv58, Sch69, SW50, Tay99, Wil51a]. Diverse [Pat84]. Diversity [Dic98, Gib92]. Divide [Has61]. Divided [Gre93b, Jo70]. Divine [McE84, Was58]. Division [ES56b]. Divorce [Goo53]. Dixon [Cop51, Llo50, Rob50]. Dmitrievich [Dre94]. Do [Adr92, Cus50, For96, Per96, Rei86, Wol91]. Doab [Pad77]. Docket [McK94]. Dockyards [Haa80]. Doctor [Cla57]. Doctors [Cor53, Mar86, Sho93]. Doctrine [Fra60b, Sta55]. Doctrines [Gar80]. Documentary [Poe56]. Documentation [Cra90, dT58a, dT58b]. Documenting [Smi71]. Documents [BB53, Chi57b, Haa50b, Pot62]. Dodd [Lop85]. Does [Hoe90, Orl71, Sil95]. Dog [For76]. Dogs [Kin54]. Dollar [Nic67]. Dolores [Ben79]. Dolphin [Lil62]. Domestic [Kin54, Leo53]. Domestication [Mün59]. Don [Neu98]. Donald [Sta98]. Done [McG96]. Doolittle [Yoc96]. Doorway [Wal93]. Doryline [Sch57b]. Double [Bar60]. Doubled [Gib95]. Dougan [Nie92]. Doughty [Wes98a]. Downfall [Sta55]. Dowsing [Rid52]. Dr [McG99]. Dr. [Bel50, But51, Coh95, CS54, Cor58, D’E66, Jef65, Lan59, Lin51, Lin94b, Rob57b, Sca87, Sio51, Was73, Wil75b]. Draft [Cha51, Swe73]. Dragon [But74]. Dragon-Killer [But74]. Drama [Pay78]. Dramatist [Ben71]. Drawings [Ver58]. Dredgings [Ste63b]. Dress [Kli73]. Drift [Els68, Kay68, Pie68]. Driftwood [Gid52]. Drinker [Chi56]. Droits [Chi54]. Droseraceae [Wil76]. Drugs [Sam99]. Dual [Eks99]. Dualism [Ste98]. Duane [Hec94]. Dubourg [Ald51]. Dubridge [RB96]. Duby [Con99]. Duché [Gar75]. Duel [Hor61]. Duhamel [Ewa59]. Dumesnil [Haa50b]. Dunlop [Ano98i, Lan98b, Lan98c]. Dupré [Zig57]. Durand [Rob57a]. during [Ald76, CR52, Cop55, For96, Fre93, GS86, McB56, Nem68a, Rei68, Ric70, Sch95a, Smi89, War85, Woo53]. Duroselle [Wri97, Nic62]. Dutch [HK77]. Duty [Bro72]. Dwight [WW70]. Dyck [Mar77]. Dying [Bev89]. Dynamical [Mac65a]. Dynamics [Eis84]. Dyspnea [Jar72]. Dzibilchaltun [And60].

E. [GHKW95, Hou90, Mil95]. Eades [Fry96]. Ear [Kam86]. Earl [Don76, Kni79, Eis50]. Earliest [Gri72a]. Early [Blo86, BG91, Chi57a, Dav60, Dav63, Dur87, DM50, Dye76, Dye89, Gre60b, Gri72b, Jac75, Ken71, Kle89a, Kub59, Moo87, Phi53, Rom56, RK82, Sim78, Sta91, Ste89, Stu65, Wil61c, Wit52, Zei88, Ano89g, Dod63, Goe53, Gra89, Hin64, How59, Joh74, Jor61, Lin51, Lin52a, Neu72, Row91, Tur71]. Earth [Hed65, Bad68, Mac65a, MD64, Pat70a, Rob76, TAT53, Whi58a, Wil68]. Earthquake [Kno93]. Earthworm [Mom50]. East [Goi71, Mal53, Spe51a, Can77, KB82, She85, oB90]. Eastern [Ada66, HC88, Jac75, Wit52]. Eatton [Reh54]. Ecclesiastical [Dow58]. Ecclesiology [Nem68a]. Echolocation [Gou65]. Eclipse [NPZ81]. École [Chi58]. Ecological [Dia94, Lik99, She61]. Ecology [Bri75, Bro75a, Rat68, Sta60]. Economic [Ald76, Cho89, Cla50, Coc58, Coc64, Cop54, Dys78, For98, Har62, Hoo55b, Kin78, Kle71, Kot51, Kuz58, Kuz67, Lan99a, McG88, Nou50, Roo72, Smi88, Sou52, Spe51b, Wei93]. Economics [Cho77, Cho89, Fuc96]. Economies [Tay99]. Economist [Fab61]. Economists [Mac65c]. Economy [Bec92, Cop56b, EWW78, Kut99, Sil95, Spe69]. Ectoprocta [Sch69]. Ecuador [Sar74]. Edgerton [Str91]. Edict [Lew54]. Edinburgh [Bel50]. Edith [Mel98]. Editing [Bur86, Lew63, Met87]. Edition [VD53]. Editions [Cra58, Phi53]. Editor [Tat54]. Editorial [But54]. Editors [Phi53]. Edmund [Mil98, Bar83]. Edoardo [Sal92]. Education [Ast82, Atk99, Bur73, Cho89, Cor50, Cre85, Dic98, Fuc98, Gre93b, Ham92, Ken60, Keo98, Lan99b, Ler77, May60, Pai54, Rav85, Rob83a, Shu53, Wel74]. Educational [Dua58, Dua66, Hea71, Ple67]. Educator [McI75]. Edward [Bow90, Gre98c, EG59, Eis97, FK96, Gal91, Gil82, Law72, Mur61b, Pet63, Ram99, VD94, Wel74]. Edwin [Bar64a, BB84, Fai92, Sea93, vH93]. Effect [Cra90, Mur55]. Effective [Ede75]. Effects [Cur63, Dew56, LVF98, Luc70, MD64, Rob79]. Egesta [Mer77]. Eggan [Fen92]. Egypt [Lan72, Lew54, Wil56]. Egyptian [Gil89, Ant63]. Ehrlich [Wit99]. Eighteenth [Ler77, SC89, Bla73b, Chi58, Com56, DM53a, For76, Frä99, Gla60, Gou89, Han53b, Han54, Kan89, Lok78, Mae80, Pac50, Rei75, Sch95a, Sch71, Smi71, Smi88, Smi89, WW70]. Eighteenth-Century [Ler77, Chi58, DM53a, Gla60, Lok78, Rei75, Sch71, Smi71, Smi88, WW70]. Eighth [Dum69, Neu62b]. Einstein [Dic81, Hol81, Mar88, Wei81, Bai89]. Eisenia [Mom50]. Election [Sto85, Ray80]. Elections [Mid85]. Electric [Hou50]. Electrical [Dav68, Mom50]. Electricity [Coh56]. Electromotive [Mac53]. Electronics [Bak77, Sei96, Zwo61]. Electronics-Promise [Zwo61]. Elemental [ES56a]. Elementary [Gou77, May60]. Elements [Jes51, Nic64]. Eleusis [Myl55]. Eleven [Kle82]. Eleventh [Dum51]. Elias [Rob57a]. Elihu [Abr72a]. Eliot [Mey98]. Elite [Sec65]. Elizabethan [Her82]. Ellen [Tra98]. Ellery [Ric53]. Elongation [ES56b]. Eloquence [Nor93]. Elsewhere [Dav52]. Elsie [Rei50]. Elso [Kno91]. Elysian [Cof86]. Embattled [GS86, Kat70]. Emblems [Nor93]. Emergence [HvW53, HvW61].

Emergency [Bre61]. Emil [Fru85, Wil94a]. Emperor [Bow99, Kae64]. Empire [Gre92, HK77, Jul87, Fah71, Han53a, Toy55]. Empirical [Mal98, Ric80]. Employment [Gin78]. Encounter [GS80]. Encourage [Cus50]. Encyclopédie [Sha70, Wat61]. End [Bev89, Cra80, Gou89, Han54, Koo90]. Endemicity [Mag56]. Enders [Rob91a]. Endowment [Dub79]. Enduring [Oli57, Oli69]. Enemy [Joh74, Ric68]. Energetics [Gre77]. Energy [Bac77, Bro81, BCGR81, Cro81, Dan71, Gib81, Gib95, Mel79, Mor77, Mum54, Pan89, Ros81, Wei80]. Energy-Progress [Dan71]. Enforceable [Ams91]. Engelmann [Kam86]. Engineer [Gra96]. England [Aul82, Bur73, Col63, Dou59, Het75, Jor61, Kim77, Puk58, Say71, Wes98b, Zak91]. English [Tur71, Bau59, Cow60, Dav85, Han54, Joh74, Ler77, Rob53, Sto99a, Wes56]. Enhancement [McG67]. Enjoyed [Chi56]. Enlightenment [Gia97, Whi70a, Whi58b]. Enlistment [Con92a]. Ensign [Jef99]. Enterprise [But54, Dub71]. Entitled [Fos71]. Entrepreneurial [Sol71]. Entropy [Sch74]. Environment [Har71b, Kee70, Sum60]. Environmental [Mor77, Whi94]. Environments [Ekv74]. Epic [Ong76, Sol68]. Epidemic [Thi93, Wol82]. Epilogue [Gol88a]. Episcopate [Dav85]. Episode [Dur62]. Episodes [Dar59b]. Epistemic [Ste98]. Epithelial [Hug69]. Equal [Spi93]. Equality [Ams91]. Equilibrium [Bea57]. Era [Sat82, Wri72, Con93, Fre93, Her82, Jor61, Was73]. Erased [Mer83]. Erasmus [Hir51]. Eric [Abr91]. Erich [Sat82]. Ernest [Bro62b, Eis98]. Ernst [Opp90, Opp82]. Errata [Ano62m, Ano63m, Ano89g]. Erratum [Ano72m, Ano79m, Ano98i]. Errors [Mal98]. Erwin [Bon94, Gel95a]. Esarhaddon [Por96]. Esau [Ste99a]. Eskimo [KB80]. Eskimos [Swa52]. Especial [Mur55, Sta91, Tob96]. Essay [Ald76, Chi57a, Gum55, Lur72]. Establishment [Pri62]. Estate [McG99, Sch95a]. Estimates [Gem89]. Estimating [Bec89]. Estivation [FGWP92]. Estranged [Kin76]. Eternal [Mac84]. Ethical [Fre71]. Ethics [Dav81]. Ethnic [Bur80, Gib92, Jul87, Swa52, McP99]. Ethnicity [Lan89]. Ethnoanatomy [SW77]. Ethnographic [Bur83, Bur88]. Ethnohistory [Klu59]. Ethnological [Jet61b, Kro57, Sny82]. Ethos [Ekv74, Spe56]. Etiquette [Mar93]. Eudoxus [BG91]. Eugene [Str91, Pow95, Wei97]. Eugenics [Gla81, GS86]. Euoenotheras [Cle54]. Europe [DL57, Ser91, oB90]. European [Anc74, Ano89g, Bec89, Blu82, Bor65, Col72, Cor72, Cza76, Fre93, Gra89, Jon99b, McN79, Par93, Pye78b]. Evaluating [Wri50b]. Evaluation [Ken52, Wel75]. Evangelist [Bes79]. Evans [Bur83, Coh95, Lan59]. Evans-Pritchard [Bur83]. Eve [Mor74a, Wil80]. Evelyn [Pat94a]. Evening [Cle64]. Every [Car89a, Car89b, Fuc96]. Everyman [Wri62]. Everything [Wol91]. Evidence [Bro59, Cla59, Gol71, Gou65, Luk81, Pie68, Vin68]. Evil [Bla88, Che54]. Evolution [Aul83, Bod63, Bur93a, Cla59, Cle54, Dob59, Gom95, Hac97, Jen71, Lap83, Mag56, Mor60, New73, Osb53, Pic54, Rai86, Riv58, Rom56, Sim59, Sjo72, Smi79a, Ste59a, Sto87, Sto99a, TZ66, WE79, Wil76, Wil79, Wol91, dD98].

Evolutionary [CBL61, Cle64, May59, Mur61b]. Evolutionism [Dob65]. Evolving [Ong76]. Ewing [CR52]. Ex [Big54, Plu77]. Exact [Neu63]. Exactions [Hof83]. Examined [Enr75, Leo53, Ray80]. Examples [Lew94]. Excavation [MO59]. Excavations [And60, Coo52b, KJF52, Wac53]. Excellence [Wil61a]. Exchange [Sou52]. Exchanges [Chi57a]. Execution [Jes51, Lan90]. Executive [Rus57]. Exercise [Wil67]. Exertion [Dub60]. Exhibition [Sel56, Ano56m]. Exhortation [Spe56]. Existence [Chi58]. Expanding [McC91, Tay54, Wri50a]. Expansion [Dur67]. Expectancy [Kan89]. Expedition [Jet61a, Ati52, Bor65, DeV55, Gil89]. Expense [Ost95]. Experience [Gre87, Gut64, Mur56, Pat76b, Put84]. Experiences [Han63, Ric68, Ste74]. Experiment [Cor76, Koy53, LC62, Mel75, Wel74]. Experimentalists [New90]. Experiments [Coh52b, Con50, Han63, Nyg70, Wil55]. Expertise [Bro75b]. Explained [Qui60]. Exploitation [Wal51a]. Explorations [Hub51, Ric50b]. Exploring [Bor65]. Explosions [TAT53]. Exposure [War63]. Expression [Jan89, Sto87]. Extensively [Ver64]. Extent [Het75]. External [Joh54]. Extinction [Hen82, Sim85]. Extra [Ken52]. Extra-Sensory [Ken52]. Extragalactic [SW50, Wri50b]. Eye [Det55, GH58, Kam86]. Eyed [Gun93]. Eyes [Boy73, Rob86].

F [Bok95, Ede97, Sca87]. Fabian [Sti59]. Fable [Wol91]. Fabricant [Lip91]. Fabrication [Gre87]. Fabricators [Sol71]. Facet [Rid52]. Facets [Lin55b]. Facilities [Dua58, Dua66]. Facsimile [Coh50]. Fact [Pac55, Par83, Sim78, War66]. Factor [Gre93a, May59, Ske75]. Factors [Alb72, Cla50, Coc58, Kin78, Spe51b]. Failure [Pur65, Woo66b]. Fairbank [Coh93, Hof91]. Fall [Gre92, dBdS54]. Fallible [Ing65]. Fallowing [Ott89]. False [Büh70]. Family [Hof75, Kir56, RK82, Smi89]. Famous [Wat78]. Faraday [Coh87]. Farm [McG99, Poe56]. Farmer [Bic65, Wal56, McG99]. Farmers [Tol51b]. Fascination [Jef65, Mac65b]. Fathers [Dys78]. Feathers [GBF60]. February [Bow93, Bre92, Cas92, Coa99, Cox94, Cra93, Dul95, Dun97, Gil95, Gol97, Hea88, Her95, HE96, Hof91, Kau95, Ken95, Lew97, Lhe93, Loc97, May97, Met93, Rob91a, Ros75, Sme93, Sme95, Swe93, VD94, Wal99, Wat99, Woo98, vH93]. Federal [Pet99, Sea66, Smi79b, Ada84, Gre87, HC88, Mag58, MS98, Sec65]. Federalism [Dou59]. Federalist [Mos87, Car89a, Car89b]. Feeding [GJ55]. Feline [GB94]. Felix [WE98, Cra93, Bla88]. Fell [Kun86]. Females [CB96, Gru89]. Fenimore [Wal52]. Fenollosa [Bro62b]. Ferdinand [NRF81]. Fermi [Mos50b]. Fernandus [For96]. Ferryman [Mur59]. Fertility [BP67, Coa67, Coa88, Kuz74, Kuz75]. Festival [Lew56a]. Fez [Min73]. Fiat [Hoo55a]. Ficta [Jac68]. Fiction [Blu82, Bré70, Par83]. Fiefs [May84]. Field [CS52a, Kar60, Mar53, Mos50a]. Fields [Cof86]. Fieldwork [Bur83]. Fifteenth [Nem80]. Fifth [Dum60, Kro57]. Fifty [Goo58, Sli98]. Fight [Mae77]. Figures [DW65, Ger74]. Fiji [Dod63]. Finance [Ker54, Mit58]. Financed [Foo65]. Financial [Joh54, Mal98, McG88, Tob99].

Financing [Bro85]. Findings [Ste77]. Fire [Dum51, Dum64]. Fire-Altar [Dum51]. Fire-God [Dum64]. Fireball [Oli57]. Firm [Saw89]. First [And76, App97, Coh56, Dum54, Dum57, Dum64, Eis54b, Kam94, Kni79, Lee81, LG78, Rai85, Roa66, Sch99b, SE98, Swe73, Wil80, WD96, Zei88, Con58, NPZ81, Sch59, Sto87]. First-Hand [Sch99b]. Fiscal [Joh54, Mus72]. Fischer [Fru85]. Fisher [Dou59]. Fishes [Edw61, Riv58]. Fission [BHT86]. Fitch [Tre74]. Fitzhugh [Dav61]. Fitzwilliam [Hof50]. Five [CB96, Dub77, See61]. Flag [CS52b]. Flame [Has80]. Fleece [Atk99]. Flexner [Spr98]. Flick [Det87]. Flies [Lam80]. Flood [Sto59]. Floor [Kit73]. Flora [Rob57a]. Florence [Dav63]. Florentine [Bec62, Dav60]. Florida [Ott86, Por74]. Flow [Wei69]. Flower [Has61]. Flowery [CS52b]. Fluids [Her93]. Fly [Det81]. Flycatchers [Ste63a]. foetida [Mom50]. Folding [Kar99]. Folio [Wat61]. Folk [For71, Han68]. Folklore [Bau79, CH81, Dor57, For59, Sny82]. Folktales [Ran53]. Following [CWH^{+}71]. Food [EJ56, Hil69, HC84, Jen51, Man65, Sch95c, Shi70, Spo51, Wal51a]. Footnotes [Ano72m, Smi72]. Force [Mac53]. Forces [Hou50, Wal80]. Foreign [Bun75, Jes51, Leo53, Mae77, She85, Sta50c, Wri56]. Forest [DG73, Ott89, Sch57c, Hij80]. Forestry [Sch57c]. Forests [Han50]. Foreword [Car90b]. Forgery [Fle86]. Forgetting [Hux86]. Forging [She85]. Forgotten [Mur59]. Form [Mor80a, New61, Per62, Sal50]. Formal [Goi71, Shr60b]. Formation [Pat76a]. Formosa [HK77]. Formulas [Car81]. Formulating [Bur60b]. Formulation [Jes51]. Forster [Kah58]. Fortes [Zan92]. Fortifications [Web76]. Fortune [Kit73]. Forty [Kra73]. Forz [Tur71]. Fossil [Kee70, New59, SESB65, Tre74]. Fossils [Ric51a, Rom68]. Fothergill [Boo78, CS54, Cor58]. Fougeroux [Ewa59]. Found [Sim55]. Foundation [Alp55, Dic81, Dub77, Fin82, McC85]. Foundations [And61, Kot51, Moe58, Saw89]. Founders [Kam87]. Founding [Dup57, Kro58]. Four [App84, Han53a, Hij80]. Fourteenth [Boa54, Gol88c, Nem80, Str61]. Fourteenth-Century [Boa54]. Fourth [Cor57, Dum61, Dum63a, Sch63, Con58, Oli69]. Fourth-Century [Cor57]. Fowler [Sal98]. Fra [Dav60]. Fragmented [Kle89b]. Fragonard [She83]. France [Chi57a, Kam71, KS84, Kie80, Laf75, McB56, Sha58, Som76, vD77]. Francesco [Jar72]. Francia [Wil75b]. Francis [Ahr99, And76, Blo97, Bon74, Hop61, Ken58, Mad62]. Francisque [Roa70]. Francois [Chi57a, Dar99, Sch56b, Sch57c, VD53]. Francois-André [Chi57a, Sch57c]. Frank [Gli94, Mer83, Wal51b]. Franklin [Ano60m, Rob91a, Ald50b, Ald50a, Ald50c, Ald51, AF54, Ald62, Ano56m, Bel55, Bel56, BL57, Ber93, Bry88, But50, Cha85, Chi50, Chi53, Chi55, Chi56, Cla54, Cla57, Coh50, Coh52b, Coh56, Coh87, CBA97, Cop51, Cor53, Cor58, Cor56, DM52, DM55, Ech54, Han53b, Han54, Jam54, Kah55a, KRF55, Kah55b, Kah56, Kah58, Lab56, LB57, Lab64, Lew56b, Lin50, Lin51, Lin52a, Lin55a, Lin56, Lop64, Lop77, Lop85, Mag85, Mar72, McB56, Mil55, New71, Nic56, Pac50, Pac52, Pac55, Phi53, Ril51, Rob53, Rol56, Sel50, Sel56, She83, Shy87, Spi56, Ste87, Waa74, Wil83, Wol55, dM56]. Franklinia [Plu77]. Franks [Sme95]. Franz [Fru85]. Fraud [War66]. Fred [Fen92]. Frederic [Bra92]. Frederick [Cle98, Bur75, Ewa52, Jur66]. Free [Abr62, Sto87]. Freed [McP95]. Freedom [Bai93, Car57a, For98, Fre68, Fry89, Ful68, Gid50, Gou50, Hea71, Ing57, Mil82a, Sil95, Tay57]. Freeze [Ste63b]. French [GB56, Hoi75, Ald50a, Ald51, Chi53, Far82, Gil89, Hys60, O’G78, Roa66, Rob57a, Sti77, Wil75a]. Fresh [Say71]. Freund [Cox94]. Freyre [dCTdM92]. Friedrich [Cra68]. Friend [Ric68, Waa74, Ber93]. Friends [Sny58]. Friendship [Cor58, D’E73, Goi71, Hir51, Hir53, Shk93]. Friendships [Lab64]. Fritz [Jef65, Wat95]. Frog [Leh59]. Front [Ano50m, Ano50n, Ano50o, Ano50p, Ano50q, Ano50r, Ano51m, Ano51n, Ano51o, Ano51p, Ano51q, Ano51r, Ano52m, Ano52n, Ano52o, Ano52p, Ano52q, Ano52r, Ano53m, Ano53n, Ano53o, Ano53p, Ano53q, Ano53r, Ano54m, Ano54n, Ano54o, Ano54p, Ano54q, Ano54r, Ano55n, Ano55o, Ano55p, Ano55q, Ano55r, Ano55s, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57m, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano58m, Ano58n, Ano58o, Ano58p, Ano58q, Ano58r, Ano59m, Ano59n, Ano59o, Ano59p, Ano59q, Ano59r, Ano60n, Ano60o, Ano60p, Ano60q, Ano60r, Ano60s, Ano61m, Ano61n, Ano61o, Ano61p, Ano61q, Ano61r, Ano62n, Ano62o, Ano62p, Ano62q, Ano62r, Ano62s, Ano63n, Ano63o, Ano63p, Ano63q, Ano63r, Ano63s, Ano64m, Ano64n, Ano64o, Ano64p]. Front [Ano64q, Ano64r, Ano65o, Ano65p, Ano65q, Ano65r, Ano65s, Ano65t, Ano65u, Ano66m, Ano66n, Ano66o, Ano66p, Ano66q, Ano66r, Ano67m, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68m, Ano68n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69m, Ano69n, Ano69o, Ano69p, Ano69q, Ano69r, Ano70m, Ano70n, Ano70o, Ano70p, Ano70q, Ano70r, Ano71m, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano72n, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano74m, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano76m, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano79n]. Front [Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano84i, Ano84j, Ano84k, Ano84l, Ano85i, Ano85j, Ano85k, Ano85l, Ano86i, Ano86j, Ano86k, Ano86l, Ano87i, Ano87j, Ano87k, Ano87l, Ano88i, Ano88j, Ano88k, Ano88l, Ano89h, Ano89i, Ano89j, Ano89k, Ano90e, Ano90f, Ano90g, Ano90h, Ano91f, Ano91g, Ano91h, Ano91i, Ano92g, Ano92h, Ano92i, Ano92j, Ano93f, Ano93g, Ano93h, Ano93i, Ano94e, Ano94f, Ano94g, Ano94h, Ano95f, Ano95g, Ano95h, Ano95i, Ano96i, Ano96j, Ano96k, Ano96l, Ano97i, Ano97j, Ano97k, Ano97l, Ano98j, Ano98k, Ano98l]. Front [Ano98m, Ano99i, Ano99j, Ano99k, Ano99l]. Frontier [Bic65, Jur66, Ker82]. Frontiers [Dry53, Egg69, Goo69]. Frost [Sch58].

Frozen [Beg97]. Fruits [McC84]. Frye [Bev93]. Fuel [Kee70]. Fulbright [Woo98]. Fulk [May89]. Full [Dum57, Dum59, Dum60, Dum61]. Full-Moon [Dum57, Dum59, Dum60, Dum61]. Fullerton [Her54]. Function [Esa67, Kae64, Sny61]. Functional [Gal92, Pen54]. Functions [Gre60a, Sny54]. Fundamental [Gol75]. Fundamentalisms [Mar98]. Fundamentals [BR79, Mag58]. Funding [McC85]. Funerals [Jon99a]. Fuori [KJF52]. Furet [Dar99]. Further [Dav51]. Fusataro [Kub59]. Futūwa [Sal50]. Future [Blu97, Cow93, Egg65, Fra60a, Gol86, Haa96, HC84, Hut79, Kra80, Mah94, Mar98, Rob55]. Futuristics [Jo70].

G [Ano98i, Bus97, Kid59, Lan98b, Lan98c, Mer83, Rob57a, Wal51b]. Gager [Bro51]. Gain [RD73]. Galaxies [Ree93]. Galileo [Ros54, vH96]. Gallatin [Goo58, dT59]. Gallery [Mel75]. Gambit [Old94]. Gamma [Fri97]. Gangs [Con67]. Gap [Rei70]. Garde [Nem68b]. Garden [RH58]. Gardener [RH58]. Gardens [Ewa59]. Gardiner [Fah71]. Garrett [Lan98a]. Gaston [HE96]. Gatekeeping [McK94]. Gating [Key92]. Gave [CR52]. Gazette [Ald62]. Gelb [Lei98]. Gellhorn [Hen97]. Gellner [Eis98]. Gene [DSR50]. General [Bur60b, Dri64, Mou59]. Generated [Cla87]. Generation [Ada87b, Con50, Rei70]. Generations [Fai64]. Genes [Ber98, Dem54, Dul97, Sta57]. Genesis [Bod63]. Genetic [Jen51, Rus63]. Geneticists [GS86]. Genetics [CBL61, Dun73, Gla65, Man65, Men51, Son65, Ste65]. Geneva [Wat61]. Génie [She83]. Genius [Bla88, Nis82]. Geniza [GP81]. Genome [McK97]. Genomic [Sin94]. Gentry [Kan89]. Genuine [Cla53]. Genus [CT69, Sib58]. Geoforms [Spi84]. Geographic [Sch69, Sel53]. Geographical [Riv58]. Geography [HK77, Her61, Li55]. Geologic [Bar74]. Geological [Clo80, NRF81]. Geologist [Bra70, Koh74, Lan59, Law72]. Geology [New90, RC67, Vin68, Wil80]. Geophysical [Ber57]. George [Tie98, Abr72a, Bec94, Bel92, Fra79, GS80, Kah58, McD59, McG99, Pat94a, Pom54, Tol51a, Wil98]. Georges [Con99, Jam95]. Georgia [vH81]. Gerald [Fry96]. Gerhart [Wol95b]. German [Bel55, Cor72, Cra91, Gla81, Mor74b, Puk57, Sec65, Shu53, Wok89, Wol94]. Germans [Ald50c]. Germany [Cor72, Haa50b, Wer81]. Gershom [Jod87]. Gerusalemme [Lee81]. Geyl [Nic62]. Ghetto [Fin82]. Ghiberti [Sca76]. Ghost [Bur83]. Giamatti [Pel95]. Giambatista [Pac52]. Giant [DW65]. Gibbs [Wil61b]. Gidian [Bré70]. Gift [Dic98, Ric51a]. Gifts [Ing65, Ros64]. Gigantic [Gra96]. Gilbert [Cra93]. Gilberto [dCTdM92]. Girard [Chr66, Kir66]. Girardinini [Riv58]. Girls [Osw64]. Girolami [Dav60]. Girolamo [Gra98]. Giuseppe [Gla93]. Glaciations [Mil64b]. Glaciers [Mil64a]. Gladys [Set92]. Glessner [Roy96]. Glimpse [Boy73]. Global [Eis94, Kut99, Roo94, Str88]. Globalization [Gal99, Tay99, Tob99]. Glory [Jen68]. Glutamate [Smi79a]. Glynne [Lan76]. Go [Dic98]. Goals [CWH^{+}71, Ken60, Kre71]. God [Dum64, Lat70, Par92]. Gods [Kan61]. Goes [Hir51, Hir53]. Gogol [Bro91]. Gold [Jef64, Jef65, Mac65b, Mac65b].

Goldberg [Gin92]. Golden [Atk99, Sli98]. Golein [Jac69]. Gondwanaland [Pie68, Rom68]. Good [Cha85]. GOP [Bes79]. Gordion [You63]. Gordonia [Plu77]. Gorilla [Eml62]. Görlitz [Hei94]. Gospel [Ric53]. Gospels [Cad51]. Got [Kri84]. Gove [Bla62]. Governance [Viv74]. Government [Bec62, Bok95, Dub71, Foo65, Sea66, Sym72, Zal91, Gom92]. Government-Financed [Foo65]. Government-Partners [Sea66]. Governments [Gol91]. Governor [But51, Hau75, Hop61]. Grade [LB79]. Gradual [Fen56]. Graeco [Ros66]. Grail [Roa66]. Grains [Wat78]. Grand [Sno58]. Granit [EC94]. Grant [Bar64a, BB84]. Grassland [Kle82]. Grateful [Ben91]. Gratings [Har58]. Gravestones [Hij83]. Gravitation [Dic81, See51]. Gravitational [Dic81, MD64]. Gravity [Sjo72, Tay93]. Grayson [Lin94b, Ste99b]. Great [Bor50, Gri72a, Ken91, Kni79, Lin51, Lub78, Mor58, Wol82, Bar64b, Dil67, DM53b, Hab96, Rau73, She85]. Greater [Rob67a, Pye78b]. Greco [Gum55]. Greece [Hab85]. Greek [Alb72, Ant63, Bar64b, Bie59, BG91, Cra58, Jon63b, Met87, Neu72, Nob68, Sol68, Ver58, Ver64, Ver65]. Greeks [Bor87, Ric51b]. Green [Gun93, McG99]. Green-Eyed [Gun93]. Greenberg [Rin93]. Greenewalt [Jef95]. Greenhouse [Ewa83]. Greenland [LF52]. Greeting [Pop97]. Grew [Lef71]. Grievance [Gar74, Gil74]. Grimm [Kah55a]. Griswold [Gel95a]. Ground [DW65]. Grounds [Rob67a]. Group [Chi56, Cle64, Mac73, Fru88]. Groups [Dub60, Ega71, Fru85, Kin54]. Grow [Ekv80, Hoe90]. Growth [And61, Coc58, ES56a, ES56b, Gib92, Kuz58, Kuz67, Lan99a, Mom50, Mor80b, RD73, Sch95c, Tay99, Wei60, Wil83]. Guardian [Tay57]. Guatemala [Kid61]. Guatemalan [GBN77]. Guayana [Mag56]. Guide [Ewa59, Hab85]. Guinea [Gus61]. Gulf [Edw61, Ise52]. Gun [Wil75b]. Guy [Fis94].

H [Egg65, Him99, Kah55a, Mur61b, Sch95b]. H. [She83, Win83]. Haas [Wat95]. Habilis [Tob96]. Habitation [Jac75]. Habits [Hil69, Pot75, Shi70]. Hacienda [Ben79]. Haeckel [Opp82]. Haemogregarina [Leh60]. Haiti [Mat96]. Hale [Gru67]. Half [Dry54, Hun54, Kam94]. Half-Century [Kam94]. Hall [Ano60m, Shr60b, Lin50, Pet51, Tho90]. Halliday [Sny57]. Hallock [Jef95]. Hallowell [Sch94]. Hallucinatory [Was58]. Hamburg [DL57]. Hamilton [Mal58, Kro58, Mit58, Mor58]. Hamlet [VD53]. Hammer [Par92]. Hand [Gun96, Sch99b]. Handsome [Too89]. Hanseatic [Rot77]. Happening [McN79]. Harappan [LBWD86]. Harmony [Löw91]. Harold [Boo97, Bur75, Str91]. Harpur [Boo78]. Harris [Hiż94]. Harrison [GS80, Oak95]. Harry [Bat96, Gel95b]. Hartford [Pot75]. Hartington [Ros75]. Hartley [D’E70]. Harvard [Lan55, Bir93a]. Hasbrouck [And94]. Hassler [Che94]. Hastie [Sch70, Sch99a]. Haury [Wil94a]. Hauser [Kit96]. Hayden [NRF81]. Hays [Bel75]. Hazard [Shu98]. Hazards [Bru63, Rus63]. Head [Win79]. Headed [Cho98]. Healing [Mah94]. Health [And57, Bau93, Bla62, Fuc96, Ham92, Ken53, Lan99b, Lon57, Par57, Pau52, Rus57, Sat82, Sho93, Hor95]. Health-Care [Sho93]. Heaney [Ven99]. Hear [Wol91]. Heart [Fos89, Gib65, Joh65, Kol65, Shu71]. Heat [Men61]. Heavens [Hos89]. Heaving [Sch58]. Heckewelder [Wal52]. Hedonic [Ste74]. Heidelberger [McC93]. Heinrich [Har93, Opp82]. Heinz [Roc93b]. Heir [Waa74]. Held [Tho90]. Helena [Llo50]. Heliocentrism [Eas82, Gin73]. Hellenistic [Tho53]. Heller [Tob91]. Help [Sul68]. Hemisphere [Whi54]. Hemoglobin [Pau52]. Hemoproteins [Lem66]. Hendrik [vdH94]. Henrik [LF52]. Henry [Fos71, Spo81, SC84, Adr88, Ano62m, But51, Chi56, Hea88, Hew70, Lin51, Mea78, Moo87, Roc93b, Smi62, Soz92, Zal91]. Hephaisteion [Mer75]. Her [OC81]. Herbarium [Mea78, Rob60]. Herbert [Car74]. Herculaneum [Bie62]. Hereafter [Sol68]. Hereditary [Ste59b]. Heretic [Rob62]. Heritage [Ada93, Gum55, Ing65]. Herlihy [Bow93]. Hermann [WE98]. Hermeneutics [Ste86]. Hermetic [EJ56]. Hero [Lar93, Ste76]. Herodotus [Lan84]. Heroes [Jac89]. Heroism [Jac89]. Herrlee [Roy96]. Herschel [Can61, Hos89]. Hessian [But50]. Hessians [Wil83, Ech54]. Heterodoxies [Eis84]. Hewes [Fos71]. Hewson [Bel56]. Hexagrams [Cam91]. Hexter [Him99]. Hidden [Gla81]. Hierarchies [Pol85]. Hierarchy [Sim87]. Hieroglyphs [Cla87]. High [Ekv80, Fen70, Mor77, Pan89]. High-Pasturage [Ekv80]. Higher [Ast82, Bur73, Dic98, Fuc98, Ken60, Kim93]. Highest [Gol91]. Highland [Mag56]. Highlands [GBN77]. Highmore [Mil78]. Highway [Han50]. Highways [Glu56]. Hilliard [Shi70]. Himation [Bie59]. Himself [Pey70]. Himsworth [Boo97]. Hipparchus [BG91]. Hiroshima [War77b, Wil96b]. Hispanic [Eas65]. Historian [Bus97, Eas82, Pel93, Wel89]. Historians [Bow57]. Historic [Pen50]. Historical [Aul83, BB53, Blu82, Chr66, Col65, Cra91, EB79, Gem89, HK77, Hoo55a, Ker54, Lin53, Mor80a, Mor64b, Nic64, Pat97, Smi83, Wel89, Wri50a]. Historiographical [Sto99a]. Historiography [Fin63, Hin64, Ken71, Moo87, Nic62, Sei90]. History [Ald76, Ano72m, Bai62, Ber54, Bes60, Bie57, Bie59, BC69, Boy73, Bur88, Cle64, Coc64, Coh55, Dun65, Dur62, EB79, Fah71, Fen53, Fon98, Ger62, Gib63, Gol77, Gor80, Har73, Hed65, Her61, Hij83, Hof83, Hou90, Hur60, Kae64, Kam78, Kat97, Kle89b, Kra63, Lan84, Lew99, Lud90, Mac65a, Mar77, McN79, Met87, Nic61c, Ott86, Ott89, Pal61, Pal81, Pal87, Par77, PB74, Pot62, Pro84, Rei92, Ric51a, Sch97, Shr61, Shr64, Smi72, Spe51a, Spe63, Spo81, Taf62, Toy55, Wei59, Wit50, Woo53, Mea54]. Hitchcock [Law72]. Hitherto [Tre74]. Hitler [Fry98]. Hittites [Güt84]. Hoax [Cla57]. Hoaxes [Ste87]. Hoebel [Mil95]. Hofmeister [Fru85]. Hofstadter [Her95]. Hog [Hil69, Shi70]. Hole [Whe81]. Holes [Ree93]. Holland [Gal91]. Hollis [Rob53]. Holmes [Bur75, Kat97]. Holocene [Sar74]. Hologram [MM72]. Holy [Con58]. Homage [Bra81]. Homans [Bel92]. Home [Dic98, HH51, Pet51, Sau62, Fah71]. Home-Raised [HH51]. Homeostasis [Pic54]. Homer [Lat70, Ver93]. Homing [Hit52]. Homme [Chi54]. Homo [Tob96]. Honest [Rob62]. Honor [Jan89]. Hoover [Bes79]. Hope [Tal99, Kap98]. Hopi [SW77]. Hopkins [Sch71]. Hopkinson [And76].

Horace [Gri95]. Horatio [Gru67]. Horizon [Ger78]. Hormonal [Sta50a]. Hormones [Ken53, Li72]. Horoscopes [GP81]. Horror [Ett79]. Hosack [Rob60]. Hostility [Mor66]. Hotar [Dum60, Dum62]. Hotel [Ano60m]. Hotu [Ang52, Coo52b, Dup52]. House [GJ55, Lub78, Fos82, Ros75, Vog73]. Household [Sch57a]. Houses [Cha80]. Hove [Cas92]. Hovering [Hou66]. Howard [Rob91b, Cho98, Dul95, Eag98, Jef99]. Howling [CS52a]. Hubbard [Lik99]. Hudibras [Lan76]. Huggins [For99]. Hughes [Cho98]. Huguenot [Sta55]. Human [Ang52, Bec92, BP67, Bro75a, Cha51, Cha83, Cla59, Cla66, Coa67, Dav68, Dev69, Dum63a, Har71a, Jak68, McK97, Osb53, Pen54, Per96, Rad93, Rid52, Sim93, Sin94, Spi76, Spi56, Sta71, Ste65]. Humanism [Kro63, Mar88]. Humanist [McI75, Sat82]. Humanists [McC91]. Humanitarian [Cow60]. Humanities [Kee68, Lum78, McC85, Sea66]. Humankind [McG95]. Humans [LVF98]. Humboldt [Whi60, dT58a, dT58b, dT59, dT60]. Hummingbird [GBF60, Hou66]. Humor [Kil96]. Humorous [AF54]. Humors [Dew64]. Hundred [Goo58, See61]. Hundredth [Bar64a, Coh52b]. Hunger [Ste74]. Hungry [Tol51b]. Hunsgi [Pad77]. Hunt [All66]. Hunter [McD59]. Hunting [Wit52]. Husayn [Daw60]. Husserl [Bar83]. Hutchinson [Pat94a]. Huxley [Ran70]. Hybridization [Sib58, Ste59a]. Hyperbolic [See51]. Hypnosis [Hil71]. Hypothesis [Mor64b, Smi79a, Yng60]. Hypotheticality [BG77].

Ibelin [May78]. Ibn [RT99, Daw60]. Ibn-’Ali [Daw60]. Iceland [Ein57, Rau73]. Idea [Ano56t, Gia97, Opp82, Sel80, Ste76, Whi54]. Ideal [Jam93]. Ideas [Sol68]. Identical [See51]. Identification [Gol80]. Identity [Con93, Goo90, Ske75]. Ideological [May91, Pan63]. Ideology [Hoo55b, Ken55, Lad75, Mos55a, Ree86]. Ignace [Lei98]. Ignored [For96]. Ignoring [Reu97]. II [Say71, Bur60a, Cre85, Mer78, Sto99a]. III [Roc93b, Cla78]. ILL [Kam95]. Illegal [Lan90]. Illiers [Bré68]. Illuminate [Hol90]. Illusion [Enr75, Lav75]. Illusions [Lof98]. Illustrated [Lon57]. Illustrating [DK54]. Illustrations [Ehr90, Lee81]. Image [CS52b, Kni79, Kra58, Pye78b, Rub63, Ver65]. Imagery [Scr90]. Images [KS91]. Imagination [Abr72b]. Imaging [Cha83]. Imitation [Bak84]. Immigrant [Gru89, Hea51]. Immigration [Ano89g, Gem89, Gib92, Gra89, Gru89, Lan89, Rei92, Tei92, Wok89]. Immunity [HH76]. Impact [Alb81, Amd70, Can61, Dic98, EWW78, May95, Mur51, Tay68, Win67]. Imperative [Eis94]. Imperfections [Wea60]. Imperial [Top76, Rie84, Sch53]. Imperialism [Laf75]. Imperials [Jon63b]. Imperiled [Kea76]. Implantation [SH57]. Implications [Byn98, Kno93, Kuz75, McG88, Mul72, Shr55, Stu70, Wei59, WM70, Wol59, Wri50a, Yng64]. Importance [Pat84, Pot62]. Important [Ver64, Wie71]. Importation [Sod89]. Imposing [Sch99a]. Improbable [Clo80]. Improvement [Dua58, Jen51]. Improvisation [Bau59]. Inamgaon [LBWD86]. Inanna [Kra50]. Inca [Par83]. Incarnation [Kli73]. Inception [Pom54]. Incident [Bar64b]. Incidents [Pro84]. Included [McG99]. Including [DSR50]. Incomplete [Pye78b]. Increase [Jon55]. Increased [Jen51]. Increasing [Dic98, Wal51a]. Incunabula [Büh70]. Independence [But51, Oak77]. Indeterminacy [Bro91]. India [Alb81, Bro72, Eme53, Eme54, Gor88, Ing65, LBWD86, Mor60, Mur74, Pad77, Sar78]. Indian [HC88, Lew56a, Bar50, CS53, Egg65, Fen56, Gre60b, GBN77, Jen66, Klu59, Kra58, Lin52b, Pin78, Roc93a, Sny53, Sny58, Ste52, Voe50, Voe52, VH53, Wal51c, Wit52]. Indianism [Her79]. Indians [Ald50b, Ald50c, DS56, EJ56, Fen53, Fre62, Gou66, Gus56, Han63, Hau75, HC88, Hei51, Jen68, Moc68, Par77, Sny57, Swa52, Tol63, Wal52, Was63]. Indicators [Bri75]. Individual [Mag99]. Indonesia [Bar52b]. Indoors [Nor73]. Induced [RD73]. Industrial [Bar54, Cha86, Dav51, Hea51, Hou90, Jef58, Mor60, Moy92, Spi71, Sta91]. Industry [Dav81, Gom85, Sol71]. Inedible [Sta51]. Inf [LL74]. Infancy [HHDA66]. Infant [HZ58]. Infanticide [Row91]. Infectious [Sch63]. Inference [DeV55]. Infinite [HC84]. Influence [Boa51, Irv59, Mag85, Mik58, Mur61a, Nic61a, Sch57c, Sta55]. Influences [Bev67, Esp92]. Influenza [Wol82]. Information [Ano50s, Ano51s, Ano52s, Ano53s, Ano54s, Ano55t, Ano56u, Ano57s, Ano58s, Ano59s, Ano60t, Ano61s, Ano62t, Ano63t, Ano64s, Ano65w, Ano66s, Ano67s, Ano68s, Ano69s, Ano70s, Ano71s, Ano72t, Ano73s, Ano74s, Ano75s, Ano76s, Ano77s, Ano78s, Ano79t, Ano79u, Ano80s, Ano81s, Ano82s, Ano83s, Ano84m, Ano85m, Ano86m, Ano87m, Ano88n, Ano89l, Ano90i, Ano91j, Ano92l, Ano93j, Ano94i, Ano95j, Ano96m, Ano97n, Ano98n, Ano99m, Büh70, Gom95, McC91, Sho93]. Infrared [Neu86]. Ingenhousz [CBA97]. Ingenious [Waa74]. Inheritance [Col65, Rob64]. Initiation [Bra81]. Initiatives [RT99]. Inner [Sul68, Wil92]. Inner-City [Wil92]. Innocent [Wri72]. Innovation [Cur61]. Inquiry [Sou52]. Inquisitorial [Ted87]. Insane [Sto99a]. Insects [Gre60c, Reh54, Wil79]. Insertion [Smi79a]. Insoluble [Bar87]. Inspiration [Maz55]. Institute [Mos55b, Bry88]. Institution [For80, Wol94, Car57b]. Institutional [Hea71, vD77]. Institutions [Bla57]. Instruction [Fre93]. Instructive [Gre87]. Instructor [Lin52a]. Instrument [Bry88, Bur52, Zim74]. Instruments [CS53]. Insulin [Sta50a]. Integrity [Fen71]. Intellectual [Alb72, Bou63, Bur75, HH51, LC62, Per53, Tay54]. Intellectuals [Bri55]. Intellectuelle [Cai60]. Intelligence [Fri62, Gou56]. Intelligibility [VH51]. Intentions [Kam87]. Interaction [Cot71, HK77, Rap60, Sch56a]. Intercommunal [vD86]. Interest [Chr66, Gre60b, Smi79b]. Interesting [Rid52]. Interests [CBA97, Ker82]. Interference [Mos50b]. Interior [MD64, Smi79a]. Intermediary [Opp82]. Internal [Joh54]. International [Ber57, Cha51, Cla78, Dub79, Kro57, Kuz74, May63b, Pou97, Roo72, Sch73, Sha70, Wil67, Woo53, Wri51, Ber57, Mil76, Viv74]. Internationale [Cai60]. Interpersonal [Cot71]. Interplay [Wei93]. Interpretation [Chi58]. Interpretations [Whi70a]. Interpreting [Hou90]. Interrupt [Spi83].

Interrupted [Cop56a, Jon99a]. Interruption [Det87]. Interstate [Goo53]. Interventions [Ham92]. Interview [BB84]. Intimate [Rob53]. Intravenous [RD73]. Introduction [Bar90a, Bel64b, Coh52b, Egg69, Gol92a, Gor88, Hes91, Kat85, Li55, Mou91, Nic61a, Pie95, Pie99, Smi86, Stu70]. Introductory [Bar52b, Gal74, Has79, Her55, Pie68, Pie80]. Intuitions [Wil96a]. Invade [Lan72]. Invented [Ros54]. Invention [Nic64]. Inventors [Pra68]. Inventory [Mil64a, Sch57a]. Invest [Ede97]. Investigation [KHO52]. Investigations [Bra72a, Kid61]. Invisible [Fri91a, Fri97]. Invoke [Per96]. Involuntary [Ash78]. Ion [Kar99, Key92]. Ionian [Con93]. Ionic [Gil82]. Ippolito [Jar72]. Ipsos [Gri82]. Iran [Ang52, Coo52b, Dup52]. Iraq [Ste77]. Ireland [Lub78]. Iridescent [GBF60]. Irish [Kea76, Mid85, Mor74a, Wok89]. Irishman [Car89a, Car89b]. Iron [PB74, SESB65, Luk81]. Ironworking [CWMM79]. Iroquoian [Pen94]. Iroquois [Boy73, CS53, Coo52a, Fen51, Fen53, Ran53, Sny61, Wit51]. Isidor [Eds85]. Islam [Böw97]. Islamic [Ett79, Gib63]. Island [Fra83, HK62, HR55, Mur62]. Islands [Kea76, Smi60]. Isles [Rau73]. Isolating [Ste63a]. Isolation [May59]. Issue [Hor61, Zak91]. Issues [Bau71, Dav81, Gid50, Lub78, Lud90, Taf62]. Istis [Dum54]. Istorii [Nic61c]. Italian [Lav75, Mae80, Set56]. Italy [Hir53, Ker82, Lom97, Pac50]. Itch [SC84]. Itinerant [Was73]. Itself [For98, Han76, Wil51b]. IV [Ste63b]. Ivan [CV92].

J [Bow93, Bur75, dK95, Kel92, Lei98, Sch95b, Wol95a]. J. [Him99, Kah55a, Rob57a, She83, Str97, Wil61b, Woo98]. J.-H. [She83]. Jāt [Lew56a]. Jackson [Jur66, Moo99, Ald50a, Sny57]. Jacksonian [Mil59]. Jacob [Gar75]. Jacobean [Lew94]. Jacobsen [Sjö99]. Jacques [Ald51]. Jaffa [May84]. Jamaica [Sim55]. James [FK96, And99, Bur60a, Cle98, Fos71, Fre96, Gob70, McK73, Mil82b, Moy97, Oak98]. Jamestown [Wer57]. Jan [Tob97, vdH94]. Jane [Lin51, McP98b, Sel55]. Jansenism [Mil87]. January [Ahr99, And99, Bau94, Bec94, Bev93, Bon94, Cla91, Con92b, Cor99, Dys99, Eis97, For99, FC92, Gin92, Gre94, Güt92, dK95, KH92, Kel92, Kno91, Lan98a, Lat95, Lin94a, MV94, Neu98, Pat94a, RB96, Rod92, Roy96, Shu98, Spr98, Sto99b, Str91, Wat95, Wei97, Zan92, Sel56]. Janus [Bak84]. Japan [Bro62b, Con56, Fle66, Miz79]. Japanese [Kim62, Kub59, Tho50a]. Jayne [Adr88, Gom95, Hea88, Ano79m, Blu97, Cra91, Dai95a, Gol92b, Key79, McC91]. Jealousy [Gun93]. Jean [Hoi75, Jac69, Lhe93, Olm60, Pau70, VD53, Wri97]. Jean-Baptiste [Hoi75, Wri97]. Jean-Philippe [Pau70]. Jednota [Nem68b]. Jefferson [D’E73, Bai93, Bed64, Boy58, But50, Han54, Jac80, Mat96, Par93, Pet61, Ple67, Ric51a, Ver93, dT59]. Jeffries [Mac65b]. Jehoiachin [Fel95]. Jenner [Wel74]. Jeopardy [Bar60]. Jeremiad [Ega78]. Jeremiah [Rob50]. Jerne [Mel97]. Jerome [Gil94, Sax96]. Jersey [Cop53, RH58]. Jerusalem [Con58, May78, May89]. Jewish [Kie80, Ros89]. Jews [Fin82]. Jigsaws [Spi84]. Johan [Wil94b]. Johann [Sch97]. Johannes [BR79]. John [But74, CR52, CS54, Lan59, And94, Ano63m, Ash81, Bou84, Boy73, Can61, Chr84, Coh93, CSL50, Fox68, Fry89, Gol97, GS94, Han76, Hom81, Kai70, Key99, Kin76, Kno63, Law92, Llo50, May84, Mor74b, Mur61b, Mus69, Reh54, Ric50b, Rob53, Rob91a, Rob91b, Shu98, Smi86, Wal52]. Johnson [Sno58]. Joining [Car81]. Joint [Ano92k]. Jones [Gel95b, Tra98, Can77]. Jonesport [Dor57]. Joseph [Bec94, Bel51, Boy73, Bra70, Cla91, Con50, DG73, Kel92, McE84, Rau73, Rob62, vH81, O’G78]. Josephus [Fel95]. Josiah [Kin76]. Josse [Mil87]. Journal [DS56, Sny57, Hor95]. Journalist [dM56]. Journals [Ehr90, McD59, Mou83, Rau73]. Journey [Haa50b]. Journeyman [RH58]. Journeymen [Wer81]. Jr [Bai97, Ede97, GHKW95, Jef99, dK95, Kno91]. Jr. [Bur75, Fos71]. Juan [Ben79, Par77]. Judge [Gun96, Smi86]. Judges [Wol94]. Judicial [Blu88, Fra54, Row91]. Judiciary [Ada84]. Julian [Kae64]. Julius [Joh96]. July [Ano98i, Bai97, Bat96, Bev93, Bir93b, Dar99, Dem94, Eis97, EoL99, Fry96, Gel95a, Gla93, Gol94, GHKW95, Got91, Gre98c, Güt92, Hec94, Hei98, Kam71, Key99, Lan98b, Lan98c, Mil95, Neu98, Nie92, Oak95, Rob91b, Set92, Soz92, TG91, Tho91, Vry99, Wil98, Wol95a, dCTdM92, Rau73]. June [Arn92, Bau94, Bow91b, Fed94, FK96, Gil95, Gli94, Hei98, Jef94, Jef99, Joh96, Kit95, Kno91, Mac97, May97, McC93, O’C95, Pow95, Roy96, Sch95b, Sid99, Sjö99, Ste99a, Tob91, Tob97, Wil94b, WE98, Wol95a]. Junk [Sin94]. Jurisdiction [Dow58, McK94]. Jurisdictionalism [Mil87]. Jurisdictions [Kim77]. Jurisprudence [Sto87]. Jurists [Mae80]. Just [Goo76, Rai86]. Justice [Tay68]. Justinian [Lad75]. Juvenal [Kai70].

Kāmya [Dum69]. Kānda [Dum59, Dum60, Dum61, Dum63a, Dum63b, Dum64, Dum51, Dum54, Dum62, Dum65, Dum69]. Kabira [Smi60]. Kafka [Beg97]. Kahkwas [Pen94]. Kaj [Mel97]. Kakouagoga [Pen94]. Kaminaljuyu [Kid61]. Karakoram [Jet61a]. Karl [Nic61c, Opp90]. Karyolysus [Leh59]. Kashmir [Jet61a]. Kathas [Dum51]. Katherine [Ste99a]. Kaukilī [Dum65]. Kaukilī-Sautrāmanī [Dum65]. Kautilya [Spe69]. Keats [Kro63]. Keckermann [Fre97]. Kemp [Hop61]. Kenneth [Lat95, May97, Par96, Sto99b]. Kepler [BR79, Eas82, Gin73]. Keresan [Hoe68]. Key [For96]. Killer [But74]. Kinds [Dum51, Pal87, Voe50]. King [Coh93, May89, Por96, Say71, Bev89]. Kingdom [Pal61]. Kingsley [Coa99]. Kingston [Sim55]. Kinship [Bur75]. Kinzel [Gol94]. Kirk [Ste99b]. Kissufim [RGSS72]. Kita [HR55]. Klamath [Spe56]. Knew [Wal51c]. Knight [Eri67]. Know [Fuc96]. Knowledge [Gre93b, Het75, Kle89b, Len66, Mas92, Mün59, New61, Oeh74, Orl71, Pie79, Rei75, Rob89, Wei60, Tho90]. Known [Coh50]. Knox [VK70]. Kohoutek [Whi76]. Konrad [Wal99]. Korea [Kim62]. Korean [Con56]. Kourion [Las75, Sti61]. Kramer [Moo99, Sjö94]. Krieble [Set92]. Krieger [Sch99c]. Kubler [Wil98]. Kuhn [Hei98, Bra85]. Kuhnian [Bra85]. Kurt [Güt92, Kit95]. Kuttner [Tie98].

L [Gil95, Win83]. LaBarre [Adr88, Hea88]. Labeling [Wes78]. Labor [Sha76, Taf62, Tay72]. Laboratory [Arr99, Hol90, Sok72]. Laboring [Smi89]. Labors [Kin76]. Ladner [Wol95b]. Lafayette [Lop64]. Lake [Too89]. Lamentation [Ega78]. Lamps [Bro52]. Lancaster [Sch98]. Land [Kar60, Pot75, Rom56, Sum60, Woo66a]. Land-Locked [Sum60]. Landolt [Kah55a]. Lands [Jam54, Tho66]. Landscape [Law92]. Language [Bak77, CH81, Gre93a, Hoe90, Kli73, Pye78a, Ske75, Smi79b, Swi98, Wei59, Yng60, vH96, vF56]. Languages [Lin52b, Smi79b, Smi83, Tay63, Voe50, VH51, Voe52, LH85]. Lao [Hau67]. Lao-Tze [Hau67]. Laplace [Gil72]. Lapse [Boy58]. Laski [Bur75]. Last [Neu51, Scr91, Top76, Wil61b]. Lasting [Luc70]. Late [Com56, Cur63, Per58, Smi89, Ald76, Ben79, Boy74, Her82, Heu66, Kno76, Ler77, Mur74, Pet73, SC89]. Late-Eighteenth-Century [SC89]. Later [Car89a, Ede97, Tau76, Dav91]. Latin [Cra58, Cun57, Fre62]. Latrobe [SC84]. Laurent [DK54]. Lausanne [Gre54]. Lavoisier [DK54]. Law [Bod63, Cla50, Dri53, Fel87, Gel68, Goo53, Gre98a, Hoe68, Hur60, Joh74, Kim93, Lad75, Lev88, Moe58, Pol85, Roc93a, Smi86, Spe63, Wri51, Kam95, Mil59]. Lawrence [Rav91]. Laws [Ber67, Mil67, Cow60]. Lawyers [Sym72]. Lay [Wol94]. Layers [Gri73]. Lead [Pat70a]. Leader [Ros75, Wes98b]. Leadership [Bes79, Tru59]. League [Fin82, Fos71, Bar64b, Dub79, Hew70, Rot77]. Lear [Bow90, Bev89]. Learned [Gun96, Lan76]. Learning [Bak94, Har71b, Hyd67, Kno76, Mil67, Mos58, Rei68, Rho98, Rob67b, Str64]. Leaves [Esa67]. Leconte [Reh54]. Lecture [Hea88, Key79, Adr88, Ano72m, Ano79m, Ber67, Blu97, Bok95, Cla59, Cor65, Cra91, Dai95a, Dry60, Dub63, Ede97, Fol58, Ful61, Gal66, Gol92b, Gom95, Gra98, Hux86, Lat70, Lin50, Lur72, McC91, Per69, Rob79, Sca87, Set69, Smi72, Sta57, Str64, dM56]. Lee [Pet63, RB96]. Lefebvre [Chi50]. Left [Por92]. Legacy [Ber98, Cho98]. Legal [Abr62, Bod63, CR54, Fre71, Par83, Pot62, KS84]. Legality [Gel68, Kie80]. Legend [Gar61, Pac55]. Legendary [Cha85]. Leges [Lad75]. Legged [Leh59]. Legionella [Fra86]. Legislative [Mar65]. Lenape [Bec89]. Length [BG91, Jon55, Sza72]. Lenher [Jef94]. Lenses [Tei88]. Leo [Set69, DM52, Gin92]. Léon [Cas92]. Leonard [Sch99c]. Leonardo [Win67]. Leopold [Gil86]. Less [Kuz75]. Lesseps [Lan50]. Lessing [Bar83]. Lesson [Whe81]. Let [Wat78]. Lethal [DSR50]. Letter [Ald50c, Kah56, Lin51, RR92a, RC67, Tod92]. Letters [Abr72a, Bel56, Can61, Coh56, Con51, Cop50b, Fah71, KRF55, Kah55b, Lin51, Tre74]. Levant [How59]. Level [LB79, Mor96, Sim93]. Levels [Gol91, Rap60]. Leviathan [PSP55]. Levin [Bat96]. Lewis [Bau94, DeV55, Bea96, DL57, Egg65, Ehr90, Jac80, KH92, Mou83, Mur61b, Spo81, Wea54]. Lexical [Mil72]. Lexico [Swa52]. Lexicography [Wol88]. Liability [Nic67]. Libanius [Dow59]. Libel [Fre68]. Liberal [Gre98a, Hau67, Ros75]. Liberalism [Boy74, Hea95, Par93]. Liberalization [Tay99]. Liberata [Lee81]. Liberty [Ada93, Rog51]. Librarian [Bel75]. Libraries [Haa96].

Library [DeM84, Eis54a, Gre60a, Mor74b, Set60, Shr60b, Alb60, Ano56m, Ano60m, Ano65v, Bel59, Fre62, GS94, Klu59, LL74, Lin53, Pet51, Shr60a]. Liebig [Fru88]. Liebling [Kam95]. Lieutenant [Bor65]. Life [Bro51, Dar59b, Fen70, Gou97, Her79, Kam87, Kan89, Koo90, Lop77, Men69a, OC81, Rob76, Sch99b, Sta57, WM70, Wri72, dD98]. Lifelong [Rob83a]. Lifework [Por85]. Light [Abr72b, Eks99, Mün59, Wal93]. Lightning [Coh52b]. Lights [Mer53]. Likenesses [Ver64]. Limbs [Sim58]. Limit [Pan89]. Limiting [Cla50]. Limits [D’E73, Mad62, Mor80b, Sch98, Sim77, Wig50]. Lind [Dau56]. Lindemann [Sat82]. Line [Cop56a]. Lines [Sim77]. Lingelbach [Rob57b]. Linguistic [Eme54, Fen56, Gre80, Har64, MMK^{+}86, Swi98, GS94]. Linguistics [Gre60b, Klu59, Swi98]. Links [Eis54b]. Linnaean [Rob60]. Linnaeus [Sta60]. Linus [Dun97]. Lipman [BK96]. Lisbon [Vog73]. Literary [Mil78, Pot62]. Literature [Ein57, Irv59, Kra63, Kra80, Mur56]. Litigation [Fra54]. Little [Coh50, Sel55]. Little-Known [Coh50]. Littlepage [DL57]. Lived [Kam94]. Lives [Ler77, Smi89, Wil75a]. Living [Kam87]. Llanquihue [Heu66]. Local [Cop55]. Locating [Rid52]. Locked [Sum60]. Locomotion [Gre77]. Lodge [Fos71, Hew70]. Logan [Tol51a]. Loge [Han53b]. Logic [Bak77, Dub63]. Logical [Ong76]. London [Bol62, Cor50, Dur87]. Long [Kam94, Kuz58, Lik99, Luc70, Oli57, Oli69, Ros81, Sta60, Stu70, Dil67, Mur62]. Long-Lasting [Luc70]. Long-Lived [Kam94]. Long-Term [Lik99, Stu70]. Longhouse [CS53]. Look [Gra85, Güt84, Lop85, RC99a, Say71]. Looking [Hij80, MS98]. Lookout [Ven99]. Looks [Fah71, Fra60a, Ros90]. Lord [Gle60, Ano97m, Arn92, Bel95, Dai95b, Ros75, Spe69, Wes98b]. Lords [Fos82]. Lorentz [Ive51]. Lorenz [Wal99]. Lorenzo [KJF52, Ser91]. Lost [Con69, Kah56, Plu77, Fry76, Ste76]. Loudoun [McG99]. Louis [Ree86, Spr98]. Love [Hof83, Maz55]. Loved [Con69]. Lower [Bec89, Boy74]. Lowland [Web76]. Lubbock [Mur61b]. Lubeck [Rot77]. Lucca [Dav74]. Lucien [Bro78]. Ludwig [Eds91]. Luigi [Li55]. Lunar [Dum54, Gol71, GCM94, Mul72, NPZ81, Sjo72]. Lund [LF52]. Lungfish [FGWP92]. Lungs [Gib65]. Luria [Wat99]. Luther [GS80, Lon59]. Lwoff [Ste97]. Lyell [Con51, Can61, RC67, Wil80]. Lyon [Laf75].

M [Dul95, DM55, Kit96]. M. [Bat86]. Macaca [HvW53, HvW61]. Macbeth [Fry98]. Macedon [Bor87]. Machiavellianism [Jor71]. Machine [ER90]. Machines [Mor64a, Wig69, Win67]. Machlup [Jef65]. Mackenzie [Ric50a]. MacKnight [McC66]. Macromolecular [Sch56a]. Madagascar [Gou65]. Made [Ano60m, Ric51b, Sui64, Hea88]. Madison [Mar65, dT59]. Maestlin [Gra73]. Magic [Han68, Nic61b, Ted87, Wal93, Wit99]. Magnetic [Cha83, Hou50, Rob89, Sli98, Voe50]. Magnetically [Bea57]. Magnetism [Rad93]. Magneto [Gil82]. Magneto-Ionic [Gil82]. Mahādeva [NP67]. Mahurjhari [Luk81]. Maine [Dor57]. Mainstream [Fry83, RR92b]. Maintenance [Kle71, Pat76a]. Maitrāvaruna [Dum62]. Major [Cho98]. Make [Reu97]. Makes [Dem54]. Making [Met91, New71, Pol85, Ric53, Sch71, Swe96]. Malay [Bar52b]. Male [Wil92]. Malecite [Sza72]. Malinowski [Bur83, Pay81]. Malraux [Bro70]. Malwa [Jac75]. Mamiya [HR55]. Mammalian [SH57]. Mammals [Sim78, Sim84]. Mammoth [Yoc92]. Man [Det81, Gre92, Har73, Her55, Hor95, Mur51, RD73, Sau62, Sch50, Shu71, SC89, Ski64, Sui64, Car58, Coh95, How59, Kee70, Lat70, Son65]. Man-Made [Sui64]. Manage [Wil51b]. Managed [Hor58]. Management [Ben79, McK94, Mot75, Plo78]. Managing [Str88]. Mandates [Sto85]. Mangelsdorf [TG91]. Maniculatus [She61]. Manipulation [Ker82]. Mankind [Mur61b]. Mansions [Dum54]. Manuals [Ted87]. Manufactory [Bro89]. Manufacture [Bry88, Rob72]. Manuscript [Fen53, Fre62, LL74, Mea54, Pra68, Sta60, Wad57, Sny58]. Manuscripts [Alb60, Ber54, Boe58, CR52, Ewa59, Kri76, Pac52, Sny53, Tho60, Ver58, Wea54]. Many [Sim77]. Map [Eis63, Kit73, LL74, Spi93]. Map-Projection [Eis63]. Mappae [Car81]. Mapping [Cap74, McK97]. Maps [Por74, Spi83]. Marble [Ver58, Ver64]. Marc [Faj74]. Marc-Michel [Faj74]. March [And94, Arn92, Bel95, Bor93, Bow91a, Boy98, Bra92, CV92, Che94, Dar99, EoL99, EC94, Fir91, Fis94, Fri91b, Gel95b, Gra93b, Har93, Hec94, Kau95, Kay98, Key99, Kit95, Lai98, Lat95, Led94, McP98b, Mel98, Moy97, Nie92, Par96, Ram99, Ros98, Sal98, Sch95b, Sel96, Spr98, Tho91, Tie98, Wat95, dCTdM92]. Marcus [Dem94]. Margaret [Kun86, Met93]. Margenau [Din55]. Marginality [Smi88]. Maria [OC81]. Marie [Con92b]. Marie-Thérèse [Con92b]. Marine [Ric50a]. Mariner [Pic66]. Marion [McP98b]. Market [Sil95]. Markets [Tay72]. Marking [Wes78]. Marriage [Jor71, Kra63]. Marriages [Dav91]. Mars [Pic66, Sch99b]. Marshall [Plu77]. Martin [Hof91, Ost99, Ray80, Dar79]. Marx [Nic61c]. Marxism [Bau79]. Mary [Bel56, Bla62, Pat74, Tho71, Tra98]. Maryland [CR54]. Mascon [Mul72]. Maskelyne [Llo50]. Mason [VD94, Bry88, Cop51, CR52, Llo50]. Massachusetts [Hom81]. Massanutten [Cha80]. Massive [Fow65]. Master [Han53b]. Mastodon [ST54]. Materials [Fen53]. Maternal [HHDA66]. Mathematical [Coo84, Wea54]. Mathematics [Ati93b, Bir93a, Mac73, New61, SC84]. Mathematicus [Sti92]. Matrilineal [Goo76]. Matrimony [Jor71]. Mattei [Bay69]. Matter [Ano50g, Ano50h, Ano50i, Ano50j, Ano50k, Ano50l, Ano50m, Ano50n, Ano50o, Ano50p, Ano50q, Ano50r, Ano51g, Ano51h, Ano51i, Ano51j, Ano51k, Ano51l, Ano51m, Ano51n, Ano51o, Ano51p, Ano51q, Ano51r, Ano52g, Ano52h, Ano52i, Ano52j, Ano52k, Ano52l, Ano52m, Ano52n, Ano52o, Ano52p, Ano52q, Ano52r, Ano53g, Ano53h, Ano53i, Ano53j, Ano53k, Ano53l, Ano53m, Ano53n, Ano53o, Ano53p, Ano53q, Ano53r, Ano54g, Ano54h, Ano54i, Ano54j, Ano54k, Ano54l, Ano54m, Ano54n, Ano54o, Ano54p, Ano54q, Ano54r, Ano55h, Ano55i, Ano55j, Ano55k, Ano55l, Ano55m, Ano55n, Ano55o, Ano55p, Ano55q, Ano55r, Ano55s, Ano56g, Ano56h, Ano56i, Ano56j, Ano56k, Ano56l, Ano56n, Ano56o, Ano56p, Ano56q, Ano56r, Ano56s, Ano57g, Ano57h, Ano57i, Ano57j].

Matter [Ano57k, Ano57l, Ano57m, Ano57n, Ano57o, Ano57p, Ano57q, Ano57r, Ano58g, Ano58h, Ano58i, Ano58j, Ano58k, Ano58l, Ano58m, Ano58n, Ano58o, Ano58p, Ano58q, Ano58r, Ano59g, Ano59h, Ano59i, Ano59j, Ano59k, Ano59l, Ano59m, Ano59n, Ano59o, Ano59p, Ano59q, Ano59r, Ano60g, Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60n, Ano60o, Ano60p, Ano60q, Ano60r, Ano60s, Ano61g, Ano61h, Ano61i, Ano61j, Ano61k, Ano61l, Ano61m, Ano61n, Ano61o, Ano61p, Ano61q, Ano61r, Ano62g, Ano62h, Ano62i, Ano62j, Ano62k, Ano62l, Ano62n, Ano62o, Ano62p, Ano62q, Ano62r, Ano62s, Ano63g, Ano63h, Ano63i, Ano63j, Ano63k, Ano63l, Ano63n, Ano63o, Ano63p, Ano63q, Ano63r, Ano63s, Ano64g, Ano64h, Ano64i, Ano64j, Ano64k, Ano64l, Ano64m, Ano64n]. Matter [Ano64o, Ano64p, Ano64q, Ano64r, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano65o, Ano65p, Ano65q, Ano65r, Ano65s, Ano65t, Ano65u, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano66m, Ano66n, Ano66o, Ano66p, Ano66q, Ano66r, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano67m, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano68n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano69o, Ano69p, Ano69q, Ano69r, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano70n, Ano70o, Ano70p, Ano70q, Ano70r, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano71o, Ano71p]. Matter [Ano71q, Ano71r, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72n, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano79g, Ano79h]. Matter [Ano79i, Ano79j, Ano79k, Ano79l, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j]. Matter [Ano87k, Ano87l, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano89e, Ano89f, Ano89h, Ano89i, Ano89j, Ano89k, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98e, Ano98f, Ano98g, Ano98h, Ano98j, Ano98k, Ano98l, Ano98m, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ree93, Rub88a, Rub88b]. Matthew [Gal74, Rai86]. Mattison [Sea93]. Matveevich [CV92]. Maurolico [Ros57]. Maury [Bei51]. Mawr [McP98a]. Maximization [Ada78]. May [BK96, Bat96, Bel92, Boo97, Bow93, Che94, Coh93, Cra93, Dai95b, Dre94, Eag98, Eds91, Fen92, FK96, Fre96, Gil94, Gol97, Gra93b, Gre94, Gri95, Him99, Hiż94, HC84, Joh96, LB95, Lit96, Mau99, Moo99, Moy97, Ost99, Pat94a, Rob91b, Sax96, Sel96, Sjö99, Sta98, Ste97, Swe93, Wil94a, Bol62]. Maya [Pro84, Web76, Wil90]. Mayall [Gre94]. Maynard [Con67, Lan76]. Mayor [RR92a, Tod92]. Mays [ES56a, ES56b]. McClintock [Fed94]. McGeorge [Kay98]. McMillan [Sea93]. Mcshane [FK96]. Me [Rei86, SC84]. Meaning [Sch98]. Meanings [Bro91]. Means [Kel57]. Measles [Wol82]. Measurement [Koy53, Sel53]. Measurements [Sjo72]. Measures [War63]. Measuring [Wri50b]. Meat [Hil69, Shi70]. Mecca [Daw60]. Mechanical [Gre60c, Yng64]. Mechanics [Els68, Tri80]. Mechanism [Bor64, Noy51, Sch58]. Mechanisms [Ste63a, Ste74]. Mechanistic [Lok78]. Mecom [Lin51, Sel55]. Medal [Lin50, Fox68, Rol56]. Medallic [Zig57]. Medawar [Bre92]. Mediated [HH76]. Medical [Bla57, Blu97, Bur73, Cor65, D’E66, Gor88, Han68, Kee63, Mah91, Zwo61]. Medicine [Bau71, For71, Lud90, Pep50, Rob89, Sny83]. Medico [Bol62]. Medico-Psychological [Bol62]. Medieval [Cla64, Goi71, Gra85, Kit73, LL74, Roa70, Rus62, Set60, Aul82, Kri76, Tho60]. Mediocrity [Wil61a]. Mediterranean [Bis79, LVF98]. Meet [Mel79]. Meeting [Hol70, Sny58]. Megalith [Luk81]. Megalith-Builders [Luk81]. Megalonyx [Boy58]. Megatherium [Boy58]. Meiggs [Bow91b]. Meiji [Con56]. Meiosis [For96]. Mello [dCTdM92]. Member [WD96]. Members [DM65, Wil51a]. Memoir [QC53, Sti77, DL57]. Memorial [Ano72m, Ber67, Cla59, Cor65, Dry60, Dub63, Ede97, Fol58, Ful61, Gal66, Hea88, Lat70, Lur72, Per69, Rob79, Sca87, Sch56b, Set69, Smi72, Sta57, Str64, dM56]. Memory [Boy58, Fle67, Hyd67, Lof98, McG67, Mil72, Rob67b, RC99b, Sch98, Voe50]. Men [Bie59, Bor50, May89, Wal51c]. Mencken [Win83]. Mendel [Dun65, Ste65]. Mendeleev [Raw78]. Mendelians [Stu65]. Mendelism [Dob65]. Mental [Cro70, KS91, Mal55, Pic54, Sat82]. Mentioned [Ewa59]. Merchant [Dye76]. Merchants [Oak77]. Mercury [Coh50, Neu51]. Meridian [Cop56a]. Meritt [Tho91]. Meriwether [Jac80]. Merle [Boy98]. Mesmerism [Nyg70]. Mesolithic [Cza76, Mov53]. Meson [Hal63]. Mesopotamian [Ada78, Fin63, Jac63, Kil71]. Message [Nic56]. Messenger [McC66]. Metabolism [Bur84, HPSH50]. Metal [Car81, Sol71]. Metal-Joining [Car81]. Metallurgical [Ano72m, Smi72]. Metallurgy [Eas65]. Metals [HPSH50]. Metalworking [Eas65]. Metamorphosis [Ken93]. Metaphysics [Ken93]. Metascience [Abr72b]. Meteor [Oli50, Oli57, Oli69]. Meter [Cha87, Ste69]. Meters [Sch69]. Method [Neu51, Rid52]. Methodological [Lud90]. Methodology [Mur55]. Methods [Dum51, Her55, Neu63, VH51]. Metric [Bra72b]. Metrology [Len65]. Metropolitan [Abr62, Har62]. Metternichian [Wer81]. Mexican [BC69]. Mexico [And60, Ben79, Gri72a, MO59, Wei93]. Meyer [May97, Ros98, Zan92]. Mice [Fle67]. Michael [Gra73, McC93]. Michaux [Chi57a, RH58, Sch56b, Sch57c]. Michel [Ste97, Faj74, Roa70]. Michelet [Haa50a, Haa50b]. Michigan [Hea95]. Microorganisms [HPSH50]. Microwave [Rob83b]. Mid [McG97, Fen51, Wit51]. Mid-Century [Fen51, Wit51]. Mid-Weekend [McG97]. Middle [Hin64, Men89, Bau59, Byn98, Coc64, Hin64, Neu63, O’G78, Shr64, Yoc96]. Midwestern [Wal56]. Might [McG96]. Mighty [Gre92]. Migration [Esp92, Gri73, Sod89]. Milbert [Rob57a]. Mile [See61]. Milieu [Nis82]. Militarism [Wil75b]. Military [Dur87, For80, Sec65, Wie67, vD86]. Milky [Bok52]. Millenarianism [Wei99]. Millennium [Ber98]. Millerand [Far82]. Million [Ric70]. Millions [Rob76]. Millipedes [CT69]. Mills [Ram99]. Milo [Pac82]. Milton [KWB97, Fry76, Fry89, Ong76, Pat76b, Vry99]. Mimesis [Mor80a]. Mind [Bun91a, Rid52, Sim93, Ste98]. Mind-Body [Bun91a]. Minds [Gra96]. Miner [Gri95]. Mineralogy [Gre69]. Minerals [Yod82]. Minerva [And76]. Ming [Pet73]. Mining [Wal83]. Minis [Bel75]. Minister [Mit58]. Mink [End52]. Minoan [Cas69, VW78]. Minor [Esa67, Ver58, Bun91b]. Mint [Mer75]. Minto [Eis50]. Miracle [Ada87a]. Misallocation [Bou63]. Miscellany [Sca76]. Misgivings [CC95]. Missing [CB96, Eis54b]. Mission [McC84]. Mission-Oriented [McC84]. Missionary [Fen56]. Missouri [Ric50b]. Mistress [Boo78]. Mistresses [Ler77]. Misunderstanding [Han53a]. Mitchel [McC66]. Mitchell [Gri95]. Mitchill [Mur62]. Mitigations [CC95]. Mixed [Cad51, Kut99]. Mobile [Wil68]. Mobility [Min73]. Mobs [Sch55]. Mocking [Mar77]. Mode [Lap83]. Model [Coc70, Gin73, Koh74, Rob67b, Yng60]. Models [Går99, Gre80, Hil64, Mos58]. Moderator [Rho88]. Modern [Adr56, Anc74, Bak84, Bro85, Byn98, Cha86, CBL61, Cla63, Dur67, Fry98, Jor61, Keo98, Mar61, May95, Mur61a, Nem68b, Nic61a, Por74, RC99a, Set60, Spe51a, SS57, CS53]. Moderna [Nem80]. Modernization [Miz79]. Modernizing [Pay78]. Modification [Wol94]. Moerbeke [Cla82]. Mohave [Kro65]. Mohawk [Bar52a]. Molecular [EB79, Hyd67]. Molecule [Pau52]. Molecules [Eds67, Her93, Rad93, Wes78]. Monastery [Ati52, Hut52, Wei66]. Monceau [Ewa59]. Mondai [Con56]. Monedatge [Rus62]. Monetary [Fri72, Sch73]. Money [BB84, Bur52, Fri61, New71, Wil51b]. Mongolia [Hut52]. Monieux [Gol69a]. Monitor [Gre87]. Monitoring [Bri75]. Monkey [HvW53, HvW61]. Monkeys [CS52a, HZ58, HHDA66]. Monopoly [Par83]. Monster [Gun93]. Montaigne [For72]. Montalenti [Gla93]. Montgomery [Bor93]. Montréal [Pye78b]. Monumental [Ver64]. Moon [Dum57, Dum59, Dum60, Dum61, Enr75, Mac65a, Men69a, Ure71]. Moore [Gol92b, Soz92]. Moral [Keo98, Mad62, Tay54, Wil96a]. Moralist [Pey58]. Moravian [Wal51c]. Morgan [All66, Bel51, Egg65, Mur61b, Spo81]. Morley [Kid59]. Mormon [Dys78]. Moroccan [RT99]. Morphogenetic [Mil64b]. Morphology [Leo94, Sim59]. Morphou [VW78]. Morris [Hec94, Sch56b]. Mortality [Coa88, Wil51a]. Morton [Dem94]. Mosaic [Kit73, Wei66]. Moscow [Sch95a]. Mot [Cha85]. Mothering [HHDA66]. Motif [Lan72]. Motion [Neu62b]. Motions [Wri50b]. Motivation [Bus54]. Motives [Cad51, dT60]. Motyxia [CT69]. Mount [Ati52, Wei66]. Mountains [Ott89]. Mounting [Ede75]. Movement [App97, HC88, Lon57, May63b, Nem80]. Mr [Bel51]. Mr. [Lab56, Lew56b, McC98, Poe56, Yoc92]. Mrs. [But74]. Mu [Hal63]. Much [Orl71, Sil95]. Muhammad [RT99]. Muhlenberg [Ano62m, Mea78, Smi62]. Mulatta [HvW53, HvW61]. Multiple [RC99b, Smi79a]. Multiples [Mer61]. Multum [Fra83]. Mumford [KH92]. Mundi [Pac82]. Municipal [Laf75]. Munsee [Moc68]. Mura [KJF52]. Muralt [Eds91]. Muscle [Sta50a]. Muse [Chi53]. Museum [Sel80, Fra60a, Kra80, Ric51a, Sel80, Yoc92, Zio87]. Museums [Ada91]. Mushroom [Was58]. Music [Chi56, Con94, Dri52, Kil71, Roc97, Ros66]. Musica [Jac68]. Musical [Mik58, Win67, CS53]. Musician [Pau70]. Muslim [Ros64]. Mustela [End52]. Mutate [Dem54]. Mycenae [Ben53, Myl55, Ven99, Wac53]. Mycenaean [Cas69, VW78]. Mynors [Got91]. Myriopods [Reh54]. Mystery [Man83, Sim78]. Myth [Ray80, Sch98, Sto99a, Wil83, Wil56]. Myths [CC95, Mei66].

N [Leh60, Leh59]. N. [Bus97, DSR50]. Nagasaki [War77b]. Naksatras [Dum54]. Names [Coo52a]. Napoleon [Jor71, dBdS54]. Napoleonic [Shu58]. Narobanchin [Hut52]. Narrative [Bré70]. Nathaniel [Gel95a, Bon74, Ega78, Gle60]. Nation [Kel57, Kro58, Orl71, Pen94, Shy87]. National [Alp55, Ano89g, Dub77, Gra89, Ker82, Koh74, Lan89, LF52, Par57, Pet61, Rus57, Sch73, Sei90, Ske75, Sta50b, Bei51, Dup57, Lon57, Mel75]. Nationalism [Far82, McP99]. Nationalist [Kro65]. Nationality [Her79, Opp90]. Nations [Fos71, Har71a, Jo70, Kir56, Bar64b, Dub79, Fen53, Hew70, Kot51, Rub63]. Native [Bec89, LH85]. Natives [Fen56]. Natural [EG59, Kam78, Kar99, Ken91, Leo94, Ler59, Lev88, Lok78, May91, Pen50, Pet73, Rai92, Ric51a, VH51]. Naturalist [For62, Mil82b, Mur59]. Naturalization [Car89a, Car89b]. Nature [Bea61, Can60, Eks99, Fra54, Fri66, Gol71, Lik99, Mur51, New59, Shr55, Sta57, Ste92, Wat53, Han76, Sta60]. Nauta [Sch95b]. Naval [Plo78]. Navigating [Pic66]. Navigation [Dra60, Olm60, Wee54]. Navy [Plo78]. Neanderthal [Ste77]. Near [Spe51a, Goi71, Ada66]. Nebulae [SW50, Wri50b]. Necessary [Dai95a, Wei80, Wei86]. Need [Adr92, Ful61, Orl71, Wil61c, Yod82]. Needed [Sto55]. Needs [Bar55, Coh55, Mar55, Wal51a]. Neglected [Alb72, Sti77]. Negotiations [Ken58]. Negro [Elk63]. Negroes [vH81]. Nehemiah [Lef71]. Nemesis [Mae77]. Nemours [Hou90]. Neo [CWMM79]. Neo-Assyrian [CWMM79]. Neoplasia [Rho76]. Nepal [Kar60]. Neptune [Bra81]. Nerves [Wei69]. Nervous [Gra75, Pic54]. Nestling [GJ55]. Nestor [Ble57]. Nether [Kra50]. Netherland [HK77]. Netherlands [Gey62]. Nettie [OC81]. Neuf [Han53b]. Neugebauer [Swe93]. Neuron [Luc70]. Neurons [Key92]. Neuroscientists [Bun91a]. Neurospora [DSR50]. New-Moon [Dum57, Dum59, Dum60, Dum61]. Newbold [Shu98]. Newborn [Cha85, Coh87, Mel79]. Newcomb [Eis57]. Newer [Car57a, Vis72]. News [Obe79, Tay68]. Next [Ber98, Pat94b, Spi71]. Niagara [Pen94]. Nicaragua [Lan50]. Nicholas [Gre94]. Nichols [Bla62]. Nicolas [Bra70]. Nicole [Cla64]. Nicollet [Bra70]. Niels [Mel97]. Nier [Mau99]. Night [For76]. Nineteenth [Boy74, Dav91, Coh95, DS56, Kan89, Kli73, Pat74, Ske75, Spo81, Wal83, War85]. Nineteenth-Century [Boy74, Coh95, DS56, Kan89, Kli73, Pat74, Spo81]. Nisbet [Per99]. Nitrogen [Aul70, Bur84]. NMR [Sli98]. No [Jan89, Ano88m]. Noah [Sjö94]. Nobel [Cra90, Eks99]. Nolan [LB95]. Nomadic [Ekv74, Sal67]. Nomenclature [DK54]. Nominating [Mor69]. Nomura [Con70]. Non [Col72]. Non-Western [Col72]. Non-White [Col72]. Nonconformity [Sto55]. Nonindustrial [BP67]. Noninvasive [Cha83]. Nonrecognition [Mat96]. Nonviolent [Tol63]. Nordenfalk [Wil94b]. Normal [Gib65, GH58, Par76]. Normans [May89]. North [Gum55, Ric50a, RK82, Bar74, Cle54, Cle64, CSL50, Hei51, Kay68, Kim93, Lar93, Lew56a, Swa52, Ver64, VH53]. Northeastern [BT60, Rob57a]. Northern [Ben79, HR55, Smi71, TNN71]. Northrop [Rob91b, Bev93]. Northward [Cop56a]. Northwestern [And60, Gou66]. Norton [Boy73]. Norway [Ric68]. Nostrum [Bay69]. Notable [Lin51]. Note [Bur88, Lin52b, Pen57, Rob50, Shi70, Tho67]. Notebooks [Hol90]. Notes [Bar52b, Bel72, But74, Chi54, Chi59, Dav52, Lon58, Moe61, Neu62a, NV64, Phi53, Pot62, Tru59, Wil75b, Wil61b]. Novel [Blu82]. November [Abr91, Bai97, Bea96, Blo97, Boo97, Dys99, Eis98, Gal91, Gel95a, Gin92, Gin99, Gre98b, Gri95, Her95, Kel92, Kor97, Lan98a, Lew97, Mey98, Mil95, Oak95, Oak98, Set92, Sjö94, Sta98, Ste99b, Wal99, Wei97, WE98, Wri97, vdH94, Ano60m, Bor65]. Nubian [Wil67]. Nuclear [BHT86, Ben85, Car90a, Cha83, Rob89, Sim87, Sli98, Wei80, Wil61c]. Nucleus [Col65]. Number [KB80]. Numbers [Bec89]. Numerical [Sto59]. Numismatic [Jon63b]. Nutrition [Mur55]. Nutritional [Luk81]. Nutritive [Sta51]. Nuttall [Bei60].

O [Blo97, Fai92, Mau99]. Obedience [Kam71]. Obeying [Mos50b]. Objectives [Her55]. Objects [Ver58]. Oblation [Dum64]. Oblations [Dum54]. Obligation [Hea71, Ost95]. Obligations [Mac57]. Observational [Dic81]. Observations [Big54, Fra54, Gol88c, Gre93a, Hac97, Hit52, KJF52, Lee81, Pen54]. Observatories [Bel64a]. Observatory [Plo78]. Occupance [Jak68]. Occupation [Cor72, Shu53, Wie67]. Occupational [Boy74]. Occupations [Dub60]. Occurrence [Sim58]. Ocean [Spi83]. Oceanic [Bur93a]. Ochoa [Kor97]. October [Abr91, And94, And99, Ano97m, BK96, Blo97, Bon94, Bow91b, Bre92, Cla91, Coa91, Con99, Cor99, EC94, Fai92, Gol94, Got91, Hiż94, KWB97, KH92, Lec94, Led94, Lei98, Lit96, MV94, Mel97, Mey98, Mil98, Nat91, Ost98, Rau73, Rav91, Roc93b, Sax96, Sch99c, Sid99, Sme95, Ste99b, Str97, Wes98a, Wil98, Wol96, CR52]. Ocular [SH57]. Odes [Kro63]. Odyssey [Gar75]. Oeconomy [Sta60]. Official [Kat97, Lew54]. Oglethorpe [Con67]. Ohio [Jak68, Jam54]. Old [Ekv80, Fel87, Gar80, Lew94, Lin55b, Lop85, Mel66, Nor73, Ram69, Sin99, Wil75a, Eng68, Roa66]. Old-Age [Lin55b]. Older [Bar54, Car57a]. Oliver [Bur75, Sme95, Smi86]. Olmec [GB94]. One [Bar64a, Ehr90, Goo58, Kri84, Lop77]. Oneida [CH81]. Ones [Ekv80]. Oort [vdH94]. op [Faj74]. Opaque [Gri73]. Open [Ott86, Ver64]. Open-Range [Ott86]. Opening [Shr60b, Ste93b, Wei93, Wil72]. Operation [Ste63b, Tau76]. Operations [Wal83]. Opinion [Lub78]. Oppenheimer [Bac72, McP98b]. Opponents [May84, Rei68]. Opportunities [Gib81]. Opportunity [Cro70, Yod82]. Optical [Cha83, Whi58a]. Optics [War85]. Optimism [Wri50a]. Oracle [Bel50]. Oral [Lan84, Roc93a]. Oration [Dow59]. Orbits [Oli57]. Order [Cun57, Sha76, Wil70]. Oreas [She61]. Oregon [BT60, CSL50, Lan59]. Oresme [Cla64]. Oresteia [Ven99]. Organ [Fre71, Ram69]. Organic [New73]. Organization [CS52a, Eds67, Goo69, Kno93, May63b, Pen54, Sal67]. Organize [Zie82]. Organs [Leh71, Sta71]. Oriental [Sal50, Sim77]. Orientation [vF56]. Oriented [Gri73, McC84]. Origin [Bad68, Bar99, Daw60, Dri53, Dun73, Eng68, Gla60, Gol80, Goo76, Hac97, Kot80, Lan89, Löw91, May63b, Moe61, Mor69, Riv58, Sel65, Sou52, Sto99a, Swi98, Whi54, Wol91, dD98, Sta60]. Original [Fra60b, Haa50b, McK94]. Originality [Bak84]. Origins [Bry88, Chi54, Fry89, Oak77, She85, Whi70b]. Orkney [Rau73]. Orleans [Ewa83]. Ormsby [McC66]. Orr [Fir91]. Orthodoxy [Coh52a, Ler77, Zak91]. Orton [Mil82b]. Oscar [Bir93b]. Oscillators [Gre60c]. Osteogenesis [Hug69]. Ostraca [Neu62a, NV64]. Other [Con51, Gre80, Ste74, Tat54, Voe50, Win67]. Otherwise [Går99]. Otto [Wat95, Swe93]. Ottolenghe [vH81]. Ottoman [Toy55]. ould [Ber93]. Our [Bac77, Ber98, Dil68, Ede97, Fol58, Gre93b, Hea88, Pat94b, Pet99, Roc63, Ste89, Tru59, Was63, Wei69]. Outdoors [Nor73]. Outer [Hut52]. Outfitting [Bor65]. Outlook [Han76]. Outpost [Kea76]. Output [Cro73]. Outsider [RC99a]. Overcoat [Bro91]. Overview [Ber95, Saw92]. Ovum [SH57]. Own [Mel79]. Owned [Llo50].

P. [She61]. Pachyderms [Bed81]. Pacific [She85, Spo81]. Packard [Lai98].

Pagan [Fra75]. Page [Mer83]. Pages [Mac65b, Shi70]. Paid [Sny57]. Pain [Hil71]. Paine [For98, Rob83a]. Painter [Mar77]. Painting [Cor57, Fon71]. Paintings [Mar53, Reh54]. Paix [PSP55]. Palaeoecology [Lam80]. Palaeolithic [Mov53]. Palchinsky [Gra96]. Paleo [GBN77, Wit52]. Paleo-Indian [GBN77, Wit52]. Paleoecology [Sar74, Sar78]. Paleontology [Rai92]. Paleopathology [LBWD86]. Palestine [May63a]. Palisot [Gil92]. Palliati} [Bie59]. Palm [TZ66]. Palomar [Hub51]. Pan [Her79]. Pan-Indianism [Her79]. Panama [Lan50, Woo66a]. Panathenaia [Mer88]. Panta [Wei69]. Pantodonts [RK82]. Papal [Bed81]. Paper [Jef65, Oli57, Oli69, Tri80]. Papers [Bai89, Chr66, LB57, Lin51, Lin52b, Lin55a, Mos87, Sel51, Hof50, Jon63a, Mar51]. Paperworkers [Zie82]. Papyri [Neu62a, NV64]. Parabola [Gia97]. Paradigm [Lok78, RC99a]. Paradigms [Bra85]. Paradise [Cor94, Fry76, Ste76]. Paradox [Now95, Sto85, Tri80]. Paraguay [Wil72]. Paraguayan [Wil75b]. Parasite [Leh59]. Parents [Hof75]. Paris [Bla88, Chi55, Ham84, Ric51a]. Parish [Wil72]. Parker [Her79, Zal91]. Parliamentary [Sec65]. Parliaments [Dav85]. Parsons [Rei50]. Part [Dum57, Dum59, Dum60, Dum61, Dum62]. Particle [Gol75, Pan89, Sim87]. Particles [Mos50b]. Partners [Sea66]. Party [Mid85]. Parvo [Fra83]. Passage [Sha76]. Passing [Han68]. Passivity [Hac74]. Past [Bro89, Dia94, Gol86, Haa96, Pat97, Ric70, Sch57c, Men69b, Mil55]. Pastoralists [Ekv74]. Pasturage [Ekv80]. Pathology [Luk81]. Patients [Sho93]. Patrick [Lin51]. Patronage [Bar87, Don76]. Pattern [Ber57, Dav61, Kno93, Pat84]. Patterns [Cun57, Luk81, Mag56, Sch56a]. Paul [RH58, Wei97, Cox94, Moo99, Pey70, TG91, Wes98a, Wit99]. Pauling [Dun97]. Pausanias [Hab85]. Pay [Ede97, Gal99]. Payments [Roo90]. Payne [For96]. Peace [Cla53, Ken58, Lin51, Pat70b, Sch60, Dub79, Gre54, Kim77]. Peacemaker [CS54]. Peale [Law92, Mur57, Nyg70, Poe56, Por85, Sel50, Sel51, Sel80, ST54, Yoc92]. Peasant [Bla73b]. Peasantry [Blu82]. Peculiarities [Fra86]. Peer [HHDA66]. Peerage [Mid85]. Peirce [Eis57, Eis63, Len65]. Pekeris [Gil95]. Pelagian [Wol59]. Pell [MS98]. Peloponnesian [Mer71, Mer78]. Penal [Sel65]. Penalty [Mae80]. Peninsula [Cop56a]. Peninsular [Pad77]. Penn [Woo92, Ada93, Bur60a, Cor94, Dun83a, Dun83b, DM53b, Mor83]. Penney [Arn92]. Pennsylvania [CR54, Gol92b, Sch98, Cop53, Cop55, Dun83b, Lan89, Llo50, Row91, Sch56b, SC89, Sod89, Wit50, Wit52, Ald62]. Penology [Anc74]. Penrose [Bok95, Ede97, Sca87, Ano72m, Ber67, Cla59, Cor65, Dry60, Dub63, Fol58, Ful61, Gal66, Hux86, Lat70, Lur72, Per69, Rob79, Set69, Smi72, Sta57, Str64, dM56]. People [Kim93, RR92b]. Peoples [Sal67]. Peptides [Kar99]. Perception [Can60, ER90, Gol64b, Ken52, Obe79, Wal83, vM80]. Perceptions [Kam87]. Perceval [Roa66]. Pères [Chi58]. Perfect [Wil87]. Performance [Cha87]. Performed [Chi56]. Perhaps [For96]. Perikles [Mer75]. Peril [Set69].

Period [Ald76, Ano89g, Fle66, Gra89, Rei68]. Perjury [Gor80]. Permanence [Fry88]. Permanent [Gar74, HvW61]. Permian [Fos89]. Permission [Dic98]. Peromyscus [She61]. Perpetual [Her93]. Perpétuelle [PSP55]. Perpetuity [For98]. Persevere [Poe56]. Persistence [Dri53]. Person [GH58, SS57]. Personal [App97, Coo52a, Goo90, KK91, War77a]. Personality [Dun83a, Ekv74, Han73, Kin76, Mor80a, Sum60]. Persons [Ewa59]. Perspective [Bun91a, Cra91, Gom95, Har89, McK97, Pat97]. Perspectives [Bro81, Man83, Mel66]. Pertaining [Sny58]. Peter [Bre92, DM53b, Har93, Pal78, Smi83, Swi98]. Petersen [Eag98]. Petra [Bow90]. Petroglyphs [Dav52]. Petrography [Ste63b]. Phaedrus [Maz55]. Phantom [Sim58]. Phelps [Kau95]. Phenomena [Mos50b]. Phenomenology [Gob70]. Phenomenon [Hou66, LG78, Mac61, McG88, Rai85, Sha70]. Philadelphia [Kro57, Sny58, Tho90, Ada87a, Bar90a, Car89a, Car89b, Cor76, Fen53, Gre69, Har89, Kan89, Kle89a, Lan89, Maa81, Nas68, Oak77, Pen50, Pet51, Rai92, Ros89, Sny53, War85, Wie67, Wok89, Wol52, dT58b]. Philadelphians [Gou89, Smi89]. Philanthropist [Hij80]. Philanthropy [Cho98, Cur61, Fab61, Jor61, Kar85, Koh85, Moe61]. Philip [Kit96]. Philippe [Pau70]. Philokles [Mer88]. Philosophe [Pau70]. Philosopher [Fuc96, McE84, Ple67]. Philosophers [Con51, Fra75]. Philosophia [Pac82]. Philosophical [Ano60m, Bel59, Bel64a, Bel72, Bel75, Bry88, Bun91a, Car93, Car57b, Con50, D’E73, DM50, Fre62, Gil86, GS94, Hea88, Kro57, Lin53, Lin55a, Lin56, O’M77, Pen50, Rol56, Roo94, Shr60a, Smi84, Sta91, Ste93c, Tho90, Wil51a, Wol59, WD96, Lin50, Tho90]. Philosophies [Spe51a]. Philosophy [Bla52, Boa51, Bun91a, Fre93, Lok78, Mar55, Mar93, Pet73, Win83]. Phonetic [Cun57]. Photoaffinity [Wes78]. Photochemistry [Noy51]. Photographic [Bun91b, GJ55, Mar53, Tat54]. Photography [Kar60]. Phylogenetic [Pic54]. Phylum [Sch69]. Physalis [Men51]. Physic [Bel50]. Physical [Dub60, Eds85, Hil64, Sch50, vM80]. Physician [Cor57]. Physicians [Pen57]. Physicist [Nie86]. Physics [Cow93, Gol75, Mac73, Mar61, Pan89, Per69, Pye78b, Tho60, Yan96]. Physiologic [Amd70]. Physiological [Ful51]. Physiology [Wil76]. Piast [Kno76]. Pictographs [Dav52]. Picture [Hol90, Sel55]. Piecing [Bro89]. Pierre [Per53, PSP55, Per58, RH58]. Piers [Moe58]. Pigeons [Hit52]. Pigment [Rob86]. Pilgrim [Dys78]. Pinewoods [Ott86]. Pioneer [Li55, Lon57, Roa70, Smi62, Ano62m]. Pioneers [Dys78]. Pitfalls [Gou77]. Pitt [Kni79]. Place [Bie57, Dub62, Dun65, Gre98a, Kro58, WW70, Woo53]. Placenta [Ram69]. Places [Spi83]. Plains [Bic65]. Plan [Ano56t, But50, Goo58, Voe52]. Plandome [Mur62]. Planetary [Gol88c, GCM94, Gra73, Neu54, Swe73]. Planner [Sch70]. Planning [Dua58, Har62, Nas68, Pol77, Sch98, Shr60b]. Plant [Kra73, Li55, WM70, Wil76]. Plantation [Dav61]. Plants [Jen51]. Plasma [Cow93]. Plasmapheresis [SS57]. Plat [Mil87]. Plate [Sio51, Spi84]. Plato [Che54, Spe69]. Play [Bré70]. Players [Sto62]. Pleistocene [Dup52, Heu66, How59, Mil64b, Mur74]. Plentiful [New71]. Plight [Ede97, Wil92]. Plowman [Moe58]. Pneumococcal [Aus81]. Poeciliid [Riv58]. Poem [AF54]. Poet [And76, Bak84, Bro70, LF52, Mal50, Maz55, Sto62]. Poetic [Abr72b, Maz55]. Poetry [Bak84, Blo86, Gal74, Hof83]. Point [Enr75]. Points [Nic62]. Poland [Kno76]. Polar [Spi93]. Polarity [Roc97]. Pole [Dav61]. Pole-Star [Dav61]. Police [Cra68]. Policies [Adr92, Tei92]. Policy [Atk99, Bun75, Con92a, Cop54, Esp92, Fri72, Jes51, Mae77, MS98, Moy92, Mus72, Por92, Sch53, She85, Sta55, Sta50c, Viv74, Woo66b]. Poliomyelitis [Riv54]. Political [Bat86, Boy74, CR54, DL57, Dav60, Fre68, Hoo55a, Jo70, Ken71, Ker82, Kni79, Law92, Lub78, Nas68, Nou50, Par63, Sal67, Sch55, Sec65, Sey57, Spe69, Tud69, Wei93, Wes98b, Wil96b]. Politician [Wes98b]. Politicization [Rav85]. Politics [Bor87, Bro75b, Gil72, Gin78, Gre98a, Hea95, Kam71, Mor74a, Por96, Pye78a, Sei98, Shk93, Tei88, Thi93, Wil90]. Politics-Problems [Bro75b]. Polity [Orl71]. Pollen [Heu66]. Pollock [Bur75]. Pollutants [Amd70]. Pollution [Pat90, Stu70, WM70]. Polydesmida [CT69]. Polyglot [Pel93]. Polynesia [Dod63]. Polynesians [Was63]. Polypeptide [Smi79a]. Ponceau [Smi83, Swi98]. Pont [Hou90]. Ponte [Ser91]. Pontem [Ser91]. Poor [Cow60, Ham92, Her82, Hor95, McC85, SC89]. Pope [Kam71, Set69]. Popular [Bec62]. Popularity [Bad78]. Popularization [Cre85]. Populated [Myr51, Spe51b]. Population [Bec89, BC69, CS52a, Dav51, Dur67, EWW78, Gib95, Gib92, Kuz58, Kuz67, Myr51, Ros89, Sch95c, SC89, Tho67, Wol82]. Populations [Coa88, Rap60, Ste63a]. Porada [Mel98]. Porkdom [Hil70, Shi70]. Portrait [Lav75, Ric51b, Fel95]. Portraits [Mar86, Sel50, Sel56, Ver64, dT58b]. Portraiture [Han73]. Portugal [Smi71]. Position [Leo53, Mil64a]. Possible [Coo76, Dav53, ER90]. Post [LBWD86, Moy92]. Post-Harappan [LBWD86]. Post-Industrial [Moy92]. Postglacial [Han50, Ric50a]. Postscript [Cla54]. Potash [Rob72]. Potential [HC84]. Potentials [Mom50]. Pottery [VW78]. Poverty [Gel68, Smi88]. Power [Ben85, Ber67, Bis79, Bus97, Par63, Sha76]. Powers [dK87, Ken91, Bar64b]. Pozzo [Ver58]. Pozzo-Albani [Ver58]. Practical [CC95, Cla53]. Practice [Dri53]. Pragmatic [Mar88]. Prague [BR79]. Prairie [Bic65, Rob67a]. Praise [Wat78, Dow59]. Prapāthaka [Dum59, Dum61, Dum63a, Dum63b, Dum64, Dum54, Dum60, Dum62, Dum65, Dum69]. Prapāthakas [Dum51]. Pre [Bar99, Dal66, Eas65, Jo70, Tho50b]. Pre- [Dal66]. Pre-Copernican [Tho50b]. Pre-Hispanic [Eas65]. Pre-Scientific [Bar99]. Pre-Theory [Jo70]. Precedent [Mar65]. Precessional [Wri50b]. Precision [See61]. Precordial [Jar72]. Precursors [Eas82]. Predecessors [All60, Chi57a]. Predicament [Lin94b]. Prediction [For59, Kno93, Ram77]. Preeminent [Wes98b]. Preface [Gol87, Gol88d, Gol93, Tod92]. Prefaces [Cra58]. Prefatory [C.65, Tho67]. Prehispanic [Pol77]. Prehistoric [Bra72a, DW65, Swa52]. Prehistory [Cla66, Eme54]. Prejudice [Tay68]. Prejudiced [Ker54]. Preliminary [Coo52b, Rho76, Sny53, Jet61b]. Prelude [Sei96]. Premature [Tob96]. Premier [Far82]. Preparation [Chr66]. Prerogative [Fus57]. Present [And61, Bar55, Bur88, Coh55, Cow93, Gol86, Mar55, McG99, Mil55, Mün59, Pat97, Pot72, Sch57c, Shu71, Tri80, Wit50]. Present-Day [Mün59]. Presentation [Rol56, Voe52]. Presenting [Tre74]. Preservation [Alb60, EJ56, Leh71]. Preserved [SESB65]. Presidency [Tru59]. President [Gol86, KK91, Cai60]. Presidential [Ray80]. Presidents [Hea88]. Press [Con67, Fre68, Hys60, Kni79, Som76, Tay68, dBdS54]. Pressure [Ega71, Mur55]. Pressures [Fen70]. Preston [Rod92]. Price [Neu98]. Priest [Eri67, Ros60]. Priestley [Con50, McE84, Rob62]. Priests [Boy74]. Primacy [Mae77]. Primary [ES56a, ES56b, Thu50]. Primitive [Red52, Sau62, Was58]. Primroses [Cle64]. Prince [Jor71]. Princes [Tho53]. Princess [WD96]. Principal [Anc74, RC99a]. Principle [AE62]. Principles [Hum51, Ker82, Ste89]. Printed [Coh50]. Priorities [MS98]. Prisoners [Las91]. Prisons [Las91]. Pritchard [And99, Bur83]. Privacy [War77a]. Private [Lew63, Pot62, DL57]. Prize [Cra90]. Probability [Gil72, Gou77]. Probable [Pen57]. Probably [Llo50]. Problem [All66, Bar87, Bro81, Bro59, Bun91a, Bur60b, Con56, Eis63, Fab58, Fai64, Fre91, Ful68, Gol80, Gom74, Goo53, Ker54, Man65, Nic62, Nor73, Ohe78, Ott86, Ott89, Rob64, Sch95c, Tre53]. Problems [Ams91, Bac77, Ber54, Bro75b, Cha51, Coh56, Cop50b, Cor94, Dev69, Fre71, Gid52, Hac97, Mar53, Mea54, Mit62, Pat90, Pra68, Ril88b, Sto59, Tho67, Wei59]. Procedure [Fos82]. Proceedings [Ano88m, Hol70, Shi70, Mac65b]. Process [Abr87, CBL61, Fra54, Mos58, Sas80, Wei60]. Processes [Eds67, Sch57b]. Proclivities [Mei66]. Product [Tay72]. Production [Ald76, Hil71, HC84, Leo53, Men61, Sta51]. Productivity [Dew56, Gib81]. Profession [Ben71]. Professional [Dri53, Ing57, McK73, Rus57]. Professionalism [Bur93b]. Professionals [Hol70]. Professions [Lan99b]. Professor [Rin93]. Professors [Nic62]. Program [Eis54a, Hub51, Jas62, Klu59, Mur61a]. Programmes [Pou97]. Progress [Bor50, Bre61, Coh52a, Dan71, Gib81, LB57, Mur61b, Nis75, Pin99, Sea66, Sui64]. Progressivism [Nem68b]. Prohibition [Reu97]. Project [Bai89, CR54, Ano92k, CH81]. Projected [Ewa59]. Projection [Eis63]. Projections [Spi93]. Projects [Gra96]. Projet [PSP55]. Prolegomena [Ric80]. Prologue [Gol88b]. Prometheus [Lon58]. Promise [Fuc98, Kra80, Now95, Str61, Zwo61]. Promised [Bev89]. Promoting [Len66, Mas92, Pie79, Tho90]. Proof [App84, Mac84, Put84]. Proofs [Coo84]. Proper [Spi83, Wri50b]. Property [KS84]. Prophetess [Bla62]. Proportion [Gru89]. Proposal [McG97, Wal80, Wri50b]. Proprietors [Cop55]. Prospect [Wal80]. Prospecting [Eis94]. Prospective [McK97]. Prospects [Ams91, Cla53, Rho98, Ril88b, Ros81, Tei92, Dry60]. Prosperity [Kle71]. Protective [War63]. Protein [Eds67, Fru85, Kam78]. Proteins [Per69]. Protohistoric [Dal66]. Proust [Bré68, Scr90]. Provenance [Mar51]. Province [Bic65, Heu66]. Provincial [vD77, Chi53]. Provinciale [Chi53]. Provost [CR52]. Prudent [Cas92]. Prussia [Cra80]. Psychological [Bol62, But50]. Psychologists [Bun91a]. Psychology [D’E70, Ric80, Sok72]. Psychopharmaceuticals [Kli59]. Psychophysiological [Rat76]. Psychosurgery [Ful51]. Ptolemy [Dav74]. Public [Bes60, Gom74, Maa81, Mah91, Ost95, Por96, Smi84, Ver64, vH81]. Publication [Mea54]. Publicizing [Sch99a]. Published [Pet73, Ver64]. Publishing [Ker54, Scr80]. Pueblo [Hoe68]. Puerto [Mur55]. Pulpit [Som76]. Pulsars [Tay93]. Pulse [Det87]. Punishment [Mor96, Zim74]. Pupils [Cor57]. Purcell [Ram99]. Purchase [Her54]. Purchasers [Cor94]. Pure [RD73]. Puritan [Mor66]. Puritans [Fry89, Zak91]. Purple [Rob86]. Purpose [Haa50a, McG95, Pet61]. Purposes [Mor96]. Purposiveness [Gra75]. Pursh [Ewa52]. Pursuit [Ker82]. Pushkin [Bow99]. Pyat [Bla88]. Pygmies [Gus61]. Pylos [Ble57]. Pym [Maa81]. Pyrite [SESB65].

Qanats [Eng68]. Quadrant [Cop53]. Quakers [Tol63]. Quality [Rob76]. Quantification [Sti92]. Quantum [Coo76, Tri80]. Quarry [BT60]. Quartet [Mar72]. Québec [dB79]. Quechua [Urt81]. Queen [Ati93b]. Quest [Det81, Gou66, Pat70b, Rei75]. Question [KS84, Ted87]. Questions [Kle89b]. Quincy [But74, Kai70, Mor74b, Mus69]. Quis [Gri82]. Quoting [Pra68]. Qurra [Neu62b].

R [WD96, For80]. Race [Kin78, Mur55, Woo66b]. Races [DSR50]. Racial [Cor76, Mur55]. Racism [GS86]. Radar [Eva69, Gol69b, Rob83b]. Radiation [AH75, Bru63, Cur63, Fle75, Rus63, War63]. Radiator [Pan63]. Radical [Boy74, Woo66b]. Radicalism [Dur87]. Radicals [Han54]. Radio [Fow65]. Radioactive [Bad68, Kam94]. Radioactivity [Bad78]. Radium [Bad78]. Rafinesque [Boe58]. Ragnar [EC94]. Railroad [Gar61]. Railway [Lub78]. Raised [HH51, Sar78]. Ralph [McK73]. Rameau [Pau70]. Rampant [GS86]. Ramphocelus [Sib58]. Ramsay [Mur57, Por85]. Ramus [Wal70]. Rana [Leh59, Leh60]. Ranching [Ott86]. Randall [Nie92]. Random [Chi59, Jes51]. Randomly [For96]. Range [Ott86, Mil64b]. Ranke [Gil86]. Rape [Lan72]. Rape-Motif [Lan72]. Rapid [Miz79]. Rapidly [Shi51]. Rare [Hal63]. Raspe [KRF55, Kah55b, Kah56]. Rat [Det81, Ric68]. Rat-Catcher [Ric68]. Rate [ES56a]. Rates [ES56b, Lub78, War63]. Ratification [Mar65]. Rational [Ric80]. Rationale [Kim93, SS57]. Rationalizing [Old94]. Rawson [Fah71]. Ray [Per69, Pou97]. Raymond [Led94]. Rays [Fri97]. Re [Leo53, Mel57, Ray80]. Re-Creation [Mel57]. Re-Examined [Leo53, Ray80]. Reaction [BHT86, Hug69, Noy51, Wil96b]. Reading [Alb81, Bun91b, Mer83]. Real [Woo99]. Realist [Blu82]. Reality [Bai62, Bro75a, Gar61, Gin73, HC84, Hoo55b, Ken55, Mos55a, Obe79, Wel89]. Really [Gre77]. Reappraisal [And76, Fru88]. Reason [Bla71, Boa71, Gra93a]. Reassessment [Mer83]. Recently [Chi57b]. Reception [Eis54b]. Receptors [Wes78]. Recluse [Hop61]. Recognition [HC88]. Recollections [Chi50]. Reconnaissance [Kar60]. Reconsideration [Rob98]. Reconstituting [Kro51]. Reconstructing [Man58]. Reconstruction [Pay78]. Record [Gem89, Hed65, Hol90, New59]. Recording [Voe50]. Records [EB79, Gru89, Wal51c]. Recovers [Win79]. Rectifiers [SE98]. Red [Hea95]. Redaction [O’G78]. Rediscovery [ST54]. Redrafting [Gou97]. Reduction [Hil71, Jac80]. Reevaluation [Pat90]. Reference [Mur55, Ric50a, Sta91, Swa52, Tob96, VH53]. Reflected [Eks99]. Reflection [KK91]. Reflections [Abr87, Bla65, Bri55, Fel87, Fon98, GB56, Gri82, Kam94, Rob83a, RR92b, Sto87, Blu97, GB56]. Reflexions [Pat76b]. Reform [Cho89, Cow60, Hir53, Hor95, Joh54, May63b, Nem80, Tho71]. Reformation [Pel61, Fre93]. Reformer [McI75, Sch71]. Refugee [Coh82]. regarding [DeV55]. Regency [May78]. Regime [Men89]. Regimen [Hor95]. Regiment [Zal91]. Regiments [Bur80]. Region [Sch95a, Wel89, Ald76, CSL50, Lan89, Pen94, RGSS72]. Regional [Kle89b, Pai54]. Regions [Kuz75, Sim77, Spi93]. Regularity [Hoe90]. Regulation [Cri88, Pic54, Rei92, Sta50a, Sto59]. Regulatory [Eds67]. Reichardt [Lin94a]. Reinhard [Sme93]. Reintegration [Sim55]. Reinterpretation [Dup57]. Reischauer [Fai92]. Rejections [Fry83]. Rel [Big54]. Related [Esa67, Kuz58]. Relating [Fen53, Lin51]. Relation [Car58, Coh56, Gra75, Gre60a, Mom50]. Relations [Cor72, DM50, Gri72b, Hum89, Kin78, Kra73, Mal53, McG79, Olm60, Por96, She61, Shi51, Sti77, VW78, Woo53, Wri51, vD86]. Relationship [Con60, Mor77, Nem68a, Sch53, Voe50]. Relatively [Wri72]. Relativity [Mou59]. Relevant [Mil67]. Relief [Her82]. Reliefs [Ver58]. Religion [Goo81, Gre98a, Gre59, Jac63, Sny61, Too89, Was58]. Religious [Ada93, Fry89, Hor61, McI75, Spi69, Zak91]. Reluctant [Ric68]. Remains [Ang52, Ste77]. Remarkable [Tho98]. Remarks [Ada87b, Fra55, Has79, Her55, Lin87, Pie68, Pie80, Rho76, Rho88, Sca87, Ste93a, Ste93b, Ano60m]. Rembrandt [Mar86]. Remigio [Dav60]. Remote [Bro62a]. Renaissance [Coh95, Bla57, Gun93, Jac68, Kri76, Lew94, Mei66, Neu89, Set56, Tho60]. Rene [Gre98b]. Renewal [Lad75]. Reno [Car58]. Reorganizing [Abr62]. Repeated [SS57]. Reply [Hil70, Rin93]. Report [Ano65v, Coo52b, Dan71, Jet61b, LB57, Sui64]. Repositories [Sny53]. Representation [Wri56]. Representative [Bon74, Mid85]. Reprint [Ano88m]. Reprinting [Wat61]. Reprise [Bow95]. Reproduction [Alb60, End52]. Reproductive [Mor77]. Republic [Hil70, Dav74, Hoi75, Pom54, Sec65, Shi70]. Republics [Viv74]. Reputation [Bel72, Mil78]. Require [Sil95]. Requirements [Mel79]. Rescue [Kie80]. Research [Ber57, Bla57, Bro75a, Bro89, Cra90, CSL50, Dev69, Dry54, Foo65, Fru85, Fru88, Gom85, Hou90, Jet61b, Kel57, Kra73, Lan55, LH85, Mac53, Mal98, McC84, Mor64a, Mos50a, Now95, Riv54, Roo94, Sch57c, Spo81, Sto55, Wat53, Wie71, Win67, Yng64, Yod82]. Resident [Wil51a]. Residues [Smi79a]. Resilience [Ada78, BG77]. Resistance [Kie80, May91, Sti92]. Resolution [Gom74, Row91]. Resonance [Cha83, Rob89, Sli98]. Resource [Mot75]. Resources [Bou63, Fol58, Har71a, Wal51a]. Respiratory [Amd70]. Responding [Thi93]. Response [Ada97]. Responses [Amd70, Dav68, HZ58]. Responsibilities [Tay54]. Responsibility [Bes60, Hil55, Lur72, Mad62]. Rest [McG97]. Reston [Oak98]. Restoration [Her82, McB56]. Results [Jas62, Oli50, Tau76, Jet61a]. Resumption [Wil61c]. Resurrection [Byn98]. Rétif [Chi58]. Retirement [Rob57b]. Retracing [Cap74]. Retreat [Boa71]. Retrospect [Whi76]. Retrospective [McK97]. Return [Lab56]. Revealed [Per69]. Revelle [Nie92]. Revenue [Bro85]. Reverend [CR52, Bel51, McG99, Sch71]. Review [Ure71]. Reviewed [Gra72]. Revised [Met91]. Revising [Mal50]. Revision [CT69, Ewa59, Row91]. Revisions [Ive51]. Revisited [Car89a, Gal66, Lap83, Pie68, Shi70, Wit99]. Revolt [Bla71, Daw60, Plo78, Mer84]. Revolution [Alb72, Bro62a, Lin51, Pet69, Spi71, Tol51a, Wil68, Bla73a, D’E66, Har53, Hys60, Mor74a, Rob72, Sch55, GB56]. Revolutionary [D’E70, Rei68, Sha58, Was73]. Revolutionibus [Eas82]. Revolutions [Dys96]. Rex [Bat86]. Rey [Faj74]. Rhei [Wei69]. Rhenish [Cha80]. Rhesus [HHDA66, HvW61]. Rhetor [O’M77]. Rhetoric [Bar90a, Ega78, Fre66]. Rhetorical [Bev67]. Rhoades [Dem94]. Rhodesia [Sum60]. Rhodesian [Kin78]. Rice [Bar52b]. Rich [Don76, Hor95]. Richard [Say71, Sca87, Bok95, Ede97, Sto99a]. Richelieu [Som76, Sta55]. Richer [Olm60]. Riches [Hor95]. Richmond [Edw61]. Rico [Mur55]. Riddle [Jon63b]. Riddles [Tay53]. Right [Ams91, Per96]. Rights [Aar74, Cha51, Mac57, Mah91, Per96]. Ringe [Gre93a]. Ringwalt [Sid99]. Rinsō [HR55]. Rio [Str95]. Riot [Pot75]. Risco [MO59]. Rise [Car74, Rai92]. Risk [Wal83]. Risks [Cro81]. Ritchey [Abr72a]. Rites [Bra81]. Rittenhouse [Sio51]. Ritual [Pay78]. River [Leo94, Sto59, Tin70]. Roach [Wol95a]. Road [You63, Por74]. Roads [Has61]. Robert [Ken95, Ano98i, Bac72, Don76, Her95, Hij80, Lan98b, Lan98c, Mac97, Mur59, O’G78, Per99, Sno58, Tre74, Woo92]. Roberts [Fir91]. Robertus [Cor99]. Roche [Chi50]. Rockingham [Hof50]. Rocks [Hed65]. Rococo [Smi71]. Rod [Coh52b]. Roger [Got91, Bog99, Nie92]. Role [Ada84, Bla57, Dub71, Gom92, Gor88, HPSH50, Ing57, Kle71, Pet99, Ray80, Roc63, Roo72, Sam99, Sec65, Sta91, Ste59a, Tay57]. Rolls [Pan63]. Roman [Bie59, Bow99, Dav74, Gum55, Lew54, Ric51b, Sol68, Sym60, Tay50, Ver64]. Romance [Bau59]. Romani [Bie59]. Romanian [Old94]. Romanization [Bro59]. Romans [Ric51b]. Romantic [Pey70, Zio87]. Romanticism [Bré70]. Rome [KJF52, LL74, Pal81]. Ronald [Bow91a]. Roosa [Mac97]. Roosevelt [Hau75]. Root [ES56a, ES56b]. Roots [App97, Ell83, RC99b]. Rothrock [DG73]. Rothschild [Gra98]. Rousham [Cof86]. Rousseau [Jac89]. Routinely [Coo84]. Royal [Wil56, vD77, Bel51, Cop50a, Haa80, Llo50, You63]. Royce [Hil70, Kin76, Pan63]. Rubens [Mar86, Nyg70]. Rudolf [Gra93b]. Ruhr [Cor72]. Rule [Kim62, Fah71]. Ruler [Ver65]. Ruling [Har58]. Rumford [Bro52, Tho98]. Running [Cop56a]. Runnymede [Gal66]. Rural [Bun91b, Kuz74, SC89]. Rural-Urban [Kuz74]. Rush [D’E73, But51, D’E66, D’E70, Lin51]. Russell [Fen92, Bow91b]. Russia [Ber95, Kap98, Mot75, Rie84, Sch99a, Shu58]. Russian [Bow99, DM50, DM53a, DM65, Ega71, Gra96, Gri72b, Lan55, Mos55b, Sch70]. Rutherford [Bad68]. Ryūkyū [Smi60].

S [Boe58, Chi56, Eis57, Eis63, Hiż94, Ken95, Len65, Moy97]. S. [Dav63, KJF52]. Sacre [Jac69]. Sacred [Kra63, MS67]. Sacrifice [Dum62, Dum63a]. Sacrifices [Dum57, Dum59, Dum60, Dum61, Dum69]. Sagan [Gin99]. Sagendorph [VD94]. Saint [Ing65, Per53, PSP55, Per58, Wil75a]. Saint-Pierre [Per53, PSP55, Per58]. Sake [Ree86]. Sakharov [Dre94, Mar88]. Salam [Dys99]. Sale [Ech54, Wil83]. Salisbury [Oak95, Wes98b]. Salvador [Wat99]. Salvage [Bre61]. Sam [Shi70]. Same [GH58, SS57]. Samian [Mer84]. Samos [Hab96]. Sample [VH53]. Samuel [Hei98, Sch71, Fah71, Jef94, Lan76, Mil98, Sjö94]. Santa [HK62]. Sardis [Gre92]. Satan [Ste76]. Satellite [Neu86]. Satellites [Whi58a]. Satisfactions [Gun96]. Saunier [RH58]. Sautrāmanī [Dum65]. Sauvy [Coa91]. Saw [Bel72]. Sawyer [Gol97]. Scare [Hea95]. Scaurus [Bat86]. Schapiro [Ros98]. Scheiner [vH96]. Schema [Sch74]. Schiller [Cra68]. Schinkel [Zio87]. Schlachtflotte [Mae77]. Schmitt [Blo97]. Scholar [Bar52a, Mac57]. Scholars [Cop55]. School [Con60, Gol92b]. Schools [Bel55, Bes60, Cus50, Fre93, May60, Rav85]. Schrödinger [Tri80]. Schwarzschild [Ost99]. Science [Abr72b, Adr56, Adr92, Alp55, Atk99, Bak94, Bar55, Bau93, Blu88, Blu97, Car74, CWH^{+}71, Coh55, Com56, Dav86, Dry53, Dub63, Dub77, Eas82, For59, Fry83, Går99, Gol86, Gol91, Gom92, Gor88, Gou50, Hil64, Hin64, Hol70, Hux86, Jas62, Kil96, Koh85, Koh74, Kre71, Lan99a, Lik99, Lin51, Lur72, Mac61, Mag99, Mar55, May60, McG79, Mer61, Mum61, Mur61a, New61, Nic61a, Olm60, Opp90, Orl71, Pal61, Pat74, Pel61, Por92, Rai92, Ril88a, Sam99, Sea66, Sei90, Sei98, Shr61, Sta50b, SC84, Swi98, Tei88, Thi93, Tob96, Wea60, WW70, Wig50, Zuc80, Dai95a, McG79, Pur65]. Sciences [Alp55, Ati93b, Boa51, Cra90, Fra93, Ger74, Her55, Kar85, Kro57, Neu63, Pen50, Dup57, DM65, Lum78, Rai92]. Scientific [Bad78, Bar99, Bla57, Bor50, Bor52, Bus54, Cla64, Coh50, Coh52a, CR54, EB79, Fle86, Frä99, Fru85, Gil89, Gle60, Gou56, Gre60b, Han76, Hol70, Klu58, Mer61, Per58, Pri62, Ram77, Raw78, Roo94, Scr80, Scr90, Tat54, Ver58, Waa74, Wat53, Wei99, Wel74, Wie71, Wil68, Zuc80]. Scientific/Philosophical [Roo94]. Scientist [KK91, McE84, Rus57, Wil96b]. Scientists [Con51, Hil55, Por92]. Scott [Fox68, Par72]. Scottish [Bev67, Sch70]. Scourge [McG96]. Scribner [Bai97]. Scrolls [Gol80, Tre53]. Sculpture [Ant63, GB94, Smi71, dT58b]. Sea [Beg97, Gid52, Gol80, Tre53, Woo66a, Fle66]. Seafarers [Dye76, Dye89]. Sealing [EJ56]. Seamus [Ven99]. Search [Hol81, Kri76, Spi76]. Seashore [Sau62]. Seashore-Primitive [Sau62]. Season [Lop77]. Sebastian [Sch97]. Second [Don76, Dum59, Dum64, Dum65, Mil76]. Secondary [May60]. Secrecy [Cra90]. Secret [Fri62]. Secretary [Bel75]. Sector [Gom74, Pal81]. Secular [Jon55]. Secured [Oli50]. Security [Lin55b, Sei90]. Sed [Gri82]. Sedgwick [Bar74]. See [Coh82]. Seedlings [Sch58]. Seeds [Woo66b]. Segregate [For96]. Selden [Chr84]. Selected [SW50]. Selection [EG59, Leo94, Ler59, May91, Ros75, WE79]. Selections [Cra90]. Self [Ker82, Kno93, Mae80, Oeh74, Sul68]. Self-Defense [Mae80]. Self-Determination [Ker82]. Self-Help [Sul68]. Self-Knowledge [Oeh74]. Self-Organization [Kno93]. Sellers [Sel51]. Sellin [Wol96]. Selling [Dun83b]. Semitism [Old94]. Senatu [Bat86]. Seneca [Her79, CS53, Sny82, Sny83]. Senecas [Fen56]. Senescence [Cri88]. Sensory [Ken52, Wil76]. Sent [Cop50a]. Separation [dK87]. Separatist [Moo87]. September [Ano97m, Ano98i, Bor93, Bow91a, Boy98, Cas92, CV92, Cle98, Coh93, Fai92, Fed94, Fen92, For99, FC92, Fry96, Hen97, HE96, Hof91, Jam95, Jef95, Kay98, Ken95, Kit96, Kor97, Kro57, Lai98, Lan98b, Lan98c, Lin94a, Lip91, McP98b, O’C95, Pel95, Per99, Rav91, Roa98, Rob91a, RB96, Rod92, Ros98, Sal92, Sea93, Sio51, Sjö94, Ste97, Wol96, Wol95b, Woo92, Wri97]. Sequence [Bun91b, Car58]. Serendipity [Rob86]. Serkin [Gra93b]. Sermon [Ham84]. Servant [Ati93b, Gru89, vH81]. Servants [Gru89]. Service [Hoi75, Wri56]. Services [Kee63, Roo72]. Servitude [Sel65]. Sessional [Jon63a]. Set [Gol88c]. Settings [Pat76b]. Settlement [Aar74, Ada78, Cha80]. Settlers [Lan89]. Setton [May97]. Seven [Dum63b, Gid62]. Seventeenth [Fos82, Haa80, Het75, Jen68, Nas68, Per53, Per58, Smi86, Sti92, Ted87, Ver58]. Seventeenth-Century [Het75, Per53, Per58, Smi86]. Seventh [Dum61, Dum63b, Ant63]. Severe [Fab58]. Severo [Kor97]. Sex [All66, Bod85, Bur73, For59]. Sexual [Ama59, Kot80, WE79]. Shadows [Bur88]. Shaker [Ste89]. Shakespeare [Fri62, Ben71, Bla73b, Fry88, Fry98, Lev62, Lom97, VD53, Wri62]. Shakespearean [Fry88]. Shale [Yoc96]. Shall [Ega78]. Shang [Spe69]. Shanidar [Ste77]. Shannon [Fre96]. Shape [Tin70]. Sharp [Con67]. Shaw [Sti59]. Sheep [Rob86]. Sheet [Sey57]. Shelton [Llo50]. Shenandoah [Cha80]. Shewell [Sme95]. Shield [Ver65]. Shils [Eis97]. Shingleton [Hil70]. Shining [Bel50]. Shipboard [Bra81]. Shipbuilding [Cro73]. Shippen [Cor50]. Shipping [Fle66]. Shoran [See61]. Shorapur [Pad77]. Shoreline [Ric70]. Short [McC85, Sha58]. Should [Bun91a, Fuc96]. Shrubs [Ewa59]. Shull [GS80]. Sickness [Lam80, Wol82]. Side [Mer53]. Sidelights [Bei60]. Sidereal [McC66]. Sidi [RT99]. Sidney [Sti59]. Sigismund [Mik58]. Significance [Ald76, Ama59, BB53, Fri66, Gou89, Kae64, Mac73, Rob67a]. Sikinos [Fra83]. Silicon [Sei96]. Silk [vH81]. Sill [Cla91]. Silvanus [Pal78]. Simmons [Jef99]. Simon [Eis57]. Simple [McC85]. Simpson [Lap83, Rai86]. Sin [Fra60b]. Sinai [Ati52, Wei66]. Since [Bla57, Cre85, Lew68, McK94, Mil64a, Pac50, Son54]. Sinclair [Bol62].

Singing [CS53]. Singletons [Mer61]. Sino [Shi51]. Sino-American [Shi51]. Sir [Bau94, Boo97, Bre92, Can77, Con51, Got91, McI75, Par72]. Site [Dau56, Men61, Mur74, Pen57, TNN71, Wit52]. Sites [Con58, Jac75, Mov53]. sitophila [DSR50]. Situation [Tri80]. Six [Fen53]. Sixteen [Tre74]. Sixteenth [Col72, Lav75]. Sixteenth-Century [Col72, Lav75]. Sixth [Ant63, Dum62, Dum65]. Size [Ama59, Gre77, Hof75, Ken91]. Sized [McC91]. Skeletal [Ang52, Ste77]. Skeptics [Bes79]. Sketch [Hil78]. Sketches [Mur57, Zig57]. Skinner [Bra92]. Skourou [VW78]. Slate [Dav53]. Slave [Vog73]. Slavery [Elk63, VK70]. Slaves [McP95]. Sleep [Mei66, Mor64b]. Sleeping [Lam80]. Slippery [Lur72]. Slow [Dav68, Gel68]. Small [Gra96, Gre77, Sol71]. Smith [Lec94, Bar90a, Big54, Sio51]. Smithsonian [Car57b]. Smyth [Kau95]. Sniff [Det87]. So-Called [Fry83, Jon63b]. Social [Alp55, Bau93, Bes60, Boo75, BP67, BG77, Bri75, Car74, Coc70, CS52a, Fra93, Ger74, Goo69, Goo90, Hen87, Her61, Her55, Hil55, Kar85, KS84, Kin54, Kot51, LS55, Lum78, Mac61, McG79, Miz79, Mos50a, Moy92, Mum54, Pal81, Ric53, Sha76, Spi69, Tho50a, WE79, Wil70, Wil79, Wil96b, Wyl59, vD77]. Socialism [Sti59]. Societies [Blo86, BP67, Ing57]. Society [Abr62, Ada78, Bau71, Bel59, Dav51, Hea71, Kee68, Lin53, Ree86, Ril88b, Rob55, Roc63, Sch97, SC89, Tho90, Win79, Ano60m, Bel51, Bel64a, Bel72, Bel75, Car93, Car57b, Con50, Cop50a, DM50, Fre62, Gil86, GS94, Hea88, Kro57, Lin55a, Lin56, Llo50, Oli50, Pen50, Rol56, Sel56, Shr60a, Smi84, Sta91, Ste93c, Wil51a, WD96]. Sociological [Tib54]. Sociology [Bar55, Mer61, Ree86, Car74]. Socrates [Wal70]. Sœurs [Han53b]. Soil [Wol82]. Solar [Dan71, Eva69, Neu62b]. Solids [See51]. Solinus [LL74]. Solomon [vH81, Lip91]. Solution [Mit62, Sto59]. Solutions [Pat90]. Somatic [Bru63, Rob64]. Somatology [Gus61]. Some [Bac77, Bei60, Bel50, Big54, But54, Byn98, Cha51, Coh56, Cop51, Cop55, Cun57, Dar59b, Dav86, Dua66, Fah71, Fle86, Fle66, Fra54, Fru77, Har89, Hum51, HPSH50, Jef58, KRF55, Lew94, Mea78, Neu72, New61, Nob68, Not54, Oli50, Pat84, Pat76b, Pen54, Poe56, Rob83a, Sib58, Ste63b, Tho50a, Tho60, Thu50, Tru59, Wil75b, Fol58, Gri82, Her55, Reu97]. Somerville [Pat74]. sonomae [Leh59]. Soul [Pri98]. Sound [Abr72b]. Source [Blu82, Col72, Dav53, Goo90, Spo51, Wie71]. Sources [Che54, EB79, Fow65, Fre62, Ker54, May63a, McC91, Mea54, Mea78, O’M77]. South [Sar78, Shi70, Wer81, Cla63, Dal66, Hil69, Luk81, Mil82b, Sar74, Sim78]. Southeastern [SC89, Sod89]. Southern [Cap74, GB94, Heu66, Pal81, Smi60, Bok52, Bur83, May63a, Mur74, Ott89, Sum60]. Southerner [Bon74]. Southwark [Tin70]. Southwest [EJ56]. Southwestern [Bra72a]. Sovereignty [Sch73]. Soviet [Fai64, Gri72b, Hoo55a, Hoo55b, Ken55, Mos50a, Mos55a, Mot75, Mov53, Rub63]. SP. [Leh60, Leh59]. Space [Cow93, Dry60, Dub62, Dys78, Fri96, Hub51, Jas62, Nor73, See51]. Spacecraft [Pic66]. Spaceships [Dra60]. Spacetime [Wei81]. Spain [Bro59, Car93, Dar79, Lon59]. Spanish [Han53a, Han63, Jac68, Llo89, Whi58b, Whi60, Whi70a]. Sparrows [Mor77]. Special [Aul83, Coa88, Dum51, New73, Ric50a, Swa52]. Specialized [Mou83]. Species [Gla60, Hen82, Zir59, Sta60]. Specimens [Rob60]. Speck [Wal51b, Mer83]. Spectroscope [Swi69]. Speech [Ger74]. Spencer [Car74, Her54]. Spengler [Kel92]. Sperry [Bog99]. Sphere [Neu62b]. Spicer [Gal91]. Spinoza [Col63]. Spirit [Str95]. Spiritual [Gar75, Haa50b]. Spiritualism [Raw78]. Spite [Pey70]. Spokesman [Dou59]. Spontaneous [Con50]. Spread [Dav51, Eng68]. Square [Nic61b]. St [Ati52, Pal78, Smi86]. St. [Llo50, See51, Wei66]. Stability [Ada78]. Stages [Sas80]. Stalin [Fry98, Ken71, Rob55]. Stamp [Wol55]. Stand [GS86]. Standard [Met91]. Standards [Dri53]. Standpoint [CBL61, Rob67b]. Stanley [Lec94]. Star [Dav61]. Stars [Cop56a, Fow65, Wri50b]. State [Bri61, Bro85, Cha71, Dai95a, Fen71, Fre91, KK91, Moo60, Nem68a, Pet99, Rob98, Roo72, Ste93c, Ega71, Hau75, Mal55, Viv74]. Statecraft [Fri62]. Statement [C.65, Ste89]. States [Bor65, Clo80, Hen87, Jac80, Plo78, Ano89g, Ber54, Big54, Bra72b, Bry88, Coc64, Coh82, Esp92, Gem89, Gid50, Gra89, Hea51, Hin64, Hur60, Kir56, McN79, Mur56, Nic67, Rob57a, RT99, Roo72, Ros89, Shu58, Tei92, dT58b, dT60]. Static [Wil68]. Statistic [Swa52]. Statistical [Mos87, Wil55]. Statistics [Mos50b]. Statues [Ric51b]. Statuette [Bie57]. Status [And61, Bar55, Coh55, Mar55, Shu71, Wit50]. Statute [Gor80]. Steady [Pot75]. Steel [Sto99a]. Stellar [Mey98]. Stems [TZ66]. Stephan [Tie98]. Stephen [Smi83, Swi98, Chr66, Dil67, Kir66]. Sterne [Ham84]. Sterrenberg [Kno91]. Stevens [OC81]. Stevenson [Bel56]. Stigler [Bec94]. Still [Wei86]. Stochastic [Mos58]. Stockbridge [Moc68]. Stoetzel [Lhe93]. Stone [Jac75]. Stores [Mor77]. Story [Goo76, Lin50, Riv54, Sti92]. Storyknife [Osw64]. Strait [Gid62]. Strange [Con70, GS80]. Strategic [Gar80, Yod82]. Strategies [Ada78, Ohl74]. Strategy [Sch63, Sol71, Zub76]. Stratigraphic [MO59]. Stratigraphy [How59]. Stratton [Joh96]. Stream [Ise52]. Street [Mor96, Boo78]. Street-Level [Mor96]. Strength [Nic67]. Stress [Sza72]. Stresses [See51]. Strigosas [Cle54]. String [Mar72]. Strong [Fow65, SC84]. Structural [Det55, Eme62]. Structure [Cop56b, Esa67, Fru85, Har64, Per69, SC89, Ure71, Yng60]. Struggle [Chi58, CR54, May78]. Stubbs [Sto99a]. Student [Cor50, Cor65, MS98]. Students [Bel50]. Studied [HH76]. Studies [Bes52, Bes60, Bok52, DSR50, Fen51, Fri66, GJ55, Kra63, Lik99, Mer83, Mur55, Roa70, TAT53, Tre53, dT58a, dT58b]. Study [Ben79, Ber67, Bra85, Car74, Col65, CS52a, Cop54, Cor72, Haa50b, Hij80, HPSH50, Ker82, Kim62, Kin78, Klu58, Kub59, LS55, Lin68, Mil55, Moo87, Mor74a, Mos87, Noy51, Olm60, Rho76, Roo94, Smi83, Sum60, Woo53, vH96]. Studying [Arr99]. Style [Fru85, Gri72a, Kli73, Pay81, Sch99a]. Styles [Mos87]. Submarine [Vin68]. Submarining [Bea61]. Submergence [Ric50a]. Subrahmanyan [Ost98]. Subsequent [Cor50]. Substance [Tin70]. Substances [Sta51]. Substitutes [Gib65]. Subtraction [Kam71]. Successor [Ada87b]. Sudan [Bur83]. Suggested [Fol58, Sch53]. Suggestions [Kir66]. Suits [Fis94]. Sultan [RT99]. Sumerian [Kra63, Kra80]. Summarizing [Fra55]. Summary [Ste77]. Sun [Men69b, Rob79]. Sunflower [Hei51]. Sunnyside [Sch98]. Sunshine [Bel56]. Sunspots [vH96]. Suomi [Gre98c]. Superconductors [Hou50]. Superintendent [vH81]. Supermajority [Kim93]. Superstition [Dar79]. Supplies [Rid52]. Supply [Las75]. Support [CWH^{+}71]. Supported [Bea57]. Suppressed [For96]. Supreme [Abr87, Fra54, Kat97, McK94]. Surface [Gol71, Gol69b, Pat70a]. Surgeon [Was73]. Surgery [C.65]. Surgical [Joh65]. Surprises [Sli98]. Surprising [Cho98]. Survey [EB79, Lin53, Sny53, Ver64, Clo80, CR54, NRF81]. Surveys [Cop55, Eva69]. Survival [Neu63, Sel65, Str88]. Survivals [Ver65]. Susanna [Mil78]. Susquehanna [Jen66]. Susquehannock [Jen68]. Sustaining [Str95]. Swedish [Jon99b]. Swings [EWW78, Kuz58]. Swiss [Bla73b]. Sylvanus [Kid59]. Sylvester [Pal78]. Symbiosis [Bur84]. Symbolic [Sim93]. Symbolism [Han68, Spi69]. Syme [Bow91a]. Symmetry [AE62, Roc97, Yan96]. Symposium [Bor52, BCGR81, C.65, CWH^{+}71, Egg69, Has79, Pie68, Pie80, Rho76, Stu70, Tho67, Was63, Mea54]. Syn [Dav81]. Syn-Ethics [Dav81]. Synfuels [Dav81]. Synopsis [Gey62]. Synthesis [Nou50]. System [Blu88, Bra72b, Cam91, Eva69, Gra75, Hoo55b, Ise52, Ken55, KB80, Löw91, Mac65a, Mos55a, Pet99, Pic54, Roo72, Wil76]. Systematic [Ber67, New61, She61]. Systems [BG77, Gar74, Pat84, RC99b, Sch56a, Ste86, Wil79].

T [DG73]. T. [Kam86, Ken95]. Tables [NP67, GCM94]. Tablets [Ben53]. Taccola [Pra68]. Tacit [Gol75]. Tacitus [Bow99]. Taft [Hor61]. Taille [vD77]. Taittirīya [Dum57, Dum59, Dum61, Dum63a, Dum63b, Dum64, Dum51, Dum54, Dum60, Dum62, Dum65, Dum69]. Taittirīya-Brāhmana [Dum57, Dum59, Dum61, Dum63a, Dum63b, Dum64, Dum51, Dum54, Dum60, Dum62, Dum65, Dum69]. Taittiriya [Dum61]. Taittiriya-Brāhmana [Dum61]. Takano [Kub59]. Tales [Ewa83, Osw64, Rei90]. Talking [CH81]. Talleyrand [Sti77]. Tamilnadu [Sar78]. Taming [Ett79]. Tanagers [Sib58]. Tangled [Sei96]. Tapiales [GBN77]. Tarapaca [TNN71]. Task [Dub60]. Tasso [Lee81]. Taste [Mik58]. Tatti [Gil76]. Tattoos [Dye89]. Taussig [Tau76]. Tax [Mor74a]. Taxation [Joh54, Mag58, vD77]. Taylor [Mur61b]. Teacher [Dic98, Win67]. Teaching [Bau93]. Teapot [Sch99b]. Technical [BB53]. Technique [EJ56]. Techniques [Bad68, Cha83, Kar60, Rai66]. Technological [Bre61, CR54]. Technologies [Fuc98, Sta91]. Technology [CWMM79, Dry53, Gom92, Lur72, Mum61, Sho93, Wal93]. Tectonics [Spi84]. Teeth [HvW53, HvW61]. Tehuantepec [Lan50]. Telescope [Abr72a, Ros54]. Telescopes [Hub51]. Temin [Dul95]. Temperance [App97]. Tempest [Sch99b]. Temple [And76, Pen57, Ver65, Zio87]. Tempo [Lap83]. Temporary [Gib65]. Ten [Dub60]. Tending [But74]. Tenrecidae [Gou65]. Tension [Nas68]. Tensions [Bro75b]. Tenth [Dum51]. Term [Lik99, Ros81, Stu70]. Terminal [Mil64a]. Terms [Rat76]. Territorial [Bis79, Sch53]. Territoriality [Höl79]. Territories [Lan59]. Territory [Ric50b]. Tertiary [RK82]. Testament [Met87]. Testamentary [Ben91]. Testimony [Dar79]. Testing [Tay93]. Tests [Cro70, Wil61c]. Tetrabiblos [Ril88a]. tetrasperma [DSR50]. Text [NPZ81]. Textiles [BB53]. Texts [Kra63, Lew94, Sca76]. Thãbit [Neu62b]. Thalberg [Mik58]. Theaceae [Plu77]. Theater [Sti61]. Theatre [Mor66]. Their [Bla57, Chi57a, Chr66, Cla50, Coo52a, Cop51, Cop55, Dav52, Fru85, GS86, Kam87, Kel57, LVF98, Mar53, Mea54, Mit62, Rob79, Roc97, TZ66, Wei59]. Them [Coh82]. Theme [Roa66, Sin99, Ver65]. Theocratic [Ree86]. Theodor [Bar83, WE98]. Theodore [Hau75]. Theology [Bla71, Pel61, Ste89]. Theorem [App84]. Theoretical [Sch57b]. Theorist [Dav60]. Theory [Bev67, Coo76, Cza76, Dic81, EG59, Fru85, Gil82, Gra73, Jo70, Kil71, Lad75, Leo94, Löw91, Mac73, May91, Neu54, Sch53, Sou52, Sti59, Swe73, Wes56]. There [Men89, Pan89]. Thérèse [Con92b]. Thetford [Smi86]. Thicker [McP99]. Thimann [Sto99b]. Thinking [Con94, Cus50, Thu50]. Third [Dum51, Dum54, Dum59, Dum60, Dum61, Dum62, Dum63a, Dum63b, Dum69, Hoi75, Oli57]. Thirteenth [Tur71]. Thirteenth-Century [Tur71]. Thirty [War77b]. Thither [Cop50a]. Thomas [Bea96, Coh95, Sta98, All66, Bed64, Bei60, Boy58, Hei98, Hil78, dK95, LB95, Maa81, Pet61, Ple67, Rob53, Rob83a]. Thompson [Met93, Sid99]. Thomson [Abr72a]. Thoreau [Alb81]. Thorkild [Sjö99]. Thorne [Mil98]. Thorsten [Wol96]. Thought [Ant63, Bod63, Cla64, Haa50b, Kli73, May95, Mor80a, Mur61b, O’M77, Pel61, Puk57]. Thoughts [Dav86, Fry88, Pat84, Sch97]. Threat [Zim74]. Three [Cho98, Kah55b, SW50]. Three-Headed [Cho98]. Tibet [Ekv80]. Tibetan [Ekv74]. Timber [Bor87]. Time [Bec89, Ben71, Böw97, Fry88, Ger78]. Times [Bro51, Her82, Jur66]. Timothy [WW70]. Tinbergen [Tob97]. Tissues [Cha83, Leh71]. Titian [Mur57, Por85]. Title [Bur80]. Tocqueville [Gol64a]. Tocquevilles [Pal87]. Today [Ben85, Com56, Gid50, Kee63, Lev62, Mor80b, Par72, Taf62]. Todd [Cor99, Tod92]. Together [Bro89, Coh50]. Tokugawa [Fle66]. Tolerance [Sto55]. Toleration [For72, Zak91]. Tolowa [Gou66]. Tolstoy [DM52, Mag85]. Tomb [Ver65]. Tomorrow [Kee63]. Tonality [Lev88]. Tondo [Ver65]. Tone [Löw91]. Too [Sim77]. Tool [Arr99, Mac53, Noy51]. Tools [CS53]. Topography [Pal81]. Topophysis [Rob64]. Torsional [See51]. Tory [dLL70]. Totalitarianism [Fry98]. Tou [VW78]. Toumba [VW78]. Tour [Bar74, Sno58]. Tovar [Par77]. Toxin [Van70]. Tracking [Whi58a]. Tract [Amd70]. Trade [Fle66, Jen66, Leo53, Lew68, Vog73]. Trades [Kub59]. Trades-Unionism [Kub59]. Trading [Joh74]. Tradition [Ada66, Cur61, Gra98, Gum55, Maz55, Ril88a, Ven99, Wal52]. Traditional [Bod63, Min73, Osw64]. Traditions [Ada91]. Tragedy [VD53]. Traill [Ste63a]. Trains [Oli57, Oli69]. traité [Ewa59, Jac69]. Trans [Ser91]. Trans-Pontem [Ser91]. Transfer [McG67, Sta91]. Transfiguration [Jen68]. Transfigurations [Scr91]. Transformations [Har64, Ive51, Roa66]. Transistors [Key92]. Transit [Coh50]. Transition [Sil95, Tay99]. Translation [Dum51, Dum54, Dum59, Dum60, Dum61, Dum62, Dum63a, Dum63b, Dum64, Dum65, Dum69, Kai70, Lei90, Lew80, Ric80, Swe73, Tri80, Yng64]. Translations [Cra58]. Translator [Cla82]. Transmission [Pye78b, Ste59b]. Transmutation [Opp82]. Transplantation [Bil71, Shu71, Sta71]. Transplants [Fre71]. Transport [Kar99]. Transportation [Mit62]. Trastevere [Pal81]. Traube [Eds85]. Travel [Col72, Dry60]. Traveler [DM53a]. Traveling [Wer81]. Travels [Cap74, Gra96, dT58b]. Treason [Pie65, Woo99]. Treasures [Ada91, Fen71]. Treaties [Sch60]. Treatise [Dar79, Gal74]. Treatment [BG91, Joh65, SW77, Zub76]. Treaty [Ega71]. Tree [Gou97, Sch58]. Trees [Ewa59, Hoe90]. Trend [Ada66, Bor64]. Trends [Bar54, Dal66, Mal55, Myr51, Osb53, Sul68, Tud69]. Tribe [Riv58]. Tribute [Jul87]. Trick [Nob68]. Trigrams [Cam91]. Triopion [Pen57]. Triumph [Hea95, dLL70]. Trouble [Gra93a]. True [Zal91]. Truk [Goo81]. Trusteeship [McG99]. Truth [Bro72]. Try [Wel74]. Tsar [Bow99]. Tsarevich [Bus97]. Tsetse [Lam80]. Tuberculosis [Lon57]. Tuchman [Bat96]. Tudor [McI75]. Tumor [HH76]. Turbulence [Goo90]. Turgot [Nis75, Gil72]. Turin [Kit73]. Turkish [Set69]. Turner [Jur66]. Tuzo [Str97]. Twelfth [Dum51, Bis79]. Twelve [Cha87, Ste69]. Twentieth [Woo53, Nic56]. Twenty [Car89a, Dub77, Tau76, Zei88]. Twenty-First [Zei88]. Twenty-Five [Dub77]. Twilight [Bur88]. Two [Can61, Chi59, Coh52b, Coh82, DK54, Dys96, Ehr90, Gla81, Lan76, Mae80, Mal98, Pal87, Roa66, Ros66, Ted87]. Type [Mil55]. Types [Bie57, Gre80]. Typology [Swi98]. Tyranny [Fry98]. Tze [Hau67]. Tzintzuntzan [Pol77].

U.S. [Atk99, Gom92, Hum89, Lan59, NRF81, Wal80]. U.S.S.R. [Hum89]. UFO [Con69]. Ulpian [Sym72]. Ulrich [Gre94]. Ultracentrifuge [Bea57]. Unattractive [Reu97]. Uncertainty [BG77]. Unchanging [Whi70a]. Under-Developed [Cop56b]. Underclass [Gal92, Moy92, Saw92]. Underclasses [Pat97]. Underground [Gar61, Rid52, Wil61c]. Understand [Fon71, Shi51]. Understanding [Ber98, Car90a, Det81, Her55, Rai86]. Undocumented [Esp92]. Uneasy [Dai95a]. Unemployment [Ash78, McG96, Red78]. UNESCO [Han53b]. Unexpected [McC84]. Ungava [Edw61]. Unification [Cra91]. Uniform [Kan61]. Uniformitarian [RC67]. Uniformitarianism [Can61]. Unimportant [Fri66]. Union [Bur80, Wil87, Zie82, Viv74]. Unionism [Kub59, Nem68b]. United [Rob57a, Ano89g, Ber54, Big54, Bor65, Bra72b, Bry88, Clo80, Coc64, Coh82, Esp92, Gem89, Gid50, Gra89, Hea51, Hen87, Hur60, Jac80, Kir56, Kot51, McN79, Mur56, Nic67, Plo78, RT99, Roo72, Ros89, Rub63, Shu58, Tei92, dT58b, dT60]. Universal [Rat76]. Universe [Fri91a, Fri97, Hac97, Rub88a, Rub88b, Urt81]. Universities [Adr88, Fre93, Shr61]. University [Con60, Hea71, Kat70, Keo98, Sch56b, Wie71, Wil61a, Gol92b, Lan55, Lew56b, Mos55b]. Unmasked [Con67]. Unofficial [Sok72]. Unorthodox [Rid52]. Unpublished [KRF55, Kun86, Ted87, Tre74, Wil61b]. Unrest [Pot75]. Unsettled [Goo53]. until [Her82]. upon [Mur61a, Mur51, Nic61a, Rob83a, SC84]. Upper [How59, RGSS72]. Uprising [Rot77]. Uranographia [Ash81]. Urban [Dil68, Kim77, Kuz74, Per62, Pol77]. Uru [Jul87]. Usage [Jur66]. Use [App84, Fol58, Hil64, Kar60, Mea54, Mul62, SE98, Wri50b]. Used [Cop53, Llo50]. Useful [Arr99, Len66, Mas92, Pie79, Rei75, Rob89, Tho90]. Usefulness [Ven99]. Uses [D’E70, Mar53]. Using [TAT53]. USSR [Lum78]. Ustrialov [Bus97]. Utilization [Dan71]. Utopia [Rob55].

V [Mac97, Sto99b, Vry99, Big54, Ken58]. V. [Loc97]. Vaccine [Aus81]. Vacui [Ett79]. Valencia [Rus62]. Validation [Bor52]. Valley [LVF98, Pad77, Bec89, Cha80, Cro73, Gru89, Jak68, Jam54, Jen66]. Value [Hof75]. Values [Fel87, Klu58, Lan99b, Pol85]. Valvular [Joh65]. Vandeveer [NRF81]. Variables [Kuz58, Qui60]. Variation [Dob59, Mün59, Ste59b]. Variations [Rob79]. Varieties [Cla50]. Vascular [TZ66]. Vases [Nob68, Ver58]. Vatican [Alb60]. Vaughan [GS94]. Vegetational [Kle82]. Vehicle [Kam87]. Veins [Esa67]. Velocities [GCM94]. Venant [See51]. Vendôme [Gal74]. Venerable [Han53b]. Venezuela [Gus56, Mag56]. Venice [Fin82, Mei66, MM72, Top76]. Venture [Top76]. Venus [Gol69b, Men69b, Swi69]. Veracruz [GB94]. Verbal [For71]. Vergilius [Lew54]. Vermeer [Bré68]. Verner [Gre98c]. Veronica [EoL99]. Versatile [Lef71]. Verses [Mal50]. Version [Met91, Sta60]. versus [Con67, May78, May89, Nic64, Plo78]. Vertebrate [Rai92]. Vertebrates [Rom56]. Very [Ham92]. Vichy [Kie80]. Victorian [Bay69, Sok72, Wes98b]. Victorious [Far82]. Vienna [Boy74]. View [Aar74, Bro62a, Enr75, Fab61, Gol69b, Gom95, Ker54, Ler59, Nic62, Red52, Ros64, Tib54, Zei88]. Viewed [Bla73b, Gey62, Mae80, McB56]. Views [Gri72b, KB82, Mün59, Poe56, Shu58, Smi86]. Vill [Aul82]. Villa [Gil76]. Village [Lew56a, LBWD86, Mel57]. Vincent [HE96]. Vinogradov [CV92]. Violence [Bro70, Ell83]. Viral [End57]. Virgil [Sol68]. Virgin [Wol82]. Virginia [But51, McG99, Ste52, Zie82]. Virtue [Pay88]. Viruses [Hor58, Spi76, Sta57]. Visibilities [Neu51]. Vision [Moe58, Rob55, Son65]. Visit [DS56, Llo50, Sny57, Wil80, dT60]. Vison [End52]. Visual [Fry76, KS91, Thu50, vH96]. Vitamins [Jef58]. vitro [HH76]. Vleck [And94]. Vocal [Lil62]. Voice [HC88]. Vol [Ano88m]. Volta [Pac55]. Voltage [Key92]. Voltage-Gating [Key92]. Voltaire [Wad57]. Volume [Ano50s, Ano51s, Ano52s, Ano53s, Ano54s, Ano55t, Ano56u, Ano57s, Ano58s, Ano59s, Ano60t, Ano61s, Ano62t, Ano63t, Ano64s, Ano65w, Ano66s, Ano67s, Ano68s, Ano69s, Ano70s, Ano71s, Ano72t, Ano73s, Ano74s, Ano75s, Ano76s, Ano77s, Ano78s, Ano79t, Ano79u, Ano80s, Ano81s, Ano82s, Ano83s, Ano84m, Ano85m, Ano86m, Ano87m, Ano88n, Ano89l, Ano90i, Ano91j, Ano92l, Ano93j, Ano94i, Ano95j, Ano96m, Ano97n, Ano98n, Ano99m, Mac65b, Shi70]. Voluntary [And57, Coa67, Lon57, Par57]. Volunteer [Rus57]. Voprosy [Nic61c]. Vowel [Sza72]. Voyage [Olm60, Rau73]. Voyagers [Lan89]. Voyages [Cla54]. vs [Ano63m, Kno63, Pet63, Thi93].

W [Coh95, Kam86, Tob91]. W. [RC67]. Wagon [BL57]. Wake [Lan89]. Walcott [Yoc96]. Wales [Bar74]. Wall [Kra58]. Wallace [Kot80]. Walle [Sch95b]. Walter [Par72, Wil94a, Eis50, Fir91, Hen97, Tob91]. Wampum [Fen71, Sny54, Sny61, Sny82]. Wants [Bau93]. War [Cla78, Cre85, Joh74, Kle71, Lan72, Lew68, Mer71, Mer78, Pie80, Sch63, Tal99, vD77, Ald50a, Bur80, Dur62, Fin82, Gey62, Jon63a, McP99, Nic62, Sch59]. War-II [Mer78]. Ward [Ega78]. Warfare [But50, Gar80, Mor80b]. Warren [Woo92, O’C95]. Wars [Dur62, Gla81]. Warwick [Don76]. Was [Ech54, Gre92, KS84, McP99, Men89, Ros60, Sal50]. Washington [Fra79, Dau56, Lan59, McG99, Mor58, Pom54, Shy87]. Wasted [Ano92e]. Watchdog [Maa81]. Water [Kra73, Las75, McP99, Pat94b, Rid52]. Waterman [Was73]. Waterworks [Maa81]. Watson [dK95]. Waugh [Ran53]. Wave [Sto59]. Waves [Swe96, TAT53]. Way [Hea88, Bok52]. Wayland [Mad62]. Ways [Hij80]. Wealth [Gou66, Har71a]. Wealthy [Gou89]. Weapons [Gar80, Pie80, Wil61c]. Weapons/Old [Gar80]. Weather [Rob79]. Weavers [Min73]. Webb [Sti59]. Webster [She85]. Wedgwood [EoL99]. Weekend [McG97]. Weeks [Hij80]. Weighs [McC98]. Weight [RD73]. Weitzmann [Kit95]. Welcome [Mos97, Sal97]. Welfare [Mah91, Moo60, Rob98]. Wellek [Gre98b]. Weltbild [Hol81]. Wendell [Bur75]. Wentworth [Hof50]. Were [For96]. Werner [Lin94a]. West [Cor72, Mal53, Can77, Lan89, Rau73, Sim55, Ste89, oB90]. Western [Col72, Cop56b, Dav52, Dav53, Fen56, Gor88, Gus56, Hop61, Kro51, LBWD86, McD59, Nor73, Pet73, RK82, Whi54]. Westernization [Sch95a]. Wet [Lur72]. Wethersfield [Jod87]. Wette [Puk58]. Wheel [Kit73]. Where [Has61, Llo89]. Which [CR52, Jar72, McG99, Ver64]. White [Col72, Mor77, Bun91b]. White-Crowned [Mor77]. Whither [oB90]. Whitney [Che94]. Who [Cop55, McP95, Nic61c, Ric51b, Wal51c]. Whole [Sta71]. Whose [Ost95]. Wieland [WE98]. Wiesner [Sax96]. Wigglesworth [Loc97]. Wigner [Wei97]. Wild [Car89a, Car89b]. Wilderness [Glu56]. Wilhelm [Gil82]. Wilkes [Bor65]. Will [Her64, Sho93, Wil51b]. Willard [Wil61b]. William [Bel50, Can77, Sio51, Tho71, Woo98, Ada93, Bar90a, Bro51, Bur60a, Cla82, Cor94, Cor50, Dav61, DeM84, Dun83a, Dun83b, Dur87, DM53b, For62, Fus57, Gob70, Har53, Hof91, Hos89, Kni79, Lop85, Mor83, Nat91, Por74, QC53, Rai86, Ric53, Sal98, Sch70, Sha58, Tur71, VK70, Wol95a]. Williams [KWB97, Fra79]. Williamsburg [Wer53]. Willis [Abr72a]. Willmer [Gel95b]. Willson [Law92]. Wilson [GHKW95, Str97, Bro85, Ker82, Lin68, Sey57]. Window [Rad93]. Wings [Gre60c]. Winning [Lin51]. Winter [Car90a]. Wireless [SE98]. Wit [Gay91]. Witchcraft [Dar79, Sny83, Ted87]. Witches [Sny83]. Witt [Hop61]. Wolcott [Bog99]. Wolf [vH93]. Wolff [Ric80]. Woman [WD96]. Women [Ben91, Bie62, Bur73, Ler77, Lew94]. Woodbine [Wil72]. Woodrow [Ker82, Lin68, Sey57]. Word [Cun57, Jur66]. Wordsworth [Pot72]. Work [Bro78, Coh50, Cop51, Dun65, Fen56, Gle60, Haa80, Har73, Kun86, Mal50, Reu97]. Worker [Bar54]. Workforce [Mor60]. Works [Ewa59, Ros66, Sca76, Scr91]. Workshop [TNN71]. World [Adr56, Ben85, Cas69, Com56, Dav61, Dur67, Kit73, Lew68, Man65, Mar98, McC98, Mor83, Pat70b, Red52, Sch59, Sch95c, Spi83, Spi93, Tol51b, Toy55, Wal51a, Bar99, Bor87, Cla78, Cre85, Eas65, Eng68, Gla81, Han63, Kra50, Wil66]. Worlds [PG52]. Worldwide [For80, Gol91]. Worth [Goo90]. Wrens [GJ55]. Wright [Fen56]. Writer [McK73]. Writers [Ler77]. Writing [Cra78, Lin52a, Mos87, Pot62]. Writings [Fre97, Ler77, O’M77]. Written [Kim93]. Wrote [Wal51c].

X [Fri97, Per69, Pou97, Set69]. X-Ray [Per69, Pou97]. X-Rays [Fri97]. Xanadu [Por74]. Xenia [Dun73]. XI [Dow59]. XIII [Kam71]. Xochipala [Gri72a]. Xystodesmidae [CT69].

Yacht [Cha71, Ste69]. Yachts [Cha87]. Yakov [Fri91b]. Yale [Fre66, Lew56b]. Yankee [Mil82b]. Yardstick [Bla73a, See61]. Yarrell [For62]. Year [Gom95, Kle82, Neu62b, Nor73, Ber57]. Yearly [Sny58]. Years [Car89a, Dub77, Goo58, Kra73, NRF81, Ric70, Sli98, Tau76, War77b]. Yellow [Leh59]. Yellow-Legged [Leh59]. Yesterday [Par72]. Yezo [HR55]. Yiddish [Wei59]. Yield [Jen51]. York [Fen71, Hau75, Mal55, DS56, Fen56, Sny57]. Young [Chr84, Key99, Hil78, Pet63, Pom54]. Youth [Fai64, Mus69]. Yucatan [And60]. Yuk [Osw64]. Yupa [Gus56]. Yupik [KB80].

Z [Wal99]. Zachary [Key99]. Zariski [Bir93b]. Zay [Mac65b]. Zea [ES56a, ES56b]. Zealand [Fle75, Mur51]. Zel’dovich [Fri91b]. Zellig [Hiż94]. Zibaldone [Sca76]. Zin [Glu56]. Zoning [Pol77]. Zoo [Rat68]. Zoogeographic [Sim77]. Zoogeography [Dar59a]. Zuckerman [Dai95b]. Zuni [Pan72]. Zuriel [Was73]. Zusetsu [HR55].

References

Aaron:1974:SDR

[Aar74] Benjamin Aaron. The settlement of disputes over rights: a comparative view. Proceedings of the American Philosophical Society, 118 (5):422–430, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986307.

Abrams:1962:LBR

[Abr62] Charles Abrams. The legal basis for reorganizing metropolitan areas in a free society. Proceedings of the American Philosophical Society, 106(3):177–189, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985173.

Abrahams:1972:RCA

[Abr72a] Harold J. Abrahams. The Ritchey–Chrétien aplanatic telescope: Letters from George Willis Ritchey to Elihu Thomson. Proceedings of the American Philosophical Society, 116(6):486–501, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985762.

Abrams:1972:CLS

[Abr72b] Meyer Howard Abrams. Coleridge’s “a light in sound”: Science, metascience, and poetic imagination. Proceedings of the American Philosophical Society, 116(6):458–476, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985760.

Abraham:1987:RSC

[Abr87] Henry J. Abraham. Reflections on the supreme court appointment process. Proceedings of the American Philosophical Society, 131(3): 270–283, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987022.

Abrams:1991:EAB

[Abr91] M. H. Abrams. Eric Albert Blackall (October 19, 1914–November 16, 1989). Proceedings of the American Philosophical Society, 135 (2):300–306, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987036.

Adams:1966:TTN

[Ada66] Robert McC. Adams. Trend and tradition in Near Eastern archaeology. Proceedings of the American Philosophical Society, 110(2): 105–110, April 1966. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985743.

Adams:1978:SMS

[Ada78] Robert McC. Adams. Strategies of maximization, stability, and resilience in Mesopotamian society, settlement, and agriculture. Proceedings of the American Philosophical Society, 122(5):329–335, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986687.

Adams:1984:RFJ

[Ada84] Arlin M. Adams. The role of the Federal Judiciary. Proceedings of the American Philosophical Society, 128(3):231–237, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986882.

Adams:1987:MP

[Ada87a] Arlin Adams. The miracle at Philadelphia. Proceedings of the American Philosophical Society, 131(3):239–242, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987018.

Adams:1987:RAA

[Ada87b] Arlin M. Adams. Remarks addressed to the Anglo–American Successor Generation Conference. Proceedings of the American Philosophical Society, 131(1):112–119, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987118.

Adams:1991:MBT

[Ada91] Robert McC. Adams. Museums beyond treasures and traditions. Proceedings of the American Philosophical Society, 135(3):332–338, September 1991. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986771.

Adams:1993:WPA

[Ada93] Arlin M. Adams. William Penn and the American heritage of religious liberty. Proceedings of the American Philosophical Society, 137(4):516–523, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987070.

Adams:1997:R

[Ada97] Arlin M. Adams. Response. Proceedings of the American Philosophical Society, 141(4):394–396, December 1997. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987215.

Adrian:1956:ASM

[Adr56] Lord Adrian. Academies of science in the modern world. Proceedings of the American Philosophical Society, 100(4):326–330, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143758.

Adrian:1988:HLJ

[Adr88] Richard H. Adrian. The Henry LaBarre Jayne Lecture: The crisis in British universities. Proceedings of the American Philosophical Society, 132(3):237–246, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143850.

Adrian:1992:DWN

[Adr92] Lord Adrian. Do we need science policies? Proceedings of the American Philosophical Society, 136(4):526–532, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986761.

Asch:1962:PAS

[AE62] Solomon E. Asch and Sheldon M. Ebenholtz. The principle of associative symmetry. Proceedings of the American Philosophical Society, 106(2):135–163, April 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985378.

Aldridge:1954:HPB

[AF54] Alfred Owen Aldridge and Benjamin Franklin. A humorous poem by Benjamin Franklin. Proceedings of the American Philosophical Society, 98(6):397–399, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143861.

Alpher:1975:BBC

[AH75] Ralph A. Alpher and Robert Herman. Big bang cosmology and the cosmic black-body radiation. Proceedings of the American Philosophical Society, 119(5):325–348, October 1975. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986676.

Ahrens:1999:FBA

[Ahr99] Thomas J. Ahrens. Francis Birch (22 August 1903–30 January 1992). Proceedings of the American Philosophical Society, 143(4): 652–661, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181993.

Albareda:1960:PRM

[Alb60] Dom Anselm M. Albareda. The preservation and reproduction of the manuscripts of the Vatican Library through the centuries. Proceedings of the American Philosophical Society, 104(4):413–418, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985621.

Albright:1972:NFG

[Alb72] William F. Albright. Neglected factors in the Greek intellectual revolution. Proceedings of the American Philosophical Society, 116 (3):225–242, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986117.

Albert:1981:TII

[Alb81] David H. Albert. Thoreau’s India: The impact of reading in a crisis. Proceedings of the American Philosophical Society, 125(2):104–109, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986639.

Aldridge:1950:FJF

[Ald50a] Alfred Owen Aldridge. Franklin and Jackson on the French War. Proceedings of the American Philosophical Society, 94(4):396– 397, August 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143525.

Aldridge:1950:FDI

[Ald50b] Alfred Owen Aldridge. Franklin’s deistical Indians. Proceedings of the American Philosophical Society, 94(4):398–410, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143526.

Aldridge:1950:FLI

[Ald50c] Alfred Owen Aldridge. Franklin’s letter on Indians and Germans. Proceedings of the American Philosophical Society, 94(4): 391–395, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143524.

Aldridge:1951:JBD

[Ald51] Alfred Owen Aldridge. Jacques Barbeu-Dubourg, a French disciple of Benjamin Franklin. Proceedings of the American Philosophical Society, 95(4):331–392, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143280.

Aldridge:1962:BFP

[Ald62] Alfred Owen Aldridge. Benjamin Franklin and the “Pennsylvania Gazette”. Proceedings of the American Philosophical Society, 106 (1):77–81, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985213.

Alden:1976:SCP

[Ald76] Dauril Alden. The significance of cacao production in the Amazon Region during the Late Colonial Period: an essay in comparative economic history. Proceedings of the American Philosophical Society, 120(2):103–135, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986517.

Allen:1960:PC

[All60] Don Cameron Allen. The predecessors of Champollion. Proceedings of the American Philosophical Society, 104(5):527–547, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985236.

Allen:1966:THM

[All66] Garland E. Allen. Thomas Hunt Morgan and the problem of sex determination, 1903–1910. Proceedings of the American Philosophical Society, 110(1):48–57, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986001.

Alpert:1955:SSN

[Alp55] Harry Alpert. The social sciences and the national science foundation. Proceedings of the American Philosophical Society, 99(5):332– 333, October 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143931.

Amadon:1959:SSD

[Ama59] Dean Amadon. The significance of sexual differences in size among birds. Proceedings of the American Philosophical Society, 103 (4):531–536, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985554.

Amdur:1970:IAP

[Amd70] Mary O. Amdur. The impact of air pollutants on physiologic responses of the respiratory tract. Proceedings of the American Philosophical Society, 114(1):3–8, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985718.

Amsterdam:1991:EER

[Ams91] Anthony G. Amsterdam. Equality as an enforceable right: Problems and prospects ahead. Proceedings of the American Philosophical Society, 135(1):13–21, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987145.

Ancel:1974:PAM

[Anc74] Marc Ancel. Principal aspects of modern European penology. Proceedings of the American Philosophical Society, 118(3):254–259, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986443.

Anderson:1957:VAH

[And57] Gaylord W. Anderson. The voluntary agency and the health department. Proceedings of the American Philosophical Society, 101 (2):149–152, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985506.

Andrews:1960:EDN

[And60] E. Wyllys Andrews. Excavations at Dzibilchaltun, Northwestern Yucatan, Mexico. Proceedings of the American Philosophical Society, 104(3):254–265, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985243.

Andrews:1961:GPS

[And61] F. Emerson Andrews. Growth and present status of American foundations. Proceedings of the American Philosophical Society, 105(2):157–161, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985627.

Anderson:1976:TMF

[And76] Gillian B. Anderson. “The Temple of Minerva” and Francis Hopkinson: a reappraisal of America’s first poet–composer. Proceedings of the American Philosophical Society, 120(3):166–177, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986556.

Anderson:1994:JHV

[And94] Philip W. Anderson. John Hasbrouck Van Vleck (13 March 1899– 27 October 1980). Proceedings of the American Philosophical Society, 138(1):184–187, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986714.

Anderson:1999:JBP

[And99] William P. Anderson. James Bennett Pritchard (4 October 1909–1 January 1997). Proceedings of the American Philosophical Society, 143(3):472–477, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181960.

Angel:1952:HSR

[Ang52] J. Lawrence Angel. The human skeletal remains from Hotu Cave, Iran. Proceedings of the American Philosophical Society, 96(3): 258–269, June 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143835.

Anonymous:1950:Aa

[Ano50a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(1):??, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:Ab

[Ano50b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(2):??, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:Ac

[Ano50c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(3):??, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:Ad

[Ano50d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(4):??, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:Ae

[Ano50e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(5):??, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:Af

[Ano50f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 94(6):??, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1950:BMa

[Ano50g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(1):??, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143253.

Anonymous:1950:BMb

[Ano50h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(2):??, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143218.

Anonymous:1950:BMc

[Ano50i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(3):??, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143568.

Anonymous:1950:BMd

[Ano50j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(4):??, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143527.

Anonymous:1950:BMe

[Ano50k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(5):??, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143617.

Anonymous:1950:BMf

[Ano50l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 94(6):??, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143596.

Anonymous:1950:FMa

[Ano50m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(1):??, February 27, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143244.

Anonymous:1950:FMb

[Ano50n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(2):??, April 21, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143206.

Anonymous:1950:FMc

[Ano50o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(3):??, June 20, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143552.

Anonymous:1950:FMd

[Ano50p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(4):??, August 25, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143512.

Anonymous:1950:FMe

[Ano50q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(5):??, October 19, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143607.

Anonymous:1950:FMf

[Ano50r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 94(6):??, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143591.

Anonymous:1950:VI

[Ano50s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 94(6):611–iv, December 22, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143590.

Anonymous:1951:Aa

[Ano51a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(1):??, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:Ab

[Ano51b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(2):??, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:Ac

[Ano51c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(3):??, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:Ad

[Ano51d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(4):??, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:Ae

[Ano51e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(5):??, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:Af

[Ano51f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 95(6):??, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1951:BMa

[Ano51g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(1):??, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143578.

Anonymous:1951:BMb

[Ano51h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(2):??, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143333.

Anonymous:1951:BMc

[Ano51i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(3):??, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143071.

Anonymous:1951:BMd

[Ano51j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(4):??, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143287.

Anonymous:1951:BMe

[Ano51k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(5):??, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143243.

Anonymous:1951:BMf

[Ano51l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 95(6):??, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143511.

Anonymous:1951:FMa

[Ano51m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 95(1):??, February 13, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143569.

Anonymous:1951:FMb

[Ano51n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 95(2):??, April 30, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143324.

Anonymous:1951:FMc

[Ano51o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 95(3):??, June 12, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143055.

Anonymous:1951:FMd

[Ano51p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 95(4):??, August 17, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143279.

Anonymous:1951:FMe

[Ano51q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 95(5):??, October 17, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143230.

Anonymous:1951:FMf ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143505.

Anonymous:1951:VI

[Ano51s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 95(6):676–iv, December 21, 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143504.

Anonymous:1952:Aa

[Ano52a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(1):??, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:Ab

[Ano52b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(2):??, April 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:Ac

[Ano52c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(3):??, June 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:Ad

[Ano52d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(4):??, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:Ae

[Ano52e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(5):??, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:Af

[Ano52f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 96(6):??, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1952:BMa

[Ano52g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(1):??, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143744.

Anonymous:1952:BMb

[Ano52h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(2):??, April 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143722.

Anonymous:1952:BMc

[Ano52i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(3):??, June 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143839.

Anonymous:1952:BMd

[Ano52j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(4):??, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143805.

Anonymous:1952:BMe

[Ano52k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(5):??, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143662.

Anonymous:1952:BMf

[Ano52l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 96(6):??, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143638.

Anonymous:1952:FMa

[Ano52m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 96(1):??, February 29, 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143737.

Anonymous:1952:FMb ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143717.

Anonymous:1952:FMc

[Ano52o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 96(3):??, June 20, 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143831.

Anonymous:1952:FMd

[Ano52p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 96(4):??, August 22, 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143793.

Anonymous:1952:FMe

[Ano52q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 96(5):??, October 15, 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143649.

Anonymous:1952:FMf

[Ano52r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 96(6):??, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143632.

Anonymous:1952:VI

[Ano52s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 96(6):756–iv, December 20, 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143631.

Anonymous:1953:Aa

[Ano53a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(1):??, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:Ab

[Ano53b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(2):??, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:Ac

[Ano53c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(3):??, June 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:Ad

[Ano53d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(4):??, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:Ae

[Ano53e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(5):??, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:Af

[Ano53f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 97(6):??, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1953:BMa

[Ano53g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(1):??, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143736.

Anonymous:1953:BMb

[Ano53h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(2):??, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143897.

Anonymous:1953:BMc

[Ano53i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(3):??, June 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143751.

Anonymous:1953:BMd

[Ano53j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(4):??, September 1953. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3149294.

Anonymous:1953:BMe

[Ano53k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(5):??, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3149276.

Anonymous:1953:BMf

[Ano53l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 97(6):??, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3149285.

Anonymous:1953:FMa

[Ano53m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 97(1):??, February 14, 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143723.

Anonymous:1953:FMb

[Ano53n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 97(2):??, April 30, 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143885.

Anonymous:1953:FMc

[Ano53o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 97(3):??, June 30, 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143745.

Anonymous:1953:FMd

[Ano53p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 97(4):??, September 28, 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3149286.

Anonymous:1953:FMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3149264.

Anonymous:1953:FMf

[Ano53r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 97(6):??, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3149278.

Anonymous:1953:VI

[Ano53s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 97(6):713–iv, December 16, 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3149277.

Anonymous:1954:Aa

[Ano54a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(1):??, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:Ab

[Ano54b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(2):??, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:Ac

[Ano54c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(3):??, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:Ad

[Ano54d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(4):??, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:Ae

[Ano54e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(5):??, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:Af

[Ano54f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 98(6):??, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1954:BMa

[Ano54g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(1):??, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143672.

Anonymous:1954:BMb

[Ano54h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(2):??, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143648.

Anonymous:1954:BMc

[Ano54i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(3):??, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143847.

Anonymous:1954:BMd

[Ano54j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(4):??, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143884.

Anonymous:1954:BMe

[Ano54k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(5):??, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143915.

Anonymous:1954:BMf

[Ano54l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 98(6):??, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143871.

Anonymous:1954:FMa

[Ano54m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(1):??, February 15, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143663.

Anonymous:1954:FMb

[Ano54n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(2):??, April 15, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143639.

Anonymous:1954:FMc

[Ano54o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(3):??, June 15, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143840.

Anonymous:1954:FMd

[Ano54p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(4):??, August 16, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143872.

Anonymous:1954:FMe

[Ano54q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(5):??, October 15, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143905.

Anonymous:1954:FMf

[Ano54r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 98(6):??, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143859.

Anonymous:1954:VI

[Ano54s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 98(6):495–iv, December 23, 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143858.

Anonymous:1955:Aa

[Ano55a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(1):??, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Ab

[Ano55b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(2):??, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Ac

[Ano55c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(3):??, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Ad

[Ano55d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(4):??, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Af

[Ano55e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(5):??, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Ag

[Ano55f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 99(6):??, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1955:Ae

[Ano55g] Anonymous. Appendix. Proceedings of the American Philosophical Society, 99(5):352–354, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143937.

Anonymous:1955:BMa

[Ano55h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(1):??, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143694.

Anonymous:1955:BMb

[Ano55i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(2):??, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143821.

Anonymous:1955:BMc

[Ano55j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(3):??, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143783.

Anonymous:1955:BMd

[Ano55k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(4):??, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143708.

Anonymous:1955:BMe

[Ano55l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(5):??, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143938.

Anonymous:1955:BMf

[Ano55m] Anonymous. Back matter. Proceedings of the American Philosophical Society, 99(6):??, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143928.

Anonymous:1955:FMa

[Ano55n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(1):??, January 27, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143686.

Anonymous:1955:FMb

[Ano55o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(2):??, April 15, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143812.

Anonymous:1955:FMc

[Ano55p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(3):??, June 15, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143776.

Anonymous:1955:FMd

[Ano55q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(4):??, August 30, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143695.

Anonymous:1955:FMe

[Ano55r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(5):325, October 15, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143929.

Anonymous:1955:FMf

[Ano55s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 99(6):??, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143917.

Anonymous:1955:VI

[Ano55t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 99(6):474–iv, December 15, 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143916.

Anonymous:1956:Aa

[Ano56a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(1):??, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:Ab

[Ano56b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(2):??, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:Ac

[Ano56c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(3):??, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:Ad

[Ano56d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(4):??, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:Ae

[Ano56e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(5):??, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:Af

[Ano56f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 100(6):??, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1956:BMa

[Ano56g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 100(1):??, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143811.

Anonymous:1956:BMb

[Ano56h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 100(2):??, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143716.

Anonymous:1956:BMc

[Ano56i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 100(3):??, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143775.

Anonymous:1956:BMd 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143768.

Anonymous:1956:BMe

[Ano56k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 100(5):??, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143685.

Anonymous:1956:BMf

[Ano56l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 100(6):??, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143792.

Anonymous:1956:CFE

[Ano56m] Anonymous. Catalogue of Franklin Exhibition in the Library of Congress. Proceedings of the American Philosophical Society, 100 (4):385–416, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143767.

Anonymous:1956:FMa

[Ano56n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(1):??, February 24, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143806.

Anonymous:1956:FMb

[Ano56o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(2):??, April 23, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143709.

Anonymous:1956:FMc

[Ano56p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(3):??, June 28, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143769.

Anonymous:1956:FMd

[Ano56q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(4):??, August 31, 1956. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143752.

Anonymous:1956:FMe

[Ano56r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(5):??, October 15, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143673.

Anonymous:1956:FMf

[Ano56s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 100(6):??, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143785.

Anonymous:1956:IPC

[Ano56t] Anonymous. Idea and plan of the celebation. Proceedings of the American Philosophical Society, 100(4):338–339, August 1956. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143760.

Anonymous:1956:VI

[Ano56u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 100(6):613–iv, December 17, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143784.

Anonymous:1957:Aa

[Ano57a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(1):??, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:Ab

[Ano57b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(2):??, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:Ac

[Ano57c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(3):??, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:Ad

[Ano57d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(4):??, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:Ae

[Ano57e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(5):??, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:Af

[Ano57f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 101(6):??, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1957:BMa

[Ano57g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 101(1):??, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985144.

Anonymous:1957:BMb

[Ano57h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 101(2):??, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985515.

Anonymous:1957:BMc

[Ano57i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 101(3):??, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985372.

Anonymous:1957:BMd

[Ano57j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 101(4):??, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985407.

Anonymous:1957:BMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985203.

Anonymous:1957:BMf

[Ano57l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 101(6):??, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985526.

Anonymous:1957:FMa

[Ano57m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 101(1):??, February 15, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985136.

Anonymous:1957:FMb

[Ano57n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 101(2):??, April 19, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985503.

Anonymous:1957:FMc

[Ano57o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 101(3):??, June 20, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985365.

Anonymous:1957:FMd

[Ano57p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 101(4):??, August 16, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985396.

Anonymous:1957:FMe

[Ano57q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 101(5):??, October 31, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985192.

Anonymous:1957:FMf ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985517.

Anonymous:1957:VI

[Ano57s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 101(6):589–iv, December 19, 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985516.

Anonymous:1958:Aa

[Ano58a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(1):??, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:Ab

[Ano58b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(2):??, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:Ac

[Ano58c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(3):??, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:Ad

[Ano58d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(4):??, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:Ae

[Ano58e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(5):??, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:Af

[Ano58f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 102(6):??, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1958:BMa ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985306.

Anonymous:1958:BMb

[Ano58h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 102(2):??, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985285.

Anonymous:1958:BMc

[Ano58i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 102(3):??, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985578.

Anonymous:1958:BMd

[Ano58j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 102(4):??, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985226.

Anonymous:1958:BMe

[Ano58k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 102(5):??, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985599.

Anonymous:1958:BMf

[Ano58l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 102(6):??, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985571.

Anonymous:1958:FMa

[Ano58m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 102(1):??, February 17, 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985299.

Anonymous:1958:FMb ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985272.

Anonymous:1958:FMc

[Ano58o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 102(3):??, June 24, 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985572.

Anonymous:1958:FMd

[Ano58p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 102(4):??, August 27, 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985216.

Anonymous:1958:FMe

[Ano58q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 102(5):??, October 20, 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985585.

Anonymous:1958:FMf

[Ano58r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 102(6):??, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985564.

Anonymous:1958:VI

[Ano58s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 102(6):621–iv, December 15, 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985563.

Anonymous:1959:Aa

[Ano59a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(1):??, February 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:Ab

[Ano59b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(2):??, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:Ac

[Ano59c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(3):??, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:Ad

[Ano59d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(4):??, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:Ae

[Ano59e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(5):??, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:Af

[Ano59f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 103(6):??, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1959:BMa

[Ano59g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 103(1):??, February 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985384.

Anonymous:1959:BMb

[Ano59h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 103(2):??, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985156.

Anonymous:1959:BMc

[Ano59i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 103(3):??, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985481.

Anonymous:1959:BMd 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985562.

Anonymous:1959:BMe

[Ano59k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 103(5):??, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985428.

Anonymous:1959:BMf

[Ano59l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 103(6):??, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985395.

Anonymous:1959:FMa

[Ano59m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 103(1):??, February 28, 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985380.

Anonymous:1959:FMb

[Ano59n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 103(2):??, April 23, 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985145.

Anonymous:1959:FMc

[Ano59o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 103(3):??, June 15, 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985471.

Anonymous:1959:FMd

[Ano59p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 103(4):??, August 15, 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985553.

Anonymous:1959:FMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985418.

Anonymous:1959:FMf

[Ano59r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 103(6):??, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985386.

Anonymous:1959:VI

[Ano59s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 103(6):819–iv, December 15, 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985385.

Anonymous:1960:Aa

[Ano60a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(1):??, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:Ab

[Ano60b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(2):??, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:Ac

[Ano60c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(3):??, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:Ad

[Ano60d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(4):??, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:Ae

[Ano60e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(5):??, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:Af

[Ano60f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 104(6):??, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1960:BMa

[Ano60g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(1):??, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985608.

Anonymous:1960:BMb

[Ano60h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(2):??, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985664.

Anonymous:1960:BMc

[Ano60i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(3):??, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985251.

Anonymous:1960:BMd

[Ano60j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(4):??, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985622.

Anonymous:1960:BMe

[Ano60k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(5):??, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985237.

Anonymous:1960:BMf

[Ano60l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 104(6):??, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985541.

Anonymous:1960:DLH

[Ano60m] Anonymous. Dedication of Library Hall of the American Philosophical Society, November 11, 1959. Remarks made at the dinner at the Benjamin Franklin Hotel. Proceedings of the American Philosophical Society, 104(4):357–370, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985615.

Anonymous:1960:FMa

[Ano60n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(1):??, February 15, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985600.

Anonymous:1960:FMb

[Ano60o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(2):??, April 19, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985653.

Anonymous:1960:FMc

[Ano60p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(3):??, June 15, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985241.

Anonymous:1960:FMd

[Ano60q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(4):??, August 15, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985613.

Anonymous:1960:FMe

[Ano60r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(5):??, October 17, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985227.

Anonymous:1960:FMf

[Ano60s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 104(6):??, December 1960. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985528.

Anonymous:1960:VI

[Ano60t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 104(6):681–iv, December 15, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985527.

Anonymous:1961:Aa

[Ano61a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(1):??, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:Ab

[Ano61b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(2):??, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:Ac

[Ano61c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(3):??, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:Ad

[Ano61d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(4):??, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:Ae

[Ano61e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(5):??, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:Af

[Ano61f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 105(6):??, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1961:BMa ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985358.

Anonymous:1961:BMb

[Ano61h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 105(2):??, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985633.

Anonymous:1961:BMc

[Ano61i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 105(3):??, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985461.

Anonymous:1961:BMd

[Ano61j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 105(4):??, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985502.

Anonymous:1961:BMe

[Ano61k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 105(5):??, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985552.

Anonymous:1961:BMf

[Ano61l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 105(6):??, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985169.

Anonymous:1961:FMa

[Ano61m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 105(1):??, February 28, 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985351.

Anonymous:1961:FMb ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985623.

Anonymous:1961:FMc

[Ano61o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 105(3):??, June 27, 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985452.

Anonymous:1961:FMd

[Ano61p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 105(4):??, August 15, 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985492.

Anonymous:1961:FMe

[Ano61q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 105(5):??, October 13, 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985542.

Anonymous:1961:FMf

[Ano61r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 105(6):??, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985158.

Anonymous:1961:VI

[Ano61s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 105(6):612–iv, December 15, 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985157.

Anonymous:1962:Aa

[Ano62a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(1):??, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:Ab

[Ano62b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(2):??, April 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:Ac

[Ano62c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(3):??, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:Ad

[Ano62d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(4):??, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:Ae

[Ano62e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(5):??, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:Af

[Ano62f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 106(6):??, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1962:BMa

[Ano62g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 106(1):??, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985215.

Anonymous:1962:BMb

[Ano62h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 106(2):??, April 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985379.

Anonymous:1962:BMc

[Ano62i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 106(3):??, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985181.

Anonymous:1962:BMd 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985271.

Anonymous:1962:BMe

[Ano62k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 106(5):??, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985491.

Anonymous:1962:BMf

[Ano62l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 106(6):??, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985261.

Anonymous:1962:EHM

[Ano62m] Anonymous. Errata: Henry Muhlenberg — Botanical Pioneer. Proceedings of the American Philosophical Society, 106(6):??, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985260.

See [Smi62].

Anonymous:1962:FMa

[Ano62n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(1):??, February 15, 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985204.

Anonymous:1962:FMb

[Ano62o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(2):??, April 30, 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985373.

Anonymous:1962:FMc

[Ano62p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(3):??, June 29, 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985170.

Anonymous:1962:FMd

[Ano62q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(4):??, August 22, 1962. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985262.

Anonymous:1962:FMe

[Ano62r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(5):??, October 11, 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985482.

Anonymous:1962:FMf

[Ano62s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 106(6):??, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985253.

Anonymous:1962:VI

[Ano62t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 106(6):541–iv, December 12, 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985252.

Anonymous:1963:Aa

[Ano63a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(1):??, February 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:Ab

[Ano63b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(2):??, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:Ac

[Ano63c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(3):??, June 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:Ad

[Ano63d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(4):??, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:Ae

[Ano63e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(5):??, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:Af

[Ano63f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 107(6):??, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1963:BMa

[Ano63g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(1):??, February 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985470.

Anonymous:1963:BMb

[Ano63h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(2):??, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985449.

Anonymous:1963:BMc

[Ano63i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(3):??, June 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985584.

Anonymous:1963:BMd

[Ano63j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(4):??, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985676.

Anonymous:1963:BMe

[Ano63k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(5):??, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985652.

Anonymous:1963:BMf

[Ano63l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 107(6):??, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986111.

Anonymous:1963:EJD

[Ano63m] Anonymous. Errata: John Dickinson vs. John Adams: 1774–1776. Proceedings of the American Philosophical Society, 107(5):??, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985651. See

[Kno63].

Anonymous:1963:FMa

[Ano63n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(1):??, February 15, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985462.

Anonymous:1963:FMb

[Ano63o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(2):??, April 15, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985434.

Anonymous:1963:FMc

[Ano63p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(3):??, June 19, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985579.

Anonymous:1963:FMd

[Ano63q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(4):??, August 15, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985669.

Anonymous:1963:FMe

[Ano63r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(5):??, October 15, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985642.

Anonymous:1963:FMf

[Ano63s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 107(6):??, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986105.

Anonymous:1963:VI

[Ano63t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 107(6):542–iv, December 20, 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986104.

Anonymous:1964:Aa

[Ano64a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(1):??, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:Ab

[Ano64b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(2):??, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:Ac

[Ano64c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(3):??, June 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:Ad

[Ano64d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(4):??, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:Ae

[Ano64e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(5):??, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:Af

[Ano64f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 108(6):??, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1964:BMa

[Ano64g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(1):??, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985996.

Anonymous:1964:BMb

[Ano64h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(2):??, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985896.

Anonymous:1964:BMc

[Ano64i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(3):??, June 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985683.

Anonymous:1964:BMd

[Ano64j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(4):??, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985914.

Anonymous:1964:BMe

[Ano64k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(5):??, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985820.

Anonymous:1964:BMf

[Ano64l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 108(6):??, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985868.

Anonymous:1964:FMa

[Ano64m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(1):??, February 28, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985989.

Anonymous:1964:FMb

[Ano64n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(2):??, April 15, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985885.

Anonymous:1964:FMc

[Ano64o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(3):??, June 22, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985677.

Anonymous:1964:FMd

[Ano64p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(4):??, August 27, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985906.

Anonymous:1964:FMe

[Ano64q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(5):??, October 20, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985810.

Anonymous:1964:FMf

[Ano64r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 108(6):??, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985860.

Anonymous:1964:VI

[Ano64s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 108(6):486–iv, December 15, 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985859.

Anonymous:1965:Aa

[Ano65a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(??):??, ???? 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Ab

[Ano65b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(1):??, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Ac

[Ano65c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(2):??, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Ad

[Ano65d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(3):??, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Ae

[Ano65e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(4):??, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Af

[Ano65f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(5):??, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:Ag

[Ano65g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 109(6):??, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1965:BMa

[Ano65h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(??):??, ???? 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985858.

Anonymous:1965:BMb

[Ano65i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(1):??, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985779.

Anonymous:1965:BMc

[Ano65j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(2):??, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985788.

Anonymous:1965:BMd

[Ano65k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(3):??, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986132.

Anonymous:1965:BMe

[Ano65l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(4):??, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985884.

Anonymous:1965:BMf

[Ano65m] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(5):??, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986083.

Anonymous:1965:BMg

[Ano65n] Anonymous. Back matter. Proceedings of the American Philosophical Society, 109(6):??, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986140.

Anonymous:1965:FMa

[Ano65o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(??):??, ???? 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985856.

Anonymous:1965:FMb

[Ano65p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(1):??, February 18, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985772.

Anonymous:1965:FMc

[Ano65q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(2):??, April 9, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985780.

Anonymous:1965:FMd

[Ano65r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(3):??, June 15, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986123.

Anonymous:1965:FMe

[Ano65s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(4):??, August 18, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985876.

Anonymous:1965:FMf

[Ano65t] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(5):??, October 19, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986075.

Anonymous:1965:FMg

[Ano65u] Anonymous. Front matter. Proceedings of the American Philosophical Society, 109(6):??, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986134.

Anonymous:1965:RCL

[Ano65v] Anonymous. Report of the Committee on Library for 1964. Proceedings of the American Philosophical Society, 109(??):187–198, ???? 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985857.

Anonymous:1965:VI

[Ano65w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 109(6):441–iv, December 10, 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986133.

Anonymous:1966:Aa

[Ano66a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(1):??, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:Ab

[Ano66b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(2):??, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:Ac

[Ano66c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(3):??, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:Ad

[Ano66d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(4):??, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:Ae

[Ano66e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(5):??, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:Af

[Ano66f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 110(6):??, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1966:BMa

[Ano66g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(1):??, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986004.

Anonymous:1966:BMb

[Ano66h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(2):??, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985748.

Anonymous:1966:BMc

[Ano66i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(3):??, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985988.

Anonymous:1966:BMd

[Ano66j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(4):??, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985690.

Anonymous:1966:BMe

[Ano66k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(5):??, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986025.

Anonymous:1966:BMf

[Ano66l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 110(6):??, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985797.

Anonymous:1966:FMa

[Ano66m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(1):??, February 18, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985997.

Anonymous:1966:FMb

[Ano66n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(2):??, April 22, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985740.

Anonymous:1966:FMc

[Ano66o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(3):??, June 27, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985981.

Anonymous:1966:FMd

[Ano66p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(4):??, August 23, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985684.

Anonymous:1966:FMe

[Ano66q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(5):??, October 27, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986018.

Anonymous:1966:FMf

[Ano66r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 110(6):??, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985790.

Anonymous:1966:VI

[Ano66s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 110(6):441–iv, December 16, 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985789.

Anonymous:1967:Aa

[Ano67a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(1):??, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:Ab

[Ano67b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(2):??, April 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:Ac

[Ano67c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(3):??, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:Ad

[Ano67d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(4):??, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:Ae

[Ano67e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(5):??, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:Af

[Ano67f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 111(6):??, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1967:BMa

[Ano67g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 111(1):??, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985755.

Anonymous:1967:BMb

[Ano67h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 111(2):??, April 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986039.

Anonymous:1967:BMc

[Ano67i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 111(3):??, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985715.

Anonymous:1967:BMd

[Ano67j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 111(4):??, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986011.

Anonymous:1967:BMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986048.

Anonymous:1967:BMf

[Ano67l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 111(6):??, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985733.

Anonymous:1967:FMa

[Ano67m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 111(1):??, February 17, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985749.

Anonymous:1967:FMb

[Ano67n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 111(2):??, April 14, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986034.

Anonymous:1967:FMc

[Ano67o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 111(3):??, June 22, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985709.

Anonymous:1967:FMd

[Ano67p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 111(4):??, August 24, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986005.

Anonymous:1967:FMe

[Ano67q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 111(5):??, October 16, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986040.

Anonymous:1967:FMf ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985727.

Anonymous:1967:VI

[Ano67s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 111(6):359–iii, December 11, 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985726.

Anonymous:1968:Aa

[Ano68a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(1):??, February 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:Ab

[Ano68b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(2):??, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:Ac

[Ano68c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(3):??, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:Ad

[Ano68d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(4):??, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:Ae

[Ano68e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(5):??, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:Af

[Ano68f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 112(6):??, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1968:BMa ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986103.

Anonymous:1968:BMb

[Ano68h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 112(2):??, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985930.

Anonymous:1968:BMc

[Ano68i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 112(3):??, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986164.

Anonymous:1968:BMd

[Ano68j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 112(4):??, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985875.

Anonymous:1968:BMe

[Ano68k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 112(5):??, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986056.

Anonymous:1968:BMf

[Ano68l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 112(6):??, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985942.

Anonymous:1968:FMa

[Ano68m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 112(1):??, February 15, 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986096.

Anonymous:1968:FMb ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985925.

Anonymous:1968:FMc

[Ano68o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 112(3):??, June 21, 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986157.

Anonymous:1968:FMd

[Ano68p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 112(4):??, August 15, 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985869.

Anonymous:1968:FMe

[Ano68q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 112(5):??, October 17, 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986049.

Anonymous:1968:FMf

[Ano68r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 112(6):??, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985932.

Anonymous:1968:VI

[Ano68s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 112(6):435–iv, December 9, 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985931.

Anonymous:1969:Aa

[Ano69a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(1):??, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:Ab

[Ano69b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(2):??, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:Ac

[Ano69c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(3):??, June 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:Ad

[Ano69d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(4):??, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:Ae

[Ano69e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(5):??, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:Af

[Ano69f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 113(6):??, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1969:BMa

[Ano69g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 113(1):??, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986017.

Anonymous:1969:BMb

[Ano69h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 113(2):??, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985966.

Anonymous:1969:BMc

[Ano69i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 113(3):??, June 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985739.

Anonymous:1969:BMd 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986156.

Anonymous:1969:BMe

[Ano69k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 113(5):??, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985973.

Anonymous:1969:BMf

[Ano69l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 113(6):??, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985830.

Anonymous:1969:FMa

[Ano69m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 113(1):??, February 20, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986012.

Anonymous:1969:FMb

[Ano69n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 113(2):??, April 17, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985958.

Anonymous:1969:FMc

[Ano69o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 113(3):??, June 16, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985734.

Anonymous:1969:FMd

[Ano69p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 113(4):??, August 15, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986149.

Anonymous:1969:FMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985967.

Anonymous:1969:FMf

[Ano69r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 113(6):??, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985822.

Anonymous:1969:VI

[Ano69s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 113(6):475–iv, December 15, 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985821.

Anonymous:1970:Aa

[Ano70a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(1):??, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:Ab

[Ano70b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(2):??, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:Ac

[Ano70c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(3):??, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:Ad

[Ano70d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(4):??, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:Ae

[Ano70e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(5):??, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:Af

[Ano70f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 114(6):??, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1970:BMa

[Ano70g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(1):??, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985725.

Anonymous:1970:BMb

[Ano70h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(2):??, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986033.

Anonymous:1970:BMc

[Ano70i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(3):??, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985701.

Anonymous:1970:BMd

[Ano70j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(4):??, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985957.

Anonymous:1970:BMe

[Ano70k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(5):??, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985809.

Anonymous:1970:BMf

[Ano70l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 114(6):??, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986148.

Anonymous:1970:FMa

[Ano70m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(1):??, February 16, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985716.

Anonymous:1970:FMb

[Ano70n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(2):??, April 13, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986026.

Anonymous:1970:FMc

[Ano70o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(3):??, June 18, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985691.

Anonymous:1970:FMd

[Ano70p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(4):??, August 20, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985950.

Anonymous:1970:FMe

[Ano70q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(5):??, October 20, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985798.

Anonymous:1970:FMf

[Ano70r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 114(6):??, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986142.

Anonymous:1970:VI

[Ano70s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 114(6):488–iv, December 18, 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986141.

Anonymous:1971:Aa

[Ano71a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(1):??, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:Ab

[Ano71b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(2):??, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:Ac

[Ano71c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(3):??, June 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:Ad

[Ano71d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(4):??, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:Ae

[Ano71e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(5):??, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:Af

[Ano71f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 115(6):??, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1971:BMa

[Ano71g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(1):??, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985924.

Anonymous:1971:BMb

[Ano71h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(2):??, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985855.

Anonymous:1971:BMc

[Ano71i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(3):??, June 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985980.

Anonymous:1971:BMd

[Ano71j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(4):??, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986095.

Anonymous:1971:BMe

[Ano71k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(5):??, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985771.

Anonymous:1971:BMf

[Ano71l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 115(6):??, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985845.

Anonymous:1971:FMa

[Ano71m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(1):??, February 17, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985915.

Anonymous:1971:FMb

[Ano71n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(2):??, April 22, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985846.

Anonymous:1971:FMc

[Ano71o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(3):??, June 17, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985974.

Anonymous:1971:FMd

[Ano71p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(4):??, August 20, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986084.

Anonymous:1971:FMe

[Ano71q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(5):??, October 15, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985765.

Anonymous:1971:FMf

[Ano71r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 115(6):??, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985832.

Anonymous:1971:VI

[Ano71s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 115(6):517–iv, December 30, 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985831.

Anonymous:1972:Aa

[Ano72a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(1):??, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:Ab

[Ano72b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(2):??, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:Ac

[Ano72c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(3):??, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:Ad

[Ano72d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(4):??, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:Ae

[Ano72e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(5):??, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:Af

[Ano72f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 116(6):??, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1972:BMa

[Ano72g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 116(1):??, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985708.

Anonymous:1972:BMb

[Ano72h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 116(2):??, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986173.

Anonymous:1972:BMc

[Ano72i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 116(3):??, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986122.

Anonymous:1972:BMd

[Ano72j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 116(4):??, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985905.

Anonymous:1972:BMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986074.

Anonymous:1972:BMf

[Ano72l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 116(6):??, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985764.

Anonymous:1972:EPM

[Ano72m] Anonymous. Erratum: Penrose Memorial Lecture. Metallurgical footnotes to the history of art. Proceedings of the American Philosophical Society, 116(3):??, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986121. See [Smi72].

Anonymous:1972:FMa

[Ano72n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(1):??, February 15, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985702.

Anonymous:1972:FMb

[Ano72o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(2):??, April 17, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986165.

Anonymous:1972:FMc

[Ano72p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(3):??, June 9, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986112.

Anonymous:1972:FMd

[Ano72q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(4):??, August 15, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985897.

Anonymous:1972:FMe

[Ano72r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(5):??, October 13, 1972. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986064.

Anonymous:1972:FMf

[Ano72s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 116(6):??, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985757.

Anonymous:1972:VI

[Ano72t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 116(6):532–iv, December 21, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985756.

Anonymous:1973:Aa

[Ano73a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(1):??, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:Ab

[Ano73b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(2):??, April 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:Ac

[Ano73c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(3):??, June 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:Ad

[Ano73d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(4):??, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:Ae

[Ano73e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(5):??, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:Af

[Ano73f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 117(6):??, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1973:BMa

[Ano73g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(1):??, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985949.

Anonymous:1973:BMb

[Ano73h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(2):??, April 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986063.

Anonymous:1973:BMc

[Ano73i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(3):??, June 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986543.

Anonymous:1973:BMd

[Ano73j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(4):??, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986699.

Anonymous:1973:BMe

[Ano73k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(5):??, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986610.

Anonymous:1973:BMf

[Ano73l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 117(6):??, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986464.

Anonymous:1973:FMa

[Ano73m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(1):??, February 16, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985943.

Anonymous:1973:FMb

[Ano73n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(2):??, April 10, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986057.

Anonymous:1973:FMc

[Ano73o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(3):??, June 15, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986538.

Anonymous:1973:FMd

[Ano73p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(4):??, August 15, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986692.

Anonymous:1973:FMe

[Ano73q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(5):??, October 25, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986603.

Anonymous:1973:FMf

[Ano73r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 117(6):??, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986459.

Anonymous:1973:VI

[Ano73s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 117(6):551–iii, December 31, 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986458.

Anonymous:1974:Aa

[Ano74a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(1):??, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:Ab

[Ano74b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(2):??, April 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:Ac

[Ano74c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(3):??, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:Ad

[Ano74d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(4):??, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:Ae

[Ano74e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(5):??, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:Af

[Ano74f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 118(6):??, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1974:BMa

[Ano74g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(1):??, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986438.

Anonymous:1974:BMb

[Ano74h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(2):??, April 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986667.

Anonymous:1974:BMc

[Ano74i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(3):??, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986448.

Anonymous:1974:BMd

[Ano74j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(4):??, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986283.

Anonymous:1974:BMe

[Ano74k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(5):??, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986312.

Anonymous:1974:BMf

[Ano74l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 118(6):??, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986407.

Anonymous:1974:FMa

[Ano74m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(1):??, February 28, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986433.

Anonymous:1974:FMb

[Ano74n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(2):??, April 19, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986661.

Anonymous:1974:FMc

[Ano74o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(3):??, June 7, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986439.

Anonymous:1974:FMd

[Ano74p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(4):??, September 13, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986277.

Anonymous:1974:FMe

[Ano74q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(5):??, October 15, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986303.

Anonymous:1974:FMf

[Ano74r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 118(6):??, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986398.

Anonymous:1974:VI

[Ano74s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 118(6):595–iii, December 27, 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986397.

Anonymous:1975:Aa

[Ano75a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(1):??, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:Ab

[Ano75b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(2):??, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:Ac

[Ano75c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(3):??, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:Ad

[Ano75d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(4):??, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:Ae

[Ano75e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(5):??, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:Af

[Ano75f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 119(6):??, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1975:BMa

[Ano75g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 119(1):??, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986651.

Anonymous:1975:BMb

[Ano75h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 119(2):??, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986635.

Anonymous:1975:BMc

[Ano75i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 119(3):??, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986674.

Anonymous:1975:BMd

[Ano75j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 119(4):??, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986294.

Anonymous:1975:BMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986682.

Anonymous:1975:BMf

[Ano75l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 119(6):??, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986380.

Anonymous:1975:FMa

[Ano75m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 119(1):??, February 21, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986644.

Anonymous:1975:FMb

[Ano75n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 119(2):??, April 16, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986628.

Anonymous:1975:FMc

[Ano75o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 119(3):??, June 12, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986668.

Anonymous:1975:FMd

[Ano75p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 119(4):??, August 15, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986284.

Anonymous:1975:FMe

[Ano75q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 119(5):??, October 15, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986675.

Anonymous:1975:FMf ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986374.

Anonymous:1975:VI

[Ano75s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 119(6):479–iii, December 5, 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986373.

Anonymous:1976:Aa

[Ano76a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(1):??, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:Ab

[Ano76b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(2):??, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:Ac

[Ano76c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(3):??, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:Ad

[Ano76d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(4):??, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:Ae

[Ano76e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(5):??, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:Af

[Ano76f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 120(6):??, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1976:BMa ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986353.

Anonymous:1976:BMb

[Ano76h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 120(2):??, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986519.

Anonymous:1976:BMc

[Ano76i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 120(3):??, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986563.

Anonymous:1976:BMd

[Ano76j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 120(4):??, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986323.

Anonymous:1976:BMe

[Ano76k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 120(5):??, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986270.

Anonymous:1976:BMf

[Ano76l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 120(6):??, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986602.

Anonymous:1976:FMa

[Ano76m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 120(1):??, February 5, 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986345.

Anonymous:1976:FMb ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986511.

Anonymous:1976:FMc

[Ano76o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 120(3):??, June 15, 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986554.

Anonymous:1976:FMd

[Ano76p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 120(4):??, August 13, 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986318.

Anonymous:1976:FMe

[Ano76q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 120(5):??, October 15, 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986263.

Anonymous:1976:FMf

[Ano76r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 120(6):??, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986596.

Anonymous:1976:VI

[Ano76s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 120(6):514–iii, December 29, 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986595.

Anonymous:1977:Aa

[Ano77a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(1):??, February 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:Ab

[Ano77b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(2):??, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:Ac

[Ano77c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(3):??, June 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:Ad

[Ano77d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(4):??, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:Ae

[Ano77e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(5):??, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:Af

[Ano77f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 121(6):??, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1977:BMa

[Ano77g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 121(1):??, February 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986568.

Anonymous:1977:BMb

[Ano77h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 121(2):??, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986528.

Anonymous:1977:BMc

[Ano77i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 121(3):??, June 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986414.

Anonymous:1977:BMd 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986422.

Anonymous:1977:BMe

[Ano77k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 121(5):??, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986344.

Anonymous:1977:BMf

[Ano77l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 121(6):??, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986396.

Anonymous:1977:FMa

[Ano77m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 121(1):??, February 15, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986564.

Anonymous:1977:FMb

[Ano77n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 121(2):??, April 29, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986520.

Anonymous:1977:FMc

[Ano77o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 121(3):??, June 15, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986408.

Anonymous:1977:FMd

[Ano77p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 121(4):??, August 12, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986415.

Anonymous:1977:FMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986335.

Anonymous:1977:FMf

[Ano77r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 121(6):??, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986390.

Anonymous:1977:VI

[Ano77s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 121(6):500–iii, December 1, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986389.

Anonymous:1978:Aa

[Ano78a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(1):??, February 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:Ab

[Ano78b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(2):??, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:Ac

[Ano78c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(3):??, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:Ad

[Ano78d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(4):??, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:Ae

[Ano78e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(5):??, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:Af

[Ano78f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 122(6):??, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1978:BMa

[Ano78g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(1):??, February 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986262.

Anonymous:1978:BMb

[Ano78h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(2):??, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986627.

Anonymous:1978:BMc

[Ano78i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(3):??, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986553.

Anonymous:1978:BMd

[Ano78j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(4):??, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986537.

Anonymous:1978:BMe

[Ano78k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(5):??, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986691.

Anonymous:1978:BMf

[Ano78l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 122(6):??, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986457.

Anonymous:1978:FMa

[Ano78m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(1):??, February 15, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986257.

Anonymous:1978:FMb

[Ano78n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(2):??, April 24, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986620.

Anonymous:1978:FMc

[Ano78o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(3):??, June 9, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986544.

Anonymous:1978:FMd

[Ano78p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(4):??, August 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986529.

Anonymous:1978:FMe

[Ano78q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(5):??, October 19, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986683.

Anonymous:1978:FMf

[Ano78r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 122(6):??, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986450.

Anonymous:1978:VI

[Ano78s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 122(6):427–iii, December 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986449.

Anonymous:1979:Aa

[Ano79a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(1):??, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:Ab

[Ano79b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(2):??, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:Ac

[Ano79c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(3):??, June 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:Ad

[Ano79d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(4):??, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:Ae

[Ano79e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(5):??, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:Af

[Ano79f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 123(6):??, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1979:BMa

[Ano79g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(1):??, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986472.

Anonymous:1979:BMb

[Ano79h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(2):??, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986238.

Anonymous:1979:BMc

[Ano79i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(3):??, June 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986221.

Anonymous:1979:BMd

[Ano79j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(4):??, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986584.

Anonymous:1979:BMe

[Ano79k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(5):??, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986594.

Anonymous:1979:BMf

[Ano79l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 123(6):??, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986619.

Anonymous:1979:EJL

[Ano79m] Anonymous. Erratum: Jayne Lecture. Darwin and the “Beagle”. Proceedings of the American Philosophical Society, 123(6):??, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986618.

See [Key79].

Anonymous:1979:FMa

[Ano79n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(1):??, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986466.

Anonymous:1979:FMb

[Ano79o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(2):??, April 27, 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986231.

Anonymous:1979:FMc

[Ano79p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(3):??, June 18, 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986216.

Anonymous:1979:FMd

[Ano79q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(4):??, August 30, 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986577.

Anonymous:1979:FMe

[Ano79r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(5):??, October 15, 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986585.

Anonymous:1979:FMf

[Ano79s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 123(6):??, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986612.

Anonymous:1979:VIa

[Ano79t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 123(1):i–iii, February 20, 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986465.

Anonymous:1979:VIb

[Ano79u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 123(6):410–411, December 28, 1979. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986611.

Anonymous:1980:Aa

[Ano80a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(1):??, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:Ab

[Ano80b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(2):??, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:Ac

[Ano80c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(3):??, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:Ad

[Ano80d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(4):??, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:Ae

[Ano80e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(5):??, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:Af

[Ano80f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 124(6):??, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1980:BMa

[Ano80g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(1):??, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986660.

Anonymous:1980:BMb

[Ano80h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(2):??, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986209.

Anonymous:1980:BMc

[Ano80i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(3):??, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986372.

Anonymous:1980:BMd

[Ano80j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(4):??, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986432.

Anonymous:1980:BMe

[Ano80k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(5):??, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986576.

Anonymous:1980:BMf

[Ano80l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 124(6):??, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986247.

Anonymous:1980:FMa

[Ano80m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(1):??, February 29, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986652.

Anonymous:1980:FMb

[Ano80n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(2):??, April 29, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986202.

Anonymous:1980:FMc

[Ano80o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(3):??, June 30, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986367.

Anonymous:1980:FMd

[Ano80p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(4):??, August 19, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986423.

Anonymous:1980:FMe

[Ano80q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(5):??, October 10, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986569.

Anonymous:1980:FMf

[Ano80r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 124(6):??, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986240.

Anonymous:1980:VI

[Ano80s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 124(6):i–465, December 17, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986239.

Anonymous:1981:Aa

[Ano81a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(1):??, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:Ab

[Ano81b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(2):??, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:Ac

[Ano81c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(3):??, June 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:Ad

[Ano81d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(4):??, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:Ae

[Ano81e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(5):??, October 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:Af

[Ano81f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 125(6):??, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1981:BMa

[Ano81g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(1):??, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986188.

Anonymous:1981:BMb

[Ano81h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(2):??, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986643.

Anonymous:1981:BMc

[Ano81i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(3):??, June 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986485.

Anonymous:1981:BMd

[Ano81j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(4):??, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986334.

Anonymous:1981:BMe

[Ano81k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(5):??, October 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986201.

Anonymous:1981:BMf

[Ano81l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 125(6):??, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986302.

Anonymous:1981:FMa

[Ano81m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(1):??, February 16, 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986180.

Anonymous:1981:FMb

[Ano81n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(2):??, April 30, 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986636.

Anonymous:1981:FMc

[Ano81o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(3):??, June 23, 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986480.

Anonymous:1981:FMd

[Ano81p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(4):??, August 21, 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986324.

Anonymous:1981:FMe

[Ano81q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(5):??, October 1, 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986196.

Anonymous:1981:FMf

[Ano81r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 125(6):??, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986296.

Anonymous:1981:VI

[Ano81s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 125(6):467–468, December 18, 1981. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986295.

Anonymous:1982:Aa

[Ano82a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(1):??, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:Ab

[Ano82b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(2):??, April 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:Ac

[Ano82c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(3):??, June 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:Ad

[Ano82d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(4):??, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:Ae

[Ano82e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(5):??, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:Af

[Ano82f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 126(6):??, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1982:BMa

[Ano82g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(1):??, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986479.

Anonymous:1982:BMb

[Ano82h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(2):??, April 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986359.

Anonymous:1982:BMc

[Ano82i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(3):??, June 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986510.

Anonymous:1982:BMd

[Ano82j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(4):??, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986179.

Anonymous:1982:BMe

[Ano82k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(5):??, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986215.

Anonymous:1982:BMf

[Ano82l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 126(6):??, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986256.

Anonymous:1982:FMa

[Ano82m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(1):??, February 26, 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986473.

Anonymous:1982:FMb

[Ano82n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(2):??, April 8, 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986354.

Anonymous:1982:FMc

[Ano82o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(3):??, June 8, 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986505.

Anonymous:1982:FMd

[Ano82p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(4):??, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986174.

Anonymous:1982:FMe

[Ano82q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(5):??, October 21, 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986210.

Anonymous:1982:FMf

[Ano82r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 126(6):??, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986249.

Anonymous:1982:VI

[Ano82s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 126(6):527–528, December 17, 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986248.

Anonymous:1983:Aa

[Ano83a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(1):??, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:Ab

[Ano83b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(2):??, April 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:Ac

[Ano83c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(3):??, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:Ad

[Ano83d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(4):??, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:Ae

[Ano83e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(5):??, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:Af

[Ano83f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 127(6):??, December 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1983:BMa

[Ano83g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 127(1):??, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986366.

Anonymous:1983:BMb

[Ano83h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 127(2):??, April 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986195.

Anonymous:1983:BMc

[Ano83i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 127(3):??, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986388.

Anonymous:1983:BMd

[Ano83j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 127(4):??, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986276.

Anonymous:1983:BMe ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986504.

Anonymous:1983:BMf

[Ano83l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 127(6):??, December 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986317.

Anonymous:1983:FMa

[Ano83m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 127(1):??, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986360.

Anonymous:1983:FMb

[Ano83n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 127(2):??, April 20, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986189.

Anonymous:1983:FMc

[Ano83o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 127(3):??, June 16, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986381.

Anonymous:1983:FMd

[Ano83p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 127(4):??, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986271.

Anonymous:1983:FMe

[Ano83q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 127(5):??, October 14, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986498.

Anonymous:1983:FMf ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986314.

Anonymous:1983:VI

[Ano83s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 127(6):463–464, December 30, 1983. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986313.

Anonymous:1984:Aa

[Ano84a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 128(1):??, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1984:Ab

[Ano84b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 128(2):??, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1984:Ac

[Ano84c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 128(3):??, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1984:Ad

[Ano84d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 128(4):??, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1984:BMa

[Ano84e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 128(1):??, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986497.

Anonymous:1984:BMb

[Ano84f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 128(2):??, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986230.

Anonymous:1984:BMc

[Ano84g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 128(3):??, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986888.

Anonymous:1984:BMd

[Ano84h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 128(4):??, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987101.

Anonymous:1984:FMa

[Ano84i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 128(1):??, March 30, 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986486.

Anonymous:1984:FMb

[Ano84j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 128(2):??, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986222.

Anonymous:1984:FMc

[Ano84k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 128(3):??, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986878.

Anonymous:1984:FMd

[Ano84l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 128(4):??, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987097.

Anonymous:1984:VI

[Ano84m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 128(4):410–411, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987096.

Anonymous:1985:Aa

[Ano85a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 129(1):??, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1985:Ab

[Ano85b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 129(2):??, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1985:Ac

[Ano85c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 129(3):??, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1985:Ad

[Ano85d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 129(4):??, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1985:BMa

[Ano85e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 129(1):??, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986982.

Anonymous:1985:BMb

[Ano85f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 129(2):??, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986991.

Anonymous:1985:BMc

[Ano85g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 129(3):??, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987014.

Anonymous:1985:BMd

[Ano85h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 129(4):??, December 1985. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986940.

Anonymous:1985:FMa

[Ano85i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 129(1):??, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986972.

Anonymous:1985:FMb

[Ano85j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 129(2):??, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986983.

Anonymous:1985:FMc

[Ano85k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 129(3):??, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987007.

Anonymous:1985:FMd

[Ano85l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 129(4):??, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986933.

Anonymous:1985:VI

[Ano85m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 129(4):476–477, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986932.

Anonymous:1986:Aa

[Ano86a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 130(1):??, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1986:Ab

[Ano86b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 130(2):??, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1986:Ac

[Ano86c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 130(3):??, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1986:Ad

[Ano86d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 130(4):??, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1986:BMa

[Ano86e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 130(1):??, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987095.

Anonymous:1986:BMb

[Ano86f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 130(2):??, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987184.

Anonymous:1986:BMc

[Ano86g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 130(3):??, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986836.

Anonymous:1986:BMd

[Ano86h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 130(4):??, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986931.

Anonymous:1986:FMa

[Ano86i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 130(1):??, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987087.

Anonymous:1986:FMb

[Ano86j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 130(2):??, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987178.

Anonymous:1986:FMc

[Ano86k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 130(3):??, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986826.

Anonymous:1986:FMd

[Ano86l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 130(4):??, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986922.

Anonymous:1986:VI

[Ano86m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 130(4):482–483, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986921.

Anonymous:1987:Aa

[Ano87a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 131(1):??, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1987:Ab

[Ano87b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 131(2):??, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1987:Ac

[Ano87c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 131(3):??, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1987:Ad

[Ano87d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 131(4):??, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1987:BMa

[Ano87e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 131(1):??, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987119.

Anonymous:1987:BMb

[Ano87f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 131(2):??, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986793.

Anonymous:1987:BMc

[Ano87g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 131(3):??, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987028.

Anonymous:1987:BMd

[Ano87h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 131(4):??, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987141.

Anonymous:1987:FMa

[Ano87i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 131(1):??, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987112.

Anonymous:1987:FMb

[Ano87j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 131(2):??, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986784.

Anonymous:1987:FMc

[Ano87k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 131(3):237–285, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987015.

Anonymous:1987:FMd

[Ano87l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 131(4):??, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987134.

Anonymous:1987:VI

[Ano87m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 131(4):442–445, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987133.

Anonymous:1988:Aa

[Ano88a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 132(1):??, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1988:Ab

[Ano88b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 132(2):??, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1988:Ac

[Ano88c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 132(3):??, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1988:Ad

[Ano88d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 132(4):??, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1988:BMa

[Ano88e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 132(1):??, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143830.

Anonymous:1988:BMb

[Ano88f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 132(2):??, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986868.

Anonymous:1988:BMc

[Ano88g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 132(3):??, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143857.

Anonymous:1988:BMd

[Ano88h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 132(4):444–445, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986964.

Anonymous:1988:FMa

[Ano88i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 132(1):i–iii, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143822.

Anonymous:1988:FMb

[Ano88j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 132(2):??, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986857.

Anonymous:1988:FMc

[Ano88k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 132(3):??, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143848.

Anonymous:1988:FMd

[Ano88l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 132(4):??, December 1988. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986956.

Anonymous:1988:PVN

[Ano88m] Anonymous. Proceedings, vol. 1, no. 4 [reprint]. Proceedings of the American Philosophical Society, 132(1):137–145, March 1988. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143829.

Anonymous:1988:VI

[Ano88n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 132(4):446–448, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986955.

Anonymous:1989:Aa

[Ano89a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 133(1):??, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1989:Ab

[Ano89b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 133(2):??, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1989:Ac

[Ano89c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 133(3):??, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1989:Ad

[Ano89d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 133(4):??, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1989:BMa

[Ano89e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 133(1):??, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987164.

Anonymous:1989:BMb

[Ano89f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 133(2):??, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987060.

Anonymous:1989:EEI

[Ano89g] Anonymous. Errata: European immigration to the United States in the Early National Period, 1783–1820. Proceedings of the American Philosophical Society, 133(4):589, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986877. See [Gra89].

Anonymous:1989:FMa

[Ano89h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 133(1):??, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987157.

Anonymous:1989:FMb

[Ano89i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 133(2):??, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987040.

Anonymous:1989:FMc

[Ano89j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 133(3):??, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987102.

Anonymous:1989:FMd

[Ano89k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 133(4):??, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986870.

Anonymous:1989:VI

[Ano89l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 133(4):580–588, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986869.

Anonymous:1990:Aa

[Ano90a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 134(1):??, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1990:Ab

[Ano90b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 134(2):??, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1990:Ac

[Ano90c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 134(3):??, September 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1990:Ad

[Ano90d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 134(4):??, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1990:FMa

[Ano90e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 134(1):??, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143898.

Anonymous:1990:FMb

[Ano90f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 134(2):??, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986837.

Anonymous:1990:FMc

[Ano90g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 134(3):??, September 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986965.

Anonymous:1990:FMd

[Ano90h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 134(4):337–443, December 1990. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986890.

Anonymous:1990:VI

[Ano90i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 134(4):465–473, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986889.

Anonymous:1991:Aa

[Ano91a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 135(1):??, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1991:Ab

[Ano91b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 135(2):??, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1991:Ac

[Ano91c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 135(3):??, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1991:Ad

[Ano91d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 135(4):??, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1991:BM

[Ano91e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 135(4):??, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986825.

Anonymous:1991:FMa

[Ano91f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 135(1):85–117, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987142.

Anonymous:1991:FMb

[Ano91g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 135(2):299–321, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987029.

Anonymous:1991:FMc

[Ano91h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 135(3):443–479, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986769.

Anonymous:1991:FMd

[Ano91i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 135(4):591–629, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986809.

Anonymous:1991:VI

[Ano91j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 135(4):633–642, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986808.

Anonymous:1992:Aa

[Ano92a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 136(1):??, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1992:Ab

[Ano92b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 136(2):??, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1992:Ac

[Ano92c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 136(3):??, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1992:Ad

[Ano92d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 136(4):??, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1992:ACA

[Ano92e] Anonymous. Appendix C: America’s wasted Blacks. Proceedings of the American Philosophical Society, 136(3):399–409, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986913.

Anonymous:1992:BM

[Ano92f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 136(4):??, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986768.

Anonymous:1992:FMa

[Ano92g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 136(1):127–149, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986794.

Anonymous:1992:FMb

[Ano92h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 136(2):273–311, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987165.

Anonymous:1992:FMc

[Ano92i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 136(3):353–449, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986901.

Anonymous:1992:FMd

[Ano92j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 136(4):577–607, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986756.

Anonymous:1992:JAA

[Ano92k] Anonymous. Joint announcement of the Austin Project. Proceedings of the American Philosophical Society, 136(3):378–379, September 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986910.

Anonymous:1992:VI

[Ano92l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 136(4):613–622, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986755.

Anonymous:1993:Aa

[Ano93a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 137(1):??, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1993:Ab

[Ano93b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 137(2):??, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1993:Ac

[Ano93c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 137(3):??, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1993:Ad

[Ano93d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 137(4):??, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1993:BM

[Ano93e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 137(4):709–712, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987086.

Anonymous:1993:FMa

[Ano93f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 137(1):117–173, March 1993. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986941.

Anonymous:1993:FMb

[Ano93g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 137(2):273–305, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986727.

Anonymous:1993:FMc

[Ano93h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 137(3):431–457, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986992.

Anonymous:1993:FMd

[Ano93i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 137(4):i–657, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987062.

Anonymous:1993:VI

[Ano93j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 137(4):713–719, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987061.

Anonymous:1994:Aa

[Ano94a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 138(1):??, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1994:Ab

[Ano94b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 138(2):??, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1994:Ac

[Ano94c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 138(3):??, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1994:Ad

[Ano94d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 138(4):??, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 138(1):i–183, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986700.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 138(2):317–355, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987120.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 138(3):407–463, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986740.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 138(4):517–569, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986845.

Anonymous:1994:VI

[Ano94i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 138(4):575–585, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986844.

Anonymous:1995:Aa

[Ano95a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 139(1):??, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1995:Ab

[Ano95b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 139(2):??, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1995:Ac

[Ano95c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 139(3):??, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1995:Ad

[Ano95d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 139(4):??, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1995:BM

[Ano95e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 139(4):463–466, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987244.

Anonymous:1995:FMa

[Ano95f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 139(1):67–109, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986715.

Anonymous:1995:FMb

[Ano95g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 139(2):171–211, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987259.

Anonymous:1995:FMc

[Ano95h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 139(3):265–311, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987185.

Anonymous:1995:FMd

[Ano95i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 139(4):427–451, December 1995. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987230.

Anonymous:1995:VI

[Ano95j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 139(4):467–471, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987229.

Anonymous:1996:Aa

[Ano96a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 140(1):??, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1996:Ab

[Ano96b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 140(2):??, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1996:Ac

[Ano96c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 140(3):??, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1996:Ad

[Ano96d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 140(4):??, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1996:BMa

[Ano96e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 140(1):??, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987283.

Anonymous:1996:BMb

[Ano96f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 140(2):??, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987332.

Anonymous:1996:BMc

[Ano96g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 140(3):??, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987317.

Anonymous:1996:BMd

[Ano96h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 140(4):587–592, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987295.

Anonymous:1996:FMa

[Ano96i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 140(1):77–115, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987271.

Anonymous:1996:FMb

[Ano96j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 140(2):205–253, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987318.

Anonymous:1996:FMc

[Ano96k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 140(3):265, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987309.

Anonymous:1996:FMd

[Ano96l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 140(4):557–579, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987285.

Anonymous:1996:VI

[Ano96m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 140(4):593–599, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987284.

Anonymous:1997:Aa

[Ano97a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 141(1):??, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1997:Ab

[Ano97b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 141(2):??, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1997:Ac

[Ano97c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 141(3):??, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1997:Ad

[Ano97d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 141(4):??, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1997:BMa

[Ano97e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 141(1):??, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987258.

Anonymous:1997:BMb

[Ano97f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 141(2):??, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987308.

Anonymous:1997:BMc

[Ano97g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 141(3):??, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987209.

Anonymous:1997:BMd

[Ano97h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 141(4):515–518, December 1997. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987228.

Anonymous:1997:FMa

[Ano97i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 141(1):77–115, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987245.

Anonymous:1997:FMb

[Ano97j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 141(2):217–239, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987296.

Anonymous:1997:FMc

[Ano97k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 141(3):365–383, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987200.

Anonymous:1997:FMd

[Ano97l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 141(4):463–509, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987211.

Anonymous:1997:LCO

[Ano97m] Anonymous. Lord Caradon (8 October 1907–5 September 1990). Proceedings of the American Philosophical Society, 141(4):466–470, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987222.

Anonymous:1997:VI

[Ano97n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 141(4):519–526, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987210.

Anonymous:1998:Aa

[Ano98a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 142(1):??, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1998:Ab

[Ano98b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 142(2):??, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1998:Ac

[Ano98c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 142(3):??, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1998:Ad

[Ano98d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 142(4):??, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1998:BMa

[Ano98e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 142(1):??, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152275.

Anonymous:1998:BMb

[Ano98f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 142(2):??, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152311.

Anonymous:1998:BMc

[Ano98g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 142(3):??, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152259.

Anonymous:1998:BMd

[Ano98h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 142(4):??, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152295.

Anonymous:1998:ERG

[Ano98i] Anonymous. Erratum: Robert G. Dunlop (2 July 1909–20 September 1995). Proceedings of the American Philosophical Society, 142 (3):??, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152258. See [Ano98i].

Anonymous:1998:FMa

[Ano98j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 142(1):110–161, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152260.

Anonymous:1998:FMb

[Ano98k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 142(2):263–315, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152296.

Anonymous:1998:FMc

[Ano98l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 142(3):443–511, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152239.

Anonymous:1998:FMd

[Ano98m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 142(4):643–703, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152277.

Anonymous:1998:VI

[Ano98n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 142(4):704–715, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3152276.

Anonymous:1999:Aa

[Ano99a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 143(1):??, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1999:Ab

[Ano99b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 143(2):??, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1999:Ac

[Ano99c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 143(3):??, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1999:Ad

[Ano99d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 143(4):??, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1999:BMa

[Ano99e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 143(1):??, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181983.

Anonymous:1999:BMb

[Ano99f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 143(2):299–351, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181947.

Anonymous:1999:BMc

[Ano99g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 143(3):??, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181967.

Anonymous:1999:BMd

[Ano99h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 143(4):??, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3182002.

Anonymous:1999:FMa

[Ano99i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 143(1):1–2, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3181968.

Anonymous:1999:FMb

[Ano99j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 143(2):157–297, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181928.

Anonymous:1999:FMc

[Ano99k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 143(3):357–523, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181948.

Anonymous:1999:FMd

[Ano99l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 143(4):529–735, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181985.

Anonymous:1999:VI

[Ano99m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 143(4):736–745, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181984.

Anthes:1963:ADB

[Ant63] Rudolf Anthes. Affinity and difference between Egyptian and Greek sculpture and thought in the Seventh and Sixth Centuries B.C. Proceedings of the American Philosophical Society, 107(1):60– 81, February 1963. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985469.

Appel:1984:UCP

[App84] Kenneth I. Appel. The use of the computer in the proof of the four color theorem. Proceedings of the American Philosophical Society, 128(1):35–39, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986491.

Appleby:1997:PRF

[App97] Joyce Appleby. The personal roots of the first American temperance movement. Proceedings of the American Philosophical Society, 141(2):141–159, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987298.

Arnold:1992:LPJ

[Arn92] Lorna Arnold. Lord Penney (24 June 1909–3 March 1991). Proceedings of the American Philosophical Society, 136(3):444–448, September 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986919.

Arrhenius:1999:LAU

[Arr99] Birgit Arrhenius. Laboratory archaeology: a useful tool for studying ancient cultures. Proceedings of the American Philosophical Society, 143(1):97–101, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181977.

Ashenfelter:1978:WIU

[Ash78] Orley Ashenfelter. What is involuntary unemployment? Proceedings of the American Philosophical Society, 122(3):135–138, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986547.

Ashworth:1981:JBH

[Ash81] William B. Ashworth, Jr. John Bevis and his “Uranographia” (ca. 1750). Proceedings of the American Philosophical Society, 125 (1):52–73, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986187.

Astin:1982:CAH

[Ast82] Alexander W. Astin. Contradictions in American higher education. Proceedings of the American Philosophical Society, 126(1): 6–10, February 1982. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986475.

Atiya:1952:MSC

[Ati52] Aziz S. Atiya. The Monastery of St. Catherine and the Mount Sinai Expedition. Proceedings of the American Philosophical Society, 96 (5):578–586, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143659.

Atiyah:1993:DA

[Ati93a] Michael Atiyah. Dinner address. Proceedings of the American Philosophical Society, 137(4):704–707, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987084.

Atiyah:1993:MQS

[Ati93b] Michael Atiyah. Mathematics: Queen and servant of the sciences. Proceedings of the American Philosophical Society, 137(4):527–531, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987071.

Atkinson:1999:GFS

[Atk99] Richard C. Atkinson. The golden fleece, science education, and U.S. science policy. Proceedings of the American Philosophical Society, 143(3):407–417, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3181953.

Aulie:1970:BNC

[Aul70] Richard P. Aulie. Boussingault and the nitrogen cycle. Proceedings of the American Philosophical Society, 114(6):435–479, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986146.

Ault:1982:VME

[Aul82] Warren O. Ault. The vill in Medieval England. Proceedings of the American Philosophical Society, 126(3):188–211, June 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986507.

Aulie:1983:ESC

[Aul83] Richard P. Aulie. Evolution and special creation: Historical aspects of the controversy. Proceedings of the American Philosophical Society, 127(6):418–462, December 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986316.

Austrian:1981:DPV

[Aus81] Robert Austrian. The development of pneumococcal vaccine. Proceedings of the American Philosophical Society, 125(1):46–51, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986186.

Bacher:1972:RO

[Bac72] Robert F. Bacher. Robert Oppenheimer (1904–1967). Proceedings of the American Philosophical Society, 116(4):279–293, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985898.

Bacher:1977:SCO

[Bac77] Robert F. Bacher. Some comments on our energy problems. Proceedings of the American Philosophical Society, 121(4):275–280, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986416.

Badash:1968:RBA

[Bad68] Lawrence Badash. Rutherford, Boltwood, and the age of the Earth: The origin of radioactive dating techniques. Proceedings of the American Philosophical Society, 112(3):157–169, June 1968. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986161.

Badash:1978:RRP

[Bad78] Lawrence Badash. Radium, radioactivity, and the popularity of scientific discovery. Proceedings of the American Philosophical Society, 122(3):145–154, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986549.

Baine:1962:DDH

[Bai62] Rodney M. Baine. Daniel Defoe and “the history and reality of apparitions”. Proceedings of the American Philosophical Society, 106(4):335–347, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985267.

Bailey:1989:CPA

[Bai89] Herbert Smith Bailey, Jr. On the Collected Papers of Albert Einstein: The development of the project. Proceedings of the American Philosophical Society, 133(3):347–359, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987104.

Bailyn:1993:JAF

[Bai93] Bernard Bailyn. Jefferson and the ambiguities of freedom. Proceedings of the American Philosophical Society, 137(4):498–515, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987069.

Bailey:1997:CSJ

[Bai97] Herbert S. Bailey, Jr. Charles Scribner, Jr. (13 July 1921–11 November 1995). Proceedings of the American Philosophical Society, 141(2):232–237, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987306.

Baker:1977:LLE

[Bak77] W. O. Baker. Language and logic with electronics. Proceedings of the American Philosophical Society, 121(5):360–372, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986340.

Baker:1984:PJO

[Bak84] Carlos Baker. The poet as Janus: Originality and imitation in modern poetry. Proceedings of the American Philosophical Society, 128(2):167–172, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986229.

Baker:1994:SLA

[Bak94] W. O. Baker. Science and learning for all Americans. Proceedings of the American Philosophical Society, 138(1):1–10, March 1994. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986701.

Barbeau:1950:IC

[Bar50] Marius Barbeau. Indian captivities. Proceedings of the American Philosophical Society, 94(6):522–548, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143594.

Barbeau:1952:CCM

[Bar52a] Marius Barbeau. Charles A. Cooke, Mohawk scholar. Proceedings of the American Philosophical Society, 96(4):424–426, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143799.

Bartlett:1952:BMC

[Bar52b] Harley Harris Bartlett. A Batak and Malay chant on rice cultivation, with introductory notes on bilingualism and acculturation in Indonesia. Proceedings of the American Philosophical Society, 96 (6):629–652, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143633.

Barrus:1954:ITA

[Bar54] L. S. Barrus. Industrial trends affecting the older worker. Proceedings of the American Philosophical Society, 98(1):46–49, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143669.

Barber:1955:PSN

[Bar55] Bernard Barber. Present status and needs of the sociology of science. Proceedings of the American Philosophical Society, 99(5): 338–342, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/3143933. See comments [Din55].

Barzun:1960:ACD

[Bar60] Jacques Barzun. The American college in double jeopardy. Proceedings of the American Philosophical Society, 104(6):570–571, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985533.

Bard:1964:EGC

[Bar64a] Philip Bard. Edwin Grant Conklin, 1863–1952. The one hundredth anniversary of his birth. Proceedings of the American Philosophical Society, 108(2):55–56, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985886.

Barros:1964:GBI

[Bar64b] James Barros. The Greek–Bulgarian Incident of 1925: The League of Nations and the Great Powers. Proceedings of the American Philosophical Society, 108(4):354–385, August 1964. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985913.

Barrett:1974:SDG

[Bar74] Paul H. Barrett. The Sedgwick–Darwin geologic tour of North Wales. Proceedings of the American Philosophical Society, 118(2): 146–164, April 1974. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986663.

Baron:1983:DDT

[Bar83] Lawrence Baron. Discipleship and dissent: Theodor Lessing and Edmund Husserl. Proceedings of the American Philosophical Society, 127(1):32–49, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986363.

Barzun:1987:IPP

[Bar87] Jacques Barzun. An insoluble problem: The patronage of art. Proceedings of the American Philosophical Society, 131(2):121–131, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986785.

Barone:1990:IWS

[Bar90a] Dennis Barone. An introduction to William Smith and rhetoric at the College of Philadelphia. Proceedings of the American Philosophical Society, 134(2):111–160, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986842.

Barzun:1990:CAC

[Bar90b] Jacques Barzun. Criticism: an art or a craft? Proceedings of the American Philosophical Society, 134(4):455–460, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986899.

Barr:1999:PSC

[Bar99] James Barr. Pre-scientific chronology: The Bible and the origin of the World. Proceedings of the American Philosophical Society, 143(3):379–387, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181950.

Bates:1986:RSP

[Bat86] Richard L. Bates. “Rex in Senatu”: a political biography of M. Aemilius Scaurus. Proceedings of the American Philosophical Society, 130(3):251–288, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986827.

Bate:1996:HTL

[Bat96] W. Jackson Bate. Harry Tuchman Levin (18 July 1912–29 May 1994). Proceedings of the American Philosophical Society, 140(1): 100–104, March 1996. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987280.

Baugh:1959:IME

[Bau59] Albert C. Baugh. Improvisation in the Middle English romance. Proceedings of the American Philosophical Society, 103(3):418–454, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985475.

Baumgartner:1971:MSN

[Bau71] Leona Baumgartner. Medicine and society — new issues. Proceedings of the American Philosophical Society, 115(5):335–340, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985766.

Baumol:1979:FM

[Bau79] William J. Baumol. On the folklore of Marxism. Proceedings of the American Philosophical Society, 123(2):124–128, April 1979. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986234.

Baumol:1993:SWD

[Bau93] William J. Baumol. Social wants and dismal science: The curious case of the climbing costs of health and teaching. Proceedings of the American Philosophical Society, 137(4):612–637, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987077.

Baumol:1994:SWA

[Bau94] William J. Baumol. Sir William Arthur Lewis (23 January 1915– 15 June 1991). Proceedings of the American Philosophical Society, 138(2):336–340, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987129.

Baylen:1969:MCC

[Bay69] Joseph O. Baylen. The Mattei cancer cure: a Victorian nostrum. Proceedings of the American Philosophical Society, 113(2):149–176, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985963.

Batigne:1953:STA

[BB53] René Batigne and Louisa Bellinger. The significance and technical analysis of ancient textiles as historical documents. Proceedings of the American Philosophical Society, 97(6):670–680, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149282.

Bonner:1984:WMI

[BB84] J. T. Bonner and Whitfield J. Bell, Jr. “What is money for?”: an interview with Edwin Grant Conklin, 1952. Proceedings of the American Philosophical Society, 128(1):79–84, March 1984. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986495.

Borah:1969:CPD

[BC69] Woodrow Borah and Sherburne F. Cook. Conquest and population: a demographic approach to Mexican history. Proceedings of the American Philosophical Society, 113(2):177–183, April 1969. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985964.

Brooks:1981:SED

[BCGR81] Harvey Brooks, James F. Crow, John H. Gibbons, and David Rose. Symposium on energy discussion. Proceedings of the American Philosophical Society, 125(4):274–276, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986329.

Beams:1957:MSE

[Bea57] J. W. Beams. The magnetically supported equilibrium ultracentrifuge. Proceedings of the American Philosophical Society, 101 (1):63–69, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985140.

Beach:1961:NS

[Bea61] Edward L. Beach. The nature of submarining. Proceedings of the American Philosophical Society, 105(6):513–516, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985159.

Bearn:1996:LTN

[Bea96] Alexander G. Bearn. Lewis Thomas (25 November 1913–December 1993). Proceedings of the American Philosophical Society, 140(2): 246–251, June 1996. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987330.

Becker:1962:FPG

[Bec62] Marvin B. Becker. Florentine popular government (1343–1348). Proceedings of the American Philosophical Society, 106(4):360–382, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985269.

Becker:1989:LPT

[Bec89] Marshall Joseph Becker. Lenape population at the time of European contact: Estimating native numbers in the Lower Delaware Valley. Proceedings of the American Philosophical Society, 133(2): 112–122, June 1989. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987042.

Becker:1992:HCE

[Bec92] Gary S. Becker. Human capital and the economy. Proceedings of the American Philosophical Society, 136(1):85–92, March 1992. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986801.

Becker:1994:GJS

[Bec94] Gary S. Becker. George Joseph Stigler (17 January 1911–1 December 1991). Proceedings of the American Philosophical Society, 138(2):348–354, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987131.

Bedini:1964:TJC

[Bed64] Silvio A. Bedini. Thomas Jefferson, clock designer. Proceedings of the American Philosophical Society, 108(3):163–180, June 1964.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985678.

Bedini:1981:PP

[Bed81] Silvio A. Bedini. The Papal pachyderms. Proceedings of the American Philosophical Society, 125(2):75–90, April 1981. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986637.

Begley:1997:KAF

[Beg97] Louis Begley. Kafka: The axe for the frozen sea inside us. Proceedings of the American Philosophical Society, 141(3):253–261, September 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987201.

Beik:1951:AMN

[Bei51] Paul H. Beik. Abbé Maury and the National Assembly. Proceedings of the American Philosophical Society, 95(5):546–555, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143241.

Beidleman:1960:SBS

[Bei60] Richard G. Beidleman. Some biographical sidelights on Thomas Nuttall, 1786–1859. Proceedings of the American Philosophical Society, 104(1):86–100, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985605.

Bell:1950:SAS

[Bel50] Whitfield J. Bell, Jr. Some American students of “that shining oracle of physic,” Dr. William Cullen of Edinburgh, 1755–1766. Proceedings of the American Philosophical Society, 94(3):275–281, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143563.

Bell:1951:RMJ

[Bel51] Whitfield J. Bell, Jr. The Reverend Mr. Joseph Morgan, an American correspondent of the Royal Society, 1732–1739. Proceedings of the American Philosophical Society, 95(3):254–261, June 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143063.

Bell:1955:BFG

[Bel55] Whitfield J. Bell, Jr. Benjamin Franklin and the German charity schools. Proceedings of the American Philosophical Society, 99(6): 381–387, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143919.

Bell:1956:ACS

[Bel56] Whitfield J. Bell, Jr. “All clear sunshine”: New letters of Franklin and Mary Stevenson Hewson. Proceedings of the American Philosophical Society, 100(6):521–536, December 1956. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143786.

Bell:1959:AAA

[Bel59] Whitfield J. Bell, Jr. Archives and autographs in the American Philosophical Society Library. Proceedings of the American Philosophical Society, 103(6):761–767, December 1959. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985388.

Bell:1964:AOA

[Bel64a] Whitfield J. Bell, Jr. Astronomical observatories of the American Philosophical Society, 1769–1843. Proceedings of the American Philosophical Society, 108(1):7–14, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985991.

Bell:1964:I

[Bel64b] Whitfield J. Bell, Jr. Introduction. Proceedings of the American Philosophical Society, 108(2):145–146, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985892.

Bell:1972:SUN

[Bel72] Whitfield J. Bell, Jr. As others saw us: Notes on the reputation of the American Philosophical Society. Proceedings of the American Philosophical Society, 116(3):269–278, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986120.

Bell:1975:MHS

[Bel75] Whitfield J. Bell, Jr. Minis Hays, Secretary, Librarian, and Benefactor of the American Philosophical Society. Proceedings of the American Philosophical Society, 119(5):401–411, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986681.

Bell:1992:GCH

[Bel92] Daniel Bell. George C. Homans (11 August 1910–29 May 1989). Proceedings of the American Philosophical Society, 136(4):586–593, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986764.

Bell:1995:LBM

[Bel95] Alan Bell. Lord Bessborough (29 March 1913–5 December 1993). Proceedings of the American Philosophical Society, 139(2):172–175, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987264.

Bennett:1953:MT

[Ben53] Emmett L. Bennett, Jr. The Mycenae tablets. Proceedings of the American Philosophical Society, 97(4):422–470, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3149293.

Bentley:1971:PDS

[Ben71] Gerald Eades Bentley. The profession of dramatist in Shakespeare’s time. Proceedings of the American Philosophical Society, 115(6): 480–483, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985840.

Benedict:1979:HML

[Ben79] H. Bradley Benedict. Hacienda management in Late Colonial Northern Mexico: a case study of Juan Bustamante and the hacienda of Dolores 1790–1820. Proceedings of the American Philosophical Society, 123(6):391–409, December 1979. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986617.

Benedict:1985:NPW

[Ben85] Manson Benedict. Nuclear power in the world today. Proceedings of the American Philosophical Society, 129(4):456–475, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986939.

Bentley:1991:GDA

[Ben91] Gerald Eades Bentley. The grateful dead: Actors’ testamentary bequests to women 1580–1651. Proceedings of the American Philosophical Society, 135(3):382–387, September 1991. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986775.

Berkeley:1954:HPC

[Ber54] Francis L. Berkeley, Jr. History and problems of the control of manuscripts in the United States. Proceedings of the American Philosophical Society, 98(3):171–178, June 1954. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143842.

Berkner:1957:IGY

[Ber57] Lloyd Viel Berkner. The International Geophysical Year, 1957– 1958: a pattern for international cooperation in research. Proceedings of the American Philosophical Society, 101(2):159–163, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985509.

Berle:1967:PML

[Ber67] Adolf A. Berle. Penrose Memorial Lecture. The laws of power: an approach to its systematic study. Proceedings of the American Philosophical Society, 111(5):249–253, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986041.

Berkeley:1993:BFD

[Ber93] Edmund Berkeley. Benjamin Franklin and a “dear ould friend”. Proceedings of the American Philosophical Society, 137(3):399–405, September 1993. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987000.

Bergson:1995:BBR

[Ber95] Abram Bergson. The Big Bang in Russia: an overview. Proceedings of the American Philosophical Society, 139(4):335–349, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987233.

Berg:1998:UOG

[Ber98] Paul Berg. Understanding our genes: a legacy for the next millennium. Proceedings of the American Philosophical Society, 142(3): 416–424, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152247.

Best:1952:SDC

[Bes52] Charles H. Best. Studies on diabetes and cirrhosis. Proceedings of the American Philosophical Society, 96(1):27–29, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143739.

Bestor:1960:HSS

[Bes60] Arthur Bestor. History, social studies, and citizenship: The responsibility of the public schools. Proceedings of the American Philosophical Society, 104(6):549–557, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985529.

Best:1979:EAS

[Bes79] Gary Dean Best. An evangelist among skeptics: Hoover’s bid for leadership of the GOP, 1937–1938. Proceedings of the American Philosophical Society, 123(1):1–14, February 1979. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986467.

Bevilacqua:1967:BID

[Bev67] Vincent M. Bevilacqua. Baconian influences in the development of Scottish rhetorical theory. Proceedings of the American Philosophical Society, 111(4):212–218, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986008.

Bevington:1989:PED

[Bev89] David Bevington. “Is this the promised end?” Death and dying in “King Lear”. Proceedings of the American Philosophical Society, 133(3):404–415, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987108.

Bevington:1993:NFJ

[Bev93] David Bevington. Northrop Frye (14 July 1912–23 January 1991). Proceedings of the American Philosophical Society, 137(1):126–128, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986950.

Branscomb:1977:RHC

[BG77] Lewis M. Branscomb and Denos C. Gazis. Resilience, hypotheticality and computers — designing social systems under uncertainty. Proceedings of the American Philosophical Society, 121(5):346–349, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986337.

Bowen:1991:HTE

[BG91] Alan C. Bowen and Bernard R. Goldstein. Hipparchus’ treatment of early Greek astronomy: The case of Eudoxus and the length of daytime. Proceedings of the American Philosophical Society, 135(2):233–254, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987033.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society, 130(2):196–231, June 1986. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987181.

Bicha:1965:PFP

[Bic65] Karel Denis Bicha. The plains farmer and the prairie province frontier, 1897–1914. Proceedings of the American Philosophical Society, 109(6):398–440, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986139.

Bieber:1957:BSC

[Bie57] Margarete Bieber. A bronze statuette in Cincinnati and its place in the history of the Asklepios types. Proceedings of the American Philosophical Society, 101(1):70–92, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985141.

Bieber:1959:RMG

[Bie59] Margarete Bieber. Roman men in Greek himation (Romani Palliati) a contribution to the history of copying. Proceedings of the American Philosophical Society, 103(3):374–417, June 1959. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985474.

Bieber:1962:CHW

[Bie62] Margarete Bieber. The copies of the Herculaneum women. Proceedings of the American Philosophical Society, 106(2):111–134, April 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985377.

Biggs:1954:SOC

[Big54] John Biggs, Jr. Some observations on the case of the United States ex rel. Smith v. Baldi. Proceedings of the American Philosophical Society, 98(4):240–244, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143875.

Billingham:1971:BT

[Bil71] Rupert E. Billingham. The biology of transplantation. Proceedings of the American Philosophical Society, 115(4):251–257, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986085.

Birkhoff:1993:MH

[Bir93a] Garrett Birkhoff. Mathematics at Harvard in the 1940s. Proceedings of the American Philosophical Society, 137(2):268–272, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986733.

Birkhoff:1993:OZA

[Bir93b] Garrett Birkhoff. Oscar Zariski (24 April 1899–4 July 1986). Proceedings of the American Philosophical Society, 137(2):306–320, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986739.

Bisson:1979:MTP

[Bis79] Thomas N. Bisson. Mediterranean territorial power in the Twelfth Century. Proceedings of the American Philosophical Society, 123 (2):143–150, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986237.

Bass:1996:LBM

[BK96] Hyman Bass and Irwin Kra. Lipman Bers (22 May 1914–29 October 1993). Proceedings of the American Philosophical Society, 140(2):206–219, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987326.

Bell:1957:FWA

[BL57] Whitfield J. Bell, Jr. and Leonard W. Labaree. Franklin and the “Wagon Affair,” 1755. Proceedings of the American Philosophical Society, 101(6):551–558, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985523.

Blanshard:1952:NPA

[Bla52] Brand Blanshard. The new philosophy of analysis. Proceedings of the American Philosophical Society, 96(3):227–230, June 1952. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143832.

Blake:1957:SIS

[Bla57] John B. Blake. Scientific institutions since the Renaissance: Their role in medical research. Proceedings of the American Philosophical Society, 101(1):31–62, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985139.

Blake:1962:MGN

[Bla62] John B. Blake. Mary Gove Nichols, prophetess of health. Proceedings of the American Philosophical Society, 106(3):219–234, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985177.

Blanshard:1965:CRB

[Bla65] Brand Blanshard. Critical reflections on behaviorism. Proceedings of the American Philosophical Society, 109(1):22–28, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985775.

Blanshard:1971:RAR

[Bla71] Brand Blanshard. The revolt against reason in theology. Proceedings of the American Philosophical Society, 115(2):83–87, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985849.

Black:1973:ARY

[Bla73a] Frederick R. Black. The American Revolution as “yardstick” in the debates on the Constitution, 1787–1788. Proceedings of the American Philosophical Society, 117(3):162–185, June 1973. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986541.

Blackall:1973:SVS

[Bla73b] Eric A. Blackall. Shakespeare as viewed by a Swiss peasant of the Eighteenth Century. Proceedings of the American Philosophical Society, 117(2):112–116, April 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986061.

Blaisdell:1988:FPE

[Bla88] Lowell L. Blaisdell. Félix Pyat, the “evil genius” of the Commune of Paris. Proceedings of the American Philosophical Society, 132 (4):330–370, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986959.

Blegen:1957:NP

[Ble57] Carl W. Blegen. Nestor’s Pylos. Proceedings of the American Philosophical Society, 101(4):379–385, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985405.

Bloomfield:1986:PES

[Blo86] Morton W. Bloomfield. Poetry in early societies. Proceedings of the American Philosophical Society, 130(2):247–250, June 1986. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987183.

Bloom:1997:FSN

[Blo97] Floyd E. Bloom. Francis O. Schmitt (23 November 1903–3 October 1995). Proceedings of the American Philosophical Society, 141(4):

504–508, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987226.

Blum:1982:FEP

[Blu82] Jerome Blum. Fiction and the European peasantry: The realist novel as a historical source. Proceedings of the American Philosophical Society, 126(2):122–139, April 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986356.

Blumberg:1988:SCJ

[Blu88] Baruch S. Blumberg. Science, the constitution and the judicial system. Proceedings of the American Philosophical Society, 132(3): 247–251, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143851.

Blumberg:1997:JLR

[Blu97] Baruch S. Blumberg. The Jayne Lecture. Reflections on the future of medical science. Proceedings of the American Philosophical Society, 141(2):125–140, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987297.

Boas:1951:IPS

[Boa51] George Boas. The influence of philosophy on the sciences. Proceedings of the American Philosophical Society, 95(5):528–537, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143237.

Boas:1954:FCC

[Boa54] George Boas. A Fourteenth-Century cosmology. Proceedings of the American Philosophical Society, 98(1):50–59, February 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143670.

Boas:1971:RR

[Boa71] George Boas. The retreat from reason. Proceedings of the American Philosophical Society, 115(6):432–436, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985835.

Bodde:1963:BCC

[Bod63] Derk Bodde. Basic concepts of Chinese law: The genesis and evolution of legal thought in traditional China. Proceedings of the American Philosophical Society, 107(5):375–398, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985644.

Bodde:1985:SCC

[Bod85] Derk Bodde. Sex in Chinese civilization. Proceedings of the American Philosophical Society, 129(2):161–172, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986987.

Boewe:1958:MCR

[Boe58] Charles Boewe. The manuscripts of C. S. Rafinesque (1783–1840). Proceedings of the American Philosophical Society, 102(6):590–595, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985568.

Bogen:1999:RWS

[Bog99] Joseph E. Bogen. Roger Wolcott Sperry (20 August 1913–17 April 1994). Proceedings of the American Philosophical Society, 143(3): 492–500, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181963.

Bok:1952:SSM

[Bok52] Bart J. Bok. Studies of the Southern Milky Way. Proceedings of the American Philosophical Society, 96(5):540–555, October 1952. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143655.

Bok:1995:RFP

[Bok95] Derek Bok. The Richard A. F. Penrose Lecture. America’s disenchantment with government. Proceedings of the American Philosophical Society, 139(4):326–334, December 1995. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987232.

Boll:1962:MSM

[Bol62] Theophilus E. M. Boll. May Sinclair and the medico-psychological clinic of London. Proceedings of the American Philosophical Society, 106(4):310–326, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985265.

Boney:1974:NFR

[Bon74] F. N. Boney. Nathaniel Francis, representative Antebellum Southerner. Proceedings of the American Philosophical Society, 118(5): 449–458, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986309.

Bonner:1979:BBC

[Bon79] John Tyler Bonner. The biological basis of culture. Proceedings of the American Philosophical Society, 123(4):219–221, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986581.

Bonner:1994:EBJ

[Bon94] James Bonner. Erwin Bünning (23 January 1906–4 October 1990). Proceedings of the American Philosophical Society, 138(2):318–320, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987126.

Boocock:1975:SCC

[Boo75] Sarane Spence Boocock. The social context of childhood. Proceedings of the American Philosophical Society, 119(6):419–429, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986376.

Booth:1978:AFM

[Boo78] Christopher C. Booth. Ann Fothergill: The Mistress of Harpur Street. Proceedings of the American Philosophical Society, 122 (5):340–354, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986689.

Booth:1997:SHH

[Boo97] Christopher Booth. Sir Harold Himsworth (19 May 1905–1 November 1993). Proceedings of the American Philosophical Society, 141 (1):84–87, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987252.

Boring:1950:GMS

[Bor50] Edwin G. Boring. Great men and scientific progress. Proceedings of the American Philosophical Society, 94(4):339–351, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143518.

Boring:1952:VSB

[Bor52] Edwin G. Boring. The validation of scientific belief. A conspectus of the symposium. Proceedings of the American Philosophical Society, 96(5):535–539, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143654.

Boring:1964:TTM

[Bor64] Edwin G. Boring. The trend toward mechanism. Proceedings of the American Philosophical Society, 108(6):451–454, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985861.

Borthwick:1965:OUS

[Bor65] Doris Esch Borthwick. Outfitting the United States Exploring Expedition: Lieutenant Charles Wilkes’ European assignment, August–November, 1836. Proceedings of the American Philosophical Society, 109(3):159–172, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986129.

Borza:1987:TPA

[Bor87] Eugene N. Borza. Timber and politics in the Ancient World: Macedon and the Greeks. Proceedings of the American Philosophical Society, 131(1):32–52, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987115.

Borel:1993:DMS

[Bor93] Armand Borel. Deane Montgomery (2 September 1909–15 March 1992). Proceedings of the American Philosophical Society, 137(3): 452–456, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987005.

Boulding:1963:MIR

[Bou63] Kenneth E. Boulding. The misallocation of intellectual resources. Proceedings of the American Philosophical Society, 107(2):117–120, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985441.

Bouwsma:1984:JCA

[Bou84] William J. Bouwsma. John Calvin’s anxiety. Proceedings of the American Philosophical Society, 128(3):252–256, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986886.

Bowen:1957:HC

[Bow57] Catherine Drinker Bowen. Historians courageous. Proceedings of the American Philosophical Society, 101(3):249–254, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985367.

Bowers:1973:A

[Bow73] John Z. Bowers. Acupuncture. Proceedings of the American Philosophical Society, 117(3):143–151, June 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986539.

Bowersock:1990:ELP

[Bow90] G. W. Bowersock. Edward lear in petra. Proceedings of the American Philosophical Society, 134(3):309–320, September 1990. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986969.

Bowersock:1991:RSM

[Bow91a] G. W. Bowersock. Ronald Syme (March 11, 1903–September 4, 1989). Proceedings of the American Philosophical Society, 135(1): 118–122, March 1991. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987156.

Bowersock:1991:RMO

[Bow91b] G. W. Bowersock. Russell Meiggs (October 20, 1902–June 24, 1989). Proceedings of the American Philosophical Society, 135(3): 474–477, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986782.

Bowsky:1993:DJH

[Bow93] William M. Bowsky. David J. Herlihy (8 May 1930–21 February 1991). Proceedings of the American Philosophical Society, 137(3): 446–449, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987004.

Bowen:1995:IBR

[Bow95] William G. Bowen. Inside the boardroom: a reprise. Proceedings of the American Philosophical Society, 139(4):350–357, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987234.

Bowering:1997:CTI

[Böw97] Gerhard Böwering. The concept of time in Islam. Proceedings of the American Philosophical Society, 141(1):55–66, March 1997. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987249.

Bowersock:1999:RER

[Bow99] G. W. Bowersock. The Roman Emperor as Russian Tsar: Tacitus and Pushkin. Proceedings of the American Philosophical Society, 143(1):130–147, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181981.

Boyd:1958:MMT

[Boy58] Julian P. Boyd. The megalonyx, the megatherium, and Thomas Jefferson’s lapse of memory. Proceedings of the American Philosophical Society, 102(5):420–435, October 1958. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985587.

Boyce:1973:GIC

[Boy73] Douglas W. Boyce. A glimpse of Iroquois culture history through the eyes of Joseph Brant and John Norton. Proceedings of the American Philosophical Society, 117(4):286–294, August 1973. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986696.

Boyer:1974:CPL

[Boy74] John W. Boyer. Catholic priests in Lower Austria: Antiliberalism, occupational anxiety, and radical political action in Late Nineteenth-Century Vienna. Proceedings of the American Philosophical Society, 118(4):337–369, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986281.

Boyer:1998:MCS

[Boy98] Paul Boyer. Merle Curti (15 September 1897–9 March 1996). Proceedings of the American Philosophical Society, 142(2):276–281, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3152304.

Bourgeois-Pichat:1967:SBD

[BP67] Jean Bourgeois-Pichat. Social and biological determinants of human fertility in nonindustrial societies. Proceedings of the American Philosophical Society, 111(3):160–163, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985712.

Brackenridge:1979:JKM

[BR79] J. Bruce Brackenridge and Mary Ann Rossi. Johannes Kepler’s on the more certain fundamentals of astrology, Prague 1601. Proceedings of the American Philosophical Society, 123(2):85–116, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986232.

Bray:1970:JNN

[Bra70] Martha Coleman Bray. Joseph Nicolas Nicollet, geologist. Proceedings of the American Philosophical Society, 114(1):37–59, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985723.

Braidwood:1972:PIS

[Bra72a] Robert J. Braidwood. Prehistoric investigations in Southwestern Asia. Proceedings of the American Philosophical Society, 116 (4):310–320, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985901.

Branscomb:1972:MSU

[Bra72b] Lewis M. Branscomb. The metric system in the United States. Proceedings of the American Philosophical Society, 116(4):294–300, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985899.

Brain:1981:HNS

[Bra81] James Lewton Brain. Homage to Neptune: Shipboard initiation rites. Proceedings of the American Philosophical Society, 125(2): 128–133, April 1981. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986641.

Brackenridge:1985:KPA

[Bra85] J. Bruce Brackenridge. Kuhn, paradigms, and astronomy: Astronomy as a case study of Kuhnian paradigms. Proceedings of the American Philosophical Society, 129(4):433–455, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986938.

Brady:1992:BFS

[Bra92] Joseph V. Brady. Burrhus Frederic Skinner (March 20, 1904– August 18, 1990). Proceedings of the American Philosophical Society, 136(2):312–316, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987177.

Brew:1961:EAS

[Bre61] J. O. Brew. Emergency archaeology: Salvage in advance of technological progress. Proceedings of the American Philosophical Society, 105(1):1–10, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985352.

Bree:1968:PCC

[Bré68] Germaine Brée. Proust’s Combray Church: Illiers or Vermeer? Proceedings of the American Philosophical Society, 112(1):5–7, February 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986098.

Bree:1970:BRN

[Bré70] Germaine Brée. Beyond Romanticism: Narrative play in Gidian fiction. Proceedings of the American Philosophical Society, 114(3): 163–167, June 1970. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985692.

Brent:1992:SPB

[Bre92] Leslie Brent. Sir Peter Brian Medawar (28 February 1915–2 October 1987). Proceedings of the American Philosophical Society, 136(3):438–441, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986918.

Brinton:1955:RDI

[Bri55] Crane Brinton. Reflections on the desertion of the intellectuals. Proceedings of the American Philosophical Society, 99(4):219– 223, August 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143700.

Bridenbaugh:1961:CSA

[Bri61] Carl Bridenbaugh. Church and state in America, 1689–1775. Proceedings of the American Philosophical Society, 105(6):521–524, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985161.

Brim:1975:CSI

[Bri75] Orville G. Brim, Jr. Childhood social indicators: Monitoring the ecology of development. Proceedings of the American Philosophical Society, 119(6):413–418, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986375.

Brooke:1951:LTW

[Bro51] C. F. Tucker Brooke. The life and times of William Gager (1555– 1622). Proceedings of the American Philosophical Society, 95(4): 401–431, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143282.

Brown:1952:RL

[Bro52] Sanborn C. Brown. Rumford lamps. Proceedings of the American Philosophical Society, 96(1):37–44, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143741.

Broughton:1959:RSP

[Bro59] T. Robert S. Broughton. The Romanization of Spain: The problem and the evidence. Proceedings of the American Philosophical Society, 103(5):645–651, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985423.

Brogan:1962:RRB

[Bro62a] D. W. Brogan. The remote revolution — a British view. Proceedings of the American Philosophical Society, 106(1):1–9, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985205.

Brooks:1962:EFJ

[Bro62b] Van Wyck Brooks. Ernest Fenollosa and Japan. Proceedings of the American Philosophical Society, 106(2):106–110, April 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985376.

Brombert:1970:MPV

[Bro70] Victor Brombert. Malraux: Poet of violence and destiny. Proceedings of the American Philosophical Society, 114(3):187–190, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985695.

Brown:1972:DTA

[Bro72] W. Norman Brown. Duty as truth in Ancient India. Proceedings of the American Philosophical Society, 116(3):252–268, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986119.

Bronfenbrenner:1975:RRE

[Bro75a] Urie Bronfenbrenner. Reality and research in the ecology of human development. Proceedings of the American Philosophical Society, 119(6):439–469, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986378.

Brooks:1975:EPP

[Bro75b] Harvey Brooks. Expertise and politics-problems and tensions. Proceedings of the American Philosophical Society, 119(4):257–261, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986286.

Brodhead:1978:WCL

[Bro78] Michael J. Brodhead. The work of Charles Lucien Bonaparte in America. Proceedings of the American Philosophical Society, 122 (4):198–203, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986531.

Brooks:1981:PEP

[Bro81] Harvey Brooks. Perspectives on the energy problem. Proceedings of the American Philosophical Society, 125(4):245–250, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986325.

Brownlee:1985:WFM

[Bro85] W. Elliot Brownlee. Wilson and financing the modern state: The Revenue Act of 1916. Proceedings of the American Philosophical Society, 129(2):173–210, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986988.

Brown:1989:PTP

[Bro89] Michael K. Brown. Piecing together the past: Recent research on the American China manufactory 1769–1772. Proceedings of the American Philosophical Society, 133(4):555–579, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986876.

Brombert:1991:MIG

[Bro91] Victor Brombert. Meanings and indeterminacy in Gogol’s “The Overcoat”. Proceedings of the American Philosophical Society, 135 (4):569–575, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986817.

Brues:1963:SHR

[Bru63] Austin M. Brues. Somatic hazards of radiation. Proceedings of the American Philosophical Society, 107(1):1–4, February 1963. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985463.

Bryden:1988:AMF

[Bry88] D. J. Bryden. Alva Mason, the Franklin Institute, and the origins of philosophical and chemical instrument manufacture in the United States. Proceedings of the American Philosophical Society, 132(4): 400–419, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986961.

Bryan:1960:BQN

[BT60] Alan L. Bryan and Donald R. Tuohy. A basalt quarry in Northeastern Oregon. Proceedings of the American Philosophical Society, 104 (5):485–510, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985233.

Buhler:1970:FIC

[Büh70] Curt F. Bühler. False information in the colophons of incunabula. Proceedings of the American Philosophical Society, 114(5):398–406, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985806.

Bundy:1975:DAF

[Bun75] William P. Bundy. Dictatorships and American foreign policy. Proceedings of the American Philosophical Society, 119(4):251–256, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986285.

Bundy:1979:C

[Bun79] William P. Bundy. Commentary. Proceedings of the American Philosophical Society, 123(5):276–278, October 1979. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986589.

Bunge:1991:PPM

[Bun91a] Mario Bunge. A philosophical perspective on the mind-body problem or, why neuroscientists and psychologists should care about philosophy. Proceedings of the American Philosophical Society, 135 (4):513–523, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986813.

Bunnell:1991:MWP

[Bun91b] Peter C. Bunnell. Minor White’s photographic sequence, “rural cathedrals”: a reading. Proceedings of the American Philosophical Society, 135(4):557–568, December 1991. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986816.

Burgess:1952:MID

[Bur52] W. Randolph Burgess. Money as an instrument of democracy. Proceedings of the American Philosophical Society, 96(5):566–572, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143657.

Buranelli:1960:WPJ

[Bur60a] Vincent Buranelli. William Penn and James II. Proceedings of the American Philosophical Society, 104(1):35–53, February 1960. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985603.

Burington:1960:PFP

[Bur60b] Richard Stevens Burington. The problem of formulating a problem: General considerations. Proceedings of the American Philosophical Society, 104(5):429–443, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985229.

Burstyn:1973:ESM

[Bur73] Joan N. Burstyn. Education and sex: The medical case against higher education for women in England, 1870–1900. Proceedings of the American Philosophical Society, 117(2):79–89, April 1973. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986058.

Burton:1975:IKO

[Bur75] David H. Burton. The intellectual kinship of Oliver Wendell Holmes, Jr., Frederick E. Pollock, and Harold J. Laski. Proceedings of the American Philosophical Society, 119(2):133–142, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986631.

Burton:1980:TDA

[Bur80] William L. Burton. “Title deed to America” union ethnic regiments in the Civil War. Proceedings of the American Philosophical Society, 124(6):455–463, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986246.

Burton:1983:GMS

[Bur83] John W. Burton. The ghost of Malinowski in the Southern Sudan: Evans-Pritchard and ethnographic fieldwork. Proceedings of the American Philosophical Society, 127(4):278–289, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986275.

Burris:1984:NMC

[Bur84] Robert H. Burris. Nitrogen metabolism in the coral–algal symbiosis. Proceedings of the American Philosophical Society, 128(1):85– 92, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986496.

Burkhardt:1986:ED

[Bur86] Frederick Burkhardt. Editing Darwin. Proceedings of the American Philosophical Society, 130(4):367–373, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986923.

Burton:1988:STN

[Bur88] John W. Burton. Shadows at twilight: a note on history and the ethnographic present. Proceedings of the American Philosophical Society, 132(4):420–433, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986962.

Burchfiel:1993:ECO

[Bur93a] B. C. Burchfiel. Evolution of continental and oceanic crust. Proceedings of the American Philosophical Society, 137(1):1–29, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986942.

Burger:1993:DP

[Bur93b] Warren E. Burger. The decline of professionalism. Proceedings of the American Philosophical Society, 137(4):481–487, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987067.

Bush:1954:SM

[Bus54] Vannevar Bush. Scientific motivation. Proceedings of the American Philosophical Society, 98(4):225–232, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143873.

Bushkovitch:1997:PHC

[Bus97] Paul Bushkovitch. Power and the historian: The case of Tsarevich Aleksei 1716–1718 and N. G. Ustrialov 1845–1859. Proceedings of the American Philosophical Society, 141(2):177–212, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987301.

Butterfield:1950:PWJ

[But50] Lyman H. Butterfield. Psychological warfare in 1776: The Jefferson–Franklin plan to cause Hessian desertions. Proceedings of the American Philosophical Society, 94(3):233–241, June 1950. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143556.

Butterfield:1951:DRG

[But51] Lyman H. Butterfield. Dr. Rush to Governor Henry on the Declaration of Independence and the Virginia Constitution. Proceedings of the American Philosophical Society, 95(3):250–253, June 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143062.

Butterfield:1954:AEE

[But54] Lyman H. Butterfield. Archival and editorial enterprise in 1850 and in 1950: Some comparisons and contrasts. Proceedings of the American Philosophical Society, 98(3):159–170, June 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143841.

Butterfield:1974:TDK

[But74] L. H. Butterfield. Tending a dragon-killer: Notes for the biographer of Mrs. John Quincy Adams. Proceedings of the American Philosophical Society, 118(2):165–178, April 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986664.

Bynum:1998:DRM

[Byn98] Caroline Walker Bynum. Death and resurrection in the Middle Ages: Some modern implications. Proceedings of the American Philosophical Society, 142(4):589–596, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3152283.

C:1965:SCS

[C.65] G. W. C. Symposium on cardiac surgery: Prefatory statement. Proceedings of the American Philosophical Society, 109(3):111, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986124.

Cadbury:1951:MMG

[Cad51] Henry J. Cadbury. Mixed motives in the gospels. Proceedings of the American Philosophical Society, 95(2):117–124, April 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143329.

Cain:1960:BPC

[Cai60] Julien Cain. Bergson Président de la Commission Internationale de Cooperation Intellectuelle. Proceedings of the American Philosophical Society, 104(4):404–407, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985619.

Cammann:1991:CHT

[Cam91] Schuyler Cammann. Chinese hexagrams, trigrams, and the binary system. Proceedings of the American Philosophical Society, 135(4): 576–589, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986818.

Cantril:1960:CNP

[Can60] Hadley Cantril. Concerning the nature of perception. Proceedings of the American Philosophical Society, 104(5):467–473, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985231.

Cannon:1961:IUT

[Can61] Walter F. Cannon. The impact of uniformitarianism: Two letters from John Herschel to Charles Lyell, 1836–1837. Proceedings of the American Philosophical Society, 105(3):301–314, June 1961. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985457.

Cannon:1977:SWJ

[Can77] Garland Cannon. Sir William Jones and the association between East and West. Proceedings of the American Philosophical Society, 121(2):183–187, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986526.

Cappon:1974:RMB

[Cap74] Lester J. Cappon. Retracing and mapping the Bartrams’ southern travels. Proceedings of the American Philosophical Society, 118(6): 507–513, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986400.

Carman:1957:ONC

[Car57a] Harry J. Carman. Older and newer concepts of academic freedom. Proceedings of the American Philosophical Society, 101(5):446–449, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985198.

Carmichael:1957:SIA

[Car57b] Leonard Carmichael. The Smithsonian Institution and the American Philosophical Society. Proceedings of the American Philosophical Society, 101(5):401–408, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985194.

Carter:1958:ARA

[Car58] George F. Carter. Archaeology in the Reno area in relation to Age of Man and the culture sequence in America. Proceedings of the American Philosophical Society, 102(2):174–192, April 1958. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985283.

Carneiro:1974:HSS

[Car74] Robert L. Carneiro. Herbert Spencer’s “The Study of Sociology” and the rise of social science in America. Proceedings of the American Philosophical Society, 118(6):540–554, December 1974. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986404.

Carroll:1981:AMJ

[Car81] Diane Lee Carroll. Antique metal-joining formulas in the “mappae clavicula”. Proceedings of the American Philosophical Society, 125(2):91–103, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986638.

Carter:1989:NPW

[Car89a] Edward C. Carter, II. Naturalization in Philadelphia, 1789–1806: a “wild Irishman” under every Federalist’s bed — revisited twenty years later. Proceedings of the American Philosophical Society, 133(2):175–177, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987048.

Carter:1989:WIU

[Car89b] Edward C. Carter, II. A “wild Irishman” under every Federalist’s bed: Naturalization in Philadelphia, 1789–1806. Proceedings of the American Philosophical Society, 133(2):178–189, June 1989. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987049.

Carrier:1990:NWC

[Car90a] George F. Carrier. Nuclear winter, current understanding. Proceedings of the American Philosophical Society, 134(2):83–89, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986839.

Carter:1990:F

[Car90b] Edward C. Carter, II. Foreword. Proceedings of the American Philosophical Society, 134(4):338–339, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986891.

Carlos:1993:SAP

[Car93] Juan Carlos, I. Spain and the American Philosophical Society. Proceedings of the American Philosophical Society, 137(4):475–480, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987066.

Caskey:1969:CMM

[Cas69] John L. Caskey. Crises in the Minoan–Mycenaean world. Proceedings of the American Philosophical Society, 113(6):433–449, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985826.

Casimir:1992:LCP

[Cas92] Hendrik B. G. Casimir. Léon Charles Prudent van Hove (10 February 1924–2 September 1991). Proceedings of the American Philosophical Society, 136(4):602–606, December 1992. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986766.

Coale:1996:FDM

[CB96] Ansley J. Coale and Judith Banister. Five decades of missing females in China. Proceedings of the American Philosophical Society, 140(4):421–450, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987286.

Conley:1997:FIC

[CBA97] Timothy K. Conley and Melissa Brewer-Anderson. Franklin and Ingenhousz: a correspondence of interests. Proceedings of the American Philosophical Society, 141(3):276–296, September 1997. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987203.

Chetverikov:1961:CAE

[CBL61] S. S. Chetverikov, Malina Barker, and I. Michael Lerner. On certain aspects of the evolutionary process from the standpoint of modern genetics. Proceedings of the American Philosophical Society, 105(2):167–195, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985629.

Calvin:1964:AA

[CC64] Melvin Calvin and G. J. Calvin. Atom to Adam. Proceedings of the American Philosophical Society, 108(2):73–87, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985888.

Catley-Carlson:1995:MMP

[CC95] Margaret Catley-Carlson. Myths, misgivings, and practical mitigations. Proceedings of the American Philosophical Society, 139(3): 247–255, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987190.

Campisi:1981:TBO

[CH81] Jack Campisi and Laurence M. Hauptman. Talking back: The Oneida Language and Folklore Project, 1938–1941. Proceedings of the American Philosophical Society, 125(6):441–448, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986299.

Chafee:1951:SPD

[Cha51] Zechariah Chafee, Jr. Some problems of the draft international covenant on human rights. Proceedings of the American Philosophical Society, 95(5):471–489, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143232.

Chandrasekhar:1964:CA

[Cha64] S. Chandrasekhar. The case for astronomy. Proceedings of the American Philosophical Society, 108(1):1–6, February 1964. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985990.

Chance:1971:YDS

[Cha71] Britton Chance, Jr. Yacht design: The state of the art. Proceedings of the American Philosophical Society, 115(6):477–479, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985839.

Chappell:1980:ASV

[Cha80] Edward A. Chappell. Acculturation in the Shenandoah Valley: Rhenish houses of the Massanutten Settlement. Proceedings of the American Philosophical Society, 124(1):55–89, February 1980. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986659.

Chance:1983:NBA

[Cha83] Britton Chance. A noninvasive biochemical assay and imaging of animal and human tissues by optical and nuclear magnetic resonance techniques. Proceedings of the American Philosophical Society, 127(1):1–25, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986361.

Chapin:1985:LBM

[Cha85] Seymour L. Chapin. A legendary bon mot?: Franklin’s “what is the good of a newborn baby?”. Proceedings of the American Philosophical Society, 129(3):278–290, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987011.

Chandler:1986:BMI

[Cha86] Alfred D. Chandler, Jr. The beginnings of the modern industrial corporation. Proceedings of the American Philosophical Society, 130(4):382–389, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986925.

Chance:1987:DPT

[Cha87] Britton Chance, Jr. The design and performance of twelve meter yachts. Proceedings of the American Philosophical Society, 131(4): 378–396, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987138.

Cherniss:1954:SEA

[Che54] Harold Cherniss. The sources of evil according to plato. Proceedings of the American Philosophical Society, 98(1):23–30, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143666.

Chern:1994:HWM

[Che94] Shing-Shen Chern. Hassler Whitney (23 March 1907–10 May 1989). Proceedings of the American Philosophical Society, 138(3):464–467, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986754.

Chinard:1950:ALR

[Chi50] Gilbert Chinard. Abbé Lefebvre De La Roche’s recollections of Benjamin Franklin. Proceedings of the American Philosophical Society, 94(3):214–221, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143554.

Chinard:1953:BFM

[Chi53] Gilbert Chinard. Benjamin Franklin et la muse provinciale. (French) [Benjamin Franklin and the provincial muse]. Proceedings of the American Philosophical Society, 97(5):493–510, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149266.

Chinard:1954:NAO

[Chi54] Gilbert Chinard. Notes on the American origins of the “Déclaration Des Droits De l’Homme Et Du Citoyen”. Proceedings of the American Philosophical Society, 98(6):383–396, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143860.

Chinard:1955:ABF

[Chi55] Gilbert Chinard. The apotheosis of Benjamin Franklin: Paris, 1790–1791. Proceedings of the American Philosophical Society, 99 (6):440–473, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143927.

Chinard:1956:MEF

[Chi56] Gilbert Chinard. Music enjoyed by Franklin. Performed by a group of artists under the direction of Henry S. Drinker. Proceedings of the American Philosophical Society, 100(4):331–337, August 1956. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143759.

Chinard:1957:AFA

[Chi57a] Gilbert Chinard. André and François-André Michaux and their predecessors. An essay on early botanical exchanges between America and France. Proceedings of the American Philosophical Society, 101(4):344–361, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985401.

Chinard:1957:RAB

[Chi57b] Gilbert Chinard. Recently acquired botanical documents. Proceedings of the American Philosophical Society, 101(6):508–522, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985519.

Chinard:1958:ECI

[Chi58] Gilbert Chinard. An Eighteenth-Century interpretation of the “struggle for existence.” Rétif de la Bretonne’s “École des pères”. Proceedings of the American Philosophical Society, 102(6):547–554, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985565.

Chinard:1959:RNT

[Chi59] Gilbert Chinard. Random notes on two “bagatelles”. Proceedings of the American Philosophical Society, 103(6):727–760, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985387.

Chow:1977:CE

[Cho77] Gregory C. Chow. The computer and economics. Proceedings of the American Philosophical Society, 121(5):350–354, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986338.

Chow:1989:EEE

[Cho89] Gregory C. Chow. Economics education and economic reform in China. Proceedings of the American Philosophical Society, 133(1): 64–74, March 1989. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987162.

Choppin:1998:THB

[Cho98] Purnell W. Choppin. From a three-headed bit to a major philanthropy: The surprising legacy of Howard Hughes. Proceedings of the American Philosophical Society, 142(3):425–431, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3152248.

Christman:1966:PSG

[Chr66] Webster M. Christman. The papers of Stephen Girard: Their preparation and historical interest. Proceedings of the American Philosophical Society, 110(6):383–385, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985791.

Christianson:1984:YJS

[Chr84] Paul Christianson. Young John Selden and the Ancient Constitution, ca. 1610–18. Proceedings of the American Philosophical Society, 128(4):271–315, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987098.

Clark:1950:VEL

[Cla50] John M. Clark. Varieties of economic law, and their limiting factors. Proceedings of the American Philosophical Society, 94(2): 121–126, April 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143211.

Clark:1953:PPD

[Cla53] Grenville Clark. The practical prospects for disarmament and genuine peace. Proceedings of the American Philosophical Society, 97 (6):645–651, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3149279.

Clark:1954:FPC

[Cla54] William Bell Clark. A Franklin postscript to Captain Cook’s voyages. Proceedings of the American Philosophical Society, 98(6): 400–405, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143862.

Clark:1957:HDF

[Cla57] William Bell Clark. A hoax on Doctor Franklin. Proceedings of the American Philosophical Society, 101(6):559–564, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985524.

Clark:1959:PML

[Cla59] Wilfrid E. Le Gros Clark. Penrose Memorial Lecture. The crucial evidence for human evolution. Proceedings of the American Philosophical Society, 103(2):159–172, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985146.

Clark:1963:MSC

[Cla63] Thomas D. Clark. The modern South in a changing America. Proceedings of the American Philosophical Society, 107(2):121–131, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985442.

Clagett:1964:NOM

[Cla64] Marshall Clagett. Nicole Oresme and medieval scientific thought. Proceedings of the American Philosophical Society, 108(4):298–309, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985910.

Clark:1966:PHB

[Cla66] Grahame Clark. Prehistory and human behavior. Proceedings of the American Philosophical Society, 110(2):91–99, April 1966. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985741.

Clark:1978:DIA

[Cla78] Joseph Sill Clark. Disarmament — international accord or World War III? Proceedings of the American Philosophical Society, 122 (2):59–62, April 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986621.

Clagett:1982:WMT

[Cla82] Marshall Clagett. William of Moerbeke: Translator of Archimedes. Proceedings of the American Philosophical Society, 126(5):356–366, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986212.

Clagett:1987:CGH

[Cla87] Marshall Clagett. Computer-generated hieroglyphs. Proceedings of the American Philosophical Society, 131(2):197–223, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986792.

Clark:1991:JSC

[Cla91] George R. Clark. Joseph Sill Clark (October 21, 1901–January 12, 1990). Proceedings of the American Philosophical Society, 135(1): 92–97, March 1991. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987153.

Cleland:1954:ENA

[Cle54] Ralph E. Cleland. Evolution of the North American Euoenotheras: The Strigosas. Proceedings of the American Philosophical Society, 98(3):189–203, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143845.

Cleland:1964:EHN

[Cle64] Ralph E. Cleland. The evolutionary history of the North American evening primroses of the “Biennis group”. Proceedings of the American Philosophical Society, 108(2):88–98, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985889.

Cleland:1998:JFB

[Cle98] Robert E. Cleland. James Frederick Bonner (1 September 1910–13 September 1996). Proceedings of the American Philosophical Society, 142(4):652–655, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3152288.

Cloud:1980:IBU

[Clo80] Preston Cloud. The improbable bureaucracy: The United States Geological Survey, 1879–1979. Proceedings of the American Philosophical Society, 124(3):155–167, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986368.

Coale:1967:VCH

[Coa67] Ansley J. Coale. The voluntary control of human fertility. Proceedings of the American Philosophical Society, 111(3):164–169, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985713.

Coale:1988:FMD

[Coa88] Ansley Coale. Fertility and mortality in different populations with special attention to China. Proceedings of the American Philosophical Society, 132(2):185–195, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986864.

Coale:1991:ASO

[Coa91] Ansley J. Coale. Alfred Sauvy (October 31, 1898–October 30, 1990). Proceedings of the American Philosophical Society, 135(4): 630–632, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986824.

Coale:1999:KDA

[Coa99] Ansley J. Coale. Kingsley Davis (20 August 1908–27 February 1997). Proceedings of the American Philosophical Society, 143(3): 452–454, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181957.

Cochran:1958:CFE

[Coc58] Thomas C. Cochran. Cultural factors in economic growth. Proceedings of the American Philosophical Society, 102(2):164–167, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985281.

Cochran:1964:MAA

[Coc64] Thomas C. Cochran. The Middle Atlantic Area in the economic history of the United States. Proceedings of the American Philosophical Society, 108(2):156–157, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985894.

Cochran:1970:TMS

[Coc70] Thomas C. Cochran. Toward a model for social change. Proceedings of the American Philosophical Society, 114(5):366–370, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985801.

Coffin:1986:EFR

[Cof86] David R. Coffin. The Elysian Fields of Rousham. Proceedings of the American Philosophical Society, 130(4):406–423, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986927.

Cohen:1950:BFT

[Coh50] I. Bernard Cohen. Benjamin Franklin and the transit of Mercury in 1753: Together with a facsimile of a little-known scientific work printed by Franklin. Proceedings of the American Philosophical Society, 94(3):222–232, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143555.

Cohen:1952:OSP

[Coh52a] I. Bernard Cohen. Orthodoxy and scientific progress. Proceedings of the American Philosophical Society, 96(5):505–512, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143650.

Cohen:1952:THA

[Coh52b] I. Bernard Cohen. The two hundredth anniversary of Benjamin Franklin’s two lightning experiments and the introduction of the lightning rod. Proceedings of the American Philosophical Society, 96(3):331–366, June 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143838.

Cohen:1955:PSN

[Coh55] I. Bernard Cohen. Present status and needs of the history of science. Proceedings of the American Philosophical Society, 99(5): 343–347, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/3143934. See comments [Din55].

Cohen:1956:SPR

[Coh56] I. Bernard Cohen. Some problems in relation to the dates of Benjamin Franklin’s first letters on electricity. Proceedings of the American Philosophical Society, 100(6):537–544, December 1956. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143787.

Cohen:1982:TRC

[Coh82] Seymour S. Cohen. Two refugee chemists in the United States, 1794: How we see them. Proceedings of the American Philosophical Society, 126(4):301–315, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986176.

Cohen:1987:FFN

[Coh87] I. Bernard Cohen. Faraday and Franklin’s “newborn baby”. Proceedings of the American Philosophical Society, 131(2):177–182, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986790.

Cohen:1993:JKF

[Coh93] Paul A. Cohen. John King Fairbank (24 May 1907–14 September 1991). Proceedings of the American Philosophical Society, 137(2): 280–284, June 1993. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986735.

Cohen:1995:DTW

[Coh95] D. Walter Cohen. Dr. Thomas W. Evans, a Nineteenth-Century Renaissance Man. Proceedings of the American Philosophical Society, 139(2):135–148, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987261.

Colie:1963:SE

[Col63] Rosalie L. Colie. Spinoza in England, 1665–1730. Proceedings of the American Philosophical Society, 107(3):183–219, June 1963. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985580.

Coleman:1965:CNI

[Col65] William Coleman. Cell, nucleus, and inheritance: an historical study. Proceedings of the American Philosophical Society, 109(3): 124–158, June 1965. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986128.

Cole:1972:SCT

[Col72] Richard G. Cole. Sixteenth-Century travel books as a source of European attitudes toward non-white and non-Western culture. Proceedings of the American Philosophical Society, 116(1):59–67, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985706.

Compton:1956:WSL

[Com56] Arthur H. Compton. The world of science in the late Eighteenth Century and today. Proceedings of the American Philosophical Society, 100(4):296–303, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143755.

Conklin:1950:JPA

[Con50] Edwin G. Conklin. Joseph Priestley and the American Philosophical Society: His experiments on spontaneous generation. Proceedings of the American Philosophical Society, 94(2):127–131, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143212.

Conklin:1951:LCD

[Con51] Edwin G. Conklin. Letters of Charles Darwin and other scientists and philosophers to Sir Charles Lyell, Bart. Proceedings of the American Philosophical Society, 95(3):220–222, June 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143057.

Conroy:1956:CMK

[Con56] Hilary Conroy. Chōsen Mondai: The Korean problem in Meiji Japan. Proceedings of the American Philosophical Society, 100 (5):443–454, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143678.

Conant:1958:HSJ

[Con58] Kenneth John Conant. The holy sites at Jerusalem in the First and Fourth Centuries A.D. Proceedings of the American Philosophical Society, 102(1):14–24, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985302.

Conant:1960:CAR

[Con60] James Bryant Conant. Certain ambiguities in the relationship between school, college, and university. Proceedings of the American Philosophical Society, 104(6):562–564, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985531.

Conner:1967:MUO

[Con67] Paul Conner. “Maynard” unmasked: Oglethorpe and Sharp versus the press gangs. Proceedings of the American Philosophical Society, 111(4):199–211, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986007.

Condon:1969:ULL

[Con69] E. U. Condon. UFO’s I have loved and lost. Proceedings of the American Philosophical Society, 113(6):425–427, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985824.

Conroy:1970:SDA

[Con70] Hilary Conroy. The strange diplomacy of Admiral Nomura. Proceedings of the American Philosophical Society, 114(3):205–216, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985698.

Conacher:1992:BPA

[Con92a] J. B. Conacher. British policy in the Anglo–American enlistment crisis of 1855–1856. Proceedings of the American Philosophical Society, 136(4):533–576, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986762.

Constable:1992:MTA

[Con92b] Giles Constable. Marie-Thérèse D’Alverny (25 January 1903–26 April 1991). Proceedings of the American Philosophical Society, 136(3):418–422, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986915.

Connor:1993:IEA

[Con93] W. R. Connor. The Ionian Era of Athenian civic identity. Proceedings of the American Philosophical Society, 137(2):194–206, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986729.

Cone:1994:TAM

[Con94] Edward T. Cone. Thinking (about) music. Proceedings of the American Philosophical Society, 138(4):469–475, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986846.

Constable:1999:GDO

[Con99] Giles Constable. Georges Duby (7 October 1919–3 December 1996). Proceedings of the American Philosophical Society, 143(2):300–309, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181939.

Cooke:1952:IPN

[Coo52a] Charles A. Cooke. Iroquois personal names: Their classification. Proceedings of the American Philosophical Society, 96(4):427– 438, August 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143800.

Coon:1952:EHC

[Coo52b] Carleton S. Coon. Excavations in Hotu Cave, Iran, 1951, a preliminary report. Proceedings of the American Philosophical Society, 96(3):231–249, June 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143833.

Cooper:1976:HPB

[Coo76] Leon N. Cooper. How possible becomes actual in the quantum theory. Proceedings of the American Philosophical Society, 120 (1):37–45, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986350.

Cook:1984:CCR

[Coo84] Stephen A. Cook. Can computers routinely discover mathematical proofs? Proceedings of the American Philosophical Society, 128(1):40–43, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986492.

Cope:1950:CST

[Cop50a] Thomas D. Cope. “A clock sent thither by the Royal Society”. Proceedings of the American Philosophical Society, 94(3):260–268, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143560.

Copeland:1950:PBL

[Cop50b] Thomas W. Copeland. Problems of Burke’s letters. Proceedings of the American Philosophical Society, 94(4):357–360, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143520.

Cope:1951:SCB

[Cop51] Thomas D. Cope. Some contacts of Benjamin Franklin with Mason and Dixon and their work. Proceedings of the American Philosophical Society, 95(3):232–238, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143059.

Cope:1953:JQU

[Cop53] Thomas D. Cope. The Jersey quadrant used in Pennsylvania. Proceedings of the American Philosophical Society, 97(5):565–570, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149270.

Copland:1954:EPE

[Cop54] Douglas Copland. Economic policy and economic development in Australia from 1945: a study in economic administration. Proceedings of the American Philosophical Society, 98(2):91–114, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143640.

Cope:1955:SLS

[Cop55] Thomas D. Cope. Some local scholars who counselled the proprietors of Pennsylvania and their commissioners during the boundary surveys of the 1760’s. Proceedings of the American Philosophical Society, 99(4):268–276, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143706.

Cope:1956:WSI

[Cop56a] Thomas D. Cope. When the stars interrupted the running of a meridian line northward up the Delmarva Peninsula. Proceedings of the American Philosophical Society, 100(6):557–566, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143789.

Copland:1956:UDC

[Cop56b] Douglas Copland. The under-developed countries and the structure of the Western economy. Proceedings of the American Philosophical Society, 100(5):430–434, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143675.

Corner:1950:DBE

[Cor50] Betsy Copping Corner. Day book of an education: William Shippen’s student days in London (1759–1760) and his subsequent career. Proceedings of the American Philosophical Society, 94(2): 132–136, April 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143213.

Corner:1953:BFC

[Cor53] George W. Corner. Benjamin Franklin consults the doctors. Proceedings of the American Philosophical Society, 97(5):511–512, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149267.

Corwin:1956:FC

[Cor56] Edward Samuel Corwin. Franklin and the Constitution. Proceedings of the American Philosophical Society, 100(4):283–288, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143753.

Corner:1957:PPF

[Cor57] George W. Corner. Physician and pupils in a fourth-century painting. Proceedings of the American Philosophical Society, 101(3):245– 248, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985366.

Corner:1958:DFF

[Cor58] Betsy Copping Corner. Dr. Fothergill’s friendship with Benjamin Franklin. Proceedings of the American Philosophical Society, 102 (5):413–419, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985586.

Corner:1965:PML

[Cor65] George W. Corner. Penrose Memorial Lecture. Apprenticed to Aesculapius the American medical student, 1765–1965. Proceedings of the American Philosophical Society, 109(5):249–258, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986076.

Cornebise:1972:CGC

[Cor72] Alfred E. Cornebise. Cuno, Germany, and the coming of the Ruhr occupation: a study in German–West European relations. Proceedings of the American Philosophical Society, 116(6):502–531, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985763.

Corner:1976:BCE

[Cor76] George W. Corner. The black coalition: an experiment in racial cooperation Philadelphia, 1968. Proceedings of the American Philosophical Society, 120(3):178–186, June 1976. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986557.

Corcoran:1994:WPH

[Cor94] Irma Corcoran. William Penn and his purchasers: Problems in paradise. Proceedings of the American Philosophical Society, 138 (4):476–486, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986847.

Cornforth:1999:ART

[Cor99] John Cornforth. Alexander Robertus Todd (2 October 1907–10 January 1997). Proceedings of the American Philosophical Society, 143(3):524–528, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181966.

Cottrell:1971:CBI

[Cot71] Leonard S. Cottrell, Jr. Covert behavior in interpersonal interaction. Proceedings of the American Philosophical Society, 115(6): 462–469, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985837.

Cowherd:1960:HRE

[Cow60] Raymond G. Cowherd. The humanitarian reform of the English Poor Laws from 1782 to 1815. Proceedings of the American Philosophical Society, 104(3):328–342, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985249.

Cowley:1993:SPP

[Cow93] Stanley W. H. Cowley. Space plasma physics, present and future. Proceedings of the American Philosophical Society, 137(1):30–45, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986943.

Cox:1994:PFF

[Cox94] Archibald Cox. Paul A. Freund (16 February 1908–5 February 1992). Proceedings of the American Philosophical Society, 138(2): 322–326, June 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987127.

Cameron:1979:AC

[CP79] Ewan Cameron and Linus Pauling. Ascorbate and cancer. Proceedings of the American Philosophical Society, 123(2):117–123, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986233.

Cope:1952:AMW

[CR52] Thomas D. Cope and H. W. Robinson. The astronomical manuscripts which Charles Mason gave to Provost the Reverend John Ewing during October 1786. Proceedings of the American Philosophical Society, 96(4):417–423, August 1952. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143798.

Cope:1954:WMP

[CR54] Thomas D. Cope and H. W. Robinson. When the Maryland– Pennsylvania Boundary Survey changed from a political and legal struggle into a scientific and technological project. Proceedings of the American Philosophical Society, 98(6):432–441, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143865.

Cranz:1958:PGE

[Cra58] F. Edward Cranz. The prefaces to the Greek editions and Latin translations of Alexander of Aphrodisias, 1450 to 1575. Proceedings of the American Philosophical Society, 102(5):510–546, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985598.

Craig:1968:FSP

[Cra68] Gordon A. Craig. Friedrich Schiller and the police. Proceedings of the American Philosophical Society, 112(6):367–370, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985936.

Craven:1978:WBD

[Cra78] Wesley Frank Craven. On the writing of a biographical dictionary. Proceedings of the American Philosophical Society, 122(2):69–74, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986623.

Craig:1980:EP

[Cra80] Gordon A. Craig. The end of Prussia. Proceedings of the American Philosophical Society, 124(2):97–99, April 1980. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986204.

Crawford:1990:SNP

[Cra90] Elisabeth Crawford. The secrecy of Nobel Prize selections in the sciences and its effect on documentation and research. Proceedings of the American Philosophical Society, 134(4):408–419, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986896.

Craig:1991:JLG

[Cra91] Gordon A. Craig. The Jayne Lecture. German unification in historical perspective. Proceedings of the American Philosophical Society, 135(1):49–60, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987149.

Craig:1993:FGM

[Cra93] Gordon A. Craig. Felix Gilbert (21 May 1905–14 February 1991). Proceedings of the American Philosophical Society, 137(1):130–136, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986951.

Cremin:1985:PAE

[Cre85] Lawrence A. Cremin. The popularization of American education since World War II. Proceedings of the American Philosophical Society, 129(2):113–120, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986984.

Cristofalo:1988:RCS

[Cri88] Vincent J. Cristofalo. The regulation of cell senescence. Proceedings of the American Philosophical Society, 132(2):140–147, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986859.

Cronbach:1970:MTC

[Cro70] Lee J. Cronbach. Mental tests and the creation of opportunity. Proceedings of the American Philosophical Society, 114(6):480–487, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986147.

Crowther:1973:SOD

[Cro73] Simeon J. Crowther. The shipbuilding output of the Delaware Valley, 1722–1776. Proceedings of the American Philosophical Society, 117(2):90–104, April 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986059.

Crow:1981:CER

[Cro81] James F. Crow. Comparison of energy risks. Proceedings of the American Philosophical Society, 125(4):262–268, August 1981. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986327.

Collias:1952:FSP

[CS52a] Nicholas Collias and Charles Southwick. A field study of population density and social organization in howling monkeys. Proceedings of the American Philosophical Society, 96(2):143–156, April 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143720.

Curti:1952:FFD

[CS52b] Merle Curti and John Stalker. “The flowery flag devils”: The American image in China 1840–1900. Proceedings of the American Philosophical Society, 96(6):663–690, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143636.

Conklin:1953:SIS

[CS53] Harold C. Conklin and William C. Sturtevant. Seneca Indian singing tools at Coldspring Longhouse. Musical instruments of the Modern Iroquois. Proceedings of the American Philosophical Society, 97(3):262–290, June 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143749.

Corner:1954:DJF

[CS54] Betsy Copping Corner and Dorothea Waley Singer. Dr. John Fothergill, peacemaker. Proceedings of the American Philosophical Society, 98(1):11–22, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143665.

Cressman:1950:ARJ

[CSL50] L. S. Cressman, T. D. Stewart, and William S. Laughlin. Archaeological research in the John Day Region of North Central Oregon. Proceedings of the American Philosophical Society, 94 (4):369–390, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143523.

Causey:1969:RBM

[CT69] Nell B. Causey and Darwin L. Tiemann. A revision of the bioluminescent millipedes of the genus “Motyxia” (Xystodesmidae, Polydesmida). Proceedings of the American Philosophical Society, 113 (1):14–33, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986014.

Cunningham:1957:SPA

[Cun57] Maurice P. Cunningham. Some phonetic aspects of word order patterns in Latin. Proceedings of the American Philosophical Society, 101(5):481–505, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985202.

Curti:1961:TIA

[Cur61] Merle Curti. Tradition and innovation in American philanthropy. Proceedings of the American Philosophical Society, 105(2):146–156, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985626.

Curtis:1963:LER

[Cur63] Howard J. Curtis. The late effects of radiation. Proceedings of the American Philosophical Society, 107(1):5–10, February 1963. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985464.

Cushman:1950:WSC

[Cus50] C. Leslie Cushman. What schools can do to encourage thinking. Proceedings of the American Philosophical Society, 94(4):336– 338, August 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143517.

Cassels:1992:IMV

[CV92] J. W. S. Cassels and R. C. Vaughan. Ivan Matveevich Vinogradov (14 September 1891–20 March 1983). Proceedings of the American Philosophical Society, 136(4):608–612, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986767.

Chance:1971:DFS

[CWH^{+}71] Dr. Chance, Dr. Wiesner, Dr. Handler, Dr. Haskins, Dr. Levin, Dr. Goddard, Dr. Piore, Dr. Spilhaus, and Mr. Orlans. Discussion following symposium on new directions in science goals and support: Crisis or challenge? Proceedings of the American Philosophical Society, 115(1):13–21, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985919.

Curtis:1979:NAI

[CWMM79] J. E. Curtis, T. S. Wheeler, J. D. Muhly, and R. Maddin. Neo-Assyrian ironworking technology. Proceedings of the American Philosophical Society, 123(6):369–390, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986616.

Czarnik:1976:TME

[Cza76] Stanley A. Czarnik. The theory of the Mesolithic in European archaeology. Proceedings of the American Philosophical Society, 120(1):59–66, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986352.

Dainton:1995:JLS

[Dai95a] Lord Dainton. The Jayne Lecture. Science and the state: The uneasy but necessary alliance. Proceedings of the American Philosophical Society, 139(3):219–230, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987186.

Dainton:1995:LZM

[Dai95b] Lord Dainton. Lord Zuckerman (30 May 1904–1 April 1993). Proceedings of the American Philosophical Society, 139(2):212–217, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987270.

Dales:1966:RTP

[Dal66] George F. Dales. Recent trends in the pre- and protohistoric archaeology of South Asia. Proceedings of the American Philosophical Society, 110(2):130–139, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985745.

Daniels:1971:USE

[Dan71] Farrington Daniels. Utilization of solar energy-progress report. Proceedings of the American Philosophical Society, 115(6):490–501, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985842.

Darlington:1959:DZ

[Dar59a] Philip J. Darlington, Jr. Darwin and zoogeography. Proceedings of the American Philosophical Society, 103(2):307–319, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985155.

Darwin:1959:SEL

[Dar59b] Charles Galton Darwin. Some episodes in the life of Charles Darwin. Proceedings of the American Philosophical Society, 103(5): 609–615, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985419.

Darst:1979:WST

[Dar79] David H. Darst. Witchcraft in Spain: The testimony of Martín de Castañega’s treatise on superstition and witchcraft (1529). Proceedings of the American Philosophical Society, 123(5):298–322, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986592.

Darnton:1999:FFM

[Dar99] Robert Darnton. François Furet (27 March 1927–12 July 1997). Proceedings of the American Philosophical Society, 143(2):312–316, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181940.

Daugherty:1956:ALC

[Dau56] Richard D. Daugherty. Archaeology of the Lind Coulee Site, Washington. Proceedings of the American Philosophical Society, 100(3): 223–278, June 1956. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143773.

Davis:1951:PFS

[Dav51] Kingsley Davis. Population and the further spread of industrial society. Proceedings of the American Philosophical Society, 95(1): 8–19, February 1951. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143571.

Davidson:1952:NPP

[Dav52] Daniel Sutherland Davidson. Notes on the pictographs and petroglyphs of Western Australia and a discussion of their affinities with appearances elsewhere on the continent. Proceedings of the American Philosophical Society, 96(1):76–117, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143743.

Davidson:1953:PSA

[Dav53] Daniel Sutherland Davidson. The possible source and antiquity of the slate churingas of Western Australia. Proceedings of the American Philosophical Society, 97(2):194–213, April 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143894.

Davis:1960:EFP

[Dav60] Charles T. Davis. An early Florentine political theorist: Fra Remigio de’ Girolami. Proceedings of the American Philosophical Society, 104(6):662–676, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985539.

Davis:1961:CPP

[Dav61] Richard Beale Davis. Chesapeake pattern and pole-star: William Fitzhugh in his plantation world, 1676–1701. Proceedings of the American Philosophical Society, 105(6):525–529, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985162.

Davis:1963:ECB

[Dav63] Charles T. Davis. The early collection of books of S. Croce in Florence. Proceedings of the American Philosophical Society, 107 (5):399–414, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985645.

Davis:1968:SER

[Dav68] Hallowell Davis. Slow electrical responses of the human cortex. Proceedings of the American Philosophical Society, 112(3):150–156, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986160.

Davis:1974:PLR

[Dav74] Charles T. Davis. Ptolemy of Lucca and the Roman Republic. Proceedings of the American Philosophical Society, 118(1):30–50, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986435.

David:1981:SSE

[Dav81] Edward E. David, Jr. Synfuels and syn-ethics: Issues for a synfuels industry. Proceedings of the American Philosophical Society, 125 (4):277–285, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986330.

Davies:1985:AEE

[Dav85] Richard G. Davies. The attendance of the episcopate in English parliaments, 1376–1461. Proceedings of the American Philosophical Society, 129(1):30–81, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986979.

David:1986:CSS

[Dav86] E. E. David, Jr. China and science: Some thoughts. Proceedings of the American Philosophical Society, 130(1):21–36, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987090.

Davis:1991:WAA

[Dav91] Richard W. Davis. “We are all Americans now!” Anglo–American marriages in the Later Nineteenth Century. Proceedings of the American Philosophical Society, 135(2):140–199, June 1991. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987031.

Dawn:1960:AMA

[Daw60] C. Ernest Dawn. The Amir of Mecca Al-H .usayn Ibn-’Ali and the origin of the Arab revolt. Proceedings of the American Philosophical Society, 104(1):11–34, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985602.

deBellefeuille:1979:QCA

[dB79] Pierre de Bellefeuille. Québec comes of age. Proceedings of the American Philosophical Society, 123(5):274–275, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986588.

deSauvigny:1954:APF

[dBdS54] Guillaume de Bertier de Sauvigny. The American press and the fall of Napoleon in 1814. Proceedings of the American Philosophical Society, 98(5):337–376, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143913.

deMiranda:1992:GMF

[dCTdM92] Maria do Carmo Tavares de Miranda. Gilberto de Mello Freyre (March 15, 1900–July 18, 1987). Proceedings of the American Philosophical Society, 136(1):134–136, March 1992. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986804.

deDuve:1998:COE

[dD98] Christian de Duve. Constraints on the origin and evolution of life. Proceedings of the American Philosophical Society, 142(4): 525–532, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152279.

DElia:1966:DBR

[D’E66] Donald J. D’Elia. Dr. Benjamin Rush and the American Medical Revolution. Proceedings of the American Philosophical Society, 110 (4):227–234, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985685.

DElia:1970:BRD

[D’E70] Donald J. D’Elia. Benjamin Rush, David Hartley, and the revolutionary uses of psychology. Proceedings of the American Philosophical Society, 114(2):109–118, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986029.

DElia:1973:JRL

[D’E73] Donald J. D’Elia. Jefferson, rush, and the limits of philosophical friendship. Proceedings of the American Philosophical Society, 117 (5):333–343, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986604.

Dean:1967:CC

[Dea67] Patrick Dean. The challenge of change. Proceedings of the American Philosophical Society, 111(5):254–258, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986042.

Demerec:1954:WMG

[Dem54] Milislav Demerec. What makes genes mutate? Proceedings of the American Philosophical Society, 98(5):318–322, October 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143910.

DeMolen:1984:LWC

[DeM84] Richard L. DeMolen. The library of William Camden. Proceedings of the American Philosophical Society, 128(4):326–409, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987100.

Dempsey:1994:MMR

[Dem94] Ellen Dempsey. Marcus Morton Rhoades (24 July 1903–30 December 1991). Proceedings of the American Philosophical Society, 138 (4):560–567, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986855.

Detwiler:1955:ESA

[Det55] S. R. Detwiler. The eye and its structural adaptations. Proceedings of the American Philosophical Society, 99(4):224–238, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143701.

Dethier:1981:FRM

[Det81] Vincent G. Dethier. Fly, rat, and man: The continuing quest for an understanding of behavior. Proceedings of the American Philosophical Society, 125(6):460–466, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986301.

Dethier:1987:SFP

[Det87] Vincent G. Dethier. Sniff, flick, and pulse: an appreciation of interruption. Proceedings of the American Philosophical Society, 131(2):159–176, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986789.

DeVoto:1955:IRE

[DeV55] Bernard DeVoto. An inference regarding the Expedition of Lewis and Clark. Proceedings of the American Philosophical Society, 99 (4):185–194, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143696.

Devos:1969:PRC

[Dev69] George Devos. Problems and research in comparative human behavior. Proceedings of the American Philosophical Society, 113 (5):350–366, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985972.

Dewhurst:1956:APC

[Dew56] J. Frederic Dewhurst. American productivity: Causes and effects. Proceedings of the American Philosophical Society, 100(5): 435–438, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143676.

Dews:1964:H

[Dew64] P. B. Dews. Humors. Proceedings of the American Philosophical Society, 108(6):473–477, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985865.

Dudley:1973:JRF

[DG73] Susan Dudley and David R. Goddard. Joseph T. Rothrock and forest conservation. Proceedings of the American Philosophical Society, 117(1):37–50, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985946.

Diamond:1994:ECP

[Dia94] Jared Diamond. Ecological collapses of past civilizations. Proceedings of the American Philosophical Society, 138(3):363–370, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986741.

Dicke:1981:OFE

[Dic81] Robert H. Dicke. The observational foundation of Einstein’s theory of gravitation the gravitational acceleration. Proceedings of the American Philosophical Society, 125(1):16–19, February 1981. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986182.

Dickerson:1998:IID

[Dic98] Janet Smith Dickerson. The impact of increasing diversity in higher education or, the teacher’s gift: Permission to go home. Proceedings of the American Philosophical Society, 142(2):207–217, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152299.

Dillon:1967:SLG

[Dil67] Richard H. Dillon. Stephen Long’s Great American Desert. Proceedings of the American Philosophical Society, 111(2):93–108, April 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986036.

Dilworth:1968:COU

[Dil68] Richardson Dilworth. The chaos of our urban areas. Proceedings of the American Philosophical Society, 112(6):355–357, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985933.

Dingle:1955:CMB

[Din55] Herbert Dingle. Comments on Margenau, Barber, Cohen. Proceedings of the American Philosophical Society, 99(5):348–349, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143935. See

[Mar55, Bar55, Coh55].

Duveen:1954:TBI

[DK54] Denis I. Duveen and Herbert S. Klickstein. Two broadsides illustrating Antoine Laurent Lavoisier’s “Chemical Characters Adapted to the New Nomenclature”. Proceedings of the American Philosophical Society, 98(6):466–468, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143869.

Katzenbach:1987:SPT

[dK87] Nicholas deB. Katzenbach. Separation of powers: Then and now. Proceedings of the American Philosophical Society, 131(3):243–250, September 1987. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987019.

Katzenbach:1995:TJW

[dK95] Nicholas deB. Katzenbach. Thomas J. Watson, Jr. (8 January 1914–31 December 1993). Proceedings of the American Philosophical Society, 139(2):198–201, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987268.

Davis:1957:ACE

[DL57] Curtis Carroll Davis and L. Littlepage. An American courtier in Europe: Lewis Littlepage’s “Private Political Memoir” (Hamburg, 1795). Proceedings of the American Philosophical Society, 101(3): 255–269, June 1957. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985368.

Landon:1970:BAC

[dLL70] Michael de L. Landon. The Bristol Artillery Company and the Tory triumph in Bristol 1679–1684. Proceedings of the American Philosophical Society, 114(2):155–161, April 1970. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986032.

Dvoichenko-Markov:1950:APS

[DM50] Eufrosina Dvoichenko-Markov. The American Philosophical Society and early Russian–American relations. Proceedings of the American Philosophical Society, 94(6):549–610, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143595.

Dvoichenko-Markov:1952:BFL

[DM52] Eufrosina Dvoichenko-Markov. Benjamin Franklin and Leo Tolstoy. Proceedings of the American Philosophical Society, 96(2): 119–128, April 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143718.

Dvoichenko-Markov:1953:RTE

[DM53a] Eufrosina Dvoichenko-Markov. A Russian traveler to Eighteenth-Century America. Proceedings of the American Philosophical Society, 97(4):350–355, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3149289.

Dvoichenko-Markov:1953:WPP

[DM53b] Eufrosina Dvoichenko-Markov. William Penn and Peter the Great. Proceedings of the American Philosophical Society, 97(1):12–25, February 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143726.

Dvoichenko-Markov:1955:BFC

[DM55] Eufrosina Dvoichenko-Markov. Benjamin Franklin and Count M. A. Benyowski. Proceedings of the American Philosophical Society, 99(6):405–417, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143923.

deMurville:1956:PML

[dM56] Maurice Couve de Murville. Penrose Memorial Lecture. Benjamin Franklin: The diplomat and journalist. Proceedings of the American Philosophical Society, 100(4):316–325, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143757.

Dvoichenko-Markov:1965:RMA

[DM65] Eufrosina Dvoichenko-Markov. The Russian members of the American Academy of Arts and Sciences. Proceedings of the American Philosophical Society, 109(1):53–56, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985778.

Dobzhansky:1959:VE

[Dob59] Theodosius Dobzhansky. Variation and evolution. Proceedings of the American Philosophical Society, 103(2):252–263, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985152.

Dobzhansky:1965:MDE

[Dob65] Theodosius Dobzhansky. Mendelism, Darwinism, and evolutionism. Proceedings of the American Philosophical Society, 109(4): 205–215, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985879.

Dodge:1963:EAC

[Dod63] Ernest S. Dodge. Early American contacts in Polynesia and Fiji. Proceedings of the American Philosophical Society, 107(2):102–106, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985437.

Donagan:1976:CPR

[Don76] Barbara Donagan. The clerical patronage of Robert Rich, Second Earl of Warwick, 1619–1642. Proceedings of the American Philosophical Society, 120(5):388–419, October 1976. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986269.

Dorson:1957:CFJ

[Dor57] Richard M. Dorson. Collecting folklore in Jonesport, Maine. Proceedings of the American Philosophical Society, 101(3):270–289, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985369.

Douglass:1959:FAS

[Dou59] Elisha P. Douglass. Fisher Ames, spokesman for New England federalism. Proceedings of the American Philosophical Society, 103 (5):693–715, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985426.

Downey:1958:CAE

[Dow58] Glanville Downey. The claim of Antioch to ecclesiastical jurisdiction over Cyprus. Proceedings of the American Philosophical Society, 102(3):224–228, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985575.

Downey:1959:LOP

[Dow59] Glanville Downey. Libanius’ oration in Praise of Antioch (Oration XI). Proceedings of the American Philosophical Society, 103 (5):652–686, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985424.

Draper:1960:NCS

[Dra60] Charles S. Draper. Navigation — from canoes to spaceships. Proceedings of the American Philosophical Society, 104(2):113–123, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985654.

Drell:1994:ADS

[Dre94] Sidney D. Drell. Andrei Dmitrievich Sakharov (21 May 1921–14 December 1989). Proceedings of the American Philosophical Society, 138(1):176–182, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986713.

Drinker:1952:AM

[Dri52] Henry S. Drinker. The amateur in music. Proceedings of the American Philosophical Society, 96(5):573–577, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143658.

Drinker:1953:OPS

[Dri53] Henry S. Drinker. The origin and persistence of standards of professional conduct in the practice of law. Proceedings of the American Philosophical Society, 97(6):652–657, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3149280.

Dripps:1964:GAC

[Dri64] Robert D. Dripps. General anesthesia and the circulation. Proceedings of the American Philosophical Society, 108(1):15–18, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985992.

Dryden:1953:FAS

[Dry53] Hugh L. Dryden. Frontiers of aeronautical science and technology. Proceedings of the American Philosophical Society, 97(1):56– 60, February 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143731.

Dryden:1954:HCA

[Dry54] Hugh L. Dryden. A half century of aeronautical research. Proceedings of the American Philosophical Society, 98(2):115–120, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143641.

Dryden:1960:PML

[Dry60] Hugh L. Dryden. Penrose Memorial Lecture. Prospects for space travel. Proceedings of the American Philosophical Society, 104(5): 474–484, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985232.

Deardorff:1956:NCJ

[DS56] Merle H. Deardorff and George S. Snyderman. A Nineteenth-Century journal of a visit to the Indians of New York. Proceedings of the American Philosophical Society, 100(6):582–612, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143791.

Dodge:1950:SLG

[DSR50] B. O. Dodge, J. R. Singleton, and Anita Rolnick. Studies on lethal E gene in Neurospora tetrasperma, including chromosome counts also in races of N. sitophila. Proceedings of the American Philosophical Society, 94(1):38–52, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143250.

deTerra:1958:SDAa

[dT58a] Helmut de Terra. Studies of the documentation of Alexander von Humboldt. Proceedings of the American Philosophical Society, 102 (2):136–141, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985277.

deTerra:1958:SDAb

[dT58b] Helmut de Terra. Studies of the documentation of Alexander von Humboldt: The Philadelphia abstract of Humboldt’s American travels: Humboldt portraits and sculpture in the United States. Proceedings of the American Philosophical Society, 102(6):560–589, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985567.

deTerra:1959:AHC

[dT59] Helmut de Terra. Alexander von Humboldt’s correspondence with Jefferson, Madison, and Gallatin. Proceedings of the American Philosophical Society, 103(6):783–806, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985393.

deTerra:1960:MCA

[dT60] Helmut de Terra. Motives and consequences of Alexander von Humboldt’s visit to the United States (1804). Proceedings of the American Philosophical Society, 104(3):314–316, June 1960. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985246.

Duane:1958:PIE

[Dua58] Morris Duane. The planning and improvement of educational facilities. Proceedings of the American Philosophical Society, 102 (4):328–332, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985218.

Duane:1966:SRD

[Dua66] Morris Duane. Some recent developments in educational facilities. Proceedings of the American Philosophical Society, 110(3):156–159, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985983.

Dubois:1960:ACO

[Dub60] Eugene F. Dubois. An attempt to classify occupations in ten task groups according to physical exertion or according to the amount of physical exertion demanded. Proceedings of the American Philosophical Society, 104(1):111–112, February 1960. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985607.

Dubridge:1962:PS

[Dub62] Lee A. Dubridge. The place of space. Proceedings of the American Philosophical Society, 106(5):461–466, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985490.

Dubos:1963:PML

[Dub63] René Dubos. Penrose Memorial Lecture. Logic and choices in science. Proceedings of the American Philosophical Society, 107 (5):365–374, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985643.

Dubridge:1971:RGE

[Dub71] Lee A. Dubridge. The role of government enterprise. Proceedings of the American Philosophical Society, 115(1):10–12, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985918.

Dubridge:1977:TFY

[Dub77] Lee A. Dubridge. Twenty-five years of the national science foundation. Proceedings of the American Philosophical Society, 121(3): 191–194, June 1977. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986409.

Dubin:1979:CEI

[Dub79] Martin David Dubin. The Carnegie Endowment for International Peace and the advocacy of a League of Nations, 1914–1918. Proceedings of the American Philosophical Society, 123(6):344–368, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986615.

Dulbecco:1995:HMT

[Dul95] Renato Dulbecco. Howard M. Temin (10 December 1934–9 February 1994). Proceedings of the American Philosophical Society, 139 (4):452–462, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987243.

Dulbecco:1997:GC

[Dul97] Renato Dulbecco. Genes and cancer. Proceedings of the American Philosophical Society, 141(4):425–435, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987219.

Dumont:1951:SKA

[Dum51] Paul-Emile Dumont. The special kinds of agnicayana (or special methods of building the fire-altar) according to the Kat.has in the Taittirīya-Brāhman.a: The tenth, eleventh, and twelfth Prapāt.hakas of the third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 95 (6):628–675, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143510.

Dumont:1954:INO

[Dum54] Paul-Emile Dumont. The Is.t.is to the Naks.atras (or oblations to the lunar mansions) in the Taittirīya-Brāhman.a: The first Prapāt.haka of the third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 98(3):204–223, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143846.

Dumont:1957:FMN

[Dum57] Paul-Emile Dumont. The full-moon and new-moon sacrifices in the taittirīya-brāhman.a (first part). Proceedings of the American Philosophical Society, 101(2):216–243, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985514.

Dumont:1959:FMN

[Dum59] Paul-Emile Dumont. The full-moon and new-moon sacrifices in the taittirīya-brāhman.a (second part): The third prapāt.haka of the third kān.d.a of the taittirīya-brāhman.a with translation. Proceedings of the American Philosophical Society, 103(4):584–608, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985561.

Dumont:1960:FMN

[Dum60] Paul-Emile Dumont. The full-moon and new-moon sacrifices in the Taittirīya-Brāhman.a (third part) — the part of the Hotar in the full-moon and new-moon sacrifices — the fifth Prapāt.haka of the Third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 104(1):1–10, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985601.

Dumont:1961:FMN

[Dum61] Paul-Emile Dumont. The full-moon and new-moon sacrifices in the taittiriya-brāhman.a (fourth part): The anuvākas 1–6 and 11 of the seventh prapāt.haka of the third kān.d.a of the taittirīya-brāhman.a with translation. Proceedings of the American Philosophical Society, 105(1):11–36, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985353.

Dumont:1962:AST

[Dum62] Paul-Emile Dumont. The animal sacrifice in the Taittirīya-Brāhman.a the part of the Hotar and the part of the Maitrāvarun.a in the animal sacrifice: The sixth Prapāt.haka of the third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 106(3):246–263, June 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985179.

Dumont:1963:HST

[Dum63a] Paul-Emile Dumont. The human sacrifice in the taittirīyabrāhman.a: The fourth prapāt.haka of the third kān.d.a of the taittirīya-brāhman.a with translation. Proceedings of the American Philosophical Society, 107(2):177–182, April 1963. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985448.

Dumont:1963:TYB

[Dum63b] Paul-Emile Dumont. Taittirīya-brāhman.a 3.7.7-10 and 3.7.12-14: Seven anuvākas of the seventh prapāt.haka of the third kān.d.a of the taittirīya-brāhman.a with translation. Proceedings of the American Philosophical Society, 107(5):446–460, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985650.

Dumont:1964:AFG

[Dum64] Paul-Emile Dumont. The agnihotra (or fire-god oblation) in the taittirīya-brāhman.a: The first prapāt.haka of the second kān.d.a of the taittirīya-brāhman.a with translation. Proceedings of the American Philosophical Society, 108(4):337–353, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985912.

Dumont:1965:KST

[Dum65] Paul-Emile Dumont. The Kaukilī-Sautrāman.ī in the Taittirīya-Brāhman.a: The sixth Prapāt.haka of the second Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 109(6):309–341, December 1965. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986135.

Dumont:1969:KAS

[Dum69] Paul-Emile Dumont. The Kāmya animal sacrifices in the Taittirīya-Brāhman.a: The eighth Prapāt.haka of the third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 113(1):34–66, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986015.

Dunn:1965:MHW

[Dun65] L. C. Dunn. Mendel, his work and his place in history. Proceedings of the American Philosophical Society, 109(4):189–198, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985877.

Dunn:1973:XOG

[Dun73] L. C. Dunn. Xenia and the origin of genetics. Proceedings of the American Philosophical Society, 117(2):105–111, April 1973. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986060.

Dunn:1983:PWP

[Dun83a] Mary Maples Dunn. The personality of William Penn. Proceedings of the American Philosophical Society, 127(5):316–321, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986500.

Dunn:1983:WPS

[Dun83b] Richard S. Dunn. William Penn and the selling of Pennsylvania, 1681–1685. Proceedings of the American Philosophical Society, 127 (5):322–329, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986501.

Dunitz:1997:LCP

[Dun97] Jack D. Dunitz. Linus Carl Pauling (28 February 1901–19 August 1994). Proceedings of the American Philosophical Society, 141(4): 494–501, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987225.

Dupree:1952:PAH

[Dup52] Louis B. Dupree. The Pleistocene artifacts of Hotu Cave, Iran. Proceedings of the American Philosophical Society, 96(3):250–257, June 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143834.

Dupree:1957:FNA

[Dup57] A. Hunter Dupree. The founding of the National Academy of Sciences — a reinterpretation. Proceedings of the American Philosophical Society, 101(5):434–440, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985196.

Duroselle:1962:ACW

[Dur62] J. B. Duroselle. The American Civil War — an episode in the history of civil wars. Proceedings of the American Philosophical Society, 106(1):10–32, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985206.

Durand:1967:MEW

[Dur67] John D. Durand. The modern expansion of world population. Proceedings of the American Philosophical Society, 111(3):136–159, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985711.

Durey:1987:WCM

[Dur87] Michael Durey. William Cobbett, military corruption and London radicalism in the early 1790s. Proceedings of the American Philosophical Society, 131(4):348–366, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987136.

Davis:1965:GGF

[DW65] Emma Lou Davis and Sylvia Winslow. Giant ground figures of the prehistoric deserts. Proceedings of the American Philosophical Society, 109(1):8–21, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985774.

Dye:1976:EAM

[Dye76] Ira Dye. Early American merchant seafarers. Proceedings of the American Philosophical Society, 120(5):331–360, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986266.

Dye:1989:TEA

[Dye89] Ira Dye. The tattoos of early American seafarers, 1796–1818. Proceedings of the American Philosophical Society, 133(4):520–554, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986875.

Dyson:1978:PFM

[Dys78] Freeman J. Dyson. Pilgrim fathers, Mormon pioneers, and space colonists: an economic comparison. Proceedings of the American Philosophical Society, 122(2):63–68, April 1978. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986622.

Dyson:1996:TRA

[Dys96] Freeman J. Dyson. Two revolutions in astronomy. Proceedings of the American Philosophical Society, 140(1):1–9, March 1996. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987272.

Dyson:1999:ASJ

[Dys99] Freeman J. Dyson. Abdus Salam (29 January 1926–21 November 1996). Proceedings of the American Philosophical Society, 143(2): 346–350, June 1999. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3181945.

Eagleson:1998:HCP

[Eag98] William B. Eagleson, Jr. Howard Charles Petersen (7 May 1910–28 December 1995). Proceedings of the American Philosophical Society, 142(2):294–299, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152307.

Easby:1965:PHM

[Eas65] Dudley T. Easby, Jr. Pre-Hispanic metallurgy and metalworking in the New World. Proceedings of the American Philosophical Society, 109(2):89–98, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985784.

Eastwood:1982:KHS

[Eas82] Bruce Stansfield Eastwood. Kepler as historian of science: Precursors of Copernican heliocentrism according to “De Revolutionibus,” I, 10. Proceedings of the American Philosophical Society, 126 (5):367–394, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986213.

Edsall:1979:HRS

[EB79] John T. Edsall and David Bearman. Historical records of scientific activity: The survey of sources for the history of biochemistry and molecular biology. Proceedings of the American Philosophical Society, 123(5):279–292, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986590.

Enroth-Cugell:1994:RGO

[EC94] Christina Enroth-Cugell. Ragnar Granit (30 October 1900–12 March 1991). Proceedings of the American Philosophical Society, 138(2):328–333, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987128.

Echeverria:1954:SHW

[Ech54] Durand Echeverria. “The Sale of the Hessians.” was Benjamin Franklin the author? Proceedings of the American Philosophical Society, 98(6):427–431, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143864.

Edelman:1975:MEC

[Ede75] Marian Wright Edelman. On mounting effective child advocacy. Proceedings of the American Philosophical Society, 119(6):470–478, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986379.

Edelman:1997:RFP

[Ede97] Marian Wright Edelman. The Richard A. F. Penrose, Jr., Memorial Lecture. The plight of our children: Invest now or pay later. Proceedings of the American Philosophical Society, 141(1):1–12, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987246.

Edsall:1967:OPM

[Eds67] John T. Edsall. The organization of protein molecules for regulatory processes in biology. Proceedings of the American Philosophical Society, 111(1):57–67, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985754.

Edsall:1985:ITP

[Eds85] John T. Edsall. Isidor Traube: Physical chemist, biochemist, colloid chemist and controversialist. Proceedings of the American Philosophical Society, 129(4):371–406, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986935.

Edsall:1991:ALM

[Eds91] John T. Edsall. Alexander Ludwig von Muralt (August 19, 1903– May 28, 1990). Proceedings of the American Philosophical Society, 135(4):614–621, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986822.

Edwards:1961:FRG

[Edw61] Robert L. Edwards. The fishes of Richmond Gulf, Ungava, Canada. Proceedings of the American Philosophical Society, 105(2):196–205, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985630.

Eiseley:1959:CDE

[EG59] Loren C. Eiseley and A. Grote. Charles Darwin, Edward Blyth, and the theory of natural selection. Proceedings of the American Philosophical Society, 103(1):94–158, February 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985383.

Egan:1971:PGD

[Ega71] Clifford L. Egan. Pressure groups, the Department of State, and the abrogation of the Russian–American Treaty of 1832. Proceedings of the American Philosophical Society, 115(4):328–334, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986094.

Egan:1978:LSL

[Ega78] James Egan. “This is a lamentation and shall be for a lamentation”: Nathaniel Ward and the rhetoric of the Jeremiad. Proceedings of the American Philosophical Society, 122(6):400–410, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986455.

Eggan:1965:LHM

[Egg65] Fred Eggan. Lewis H. Morgan and the future of the American Indian. Proceedings of the American Philosophical Society, 109 (5):272–276, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986079.

Eggan:1969:ISF

[Egg69] Fred R. Eggan. Introduction to symposium on frontiers of cultural anthropology. Proceedings of the American Philosophical Society, 113(5):325–328, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985968.

Ehrlich:1990:ILC

[Ehr90] George Ehrlich. The illustrations in the Lewis and Clark journals: One artist or two? Proceedings of the American Philosophical Society, 134(2):95–110, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986841.

Einarsson:1957:IL

[Ein57] Stefán Einarsson. Iceland and its literature. Proceedings of the American Philosophical Society, 101(4):375–378, August 1957. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985404.

Eisenhart:1950:WME

[Eis50] Luther P. Eisenhart. Walter Minto and the Earl of Buchan. Proceedings of the American Philosophical Society, 94(3):282–294, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143564.

Eiseley:1954:PDC

[Eis54a] Loren C. Eiseley. The program on the Darwin Collection in the library. Proceedings of the American Philosophical Society, 98(6): 449–452, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143867.

Eiseley:1954:RFM

[Eis54b] Loren C. Eiseley. The reception of the first missing links. Proceedings of the American Philosophical Society, 98(6):453–465, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143868.

Eisele:1957:CPS

[Eis57] Carolyn Eisele. The Charles S. Peirce–Simon Newcomb correspondence. Proceedings of the American Philosophical Society, 101(5): 409–433, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985195.

Eisele:1963:CPP

[Eis63] Carolyn Eisele. Charles S. Peirce and the problem of mapprojection. Proceedings of the American Philosophical Society, 107 (4):299–307, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985673.

Eisenstadt:1984:HDC

[Eis84] Shmuel N. Eisenstadt. Heterodoxies and dynamics of civilizations. Proceedings of the American Philosophical Society, 128(2):104–113, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986224.

Eisner:1994:CPG

[Eis94] Thomas Eisner. Chemical prospecting: a global imperative. Proceedings of the American Philosophical Society, 138(3):385–393, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986744.

Eisenstadt:1997:ESJ

[Eis97] S. N. Eisenstadt. Edward Shils (1 July 1910–23 January 1995). Proceedings of the American Philosophical Society, 141(3):366–373, September 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987206.

Eisenstadt:1998:EGD

[Eis98] S. N. Eisenstadt. Ernest Gellner (9 December 1925–5 November 1995). Proceedings of the American Philosophical Society, 142(1): 134–137, March 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152270.

Euler:1956:HST

[EJ56] Robert C. Euler and Volney H. Jones. Hermetic sealing as a technique of food preservation among the Indians of the American Southwest. Proceedings of the American Philosophical Society, 100 (1):87–99, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143809.

Ekspong:1999:DNL

[Eks99] Gösta Ekspong. The dual nature of light, as reflected in the Nobel Archive. Proceedings of the American Philosophical Society, 143(1):42–49, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181972.

Ekvall:1974:TNP

[Ekv74] Robert B. Ekvall. Tibetan nomadic pastoralists: Environments, personality, and ethos. Proceedings of the American Philosophical Society, 118(6):519–537, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986402.

Ekvall:1980:HPO

[Ekv80] Robert B. Ekvall. The high-pasturage ones of Tibet also grow old. Proceedings of the American Philosophical Society, 124(6):429–437, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986244.

Elkins:1963:CCN

[Elk63] Stanley M. Elkins. Culture contacts and Negro slavery. Proceedings of the American Philosophical Society, 107(2):107–109, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985438.

Elliott:1983:BRV

[Ell83] Frank A. Elliott. Biological roots of violence. Proceedings of the American Philosophical Society, 127(2):84–94, April 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986191.

Elsasser:1968:MCD

[Els68] Walter M. Elsasser. The mechanics of continental drift. Proceedings of the American Philosophical Society, 112(5):344–353, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986055.

Emeneau:1953:CBI

[Eme53] Murray B. Emeneau. The composite bow in India. Proceedings of the American Philosophical Society, 97(1):77–87, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143734.

Emeneau:1954:LPI

[Eme54] Murray B. Emeneau. Linguistic prehistory of India. Proceedings of the American Philosophical Society, 98(4):282–292, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143883.

Emeneau:1962:BSB

[Eme62] M. B. Emeneau. Bilingualism and structural borrowing. Proceedings of the American Philosophical Society, 106(5):430–442, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985488.

Emlen:1962:DG

[Eml62] John T. Emlen. The display of the gorilla. Proceedings of the American Philosophical Society, 106(6):516–519, December 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985257.

Enders:1952:RMM

[End52] Robert K. Enders. Reproduction in the mink (Mustela Vison). Proceedings of the American Philosophical Society, 96(6):691–755, December 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143637.

Enders:1957:CVC

[End57] John F. Enders. Comments on viral cytopathogenicity. Proceedings of the American Philosophical Society, 101(4):325–329, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985398.

English:1968:OSQ

[Eng68] Paul Ward English. The origin and spread of qanats in the Old World. Proceedings of the American Philosophical Society, 112(3): 170–181, June 1968. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986162.

Enright:1975:MIE

[Enr75] J. T. Enright. The Moon illusion examined from a new point of view. Proceedings of the American Philosophical Society, 119(2): 87–107, April 1975. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986629.

Elizabeth:1999:DVW

[EoL99] Elizabeth and Countess of Longford. Dame Veronica Wedgwood (20 July 1910–9 March 1997). Proceedings of the American Philosophical Society, 143(2):352–356, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181946.

Edelman:1990:IPC

[ER90] Gerald M. Edelman and George N. Reeke, Jr. Is it possible to construct a perception machine? Proceedings of the American Philosophical Society, 134(1):36–73, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143904.

Erickson:1967:DBP

[Eri67] Norma N. Erickson. A dispute between a priest and a knight. Proceedings of the American Philosophical Society, 111(5):288–309, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986046.

Erickson:1956:EGR

[ES56a] Ralph O. Erickson and Katharine B. Sax. Elemental growth rate of the primary root of Zea Mays. Proceedings of the American Philosophical Society, 100(5):487–498, October 1956. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143682.

Erickson:1956:RCD

[ES56b] Ralph O. Erickson and Katharine B. Sax. Rates of cell division and cell elongation in the growth of the primary root of Zea Mays. Proceedings of the American Philosophical Society, 100(5): 499–514, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143683.

Esau:1967:MVB

[Esa67] Katherine Esau. Minor veins in beta leaves: Structure related to function. Proceedings of the American Philosophical Society, 111 (4):219–233, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986009.

Espenshade:1992:PIU

[Esp92] Thomas J. Espenshade. Policy influences on undocumented migration to the United States. Proceedings of the American Philosophical Society, 136(2):188–207, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987168.

Ettinghausen:1979:THV

[Ett79] Richard Ettinghausen. The taming of the horror vacui in Islamic art. Proceedings of the American Philosophical Society, 123(1):15– 28, February 1979. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986468.

Evans:1969:RSS

[Eva69] John V. Evans. Radar surveys of the solar system. Proceedings of the American Philosophical Society, 113(3):203–223, June 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985736.

Ewan:1952:FPH

[Ewa52] Joseph Ewan. Frederick Pursh, 1774–1820, and his botanical associates. Proceedings of the American Philosophical Society, 96 (5):599–628, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143661.

Ewan:1959:FBH

[Ewa59] Joseph Ewan. Fougeroux de Bondaroy (1732–1789) and his projected revision of Duhamel du Monceau’s “traité” (1755) on trees and shrubs: I. An analytical guide to persons, gardens, and works mentioned in the manuscripts. Proceedings of the American Philosophical Society, 103(6):807–818, December 1959. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985394.

Ewan:1983:CNO

[Ewa83] Joseph Ewan. From Calcutta and New Orleans, or, tales from Barton’s greenhouse. Proceedings of the American Philosophical Society, 127(3):125–134, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986382.

Easterlin:1978:CIP

[EWW78] Richard A. Easterlin, Michael L. Wachter, and Susan M. Wachter. The changing impact of population swings on the American economy. Proceedings of the American Philosophical Society, 122(3): 119–130, June 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986545.

Fabricant:1958:PSB

[Fab58] Solomon Fabricant. The problem of severe business contractions. Proceedings of the American Philosophical Society, 102(4):400–403, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985224.

Fabricant:1961:EVP

[Fab61] Solomon Fabricant. An economist’s view of philanthropy. Proceedings of the American Philosophical Society, 105(2):162–166, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985628.

Fahey:1971:SRG

[Fah71] David M. Fahey. Samuel Rawson Gardiner looks at history, Home Rule, and Empire: Some Bodleian letters, 1886–1899. Proceedings of the American Philosophical Society, 115(2):125–130, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985852.

Fainsod:1964:SYP

[Fai64] Merle Fainsod. Soviet youth and the problem of the generations. Proceedings of the American Philosophical Society, 108(5):429–436, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985817.

Fairbank:1992:ERO

[Fai92] John K. Fairbank. Edwin O. Reischauer (15 October 1910–1 September 1990). Proceedings of the American Philosophical Society, 136(3):450–454, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986920.

Fajn:1974:MMR

[Faj74] Max Fajn. Marc-Michel Rey: Boekhandelaar op de Bloemmark (Amsterdam). Proceedings of the American Philosophical Society, 118(3):260–268, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986444.

Farrar:1982:VNB

[Far82] Marjorie M. Farrar. Victorious nationalism beleaguered: Alexandre Millerand as French Premier in 1920. Proceedings of the American Philosophical Society, 126(6):481–519, December 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986254.

Fowler:1992:CDA

[FC92] William A. Fowler and Eugene W. Cowan. Carl David Anderson (September 3, 1905–January 11, 1991). Proceedings of the American Philosophical Society, 136(2):274–278, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987172.

Fedoroff:1994:BMJ

[Fed94] Nina Fedoroff. Barbara McClintock (16 June 1902–2 September 1992). Proceedings of the American Philosophical Society, 138(3): 432–445, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986751.

Feldman:1987:RAL

[Fel87] Franklin Feldman. Reflections on art and the law: Old concepts, new values. Proceedings of the American Philosophical Society, 131(2):141–147, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986787.

Feldman:1995:JPJ

[Fel95] Louis H. Feldman. Josephus’ Portrait of Jehoiachin. Proceedings of the American Philosophical Society, 139(1):11–31, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986717.

Fenton:1951:ISM

[Fen51] William N. Fenton. Iroquois studies at the Mid-Century. Proceedings of the American Philosophical Society, 95(3):296–310, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143068.

Fenton:1953:CMM

[Fen53] William Nelson Fenton. A calendar of manuscript materials relating to the history of the Six Nations or Iroquois Indians in depositories outside of Philadelphia, 1750–1850. Proceedings of the American Philosophical Society, 97(5):578–595, October 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3149272.

Fenton:1956:TGC

[Fen56] William N. Fenton. Toward the gradual civilization of the Indian natives: The missionary and linguistic work of Asher Wright (1803–1875) among the Senecas of Western New York. Proceedings of the American Philosophical Society, 100(6):567–581, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143790.

Fenn:1970:LUH

[Fen70] Wallace O. Fenn. Life under high pressures. Proceedings of the American Philosophical Society, 114(3):191–197, June 1970. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985696.

Fenton:1971:NYS

[Fen71] William N. Fenton. The New York State Wampum Collection: The case for the integrity of cultural treasures. Proceedings of the American Philosophical Society, 115(6):437–461, December 1971. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985836.

Fenton:1992:FRE

[Fen92] William N. Fenton. Fred Russell Eggan (12 September 1906–7 May 1991). Proceedings of the American Philosophical Society, 136(3): 432–435, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986917.

Fishman:1992:EAL

[FGWP92] A. P. Fishman, R. J. Galante, A. Winokur, and A. I. Pack. Estivation in the African lungfish. Proceedings of the American Philosophical Society, 136(1):61–72, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986798.

Finkelstein:1963:MH

[Fin63] J. J. Finkelstein. Mesopotamian historiography. Proceedings of the American Philosophical Society, 107(6):461–472, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986106.

Finlay:1982:FGV

[Fin82] Robert Finlay. The foundation of the Ghetto: Venice, the Jews, and the War of the League of Cambrai. Proceedings of the American Philosophical Society, 126(2):140–154, April 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986357.

Firor:1991:WOR

[Fir91] John Firor. Walter Orr Roberts (August 20, 1915–March 12, 1990). Proceedings of the American Philosophical Society, 135(4):624–628, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986823.

Fiske:1994:CGS

[Fis94] Milan D. Fiske. C. Guy Suits (12 March 1905–14 August 1991). Proceedings of the American Philosophical Society, 138(2):356–362, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987132.

Fleming:1996:EJM

[FK96] Wendell Fleming and Victor Klee. Edward James Mcshane (10 May 1904–1 June 1989). Proceedings of the American Philosophical Society, 140(4):572–578, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987293.

Flershem:1966:SAJ

[Fle66] Robert G. Flershem. Some aspects of Japan Sea shipping and trade in the Tokugawa Period, 1603–1867. Proceedings of the American Philosophical Society, 110(3):182–226, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985987.

Flexner:1967:DMM

[Fle67] Louis B. Flexner. Dissection of memory in mice with antibiotics. Proceedings of the American Philosophical Society, 111(6):343–346, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985730.

Fleming:1975:ARN

[Fle75] C. A. Fleming. Adaptive radiation in New Zealand cicadas. Proceedings of the American Philosophical Society, 119(4):298–306, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986291.

Fleming:1986:AFS

[Fle86] Stuart Fleming. Art forgery: Some scientific defenses. Proceedings of the American Philosophical Society, 130(2):175–195, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987180.

Folsom:1958:PML

[Fol58] Marion B. Folsom. Penrose Memorial Lecture. Some suggested adjustments in the use of our resources. Proceedings of the American Philosophical Society, 102(4):321–327, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985217.

Fong:1971:HUC

[Fon71] Wen C. Fong. How to understand Chinese painting. Proceedings of the American Philosophical Society, 115(4):282–292, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986090.

Fong:1998:RCA

[Fon98] Wen C. Fong. Reflections on Chinese art history. Proceedings of the American Philosophical Society, 142(1):47–59, March 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152263.

Foote:1965:GFR

[Foo65] Paul D. Foote. Government-financed research, basic and applied. Proceedings of the American Philosophical Society, 109(2):57–62, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985781.

Forbes:1959:PSF

[For59] Thomas R. Forbes. The prediction of sex: Folklore and science. Proceedings of the American Philosophical Society, 103(4):537–544, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985555.

Forbes:1962:WYB

[For62] Thomas R. Forbes. William Yarrell, British naturalist. Proceedings of the American Philosophical Society, 106(6):505–515, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985256.

Forbes:1971:VCB

[For71] Thomas R. Forbes. Verbal charms in British folk medicine. Proceedings of the American Philosophical Society, 115(4):293–316, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986091.

Ford:1972:DTC

[For72] Franklin L. Ford. Dimensions of toleration: Castellio, Bodin, Montaigne. Proceedings of the American Philosophical Society, 116(2): 136–139, April 1972. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986167.

Ford:1976:AEC

[For76] Franklin L. Ford. Assassination in the Eighteenth Century: The dog that did not bark in the night. Proceedings of the American Philosophical Society, 120(3):211–215, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986560.

Forsberg:1980:MRD

[For80] Randall Forsberg. Military R and D: a worldwide institution? Proceedings of the American Philosophical Society, 124(4):266–270, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986428.

Forer:1996:DCS

[For96] Arthur Forer. Do chromosomes segregate randomly during meiosis?: Key articles by Fernandus Payne were ignored, and perhaps “suppressed”. Proceedings of the American Philosophical Society, 140(2):148–163, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987321.

Ford:1998:CDB

[For98] Karen M. Ford. Can a democracy bind itself in perpetuity? Paine, the bank crisis, and the concept of economic freedom. Proceedings of the American Philosophical Society, 142(4):557–577, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152281.

Forster:1999:CBH

[For99] Robert E. Forster, II. Charles Brenton Huggins (22 September 1901–12 January 1997). Proceedings of the American Philosophical Society, 143(2):326–331, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3181942.

Fosdick:1971:CAJ

[Fos71] Raymond B. Fosdick. Comment on article by James E. Hewes, Jr., entitled “Henry Cabot Lodge and the League of Nations”. Proceedings of the American Philosophical Society, 115(1):65–66, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985923. See

[Hew70].

Foster:1982:PHL

[Fos82] Elizabeth Read Foster. Procedure and the House of Lords in the Seventeenth Century. Proceedings of the American Philosophical Society, 126(3):183–187, June 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986506.

Foster:1989:PCA

[Fos89] Mike Foster. The Permian Controversy of 1858: an affair of the heart. Proceedings of the American Philosophical Society, 133(3): 370–390, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987106.

Fowler:1965:MSS

[Fow65] William A. Fowler. Massive stars and strong radio sources. Proceedings of the American Philosophical Society, 109(3):181–188, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986131.

Fox:1968:JSM

[Fox68] Robert Fox. The John Scott Medal. Proceedings of the American Philosophical Society, 112(6):416–430, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985940.

Frankfurter:1954:SON

[Fra54] Felix Frankfurter. Some observations on the nature of the judicial process of supreme court litigation. Proceedings of the American Philosophical Society, 98(4):233–239, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143874.

Frank:1955:SR

[Fra55] Philipp G. Frank. Summarizing remarks. Proceedings of the American Philosophical Society, 99(5):350–351, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143936.

Francis:1960:BML

[Fra60a] F. C. Francis. The British Museum looks towards the future. Proceedings of the American Philosophical Society, 104(4):408–412, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985620.

Francis:1960:DOA

[Fra60b] Thomas Francis, Jr. On the doctrine of original antigenic sin. Proceedings of the American Philosophical Society, 104(6):572–578, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985534.

Frantz:1975:PPC

[Fra75] Alison Frantz. Pagan philosophers in Christian Athens. Proceedings of the American Philosophical Society, 119(1):29–38, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986648.

Franklin:1979:AAB

[Fra79] John Hope Franklin. Afro–American biography: The case of George Washington Williams. Proceedings of the American Philosophical Society, 123(3):160–163, June 1979. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986218.

Frantz:1983:MPA

[Fra83] M. Alison Frantz. Multum in parvo: The Aegean island of Sikinos. Proceedings of the American Philosophical Society, 127(2):71–83, April 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986190.

Fraser:1986:PL

[Fra86] David W. Fraser. The peculiarities of “Legionella”. Proceedings of the American Philosophical Society, 130(3):330–335, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986831.

Franklin:1993:SSN

[Fra93] John Hope Franklin. The social sciences for the new century. Proceedings of the American Philosophical Society, 137(4):648–655, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987079.

Frangsmyr:1999:NAS

[Frä99] Tore Frängsmyr. The new academies and the scientific climate of the Eighteenth Century. Proceedings of the American Philosophical Society, 143(1):109–115, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181979.

Freeman:1962:MSL

[Fre62] John Finley Freeman. Manuscript sources on Latin American Indians in the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 106(6):530–540, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985259.

Freimarck:1966:RY

[Fre66] Vincent Freimarck. Rhetoric at Yale in 1807. Proceedings of the American Philosophical Society, 110(4):235–255, August 1966. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985686.

Freund:1968:PLF

[Fre68] Paul A. Freund. Political libel and freedom of the press. Proceedings of the American Philosophical Society, 112(2):117–120, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985928.

Freund:1971:OTE

[Fre71] Paul A. Freund. Organ transplants: Ethical and legal problems. Proceedings of the American Philosophical Society, 115(4):276–281, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986089.

Freund:1991:SAP

[Fre91] Paul A. Freund. The “state action” problem. Proceedings of the American Philosophical Society, 135(1):3–12, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987144.

Freedman:1993:ACP

[Fre93] Joseph S. Freedman. Aristotle and the content of philosophy instruction at Central European schools and universities during the Reformation Era (1500–1650). Proceedings of the American Philosophical Society, 137(2):213–253, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986731.

Fredrickson:1996:JAS

[Fre96] Donald S. Fredrickson. James Augustine Shannon (9 August 1904—20 May 1994). Proceedings of the American Philosophical Society, 140(1):106–113, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987281.

Freedman:1997:CWB

[Fre97] Joseph S. Freedman. The career and writings of Bartholomew Keckermann (d. 1609). Proceedings of the American Philosophical Society, 141(3):305–364, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987205.

Friedman:1961:DM

[Fri61] Milton Friedman. The demand for money. Proceedings of the American Philosophical Society, 105(3):259–264, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985455.

Friedman:1962:SSI

[Fri62] William F. Friedman. Shakespeare, secret intelligence, and statecraft. Proceedings of the American Philosophical Society, 106(5): 401–411, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985484.

Friedmann:1966:SUS

[Fri66] Herbert Friedmann. The significance of the unimportant in studies of nature and of art. Proceedings of the American Philosophical Society, 110(4):256–260, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985687.

Friedman:1972:MP

[Fri72] Milton Friedman. Monetary policy. Proceedings of the American Philosophical Society, 116(3):183–196, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986113.

Friedman:1991:IU

[Fri91a] Herbert Friedman. The invisible universe. Proceedings of the American Philosophical Society, 135(3):339–345, September 1991. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986772.

Friedman:1991:YBZ

[Fri91b] Herbert Friedman. Yakov Borisovich Zel’dovich (March 8, 1914– December 2, 1987). Proceedings of the American Philosophical Society, 135(2):322–326, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987039.

Friedman:1996:AS

[Fri96] Herbert Friedman. Adrift in space. Proceedings of the American Philosophical Society, 140(1):10–21, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987273.

Friedman:1997:IUX

[Fri97] Herbert Friedman. The invisible universe of X-rays and gamma rays. Proceedings of the American Philosophical Society, 141(4): 403–416, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987217.

Fruton:1977:SAB

[Fru77] Joseph S. Fruton. Some aspects of biochemical catalysis. Proceedings of the American Philosophical Society, 121(4):309–315, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986419.

Fruton:1985:CSS

[Fru85] Joseph S. Fruton. Contrasts in scientific style. Emil Fischer and Franz Hofmeister: Their research groups and their theory of protein structure. Proceedings of the American Philosophical Society, 129(4):313–370, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986934.

Fruton:1988:LRG

[Fru88] Joseph S. Fruton. The Liebig Research Group: a reappraisal. Proceedings of the American Philosophical Society, 132(1):1–66, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143824.

Frye:1976:MPL

[Fry76] Roland Mushat Frye. Milton’s “Paradise Lost” and the visual arts. Proceedings of the American Philosophical Society, 120(4):233–244, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986319.

Frye:1983:CCS

[Fry83] Roland Mushat Frye. So-called “creation-science” and mainstream Christian rejections. Proceedings of the American Philosophical Society, 127(1):61–70, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986365.

Frye:1988:AAT

[Fry88] Roland Mushat Frye. “Not of an age, but for all time”: a Shakespearean’s thoughts on Shakespeare’s permanence. Proceedings of the American Philosophical Society, 132(3):223–236, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143849.

Frye:1989:DDO

[Fry89] Roland Mushat Frye. The dissidence of dissent and the origins of religious freedom in America: John Milton and the Puritans. Proceedings of the American Philosophical Society, 133(4):475–488, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986872.

Frye:1996:GEB

[Fry96] Roland Mushat Frye. Gerald Eades Bentley (15 September 1901– 25 July 1994). Proceedings of the American Philosophical Society, 140(1):78–85, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987277.

Frye:1998:HSS

[Fry98] Roland Mushat Frye. Hitler, Stalin, and Shakespeare’s Macbeth: Modern totalitarianism and ancient tyranny. Proceedings of the American Philosophical Society, 142(1):81–109, March 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152266.

Fuchs:1996:WEP

[Fuc96] Victor R. Fuchs. What every philosopher should know about health economics. Proceedings of the American Philosophical Society, 140 (2):186–196, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987324.

Fuchs:1998:PCN

[Fuc98] Ira H. Fuchs. The promise and challenge of new technologies in higher education. Proceedings of the American Philosophical Society, 142(2):191–206, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152298.

Fulton:1951:PBP

[Ful51] John F. Fulton. The physiological basis of psychosurgery. Proceedings of the American Philosophical Society, 95(5):538–541, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143238.

Fulbright:1961:PML

[Ful61] J. William Fulbright. Penrose Memorial Lecture. In need of a consensus. Proceedings of the American Philosophical Society, 105 (4):349–353, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985493.

Fuller:1968:FPA

[Ful68] Lon L. Fuller. Freedom as a problem of allocating choice. Proceedings of the American Philosophical Society, 112(2):101–106, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985926.

Fussner:1957:WCD

[Fus57] Frank Smith Fussner. William Camden’s “Discourse concerning the Prerogative of the Crown”. Proceedings of the American Philosophical Society, 101(2):204–215, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985513.

Galbraith:1966:PML

[Gal66] V. H. Galbraith. Penrose Memorial Lecture. Runnymede revisited. Proceedings of the American Philosophical Society, 110(5):307–317, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986019.

Gallo:1974:MVI

[Gal74] Ernest Gallo. Matthew of Vendôme: Introductory treatise on the art of poetry. Proceedings of the American Philosophical Society, 118(1):51–92, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986436.

Gallaher:1991:EHS

[Gal91] Art Gallaher, Jr. Edward Holland Spicer (November 29, 1906– April 5, 1983). Proceedings of the American Philosophical Society, 135(3):480–484, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986783.

Galbraith:1992:FU

[Gal92] John Kenneth Galbraith. The functional underclass. Proceedings of the American Philosophical Society, 136(3):411–415, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986914.

Galbraith:1999:GP

[Gal99] James K. Galbraith. Globalization and pay. Proceedings of the American Philosophical Society, 143(2):178–186, June 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181932.

Gara:1961:URL

[Gar61] Larry Gara. The underground railroad: Legend or reality? Proceedings of the American Philosophical Society, 105(3):334–339, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985459.

Garrett:1974:PSG

[Gar74] Sylvester Garrett. Permanent systems for grievance arbitration. Proceedings of the American Philosophical Society, 118(5):417–421, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986306.

Garrett:1975:SOJ

[Gar75] Clarke Garrett. The spiritual odyssey of Jacob Duché. Proceedings of the American Philosophical Society, 119(2):143–155, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986632.

Garwin:1980:NWO

[Gar80] Richard L. Garwin. New weapons/old doctrines: Strategic warfare in the 1980s. Proceedings of the American Philosophical Society, 124(4):261–265, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986427.

Gaarding:1999:MSO

[Går99] Lars Gårding. Models in science and otherwise. Proceedings of the American Philosophical Society, 143(1):3–11, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181969.

Gay:1991:BW

[Gay91] Peter Gay. The bite of wit. Proceedings of the American Philosophical Society, 135(3):327–331, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986770.

Gottschalk:1956:RBR

[GB56] Louis Gottschalk and Edmund Burke. Reflections on Burke’s “Reflections on the French Revolution”. Proceedings of the American Philosophical Society, 100(5):417–429, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143674.

Guillen:1994:OFS

[GB94] Ann Cyphers Guillén and Fernando Botas. An Olmec feline sculpture from El Azuzul, Southern Veracruz. Proceedings of the American Philosophical Society, 138(2):273–283, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987124.

Greenewalt:1960:ICH

[GBF60] Crawford H. Greenewalt, Werner Brandt, and D. D. Friel. The iridescent colors of hummingbird feathers. Proceedings of the American Philosophical Society, 104(3):249–253, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985242.

Gruhn:1977:TPI

[GBN77] Ruth Gruhn, Alan Lyle Bryan, and Jack D. Nance. Los Tapiales: a Paleo-Indian campsite in the Guatemalan Highlands. Proceedings of the American Philosophical Society, 121(3):235–273, June 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986413.

Goldstein:1994:PLV

[GCM94] Bernard R. Goldstein, José Chabás, and José Luis Mancha. Planetary and lunar velocities in the Castilian Alfonsine Tables. Proceedings of the American Philosophical Society, 138(1):61–95, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986706.

Gellhorn:1968:PLL

[Gel68] Walter Gellhorn. Poverty and legality: The law’s slow awakening. Proceedings of the American Philosophical Society, 112(2):107–116, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985927.

Gellhorn:1995:ENG

[Gel95a] Walter Gellhorn. Erwin Nathaniel Griswold (14 July 1904–19 November 1994). Proceedings of the American Philosophical Society, 139(4):428–432, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987239.

Gellhorn:1995:HWJ

[Gel95b] Walter Gellhorn. Harry Willmer Jones (4 March 1911–6 April 1993). Proceedings of the American Philosophical Society, 139(1): 110–113, March 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986726.

Gemery:1989:DHR

[Gem89] Henry A. Gemery. Disarray in the historical record: Estimates of immigration to the United States, 1700–1860. Proceedings of the American Philosophical Society, 133(2):123–127, June 1989. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987043.

Gerschenkron:1962:CCH

[Ger62] Alexander Gerschenkron. On the concept of continuity in history. Proceedings of the American Philosophical Society, 106(3):195–209, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985175.

Gerschenkron:1974:FSS

[Ger74] Alexander Gerschenkron. Figures of speech in social sciences. Proceedings of the American Philosophical Society, 118(5):431–448, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986308.

Gerschenkron:1978:THB

[Ger78] Alexander Gerschenkron. Time horizon in Balzac and others. Proceedings of the American Philosophical Society, 122(2):75–91, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986624.

Geyl:1962:SAC

[Gey62] Pieter Geyl. Synopsis: The American Civil War viewed from The Netherlands. Proceedings of the American Philosophical Society, 106(1):33–35, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985207.

Graham:1958:DNC

[GH58] Clarence H. Graham and Yun Hsia. The discriminations of a normal and a color-blind eye in the same person. Proceedings of the American Philosophical Society, 102(2):168–173, April 1958. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985282.

Gordon:1995:BWJ

[GHKW95] Roy Gordon, Dudley Herschbach, William Klemperer, and Frank Westheimer. E. Bright Wilson, Jr. (18 December 1908–12 July 1992). Proceedings of the American Philosophical Society, 139(3): 312–315, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987199.

Giarrizzo:1997:EPI

[Gia97] Giuseppe Giarrizzo. Enlightenment: The parabola of an idea. Proceedings of the American Philosophical Society, 141(4):436–453, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987220.

Gibb:1963:CIH

[Gib63] Hamilton A. R. Gibb. The community in Islamic history. Proceedings of the American Philosophical Society, 107(2):173–176, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985447.

Gibbon:1965:NHT

[Gib65] John H. Gibbon, Jr. The normal heart and temporary substitutes for the heart and lungs. Proceedings of the American Philosophical Society, 109(3):112–116, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986125.

Gibbons:1981:EPO

[Gib81] John H. Gibbons. Energy productivity: Opportunities, constraints, and progress. Proceedings of the American Philosophical Society, 125(4):251–261, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986326.

Gibson:1992:CIG

[Gib92] Campbell Gibson. The contribution of immigration to the growth and ethnic diversity of the American population. Proceedings of the American Philosophical Society, 136(2):157–175, June 1992. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987166.

Gibbons:1995:EDP

[Gib95] John H. Gibbons. Energy for a doubled population. Proceedings of the American Philosophical Society, 139(3):256–264, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987191.

Gideonse:1950:CIA

[Gid50] Harry D. Gideonse. Changing issues in academic freedom in the United States today. Proceedings of the American Philosophical Society, 94(2):91–104, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143207.

Giddings:1952:DPA

[Gid52] J. L. Giddings, Jr. Driftwood and problems of Arctic sea currents. Proceedings of the American Philosophical Society, 96(2):129–142, April 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143719.

Giddings:1962:SDB

[Gid62] J. L. Giddings. Seven discoveries of Bering Strait. Proceedings of the American Philosophical Society, 106(2):89–93, April 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985374.

Gillispie:1972:PPL

[Gil72] Charles Coulston Gillispie. Probability and politics: Laplace, Condorcet, and Turgot. Proceedings of the American Philosophical Society, 116(1):1–20, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985703.

Gill:1974:GA

[Gil74] Lewis M. Gill. Grievance arbitration. Proceedings of the American Philosophical Society, 118(5):415–416, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986305.

Gilmore:1976:BVT

[Gil76] Myron P. Gilmore. The Berensons and Villa I Tatti. Proceedings of the American Philosophical Society, 120(1):7–12, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986347.

Gillmor:1982:WAE

[Gil82] C. Stewart Gillmor. Wilhelm Altar, Edward Appleton, and the magneto-ionic theory. Proceedings of the American Philosophical Society, 126(5):395–440, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986214.

Gilbert:1986:LRA

[Gil86] Felix Gilbert. Leopold von Ranke and the American Philosophical Society. Proceedings of the American Philosophical Society, 130(3): 362–366, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986835.

Gillispie:1989:SAF

[Gil89] Charles Coulston Gillispie. Scientific aspects of the French Egyptian Expedition 1798–1801. Proceedings of the American Philosophical Society, 133(4):447–474, December 1989. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986871.

Gillispie:1992:PBA

[Gil92] Charles C. Gillispie. Palisot de Beauvois on the Americans. Proceedings of the American Philosophical Society, 136(1):33–50, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986796.

Gillispie:1994:JBA

[Gil94] Charles C. Gillispie. Jerome Blum (27 April 1913–7 May 1993). Proceedings of the American Philosophical Society, 138(3):408–412, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986747.

Gillis:1995:CPJ

[Gil95] J. Gillis. Chaim L. Pekeris (15 June 1908–25 February 1993). Proceedings of the American Philosophical Society, 139(2):178–181, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987265.

Gingerich:1973:CKH

[Gin73] Owen Gingerich. From Copernicus to Kepler: Heliocentrism as model and as reality. Proceedings of the American Philosophical Society, 117(6):513–522, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986462.

Ginzberg:1978:PE

[Gin78] Eli Ginzberg. The politics of employment. Proceedings of the American Philosophical Society, 122(3):131–134, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986546.

Gingerich:1992:LGJ

[Gin92] Owen Gingerich. Leo Goldberg (January 26, 1913–November 1, 1987). Proceedings of the American Philosophical Society, 136(2): 296–301, June 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987175.

Gingerich:1999:CSN

[Gin99] Owen Gingerich. Carl Sagan (9 November 1934–20 December 1996). Proceedings of the American Philosophical Society, 143(4): 712–716, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181999.

Greenewalt:1955:PSF

[GJ55] Crawford H. Greenewalt and Frank Morton Jones. Photographic studies of the feeding of nestling house wrens. Proceedings of the American Philosophical Society, 99(4):200–204, August 1955. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143698.

Glass:1960:ECC

[Gla60] Bentley Glass. Eighteenth-Century concepts of the origin of species. Proceedings of the American Philosophical Society, 104 (2):227–234, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985661.

Glass:1965:CBG

[Gla65] Bentley Glass. A century of biochemical genetics. Proceedings of the American Philosophical Society, 109(4):227–236, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985881.

Glass:1981:HCG

[Gla81] Bentley Glass. A hidden chapter of German eugenics between the two World Wars. Proceedings of the American Philosophical Society, 125(5):357–367, October 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986198.

Glass:1993:GMD

[Gla93] Bentley Glass. Giuseppe Montalenti (13 December 1904–2 July 1990). Proceedings of the American Philosophical Society, 137(2): 294–298, June 1993. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986737.

Gleason:1960:SWN

[Gle60] Henry A. Gleason. The scientific work of Nathaniel Lord Britton. Proceedings of the American Philosophical Society, 104(2):205–226, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985660.

Glickman:1994:FBA

[Gli94] Stephen E. Glickman. Frank A. Beach (13 April 1911–15 June 1988). Proceedings of the American Philosophical Society, 138(1): 160–164, March 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986710.

Glueck:1956:AHW

[Glu56] Nelson Glueck. Ancient highways in the wilderness of zin. Proceedings of the American Philosophical Society, 100(2):150–155, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143715.

Gobar:1970:PWJ

[Gob70] Ash Gobar. The phenomenology of William James. Proceedings of the American Philosophical Society, 114(4):294–309, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985954.

Goetze:1953:CEA

[Goe53] Albrecht Goetze. The cultures of Early Anatolia. Proceedings of the American Philosophical Society, 97(2):214–221, April 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143895.

Goitein:1971:FFM

[Goi71] Shelomo D. Goitein. Formal friendship in the medieval near east. Proceedings of the American Philosophical Society, 115(6):484–489, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985841.

Goldstein:1964:ATC

[Gol64a] Doris S. Goldstein. Alexis de Tocqueville’s concept of citizenship. Proceedings of the American Philosophical Society, 108(1):39–53, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985995.

Goldstine:1964:CP

[Gol64b] Herman H. Goldstine. Computers and perception. Proceedings of the American Philosophical Society, 108(4):282–290, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985908.

Golb:1969:M

[Gol69a] Norman Golb. Monieux. Proceedings of the American Philosophical Society, 113(1):67–94, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986016.

Goldstein:1969:RVS

[Gol69b] Richard M. Goldstein. A radar view of the surface of Venus. Proceedings of the American Philosophical Society, 113(3):224–228, June 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985737.

Gold:1971:NLS

[Gol71] Thomas Gold. The nature of the lunar surface: Recent evidence. Proceedings of the American Philosophical Society, 115(2):74–82, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985848.

Goldhaber:1975:TAF

[Gol75] Maurice Goldhaber. Tacit assumptions in fundamental particle physics. Proceedings of the American Philosophical Society, 119 (1):24–28, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986647.

Goldstine:1977:BHC

[Gol77] Herman H. Goldstine. A brief history of the computer. Proceedings of the American Philosophical Society, 121(5):339–345, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986336.

Golb:1980:POI

[Gol80] Norman Golb. The problem of origin and identification of the dead sea scrolls. Proceedings of the American Philosophical Society, 124 (1):1–24, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986653.

Golden:1986:SAP

[Gol86] William T. Golden. Science advice to the president: Past, present, future. Proceedings of the American Philosophical Society, 130(3): 325–329, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986830.

Goldstine:1987:P

[Gol87] Herman H. Goldstine. Preface. Proceedings of the American Philosophical Society, 131(3):vii, September 1987. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987016.

Golden:1988:E

[Gol88a] William T. Golden. Epilogue. Proceedings of the American Philosophical Society, 132(2):221–222, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986867.

Golden:1988:P

[Gol88b] William T. Golden. Prologue. Proceedings of the American Philosophical Society, 132(2):172, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986862.

Goldstein:1988:NSF

[Gol88c] Bernard R. Goldstein. A new set of Fourteenth Century planetary observations. Proceedings of the American Philosophical Society, 132(4):371–399, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986960.

Goldstine:1988:P

[Gol88d] Herman H. Goldstine. Preface. Proceedings of the American Philosophical Society, 132(1):v, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143823.

Golden:1991:WSA

[Gol91] William T. Golden. Worldwide science advising to the highest levels of governments. Proceedings of the American Philosophical Society, 135(1):41–48, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987148.

Golden:1992:I

[Gol92a] William T. Golden. Introduction. Proceedings of the American Philosophical Society, 136(3):319, September 1992. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986902.

Goldstine:1992:JLC

[Gol92b] Herman H. Goldstine. The Jayne Lecture. Computers at the University of Pennsylvania’s Moore School, 1943–1946. Proceedings of the American Philosophical Society, 136(1):73–78, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986799.

Goldstine:1993:P

[Gol93] Herman H. Goldstine. Preface. Proceedings of the American Philosophical Society, 137(4):iv–vi, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987063.

Golden:1994:ABK

[Gol94] William T. Golden. Augustus B. Kinzel (26 July 1900–23 October 1987). Proceedings of the American Philosophical Society, 138 (1):166–168, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986711.

Golden:1997:JSM

[Gol97] William T. Golden. John E. Sawyer (5 May 1917–7 February 1995). Proceedings of the American Philosophical Society, 141(1):104–106, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987255.

Gomberg:1974:PAR

[Gom74] William Gomberg. The problem of arbitration — the resolution of public sector disputes. Proceedings of the American Philosophical Society, 118(5):409–414, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986304.

Gomory:1985:RI

[Gom85] Ralph E. Gomory. Research in industry. Proceedings of the American Philosophical Society, 129(1):26–29, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986978.

Gomory:1992:UGR

[Gom92] R. E. Gomory. The U.S. Government’s role in science & technology. Proceedings of the American Philosophical Society, 136(1):79–84, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986800.

Gomory:1995:JLI

[Gom95] Ralph E. Gomory. 1991 Jayne Lecture. The information evolution in perspective: a view from the year 2091. Proceedings of the American Philosophical Society, 139(4):420–426, December 1995. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987238.

Goodrich:1953:UPI

[Goo53] Herbert F. Goodrich. An unsettled problem in interstate divorce law. Proceedings of the American Philosophical Society, 97(2):147– 150, April 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143887.

Goodrich:1958:GPA

[Goo58] Carter Goodrich. The Gallatin Plan after one hundred and fifty years. Proceedings of the American Philosophical Society, 102(5): 436–441, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985588.

Goodenough:1969:FCA

[Goo69] Ward H. Goodenough. Frontiers of cultural anthropology: Social organization. Proceedings of the American Philosophical Society, 113(5):329–335, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985969.

Goodenough:1976:OMC

[Goo76] Ward H. Goodenough. On the origin of matrilineal clans: a “just so” story. Proceedings of the American Philosophical Society, 120 (1):21–36, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986349.

Goodenough:1981:DRT

[Goo81] Ward H. Goodenough. On describing religion in truk: an anthropological dilemma. Proceedings of the American Philosophical Society, 125(6):411–415, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986297.

Goodenough:1990:CSI

[Goo90] Ward H. Goodenough. Concern with social identity and personal worth as a source of social turbulence. Proceedings of the American Philosophical Society, 134(3):328–334, September 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986971.

Gordon:1980:PSC

[Gor80] Michael D. Gordon. The perjury statute of 1563: a case history of confusion. Proceedings of the American Philosophical Society, 124 (6):438–454, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986245.

Gorman:1988:IWS

[Gor88] Mel Gorman. Introduction of Western science into Colonial India: Role of the Calcutta Medical College. Proceedings of the American Philosophical Society, 132(3):276–298, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143855.

Gotoff:1991:SRA

[Got91] Harold C. Gotoff. Sir Roger Aubrey Baskerville Mynors (July 28, 1903–October 17, 1989). Proceedings of the American Philosophical Society, 135(2):308–311, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987037.

Goudsmit:1950:FS

[Gou50] S. A. Goudsmit. Freedom of science. Proceedings of the American Philosophical Society, 94(2):111–113, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143209.

Goudsmit:1956:SI

[Gou56] Samuel A. Goudsmit. Scientific intelligence. Proceedings of the American Philosophical Society, 100(1):100–103, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143810.

Gould:1965:EET

[Gou65] Edwin Gould. Evidence for echolocation in the Tenrecidae of Madagascar. Proceedings of the American Philosophical Society, 109(6):352–360, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986137.

Gould:1966:WQA

[Gou66] Richard A. Gould. The wealth quest among the Tolowa Indians of Northwestern California. Proceedings of the American Philosophical Society, 110(1):67–89, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986003.

Goudsmit:1977:PEP

[Gou77] Samuel A. Goudsmit. Pitfalls in elementary probability. Proceedings of the American Philosophical Society, 121(2):188–189, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986527.

Gough:1989:SDC

[Gou89] Robert J. Gough. The significance of the demographic characteristics of wealthy Philadelphians at the end of the Eighteenth Century. Proceedings of the American Philosophical Society, 133(2):305–311, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987056.

Gould:1997:RTL

[Gou97] Stephen Jay Gould. Redrafting the tree of life. Proceedings of the American Philosophical Society, 141(1):30–54, March 1997. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987248.

Goldstein:1981:MHC

[GP81] Bernard R. Goldstein and David Pingree. More horoscopes from the Cairo Geniza. Proceedings of the American Philosophical Society, 125(3):155–189, June 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986481.

Gray:1972:BCR

[Gra72] Charles M. Gray. Bonham’s case reviewed. Proceedings of the American Philosophical Society, 116(1):35–58, February 1972. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985705.

Grafton:1973:MMA

[Gra73] Anthony Grafton. Michael Maestlin’s account of Copernican planetary theory. Proceedings of the American Philosophical Society, 117(6):523–550, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986463.

Granit:1975:ANS

[Gra75] Ragnar Granit. Adaptability of the nervous system in relation to chance, purposiveness, and causality. Proceedings of the American Philosophical Society, 119(4):262–266, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986287.

Grant:1985:NLM

[Gra85] Edward Grant. A new look at medieval cosmology, 1200–1687. Proceedings of the American Philosophical Society, 129(4):417–432, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986937.

Grabbe:1989:EIU

[Gra89] Hans-Jürgen Grabbe. European immigration to the United States in the Early National Period, 1783–1820. Proceedings of the American Philosophical Society, 133(2):190–214, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987050. See errata [Ano89g].

Graff:1993:RT

[Gra93a] Gerald Graff. Is reason in trouble? Proceedings of the American Philosophical Society, 137(4):680–688, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987082.

Graffman:1993:RSM

[Gra93b] Gary Graffman. Rudolf Serkin (28 March 1903–8 May 1991). Proceedings of the American Philosophical Society, 137(1):168–171, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986953.

Graham:1996:PTR

[Gra96] Loren R. Graham. Palchinsky’s travels: a Russian engineer’s adventures among gigantic projects and small minds. Proceedings of the American Philosophical Society, 140(2):175–185, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987323.

Grafton:1998:GCT

[Gra98] Anthony Grafton. Girolamo Cardano and the tradition of classical astrology: the Rothschild Lecture, 1995. Proceedings of the American Philosophical Society, 142(3):323–354, September 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152240.

Grew:1954:PCL

[Gre54] Joseph C. Grew. The Peace Conference of Lausanne, 1922– 1923. Proceedings of the American Philosophical Society, 98(1): 1–10, February 1954. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143664.

Greene:1959:DR

[Gre59] John C. Greene. Darwin and religion. Proceedings of the American Philosophical Society, 103(5):716–725, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985427.

Greenaway:1960:RLB

[Gre60a] Emerson Greenaway. The relation of library buildings to library functions. Proceedings of the American Philosophical Society, 104 (4):391–397, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985617.

Greene:1960:ESI

[Gre60b] John C. Greene. Early scientific interest in the American Indian: Comparative linguistics. Proceedings of the American Philosophical Society, 104(5):511–517, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985234.

Greenewalt:1960:WIB

[Gre60c] Crawford H. Greenewalt. The wings of insects and birds as mechanical oscillators. Proceedings of the American Philosophical Society, 104(6):605–611, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985536.

Greene:1969:DMP

[Gre69] John C. Greene. The development of mineralogy in Philadelphia, 1780–1820. Proceedings of the American Philosophical Society, 113 (4):283–295, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986153.

Greenewalt:1977:ELS

[Gre77] Crawford H. Greenewalt. The energetics of locomotion — is small size really disadvantageous? Proceedings of the American Philosophical Society, 121(2):100–106, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986522.

Greenberg:1980:TLM

[Gre80] Joseph H. Greenberg. Types of linguistic models in other disciplines. Proceedings of the American Philosophical Society, 124 (1):35–40, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986655.

Greene:1987:IME

[Gre87] Jack P. Greene. “An instructive monitor”: Experience and the fabrication of the Federal Constitution. Proceedings of the American Philosophical Society, 131(3):298–307, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987025.

Greenewalt:1992:WME

[Gre92] Crawford H. Greenewalt, Jr. When a mighty empire was destroyed: The common man at the fall of Sardis, ca. 546 B.C. Proceedings of the American Philosophical Society, 136(2):247–271, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987171.

Greenberg:1993:OCR

[Gre93a] Joseph H. Greenberg. Observations concerning Ringe’s “Calculating the Factor of Chance in Language Comparison”. Proceedings of the American Philosophical Society, 137(1):79–90, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986946.

Gregorian:1993:EOD

[Gre93b] Vartan Gregorian. Education and our divided knowledge. Proceedings of the American Philosophical Society, 137(4):605–611, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987076.

Greenstein:1994:NUM

[Gre94] Jesse L. Greenstein. Nicholas Ulrich Mayall (9 May 1906–5 January 1993). Proceedings of the American Philosophical Society, 138(4): 546–551, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986853.

Greenawalt:1998:RPP

[Gre98a] Kent Greenawalt. Has religion any place in the politics and law of liberal democracy? Proceedings of the American Philosophical Society, 142(3):378–387, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152243.

Greene:1998:RWA

[Gre98b] Thomas M. Greene. Rene Wellek (22 August 1903–10 November 1995). Proceedings of the American Philosophical Society, 142(2): 308–313, June 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152309.

Gregory:1998:VES

[Gre98c] Terri Gregory. Verner Edward Suomi (6 December 1915–30 July 1995). Proceedings of the American Philosophical Society, 142(3): 500–509, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152256.

Griffin:1972:XEG

[Gri72a] Gillett G. Griffin. Xochipala, the earliest great art style in Mexico. Proceedings of the American Philosophical Society, 116(4):301–309, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985900.

Griffiths:1972:SVE

[Gri72b] David M. Griffiths. Soviet views of early Russian–American relations. Proceedings of the American Philosophical Society, 116 (2):148–156, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986170.

Griffin:1973:OBM

[Gri73] Donald R. Griffin. Oriented bird migration in or between opaque cloud layers. Proceedings of the American Philosophical Society, 117(2):117–141, April 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986062.

Griswold:1982:SQC

[Gri82] Erwin N. Griswold. Sed quis custodiet ipsos custodies? Some reflections on ABSCAM. Proceedings of the American Philosophical Society, 126(6):452–460, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986251.

Griffin:1995:HMM

[Gri95] James B. Griffin. Horace Mitchell Miner (26 May 1912–26 November 1993). Proceedings of the American Philosophical Society, 139 (3):288–292, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987195.

Gruber:1967:HHD

[Gru67] Jacob W. Gruber. Horatio Hale and the development of American anthropology. Proceedings of the American Philosophical Society, 111(1):5–37, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985751.

Grubb:1989:SAR

[Gru89] Farley Grubb. Servant auction records and immigration into the Delaware Valley, 1745–1831: The proportion of females among immigrant servants. Proceedings of the American Philosophical Society, 133(2):154–169, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987046.

Glass:1980:SEL

[GS80] Bentley Glass and Geo. H. Shull. The strange encounter of Luther Burbank and George Harrison Shull. Proceedings of the American Philosophical Society, 124(2):133–153, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986208.

Glass:1986:GET

[GS86] Bentley Glass and Curt Stern. Geneticists embattled: Their stand against rampant eugenics and racism in America during the 1920s and 1930s. Proceedings of the American Philosophical Society, 130 (1):130–154, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987094.

Goodman:1994:JVL

[GS94] Roy Goodman and Pierre Swiggers. John Vaughan (1756–1841) and the Linguistic Collection in the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 138(2):251–272, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987123.

Gummere:1955:HCC

[Gum55] Richard M. Gummere. The heritage of the classics in Colonial North America. An essay on the Greco–Roman tradition. Proceedings of the American Philosophical Society, 99(2):68–78, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143817.

Gundersheimer:1993:GEM

[Gun93] Werner Gundersheimer. “The green-eyed monster”: Renaissance conceptions of jealousy. Proceedings of the American Philosophical Society, 137(3):321–331, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986993.

Gunther:1996:JLH

[Gun96] Gerald Gunther. Judge learned hand: The choices and satisfactions of a biographer. Proceedings of the American Philosophical Society, 140(4):451–458, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987287.

Gusinde:1956:YIW

[Gus56] Martin Gusinde. The Yupa Indians in Western Venezuela. Proceedings of the American Philosophical Society, 100(3):197–222, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143772.

Gusinde:1961:SAP

[Gus61] Martin Gusinde. Somatology of the Ayom Pygmies of New Guinea. Proceedings of the American Philosophical Society, 105(4):394–411, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985497.

Guttman:1964:E

[Gut64] Norman Guttman. Experience. Proceedings of the American Philosophical Society, 108(6):459–463, December 1964. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985863.

Guterbock:1984:HAN

[Güt84] Hans G. Güterbock. Hittites and Akhaeans: a new look. Proceedings of the American Philosophical Society, 128(2):114–122, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986225.

Guterbock:1992:KBJ

[Güt92] Hans G. Güterbock. Kurt Bittel (5 July 1907–30 January 1991). Proceedings of the American Philosophical Society, 136(4):578–583, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986763.

Haac:1950:MP

[Haa50a] Oscar A. Haac. Michelet’s purpose. Proceedings of the American Philosophical Society, 94(5):494–501, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143615.

Haac:1950:SJM

[Haa50b] Oscar A. Haac. A spiritual journey: Michelet in Germany 1842. A study of Michelet’s thought in 1842 according to original documents, diaries of Alfred Dumesnil and Adèle Michelet. Proceedings of the American Philosophical Society, 94(5):502–509, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143616.

Haas:1980:WAR

[Haa80] J. M. Haas. Work and authority in the Royal Dockyards from the Seventeenth Century to 1870. Proceedings of the American Philosophical Society, 124(6):419–428, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986243.

Haas:1996:ALD

[Haa96] Warren J. Haas. America’s libraries: Distinguished past, difficult future. Proceedings of the American Philosophical Society, 140(2):125–134, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987319.

Habicht:1985:ABH

[Hab85] Christian Habicht. An ancient Baedeker and his critics: Pausanias’ “Guide to Greece”. Proceedings of the American Philosophical Society, 129(2):220–224, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986990.

Habicht:1996:ASA

[Hab96] Christian Habicht. Athens, Samos, and Alexander the Great. Proceedings of the American Philosophical Society, 140(3):397–405, September 1996. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987315.

Hackerman:1974:P

[Hac74] Norman Hackerman. On passivity. Proceedings of the American Philosophical Society, 118(4):334–336, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986280.

Hack:1997:OEU

[Hac97] Margherita Hack. Origin and evolution of the universe: Observations and problems. Proceedings of the American Philosophical Society, 141(4):397–402, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987216.

Hall:1955:BC

[Hal55] H. Duncan Hall. The British Commonwealth. Proceedings of the American Philosophical Society, 99(4):251–258, August 1955. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143704.

Halpern:1963:RDM

[Hal63] Jules Halpern. Rare decays of the mu meson. Proceedings of the American Philosophical Society, 107(5):427–430, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985647.

Hamilton:1984:SSP

[Ham84] Harlan W. Hamilton. Sterne’s sermon in Paris and its background. Proceedings of the American Philosophical Society, 128(4):316–325, December 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987099.

Hamburg:1992:IEH

[Ham92] David A. Hamburg. Interventions for education and health in very poor communities. Proceedings of the American Philosophical Society, 136(3):333–346, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986905.

Hansen:1950:PFA

[Han50] Henry P. Hansen. Postglacial forests along the Alaska Highway in British Columbia. Proceedings of the American Philosophical Society, 94(5):411–421, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143608.

Hanke:1953:BCS

[Han53a] Lewis Hanke. Bartolomé De Las Casas and the Spanish Empire in America: Four centuries of misunderstanding. Proceedings of the American Philosophical Society, 97(1):26–30, February 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143727.

Hans:1953:UEC

[Han53b] Nicholas Hans. UNESCO of the Eighteenth Century. La Loge des Neuf Sœurs and its venerable master, Benjamin Franklin. Proceedings of the American Philosophical Society, 97(5):513–524, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149268.

Hans:1954:FJE

[Han54] Nicholas Hans. Franklin, Jefferson, and the English radicals at the end of the Eighteenth Century. Proceedings of the American Philosophical Society, 98(6):406–426, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143863.

Hanke:1963:DCA

[Han63] Lewis Hanke. The dawn of conscience in America: Spanish experiments and experiences with Indians in the New World. Proceedings of the American Philosophical Society, 107(2):83–92, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985435.

Hand:1968:PTF

[Han68] Wayland D. Hand. “Passing through”: Folk medical magic and symbolism. Proceedings of the American Philosophical Society, 112 (6):379–402, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985938.

Hanfmann:1973:PPA

[Han73] George M. A. Hanfmann. Personality and portraiture in ancient art. Proceedings of the American Philosophical Society, 117(4):259– 285, August 1973. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986695.

Handler:1976:NIA

[Han76] Edward Handler. “Nature itself is all arcanum”: The scientific outlook of John Adams. Proceedings of the American Philosophical Society, 120(3):216–229, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986561.

Harper:1953:WBA

[Har53] Francis Harper. William Bartram and the American Revolution. Proceedings of the American Philosophical Society, 97(5): 571–577, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3149271.

Harrison:1958:CRD

[Har58] George R. Harrison. The controlled ruling of diffraction gratings. Proceedings of the American Philosophical Society, 102(5):483–491, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985595.

Harris:1962:EAM

[Har62] Britton Harris. Economic aspects of metropolitan planning. Proceedings of the American Philosophical Society, 106(3):165–169, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985171.

Harris:1964:TLS

[Har64] Zellig S. Harris. Transformations in linguistic structure. Proceedings of the American Philosophical Society, 108(5):418–422, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985815.

Harbison:1971:HRW

[Har71a] Frederick H. Harbison. Human resources as the wealth of nations. Proceedings of the American Philosophical Society, 115(6):426–431, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985834.

Harnwell:1971:EL

[Har71b] Gaylord P. Harnwell. An environment for learning. Proceedings of the American Philosophical Society, 115(3):170–186, June 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985976.

Hartner:1973:CMW

[Har73] Willy Hartner. Copernicus, the man, the work, and its history. Proceedings of the American Philosophical Society, 117(6):413–422, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986460.

Harris:1989:DDC

[Har89] P. M. G. Harris. The demographic development of Colonial Philadelphia in some comparative perspective. Proceedings of the American Philosophical Society, 133(2):262–304, June 1989. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987055.

Harrison:1993:PHB

[Har93] Evelyn B. Harrison. Peter Heinrich von Blanckenhagen (21 March 1909–6 March 1990). Proceedings of the American Philosophical Society, 137(2):274–277, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986734.

Haskins:1961:FWR

[Has61] Caryl P. Haskins. A flower where the roads divide. Proceedings of the American Philosophical Society, 105(4):452–458, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985501.

Haskins:1979:IRS

[Has79] Caryl P. Haskins. Introductory remarks for the symposium on animal communities. Proceedings of the American Philosophical Society, 123(4):203, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986578.

Haskins:1980:FA

[Has80] Caryl P. Haskins. The flame of the amateur. Proceedings of the American Philosophical Society, 124(5):313–319, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986570.

Hauser:1967:LTC

[Hau67] Philip M. Hauser. Lao-Tze, Confucius, and the conservative-liberal debate. Proceedings of the American Philosophical Society, 111 (5):259–267, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986043.

Hauptman:1975:GTR

[Hau75] Laurence M. Hauptman. Governor Theodore Roosevelt and the Indians of New York State. Proceedings of the American Philosophical Society, 119(1):1–7, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986645.

Hodges:1984:FFP

[HC84] Carl N. Hodges and Wayne L. Collins. Future food production: The potential is infinite, the reality may not be. Proceedings of the American Philosophical Society, 128(1):27–30, March 1984. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986489.

Hauptman:1988:VEI

[HC88] Laurence M. Hauptman and Jack Campisi. The voice of Eastern Indians: The American Indian Chicago Conference of 1961 and the movement for Federal recognition. Proceedings of the American Philosophical Society, 132(4):316–329, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986958.

Hildebrand:1996:VGD

[HE96] John G. Hildebrand and Thomas Eisner. Vincent Gaston Dethier (20 February 1915–8 September 1993). Proceedings of the American Philosophical Society, 140(2):222–228, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987327.

Heaton:1951:IIU

[Hea51] Herbert Heaton. The industrial immigrant in the United States, 1783–1812. Proceedings of the American Philosophical Society, 95 (5):519–527, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143236.

Heard:1971:USE

[Hea71] Alexander Heard. University and society: Educational freedom and institutional obligation. Proceedings of the American Philosophical Society, 115(6):423–425, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985833.

Heard:1988:HLJ

[Hea88] Alexander Heard. The Henry LaBarre Jayne Memorial Lecture of the American Philosophical Society: 25 February 1988: Made in America: The way we choose our Presidents. Proceedings of the American Philosophical Society, 132(4):305–315, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986957.

Heale:1995:TLR

[Hea95] M. J. Heale. The triumph of liberalism? Red scare politics in Michigan, 1938–1954. Proceedings of the American Philosophical Society, 139(1):44–66, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986719.

Heckscher:1994:MDM

[Hec94] Maurice Heckscher. Morris Duane (20 March 1901–18 July 1992). Proceedings of the American Philosophical Society, 138(3):422–424, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986749.

Hedberg:1965:EHR

[Hed65] Hollis D. Hedberg. Earth history and the record in the rocks. Proceedings of the American Philosophical Society, 109(2):99–104, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985785.

Heiser:1951:SAN

[Hei51] Charles B. Heiser, Jr. The sunflower among the North American Indians. Proceedings of the American Philosophical Society, 95 (4):432–448, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143283.

Heilbron:1994:ACG

[Hei94] J. L. Heilbron. The affair of the Countess Görlitz. Proceedings of the American Philosophical Society, 138(2):284–316, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987125.

Heilbron:1998:TSK

[Hei98] J. L. Heilbron. Thomas Samuel Kuhn (18 July 1922–17 June 1996). Proceedings of the American Philosophical Society, 142(4): 678–686, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152292.

Henderson:1982:DES

[Hen82] D. A. Henderson. The deliberate extinction of a species. Proceedings of the American Philosophical Society, 126(6):461–471, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986252.

Henkin:1987:USC

[Hen87] Louis Henkin. The United States Constitution as social compact. Proceedings of the American Philosophical Society, 131(3):261–269, September 1987. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987021.

Henkin:1997:WGS

[Hen97] Louis Henkin. Walter Gellhorn (18 September 1906–9 December 1995). Proceedings of the American Philosophical Society, 141(4): 472–476, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987223.

Herber:1954:SFB

[Her54] Elmer C. Herber. Spencer Fullerton Baird and the purchase of Alaska. Proceedings of the American Philosophical Society, 98(2): 139–143, April 1954. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143644.

Herring:1955:TUM

[Her55] Pendleton Herring. Toward an understanding of man. Some introductory remarks about the objectives and methods of the social sciences. Proceedings of the American Philosophical Society, 99(2): 39–45, April 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143813.

Herbst:1961:SDH

[Her61] Jurgen Herbst. Social Darwinism and the history of American geography. Proceedings of the American Philosophical Society, 105 (6):538–544, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985164.

Herrnstein:1964:W

[Her64] R. J. Herrnstein. Will. Proceedings of the American Philosophical Society, 108(6):455–458, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985862.

Hertzberg:1979:NAP

[Her79] Hazel Whitman Hertzberg. Nationality, anthropology, and pan-Indianism in the life of Arthur C. Parker (Seneca). Proceedings of the American Philosophical Society, 123(1):47–72, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986470.

Herlan:1982:RPB

[Her82] Ronald W. Herlan. Relief of the poor in Bristol from Late Elizabethan times until the Restoration Era. Proceedings of the American Philosophical Society, 126(3):212–228, June 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986508.

Herschbach:1993:DFM

[Her93] Dudley Herschbach. Dense fluids: Molecules in perpetual collision. Proceedings of the American Philosophical Society, 137(4):532–544, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987072.

Herman:1995:RHF

[Her95] Robert Herman. Robert Hofstadter (5 February 1915–17 November 1990). Proceedings of the American Philosophical Society, 139(3): 276–285, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987194.

Hesburgh:1991:I

[Hes91] Theodore M. Hesburgh. Introduction. Proceedings of the American Philosophical Society, 135(1):1–2, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987143.

Hetherington:1975:AEK

[Het75] Norriss S. Hetherington. Almanacs and the extent of knowledge of the new astronomy in Seventeenth-Century England. Proceedings of the American Philosophical Society, 119(4):275–279, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986289.

Heusser:1966:LPP

[Heu66] Calvin J. Heusser. Late–Pleistocene pollen diagrams from the Province of Llanquihue, Southern Chile. Proceedings of the American Philosophical Society, 110(4):269–305, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985689.

Hewes:1970:HCL

[Hew70] James E. Hewes, Jr. Henry Cabot Lodge and the League of Nations. Proceedings of the American Philosophical Society, 114 (4):245–255, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/985951. See comment [Fos71].

Hayes:1951:IDH

[HH51] Keith J. Hayes and Catherine Hayes. The intellectual development of a home-raised chimpanzee. Proceedings of the American Philosophical Society, 95(2):105–109, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143327.

Hellstrom:1976:CMI

[HH76] Karl Erik Hellström and Ingegerd Hellström. Cell-mediated immunity to tumor antigens as studied in vitro. Proceedings of the American Philosophical Society, 120(2):93–99, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986515.

Harlow:1966:MBR

[HHDA66] H. F. Harlow, M. K. Harlow, R. O. Dodsworth, and G. L. Arling. Maternal behavior of rhesus monkeys deprived of mothering and peer associations in infancy. Proceedings of the American Philosophical Society, 110(1):58–66, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986002.

Hijiya:1980:FWL

[Hij80] James A. Hijiya. Four ways of looking at a philanthropist: a study of Robert Weeks de Forest. Proceedings of the American Philosophical Society, 124(6):404–418, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986242.

Hijiya:1983:AGA

[Hij83] James A. Hijiya. American gravestones and attitudes toward death: a brief history. Proceedings of the American Philosophical Society, 127(5):339–363, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986503.

Hildebrand:1955:SRS

[Hil55] Joel H. Hildebrand. The social responsibility of scientists. Proceedings of the American Philosophical Society, 99(2):46–50, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143814.

Hildebrand:1964:UMP

[Hil64] Joel H. Hildebrand. The use of models in physical science. Proceedings of the American Philosophical Society, 108(5):411–417, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985814.

Hilliard:1969:HMC

[Hil69] Sam Hilliard. Hog meat and cornpone: Food habits in the Ante-Bellum South. Proceedings of the American Philosophical Society, 113(1):1–13, February 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/986013. See comment [Shi70].

Hilliard:1970:RPR

[Hil70] Sam Hilliard. “The Republic of Porkdom”: a reply to article by Royce Shingleton. Proceedings of the American Philosophical Society, 114(5):409–410, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor. org/stable/985808. See comment [Shi70].

Hilgard:1971:PRP

[Hil71] Ernest R. Hilgard. Pain: Its reduction and production under hypnosis. Proceedings of the American Philosophical Society, 115(6): 470–476, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985838.

Hilts:1978:TYA

[Hil78] Victor L. Hilts. Thomas Young’s “autobiographical sketch”. Proceedings of the American Philosophical Society, 122(4):248–260, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986535.

Himmelfarb:1999:JHH

[Him99] Gertrude Himmelfarb. J. H. Hexter (25 May 1910–8 December 1996). Proceedings of the American Philosophical Society, 143(2): 318–324, June 1999. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3181941.

Hindle:1964:HSM

[Hin64] Brooke Hindle. The historiography of science in the Middle Colonies and the Early Middle States. Proceedings of the American Philosophical Society, 108(2):158–161, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985895.

Hirsch:1951:FED

[Hir51] Elisabeth Feist Hirsch. The friendship of Erasmus and Damião De Goes. Proceedings of the American Philosophical Society, 95 (5):556–568, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143242.

Hirsch:1953:FRC

[Hir53] Elisabeth Feist Hirsch. The friendship of the “reform” cardinals in Italy with Damião de Goes. Proceedings of the American Philosophical Society, 97(2):173–183, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143892.

Hitchcock:1952:AOH

[Hit52] Harold Bradford Hitchcock. Airplane observations of homing pigeons. Proceedings of the American Philosophical Society, 96(3): 270–289, June 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143836.

Hiz:1994:ZHO

[Hiż94] Henry Hiż. Zellig S. Harris (23 October 1909–22 May 1992). Proceedings of the American Philosophical Society, 138(4):518–527, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986850.

Heizer:1962:BBC

[HK62] Robert F. Heizer and Harper Kelley. Burins and bladelets in the Cessac Collection from Santa Cruz Island, California. Proceedings of the American Philosophical Society, 106(2):94–105, April 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985375.

Hauptman:1977:DAI

[HK77] Laurence M. Hauptman and Ronald G. Knapp. Dutch–Aboriginal interaction in New Netherland and Formosa: an historical geography of empire. Proceedings of the American Philosophical Society, 121(2):166–182, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986525.

Hoebel:1968:CKP

[Hoe68] E. Adamson Hoebel. The character of Keresan Pueblo law. Proceedings of the American Philosophical Society, 112(3):127–130, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986158.

Hoenigswald:1990:DLG

[Hoe90] Henry M. Hoenigswald. Does language grow on trees? Ancestry, descent, regularity. Proceedings of the American Philosophical Society, 134(1):10–18, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143900.

Hoffman:1950:WPB

[Hof50] Ross J. S. Hoffman. The Wentworth Papers of Burke, Rockingham, and Fitzwilliam. Proceedings of the American Philosophical Society, 94(4):352–356, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143519.

Hoffman:1975:VCP

[Hof75] Lois Wladis Hoffman. The value of children to parents and the decrease in family size. Proceedings of the American Philosophical Society, 119(6):430–438, December 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986377.

Hoffman:1983:ECH

[Hof83] Daniel Hoffman. The exactions of clio: History into poetry in “brotherly love”. Proceedings of the American Philosophical Society, 127(5):330–338, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986502.

Hofstadter:1991:WMF

[Hof91] R. Hofstadter. William Martin Fairbank (February 24, 1917– September 30, 1989). Proceedings of the American Philosophical Society, 135(3):460–465, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986780.

Hoisington:1975:STF

[Hoi75] William A. Hoisington, Jr. In the service of the Third French Republic: Jean-Baptiste Charcot (1867–1936) and the Antarctic. Proceedings of the American Philosophical Society, 119(4):315–324, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986293.

Holmfeld:1970:APA

[Hol70] John D. Holmfeld. From amateurs to professionals in American science: The controversy over the proceedings of an 1853 scientific meeting. Proceedings of the American Philosophical Society, 114 (1):22–36, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985722.

Holldobler:1979:TA

[Höl79] Bert Hölldobler. Territoriality in ants. Proceedings of the American Philosophical Society, 123(4):211–218, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986580.

Holton:1981:ESW

[Hol81] Gerald Holton. Einstein’s search for the “Weltbild”. Proceedings of the American Philosophical Society, 125(1):1–15, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986181.

Holmes:1990:LNC

[Hol90] Frederic L. Holmes. Laboratory notebooks: Can the daily record illuminate the broader picture? Proceedings of the American Philosophical Society, 134(4):349–366, December 1990. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986893.

Homans:1981:JAC

[Hom81] George C. Homans. John Adams and the Constitution of Massachusetts. Proceedings of the American Philosophical Society, 125 (4):286–291, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986331.

Hook:1955:HDP

[Hoo55a] Sidney Hook. Historical determinism and political fiat in Soviet communism. Proceedings of the American Philosophical Society, 99(1):1–10, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143687.

Hoover:1955:IRS

[Hoo55b] Calvin Bryce Hoover. Ideology and reality in the Soviet system: Economic aspects. Proceedings of the American Philosophical Society, 99(1):22–28, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143689.

Hopkins:1961:GWR

[Hop61] Vivian C. Hopkins. The Governor and the Western recluse: De Witt Clinton and Francis Adrian van der Kemp. Proceedings of the American Philosophical Society, 105(3):315–333, June 1961. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985458.

Horsfall:1958:CVM

[Hor58] Frank L. Horsfall, Jr. Can viruses be managed? Proceedings of the American Philosophical Society, 102(5):442–447, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985589.

Hornig:1961:RIT

[Hor61] Edgar Albert Hornig. The religious issue in the Taft–Bryan duel of 1908. Proceedings of the American Philosophical Society, 105 (6):530–537, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985163.

Horgan:1965:AA

[Hor65] Paul Horgan. The abdication of the artist. Proceedings of the American Philosophical Society, 109(5):267–271, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986078.

Horrocks:1995:PMR

[Hor95] Thomas A. Horrocks. “The poor man’s riches, the rich man’s bliss”: Regimen, reform, and the “Journal of Health”, 1829–1833. Proceedings of the American Philosophical Society, 139(2):115–134, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987260.

Hoskin:1989:WHC

[Hos89] Michael Hoskin. William Herschel and the construction of the heavens. Proceedings of the American Philosophical Society, 133(3): 427–433, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987110.

Houston:1950:EMF

[Hou50] William V. Houston. Electric and magnetic forces on superconductors. Proceedings of the American Philosophical Society, 94 (5):453–458, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143612.

Houghton:1966:HHA

[Hou66] Gerald Houghton. Hummingbird hovering, an aero-biophysical phenomenon. Proceedings of the American Philosophical Society, 110(3):165–173, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985985.

Hounshell:1990:IHI

[Hou90] David A. Hounshell. Interpreting the history of industrial research and development: The case of E. I. du Pont de Nemours & Co. Proceedings of the American Philosophical Society, 134(4):387–407, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986895.

Howell:1959:UPS

[How59] F. Clark Howell. Upper Pleistocene stratigraphy and Early Man in the Levant. Proceedings of the American Philosophical Society, 103(1):1–65, February 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985381.

Hutner:1950:SAS

[HPSH50] S. H. Hutner, L. Provasoli, Albert Schatz, and C. P. Haskins. Some approaches to the study of the role of metals in the metabolism of microorganisms. Proceedings of the American Philosophical Society, 94(2):152–170, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143215.

Harrison:1955:KYZ

[HR55] John A. Harrison and Mamiya Rinsō. Kita Yezo zusetsu or a description of the Island of Northern Yezo by Mamiya Rinsō. Proceedings of the American Philosophical Society, 99(2):93–117, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143820.

Hubble:1951:ESC

[Hub51] Edwin P. Hubble. Explorations in space the cosmological program for the Palomar telescopes. Proceedings of the American Philosophical Society, 95(5):461–470, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143231.

Huggins:1969:EOB

[Hug69] Charles Huggins. Epithelial osteogenesis: a biological chain reaction. Proceedings of the American Philosophical Society, 113(6): 458–463, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985828.

Hummel:1951:SBP

[Hum51] Arthur W. Hummel. Some basic principles in Chinese culture. Proceedings of the American Philosophical Society, 95(4):453–456, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143285.

Hummel:1989:CCR

[Hum89] Arthur W. Hummel, Jr. China’s changing relations with the U.S. and U.S.S.R. Proceedings of the American Philosophical Society, 133(1):75–83, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987163.

Hunsaker:1954:HCA

[Hun54] Jerome C. Hunsaker. A half century of aeronautical development. Proceedings of the American Philosophical Society, 98(2):121–130, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143642.

Hurst:1960:LUS

[Hur60] Willard Hurst. The law in United States history. Proceedings of the American Philosophical Society, 104(5):518–526, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985235.

Hutukhtu:1952:NMO

[Hut52] The Dilowa Hutukhtu. The Narobanchin Monastery in Outer Mongolia. Proceedings of the American Philosophical Society, 96(5): 587–598, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143660.

Hutchison:1979:FC

[Hut79] W. Bruce Hutchison. The future of Canada. Proceedings of the American Philosophical Society, 123(5):270–273, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986587.

Huxley:1986:PLD

[Hux86] Andrew F. Huxley. The Penrose Lecture. Discovery and forgetting in science. Proceedings of the American Philosophical Society, 130 (4):475–481, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986930.

Huxtable:1990:AC

[Hux90] Ada Louise Huxtable. Architecture criticism. Proceedings of the American Philosophical Society, 134(4):461–464, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986900.

Hurme:1953:BDE

[HvW53] V. O. Hurme and G. van Wagenen. Basic data on the emergence of deciduous teeth in the monkey (Macaca Mulatta). Proceedings of the American Philosophical Society, 97(3):291–315, June 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143750.

Hurme:1961:BDE

[HvW61] V. O. Hurme and G. van Wagenen. Basic data on the emergence of permanent teeth in the rhesus monkey “(Macaca Mulatta)”. Proceedings of the American Philosophical Society, 105(1):105–140, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985357.

Hyden:1967:BMA

[Hyd67] Holger Hydén. Biochemical and molecular aspects of learning and memory. Proceedings of the American Philosophical Society, 111 (6):326–342, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985729.

Hyslop:1960:APF

[Hys60] Beatrice F. Hyslop. The American press and the French Revolution of 1789. Proceedings of the American Philosophical Society, 104 (1):54–85, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985604.

Harlow:1958:DAR

[HZ58] Harry F. Harlow and Robert R. Zimmermann. The development of affectional responses in infant monkeys. Proceedings of the American Philosophical Society, 102(5):501–509, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985597.

Ingraham:1957:AFR

[Ing57] Mark H. Ingraham. Academic freedom — the role of the professional societies. Proceedings of the American Philosophical Society, 101(5):441–445, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985197.

Ingalls:1965:HFS

[Ing65] Daniel H. H. Ingalls. The heritage of a fallible saint: Annie Besant’s gifts to India. Proceedings of the American Philosophical Society, 109(2):85–88, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985783.

Irvine:1959:IDL

[Irv59] William Irvine. The influence of Darwin on literature. Proceedings of the American Philosophical Society, 103(5):616–628, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985420.

Iselin:1952:GSS

[Ise52] C. O’D. Iselin. The Gulf Stream system. Proceedings of the American Philosophical Society, 96(6):660–662, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143635.

Ives:1951:RLT

[Ive51] Herbert E. Ives. Revisions of the Lorentz transformations. Proceedings of the American Philosophical Society, 95(2):125–131, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143330.

Jacobsen:1963:AMR

[Jac63] Thorkild Jacobsen. Ancient Mesopotamian religion: The central concerns. Proceedings of the American Philosophical Society, 107 (6):473–484, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986107.

Jacobs:1968:SRD

[Jac68] Charles Jacobs. Spanish Renaissance discussion of Musica Ficta. Proceedings of the American Philosophical Society, 112(4):277–298, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985873.

Jackson:1969:TDS

[Jac69] Richard A. Jackson. The “Traité du Sacre” of Jean Golein. Proceedings of the American Philosophical Society, 113(4):305–324, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986155.

Jacobson:1975:ESA

[Jac75] Jerome Jacobson. Early stone age habitation sites in Eastern Malwa. Proceedings of the American Philosophical Society, 119 (4):280–297, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986290.

Jackson:1980:JML

[Jac80] Donald Jackson. Jefferson, Meriwether Lewis, and the reduction of the United States Army. Proceedings of the American Philosophical Society, 124(2):91–96, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986203.

Jackson:1989:RDH

[Jac89] M. W. Jackson. Rousseau’s discourse on heroes and heroism. Proceedings of the American Philosophical Society, 133(3):434–446, September 1989. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987111.

Jakle:1968:ABH

[Jak68] John A. Jakle. The American bison and the human occupance of the Ohio Valley. Proceedings of the American Philosophical Society, 112(4):299–305, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985874.

James:1954:BFO

[Jam54] Alfred P. James. Benjamin Franklin’s Ohio Valley lands. Proceedings of the American Philosophical Society, 98(4):255–265, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143878.

Jamieson:1993:DDI

[Jam93] Kathleen Hall Jamieson. Discourse and the democratic ideal. Proceedings of the American Philosophical Society, 137(3):332–338, September 1993. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986994.

Jameson:1995:GDS

[Jam95] Michael H. Jameson. Georges Daux (21 September 1899–23 December 1988). Proceedings of the American Philosophical Society, 139(1):78–80, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986722.

Jansson:1989:NEH

[Jan89] Maija Jansson. The hat is no expression of honor. Proceedings of the American Philosophical Society, 133(1):26–34, March 1989. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987159.

Jarcho:1972:AIF

[Jar72] Saul Jarcho. The “Animadversiones” of Ippolito Francesco Albertini (1748) on cardiac and precordial diseases which cause dyspnea. Proceedings of the American Philosophical Society, 116(2):163–181, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986172.

Jastrow:1962:RSS

[Jas62] Robert Jastrow. Results of the space science program. Proceedings of the American Philosophical Society, 106(1):82–87, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985214.

Jeffries:1958:SIV

[Jef58] Zay Jeffries. Some industrial vitamins. Proceedings of the American Philosophical Society, 102(4):396–399, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985223.

Jeffries:1964:G

[Jef64] Zay Jeffries. Gold. Proceedings of the American Philosophical Society, 108(5):437–442, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor. org/stable/985818. See comments [Mac65b].

Jeffries:1965:CDF

[Jef65] Zay Jeffries. Comments on Dr. Fritz Machlup’s paper “The fascination of gold”. Proceedings of the American Philosophical Society, 109(2):108–109, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985787. See [Mac65b].

Jefferson:1994:SLJ

[Jef94] Edward G. Jefferson. Samuel Lenher (19 June 1905–17 December 1992). Proceedings of the American Philosophical Society, 138(4): 538–544, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986852.

Jefferson:1995:CHG

[Jef95] Edward G. Jefferson. Crawford Hallock Greenewalt (16 August 1902–27 September 1993). Proceedings of the American Philosophical Society, 139(1):90–104, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986724.

Jefferson:1999:HES

[Jef99] Edward G. Jefferson. Howard Ensign Simmons, Jr. (17 June 1929– 26 April 1997). Proceedings of the American Philosophical Society, 143(4):718–726, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3182000.

Jenkins:1951:GIF

[Jen51] Merle T. Jenkins. Genetic improvement of food plants for increased yield. Proceedings of the American Philosophical Society, 95(1):84– 91, February 1951. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143577.

Jennings:1966:ITS

[Jen66] Francis Jennings. The Indian trade of the Susquehanna Valley. Proceedings of the American Philosophical Society, 110(6):406–424, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985794.

Jennings:1968:GDT

[Jen68] Francis Jennings. Glory, death, and transfiguration: The Susquehannock Indians in the Seventeenth Century. Proceedings of the American Philosophical Society, 112(1):15–53, February 1968. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986100.

Jennings:1971:CEC

[Jen71] Francis Jennings. The Constitutional evolution of the Covenant Chain. Proceedings of the American Philosophical Society, 115(2): 88–96, April 1971. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985850.

Jessup:1951:REF

[Jes51] Philip C. Jessup. Random elements in the formulation and execution of foreign policy. Proceedings of the American Philosophical Society, 95(2):93–99, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143325.

Jettmar:1961:BAK

[Jet61a] Karl Jettmar. Bronze axes from the Karakoram. Results of the 1958 expedition in Azad Kashmir. Proceedings of the American Philosophical Society, 105(1):98–104, February 1961. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985356.

Jettmar:1961:ERD

[Jet61b] Karl Jettmar. Ethnological research in Dardistan 1958. Preliminary report. Proceedings of the American Philosophical Society, 105(1):79–97, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985355.

Jo:1970:PFD

[Jo70] Yung-Hwan Jo. Political futuristics for the divided nations in Asia: a pre-theory. Proceedings of the American Philosophical Society, 114(5):384–388, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985803.

Jodziewicz:1987:DGB

[Jod87] Thomas W. Jodziewicz. The 1699 diary of Gershom Bulkeley of Wethersfield, Connecticut. Proceedings of the American Philosophical Society, 131(4):425–441, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987140.

Johnson:1954:IFR

[Joh54] Victor L. Johnson. Internal financial reform or external taxation: Britain’s fiscal choice, 1763. Proceedings of the American Philosophical Society, 98(1):31–37, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143667.

Johnson:1965:STA

[Joh65] Julian Johnson. The surgical treatment of acquired valvular heart disease. Proceedings of the American Philosophical Society, 109(3): 119–123, June 1965. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986127.

Johnson:1974:BWT

[Joh74] Ludwell H. Johnson III. The business of war: Trading with the enemy in English and Early American law. Proceedings of the American Philosophical Society, 118(5):459–470, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986310.

Johnson:1996:JAS

[Joh96] Howard W. Johnson. Julius Adams Stratton (18 May 1901–22 June 1994). Proceedings of the American Philosophical Society, 140(1):116–121, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987282.

Jones:1955:SIL

[Jon55] Harold Spencer Jones. The secular increase in the length of the day. Proceedings of the American Philosophical Society, 99(4):195– 199, August 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143697.

Jones:1963:ACW

[Jon63a] Robert Huhn Jones. The American Civil War in the British Sessional Papers: Catalogue and commentary. Proceedings of the American Philosophical Society, 107(5):415–426, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985646.

Jones:1963:NRC

[Jon63b] Tom B. Jones. A numismatic riddle: The so-called Greek imperials. Proceedings of the American Philosophical Society, 107(4):308–347, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985674.

Jones:1967:CD

[Jon67] Harry W. Jones. Civil disobedience. Proceedings of the American Philosophical Society, 111(4):195–198, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986006.

Jones:1999:IF

[Jon99a] Christopher Jones. Interrupted funerals. Proceedings of the American Philosophical Society, 143(4):588–600, December 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181990.

Jonsson:1999:SCE

[Jon99b] Inge Jonsson. Swedish culture in a European context. Proceedings of the American Philosophical Society, 143(1):148–156, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181982.

Jordan:1961:DPE

[Jor61] W. K. Jordan. The development of philanthropy in England in the Early Modern Era. Proceedings of the American Philosophical Society, 105(2):145, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985625.

Jordan:1971:MMM

[Jor71] James E. Jordan. Matrimony and Machiavellianism: The marriage of Prince Napoleon. Proceedings of the American Philosophical Society, 115(4):317–323, August 1971. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986092.

Julien:1987:UTC

[Jul87] Catherine J. Julien. The Uru tribute category; ethnic boundaries and empire in the Andes. Proceedings of the American Philosophical Society, 131(1):53–91, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987116.

Juricek:1966:AUW

[Jur66] John T. Juricek. American usage of the word “frontier” from Colonial times to Frederick Jackson Turner. Proceedings of the American Philosophical Society, 110(1):10–34, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985999.

Kaegi:1964:EJA

[Kae64] Walter Emil Kaegi. The Emperor Julian’s assessment of the significance and function of history. Proceedings of the American Philosophical Society, 108(1):29–38, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985994.

Kahn:1955:FGJ

[Kah55a] Robert L. Kahn. Franklin, Grimm, and J. H. Landolt. Proceedings of the American Philosophical Society, 99(6):401–404, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143922.

Kahn:1955:TFR

[Kah55b] Robert L. Kahn. Three Franklin–Raspe letters. Proceedings of the American Philosophical Society, 99(6):397–400, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143921.

Kahn:1956:ACL

[Kah56] Robert L. Kahn. Addendum concerning a lost Franklin–Raspe letter. Proceedings of the American Philosophical Society, 100(3):279, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143774.

Kahn:1958:GFB

[Kah58] Robert L. Kahn. George Forster and Benjamin Franklin. Proceedings of the American Philosophical Society, 102(1):1–6, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985300.

Kaiser:1970:JQA

[Kai70] Leo M. Kaiser. John Quincy Adams and his translation of Juvenal 13. Proceedings of the American Philosophical Society, 114 (4):272–293, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985953.

Kaminsky:1971:PFS

[Kam71] Howard Kaminsky. The politics of France’s subtraction of obedience from Pope Benedict XIII, 27 July, 1398. Proceedings of the American Philosophical Society, 115(5):366–397, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985769.

Kamen:1978:NHA

[Kam78] Martin D. Kamen. The natural history of an ancient protein. Proceedings of the American Philosophical Society, 122(4):214–221, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986533.

Kamen:1986:CEE

[Kam86] Martin D. Kamen. On creativity of eye and ear: a commentary on the career of T. W. Engelmann. Proceedings of the American Philosophical Society, 130(2):232–246, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987182.

Kammen:1987:VLF

[Kam87] Michael Kammen. “A vehicle of life”: The Founders’ intentions and American perceptions of their living Constitution. Proceedings of the American Philosophical Society, 131(3):325–340, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987027.

Kamen:1994:RFH

[Kam94] Martin D. Kamen. Reflections on the first half-century of long-lived radioactive carbon (^{14}c). Proceedings of the American Philosophical Society, 138(1):48–60, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986705.

Kamen:1995:LLI

[Kam95] Martin D. Kamen. Liebling’s Law (“ILL”). Proceedings of the American Philosophical Society, 139(4):358–367, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987235.

Kantorowicz:1961:GU

[Kan61] Ernst H. Kantorowicz. Gods in uniform. Proceedings of the American Philosophical Society, 105(4):368–393, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985496.

Kantrow:1989:LEG

[Kan89] Louise Kantrow. Life expectancy of the gentry in Eighteenth and Nineteenth-Century Philadelphia. Proceedings of the American Philosophical Society, 133(2):312–327, June 1989. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987057.

Kaplan:1998:CCC

[Kap98] Herbert H. Kaplan. Commerce, consumption, and culture: Hope & Co. and Baring Brothers & Co. and Russia. Proceedings of the American Philosophical Society, 142(2):258–262, June 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152302.

Karan:1960:LUR

[Kar60] Pradyumna Prasad Karan. A land use reconnaissance in Nepal by aero-field techniques and photography. Proceedings of the American Philosophical Society, 104(2):172–187, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985657.

Karl:1985:PSS

[Kar85] Barry D. Karl. Philanthropy and the social sciences. Proceedings of the American Philosophical Society, 129(1):14–19, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986976.

Karle:1999:NDP

[Kar99] Isabella L. Karle. Natural and designer peptides: Folding, aggregation, ion transport. Proceedings of the American Philosophical Society, 143(1):28–41, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181971.

Katz:1970:AEU

[Kat70] Milton Katz. Aspects of the “embattled university”. Proceedings of the American Philosophical Society, 114(5):343–346, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985799.

Katz:1985:I

[Kat85] Stanley N. Katz. Introduction. Proceedings of the American Philosophical Society, 129(1):1–2, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986973.

Katz:1997:OHH

[Kat97] Stanley N. Katz. Official history: The Holmes Devise history of the Supreme Court. Proceedings of the American Philosophical Society, 141(3):297–304, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987204.

Kauzmann:1995:CPS

[Kau95] Walter Kauzmann. Charles Phelps Smyth (10 February 1895–18 March 1990). Proceedings of the American Philosophical Society, 139(3):306–309, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987198.

Kay:1968:NAC

[Kay68] Marshall Kay. North Atlantic continental drift. Proceedings of the American Philosophical Society, 112(5):321–324, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986052.

Kaysen:1998:MBM

[Kay98] Carl Kaysen. McGeorge Bundy (30 March 1919–16 September 1996). Proceedings of the American Philosophical Society, 142(3):

458–465, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152251.

Koo:1980:NSY

[KB80] John H. Koo and Gust Bartman. The number system in Yupik Eskimo. Proceedings of the American Philosophical Society, 124 (1):48–51, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986657.

Kennan:1982:VAE

[KB82] George F. Kennan and Whitfield J. Bell, Jr. Views from another East Berlin. Proceedings of the American Philosophical Society, 126(6):472–480, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986253.

Kearns:1976:AII

[Kea76] Kevin C. Kearns. The Aran Islands: an imperiled Irish outpost. Proceedings of the American Philosophical Society, 120(6):421–438, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986597.

Keefer:1963:MCS

[Kee63] Chester Scott Keefer. Medical care services today and tomorrow. Proceedings of the American Philosophical Society, 107(4):294–298, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985672.

Keeney:1968:HAS

[Kee68] Barnaby C. Keeney. The humanities in American society. Proceedings of the American Philosophical Society, 112(1):1–4, February 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986097.

Keeling:1970:CDF

[Kee70] Charles D. Keeling. Is carbon dioxide from fossil fuel changing Man’s environment? Proceedings of the American Philosophical Society, 114(1):10–17, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985720.

Kelly:1957:NRD

[Kel57] Mervin J. Kelly. The nation’s research and development — their deficiencies and means for correction. Proceedings of the American Philosophical Society, 101(4):386–391, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985406.

Kelley:1992:JJS

[Kel92] Allen C. Kelley. Joseph J. Spengler (November 19, 1902–January 2, 1991). Proceedings of the American Philosophical Society, 136 (1):142–147, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986806.

Kennedy:1952:EES

[Ken52] John L. Kennedy. An evaluation of extra-sensory perception. Proceedings of the American Philosophical Society, 96(5):513–518, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143651.

Kendall:1953:HAC

[Ken53] Edward C. Kendall. Hormones of the adrenal cortex in health and disease. Proceedings of the American Philosophical Society, 97(1):8–11, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143725.

Kennan:1955:CIR

[Ken55] George F. Kennan. Comments on ideology and reality in the Soviet system. Proceedings of the American Philosophical Society, 99(1):29–30, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143690.

Keniston:1958:PNB

[Ken58] Hayward Keniston. Peace negotiations between Charles V and Francis I (1537–1538). Proceedings of the American Philosophical Society, 102(2):142–147, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985278.

Keniston:1960:GHE

[Ken60] Hayward Keniston. The goals of higher education. Proceedings of the American Philosophical Society, 104(6):565–569, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985532.

Kennan:1971:HEP

[Ken71] George F. Kennan. The historiography of the early political career of Stalin. Proceedings of the American Philosophical Society, 115(3):165–169, June 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985975.

Kennedy:1991:NSG

[Ken91] Paul Kennedy. On the “natural size” of great powers. Proceedings of the American Philosophical Society, 135(4):485–489, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986810.

Kenny:1993:MM

[Ken93] Anthony Kenny. The metamorphosis of metaphysics. Proceedings of the American Philosophical Society, 137(4):669–679, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987081.

Kennedy:1995:RBF

[Ken95] George A. Kennedy. T. Robert S. Broughton (17 February 1900– 17 September 1993). Proceedings of the American Philosophical Society, 139(1):74–76, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986721.

Keohane:1998:MEM

[Keo98] Nannerl O. Keohane. Moral education in the modern university. Proceedings of the American Philosophical Society, 142(2):244–257, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3152301.

Kerr:1954:PHS

[Ker54] Chester Kerr. Publishing historical sources: a prejudiced view of the problem of finance. Proceedings of the American Philosophical Society, 98(4):273–278, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143881.

Kernek:1982:WWN

[Ker82]  Sterling J. Kernek.  Woodrow Wilson and national self-

determination along Italy’s frontier: a study of the manipulation of principles in the pursuit of political interests. Proceedings of the American Philosophical Society, 126(4):243–300, August 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986175.

Keynes:1979:JLD

[Key79] R. D. Keynes. Jayne Lecture. Darwin and the “Beagle”. Proceedings of the American Philosophical Society, 123(5):323–335, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986593. See

erratum [Ano79m].

Keynes:1992:BTV

[Key92] Richard D. Keynes. Biological transistors: The voltage-gating of ion currents in neurons. Proceedings of the American Philosophical Society, 136(1):51–60, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986797.

Keynes:1999:JZY

[Key99] Richard Keynes. John Zachary Young (18 March 1907–4 July 1997). Proceedings of the American Philosophical Society, 143(4): 728–732, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3182001.

Kazin:1992:LMO

[KH92] Alfred Kazin and Ada Louise Huxtable. Lewis Mumford (October 19, 1895–January 26, 1990). Proceedings of the American Philosophical Society, 136(2):304–309, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987176.

Kohler:1952:ICC

[KHO52] Wolfgang Köhler, Richard Held, and Donald Neil O’Connell. An investigation of cortical currents. Proceedings of the American Philosophical Society, 96(3):290–330, June 1952. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143837.

Kidder:1959:DSG

[Kid59] Alfred V. Kidder. The diary of Sylvanus G. Morley. Proceedings of the American Philosophical Society, 103(6):778–782, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985392.

Kidder:1961:AIK

[Kid61] Alfred V. Kidder. Archaeological investigations at Kaminaljuyu, Guatemala. Proceedings of the American Philosophical Society, 105 (6):559–570, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985166.

Kieval:1980:LRV

[Kie80] Hillel J. Kieval. Legality and resistance in Vichy France: The rescue of Jewish children. Proceedings of the American Philosophical Society, 124(5):339–366, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986573.

Kilmer:1971:DAM

[Kil71] Anne Draffkorn Kilmer. The discovery of an Ancient Mesopotamian theory of music. Proceedings of the American Philosophical Society, 115(2):131–149, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985853.

Kilbourne:1996:HS

[Kil96] Edwin D. Kilbourne. Humor in science. Proceedings of the American Philosophical Society, 140(3):338–349, September 1996. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987312.

Kim:1962:JRK

[Kim62] C. I. Eugene Kim. Japanese rule in Korea (1905–1910): a case study. Proceedings of the American Philosophical Society, 106 (1):53–59, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985211.

Kimball:1977:CPU

[Kim77] Elisabeth G. Kimball. Commissions of the Peace for urban jurisdictions in England, 1327–1485. Proceedings of the American Philosophical Society, 121(6):448–474, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986393.

Kim:1993:WSP

[Kim93] Sung Hui Kim. “We (the supermajority of) the people”: The development of a rationale for written higher law in North American Constitutions. Proceedings of the American Philosophical Society, 137(3):364–389, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986998.

King:1954:CSG

[Kin54] John A. King. Closed social groups among domestic dogs. Proceedings of the American Philosophical Society, 98(5):327–336, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143912.

King:1976:LEP

[Kin76] John Owen King. Labors of the estranged personality: Josiah Royce on “the case of John Bunyan”. Proceedings of the American Philosophical Society, 120(1):46–58, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986351.

Kinloch:1978:CRR

[Kin78] Graham C. Kinloch. Changing Rhodesian race relations: a study of demographic and economic factors. Proceedings of the American Philosophical Society, 122(1):18–24, February 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986260.

Kirk:1956:USF

[Kir56] Grayson Kirk. The United States in the family of nations. Proceedings of the American Philosophical Society, 100(4):289–295, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143754.

Kirkland:1966:SGS

[Kir66] Edward Chase Kirkland. Stephen Girard: Suggestions for a biography. Proceedings of the American Philosophical Society, 110 (6):386–391, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985792.

Kitzinger:1973:WMF

[Kit73] Ernst Kitzinger. World map and fortune’s wheel: a medieval mosaic floor in turin. Proceedings of the American Philosophical Society, 117(5):344–373, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986605.

Kitzinger:1995:KWM

[Kit95] Ernst Kitzinger. Kurt Weitzmann (7 March 1904–7 June 1993). Proceedings of the American Philosophical Society, 139(2):204–209, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987269.

Kitagawa:1996:PMH

[Kit96] Evelyn M. Kitagawa. Philip M. Hauser (27 September 1909–13 December 1994). Proceedings of the American Philosophical Society, 140(2):240–244, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987329.

Krautheimer:1952:LFM

[KJF52] Richard Krautheimer, Enrico Josi, and Wolfgang Frankl. S. Lorenzo Fuori Le Mura in Rome: Excavations and observations. Proceedings of the American Philosophical Society, 96(1):1–26, February 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143738.

Katchalski-Katzir:1991:SSP

[KK91] Ephraim Katchalski-Katzir. A scientist as state president: a personal reflection. Proceedings of the American Philosophical Society, 135(3):346–367, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986773.

Klein:1971:RWM

[Kle71] Lawrence R. Klein. The role of war in the maintenance of American economic prosperity. Proceedings of the American Philosophical Society, 115(6):507–516, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985844.

Kleiner:1982:EYV

[Kle82] Edgar F. Kleiner. Eleven year vegetational comparison in an arid grassland. Proceedings of the American Philosophical Society, 126 (6):520–526, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986255.

Klepp:1989:DEP

[Kle89a] Susan E. Klepp. Demography in early Philadelphia, 1690–1860. Proceedings of the American Philosophical Society, 133(2):85–111, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987041.

Klepp:1989:FKQ

[Kle89b] Susan E. Klepp. Fragmented knowledge: Questions in regional demographic history. Proceedings of the American Philosophical Society, 133(2):223–233, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987053.

Kline:1959:CP

[Kli59] Nathan S. Kline. The challenge of the psychopharmaceuticals. Proceedings of the American Philosophical Society, 103(3):455–462, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985476.

Kligerman:1973:DIT

[Kli73] Jack Kligerman. “Dress” or “incarnation” of thought: Nineteenth-Century American attitudes towards language and style. Proceedings of the American Philosophical Society, 117(1):51–58, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985947.

Kluckhohn:1958:SSV

[Klu58] Clyde Kluckhohn. The scientific study of values and contemporary civilization. Proceedings of the American Philosophical Society, 102 (5):469–476, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985593.

Kluckhohn:1959:LNP

[Klu59] Clyde Kluckhohn. The Library’s new program in American Indian linguistics and ethnohistory. Proceedings of the American Philosophical Society, 103(6):768–769, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985389.

Knight:1979:CGC

[Kni79] Carol Lynn H. Knight. A certain great commoner: The political image of William Pitt, First Earl of Chatham, in the Colonial press. Proceedings of the American Philosophical Society, 123(2):131–142, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986236.

Knollenberg:1963:JDV

[Kno63] Bernhard Knollenberg. John Dickinson vs. John Adams: 1774– 1776. Proceedings of the American Philosophical Society, 107(2): 138–144, April 1963. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/

985444. See erratum [Ano63m].

Knoll:1976:LLP

[Kno76] Paul W. Knoll. Learning in Late Piast Poland. Proceedings of the American Philosophical Society, 120(2):136–157, April 1976. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986518.

Knoll:1991:ESB

[Kno91] Andrew H. Knoll. Elso Sterrenberg Barghoorn, Jr. (June 15, 1915– January 22, 1984). Proceedings of the American Philosophical Society, 135(1):86–90, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987152.

Knopoff:1993:SOD

[Kno93] Leon Knopoff. Self-organization and the development of pattern: Implications for earthquake prediction. Proceedings of the American Philosophical Society, 137(3):339–349, September 1993. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986995.

Kohlstedt:1974:GMN

[Koh74] Sally Gregory Kohlstedt. The geologists’ model for national science, 1840–1847. Proceedings of the American Philosophical Society, 118(2):179–195, April 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986665.

Kohler:1985:PS

[Koh85] Robert E. Kohler. Philanthropy and science. Proceedings of the American Philosophical Society, 129(1):9–13, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986975.

Kolff:1965:AHI

[Kol65] Willem J. Kolff. Artificial heart inside the chest. Proceedings of the American Philosophical Society, 109(3):117–118, June 1965. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986126.

Koop:1990:DEL

[Koo90] C. Everett Koop. Decisions at the end of life. Proceedings of the American Philosophical Society, 134(3):230–242, September 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986967.

Kornberg:1997:SOS

[Kor97] Arthur Kornberg. Severo Ochoa (24 September 1905–1 November 1993). Proceedings of the American Philosophical Society, 141(4): 478–491, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987224.

Kotschnig:1951:ESF

[Kot51] Walter M. Kotschnig. Economic and social foundations of the United Nations. Proceedings of the American Philosophical Society, 95(5):512–518, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143235.

Kottler:1980:DWO

[Kot80] Malcolm Jay Kottler. Darwin, Wallace, and the origin of sexual dimorphism. Proceedings of the American Philosophical Society, 124(3):203–226, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986370.

Koyre:1953:EM

[Koy53] Alexandre Koyré. An experiment in measurement. Proceedings of the American Philosophical Society, 97(2):222–237, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143896.

Kramer:1950:IDN

[Kra50] Samuel Noah Kramer. “Inanna’s descent to the Nether World” continued. Proceedings of the American Philosophical Society, 94 (4):361–363, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143521.

Kramrisch:1958:WII

[Kra58] Stella Kramrisch. Wall and image in Indian art. Proceedings of the American Philosophical Society, 102(1):7–13, February 1958. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985301.

Kramer:1963:CSH

[Kra63] Samuel Noah Kramer. Cuneiform studies and the history of literature: The Sumerian sacred marriage texts. Proceedings of the American Philosophical Society, 107(6):485–527, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986108.

Kramer:1973:FYR

[Kra73] Paul J. Kramer. Forty years of research on plant water relations. Proceedings of the American Philosophical Society, 117(5):381–387, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986607.

Kramer:1980:SLB

[Kra80] Samuel Noah Kramer. Sumerian literature and the British Museum: The promise of the future. Proceedings of the American Philosophical Society, 124(4):295–312, August 1980. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986431.

Krebs:1971:GS

[Kre71] Hans A. Krebs. The goals of science. Proceedings of the American Philosophical Society, 115(1):1–3, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985916.

Kahn:1955:SUR

[KRF55] Robert L. Kahn, Rudolf Erich Raspe, and Benjamin Franklin. Some unpublished Raspe–Franklin letters. Proceedings of the American Philosophical Society, 99(3):127–132, June 1955. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143778.

Kristeller:1976:SMR

[Kri76] Paul Oskar Kristeller. The search for Medieval and Renaissance manuscripts. Proceedings of the American Philosophical Society, 120(5):307–310, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986264.

Krieger:1984:OG

[Kri84] Leonard Krieger. The one that got away. Proceedings of the American Philosophical Society, 128(3):248–251, September 1984. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986885.

Kroeber:1951:WCD

[Kro51] Alfred L. Kroeber. Is Western civilization disintegrating or reconstituting? Proceedings of the American Philosophical Society, 95(2):100–104, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143326.

Kroeber:1957:APS

[Kro57] A. L. Kroeber. The American Philosophical Society. For the Fifth International Congress of Anthropological and Ethnological Sciences, Philadelphia, September 4, 1956. Proceedings of the American Philosophical Society, 101(1):1–3, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985137.

Krout:1958:AHP

[Kro58] John A. Krout. Alexander Hamilton’s place in the founding of the nation. Proceedings of the American Philosophical Society, 102 (2):124–128, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985275.

Kroeber:1963:NHK

[Kro63] Karl Kroeber. The new humanism of Keats’s odes. Proceedings of the American Philosophical Society, 107(3):263–271, June 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985583.

Kroeber:1965:MN

[Kro65] Clifton B. Kroeber. The Mohave as nationalist, 1859–1874. Proceedings of the American Philosophical Society, 109(3):173–180, June 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986130.

Kelley:1984:WWP

[KS84] Donald R. Kelley and Bonnie G. Smith. What was property? Legal dimensions of the social question in France (1789–1848). Proceedings of the American Philosophical Society, 128(3):200–230, September 1984. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986881.

Kosslyn:1991:VMI

[KS91] Stephen M. Kosslyn and Lisa M. Shin. Visual mental images in the brain. Proceedings of the American Philosophical Society, 135 (4):524–532, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986814.

Kublin:1959:TFS

[Kub59] Hyman Kublin. Takano Fusataro: a study in early Japanese tradesunionism. Proceedings of the American Philosophical Society, 103 (4):571–583, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985560.

Kunze:1986:UWM

[Kun86] Bonnelyn Young Kunze. An unpublished work by Margaret Fell. Proceedings of the American Philosophical Society, 130(4):424–452, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986928.

Kuttner:1999:CGE

[Kut99] Robert Kuttner. Can the global economy be a mixed economy? Proceedings of the American Philosophical Society, 143(2):168–177, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181931.

Kuznets:1958:LSG

[Kuz58] Simon Kuznets. Long swings in the growth of population and in related economic variables. Proceedings of the American Philosophical Society, 102(1):25–52, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985303.

Kuznets:1967:PEG

[Kuz67] Simon Kuznets. Population and economic growth. Proceedings of the American Philosophical Society, 111(3):170–193, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985714.

Kuznets:1974:RUD

[Kuz74] Simon Kuznets. Rural-urban differences in fertility: an international comparison. Proceedings of the American Philosophical Society, 118(1):1–29, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986434.

Kuznets:1975:FDB

[Kuz75] Simon Kuznets. Fertility differentials between less developed and developed regions: Components and implications. Proceedings of the American Philosophical Society, 119(5):363–396, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986679.

Kafatos:1997:CMW

[KWB97] F. C. Kafatos, E. O. Wilson, and Daniel Branton. Carroll Milton Williams (2 December 1916–11 October 1991). Proceedings of the American Philosophical Society, 141(1):116–121, March 1997. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987257.

Labaree:1956:RMF

[Lab56] Leonard W. Labaree. The return of Mr. Franklin. Proceedings of the American Philosophical Society, 100(4):342–345, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143762.

Labaree:1964:BFB

[Lab64] Leonard W. Labaree. Benjamin Franklin’s British friendships. Proceedings of the American Philosophical Society, 108(5):423–428, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985816.

Ladner:1975:JTL

[Lad75] Gerhart B. Ladner. Justinian’s theory of law and the renewal ideology of the “Leges Barbarorum”. Proceedings of the American Philosophical Society, 119(3):191–200, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986669.

Laffey:1975:MIF

[Laf75] John F. Laffey. Municipal imperialism in France: The Lyon Chamber of Commerce, 1900–1914. Proceedings of the American Philosophical Society, 119(1):8–23, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986646.

Laird:1998:DPS

[Lai98] Melvin R. Laird. David Packard (7 September 1912–26 March 1996). Proceedings of the American Philosophical Society, 142(1): 146–152, March 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152272.

Lambrecht:1980:PTF

[Lam80] Frank L. Lambrecht. Palaeoecology of tsetse flies and sleeping sickness in Africa. Proceedings of the American Philosophical Society, 124(5):367–385, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986574.

Lancaster:1950:LPN

[Lan50] H. Carrington Lancaster. De Lesseps on Panama, Nicaragua, and Tehuantepec. Proceedings of the American Philosophical Society, 94(3):258–259, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143559.

Langer:1955:RRC

[Lan55] William L. Langer. The Russian Research Center at Harvard University. Proceedings of the American Philosophical Society, 99(1): 34–35, January 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143692.

Lange:1959:DJE

[Lan59] Erwin F. Lange. Dr. John Evans, U.S. geologist to the Oregon and Washington Territories. Proceedings of the American Philosophical Society, 103(3):476–484, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985479.

Lang:1972:WRM

[Lan72] Mabel L. Lang. War and the rape-motif, or why did Cambyses invade Egypt? Proceedings of the American Philosophical Society, 116(5):410–414, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986072.

Landon:1976:LGM

[Lan76] Michael Landon. “The learned Glynne and Maynard” — two characters “dashed out” of Samuel Butler’s “Hudibras”. Proceedings of the American Philosophical Society, 120(3):205–210, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986559.

Lang:1984:HOH

[Lan84] Mabel L. Lang. Herodotus: Oral history with a difference. Proceedings of the American Philosophical Society, 128(2):93–103, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986223.

Landsman:1989:ENO

[Lan89] Ned C. Landsman. Ethnicity and national origin among British settlers in the Philadelphia Region: Pennsylvania immigration in the wake of “voyagers to the West”. Proceedings of the American Philosophical Society, 133(2):170–174, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987047.

Lang:1990:IEA

[Lan90] Mabel L. Lang. Illegal execution in Ancient Athens. Proceedings of the American Philosophical Society, 134(1):24–29, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143902.

Lane:1998:GBJ

[Lan98a] Saunders Mac Lane. Garrett Birkhoff (10 January 1911–22 November 1996). Proceedings of the American Philosophical Society, 142 (4):646–649, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152287.

Langfitt:1998:RGDa

[Lan98b] Thomas W. Langfitt. Robert G. Dunlop (2 July 1909–20 September 1995). Proceedings of the American Philosophical Society, 142 (2):284–286, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152305.

Langfitt:1998:RGDb

[Lan98c] Thomas W. Langfitt. Robert G. Dunlop (2 July 1909–20 September 1995). Proceedings of the American Philosophical Society, 142(3): 512–514, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/3152257. See erratum [Ano98i].

Landau:1999:SEG

[Lan99a] Ralph Landau. From science to economic growth. Proceedings of the American Philosophical Society, 143(2):194–207, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181934.

Langfitt:1999:VHC

[Lan99b] T. W. Langfitt. Values in health care and health professions education. Proceedings of the American Philosophical Society, 143(3): 400–406, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181952.

Laporte:1983:STM

[Lap83] Léo F. Laporte. Simpson’s “tempo and mode in evolution” revisited. Proceedings of the American Philosophical Society, 127(6): 365–417, December 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986315.

Larner:1993:NAH

[Lar93] John P. Larner. North American hero? Christopher Columbus 1702–2002. Proceedings of the American Philosophical Society, 137(1):46–63, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986944.

Last:1975:KAW

[Las75] Joseph S. Last. Kourion: The ancient water supply. Proceedings of the American Philosophical Society, 119(1):39–72, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986649.

Lasker:1991:APP

[Las91] Morris E. Lasker. American prisons and prisoners in 1990. Proceedings of the American Philosophical Society, 135(1):30–40, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987147.

Lattimore:1962:WDD

[Lat62] Richmond Lattimore. Why the devil is the devil. Proceedings of the American Philosophical Society, 106(5):427–429, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985487.

Lattimore:1970:PML

[Lat70] Richmond Lattimore. Penrose Memorial Lecture. Man and God in Homer. Proceedings of the American Philosophical Society, 114 (6):411–422, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986143.

Latzko:1995:KBJ

[Lat95] David A. Latzko. Kenneth E. Boulding (18 January 1910–19 March 1993). Proceedings of the American Philosophical Society, 139(1): 68–72, March 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986720.

Lavin:1975:IAI

[Lav75] Irving Lavin. On illusion and allusion in Italian Sixteenth-Century portrait busts. Proceedings of the American Philosophical Society, 119(5):353–362, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986678.

Lawrence:1972:EHC

[Law72] Philip J. Lawrence. Edward Hitchcock: The Christian geologist. Proceedings of the American Philosophical Society, 116(1):21–34, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985704.

Lawson:1992:CWP

[Law92] Karol Ann Peard Lawson. Charles Willson Peale’s “John Dickinson”: an American landscape as political allegory. Proceedings of the American Philosophical Society, 136(4):455–486, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986757.

Labaree:1957:PBF

[LB57] Leonard W. Labaree and Whitfield J. Bell, Jr. The papers of Benjamin Franklin: a progress report. Proceedings of the American Philosophical Society, 101(6):532–534, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985521.

Leopold:1979:BLA

[LB79] Luna B. Leopold and William B. Bull. Base level, aggradation, and grade. Proceedings of the American Philosophical Society, 123(3): 168–202, June 1979. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986220.

Leopold:1995:TBN

[LB95] Luna B. Leopold and Arthur Baker. Thomas Brennan Nolan (21 May 1901–1 August 1992). Proceedings of the American Philosophical Society, 139(3):294–297, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987196.

Lukacs:1986:PIP

[LBWD86] John R. Lukacs, Rebecca K. Bogorad, Subhash R. Walimbe, and Donald C. Dunbar. Paleopathology at Inamgaon: a post-Harappan agrarian village in Western India. Proceedings of the American Philosophical Society, 130(3):289–311, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986828.

Lewis:1962:EIC

[LC62] John F. Lewis, Jr. and Eleanor Este Campion. An experiment in intellectual cooperation. Proceedings of the American Philosophical Society, 106(4):301–303, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985263.

Lechtman:1994:CSS

[Lec94] Heather Lechtman. Cyril Stanley Smith (4 October 1903–18 August 1992). Proceedings of the American Philosophical Society, 138 (4):570–573, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986856.

Lederman:1992:CC

[Led92] Leon M. Lederman. To change a city. Proceedings of the American Philosophical Society, 136(3):326–332, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986904.

Ledeen:1994:RAM

[Led94] Michael A. Ledeen. Raymond Aron (14 March 1905–17 October 1983). Proceedings of the American Philosophical Society, 138(1): 150–157, March 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986709.

Lee:1981:OFI

[Lee81] Rensselaer W. Lee. Observations on the first illustrations of Tasso’s “Gerusalemme Liberata”. Proceedings of the American Philosophical Society, 125(5):329–356, October 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986197.

Lefanu:1971:VNG

[Lef71] William R. Lefanu. The versatile Nehemiah grew. Proceedings of the American Philosophical Society, 115(6):502–506, December 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985843.

Lehmann:1959:KSS

[Leh59] Donald L. Lehmann. Karyolysus sonomae n. sp., a blood parasite from the California yellow-legged frog, Rana boyli boyli. Proceedings of the American Philosophical Society, 103(4):545–547, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985556.

Lehmann:1960:HAS

[Leh60] Donald L. Lehmann. Haemogregarina Aurorae N. SP. from Rana A. Aurora. Proceedings of the American Philosophical Society, 104 (2):202–204, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985659.

Lehr:1971:PTO

[Leh71] Herndon B. Lehr. The preservation of tissues and organs. Proceedings of the American Philosophical Society, 115(4):271–275, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986088.

Leighton:1990:TDA

[Lei90] Lauren G. Leighton. Translation as a derived art. Proceedings of the American Philosophical Society, 134(4):445–454, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986898.

Leichty:1998:IJG

[Lei98] Erle Leichty. Ignace J. Gelb (14 October 1907–22 December 1985). Proceedings of the American Philosophical Society, 142(4):668–670, December 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152290.

Lemberg:1966:H

[Lem66] Rudolf Lemberg. Hemoproteins. Proceedings of the American Philosophical Society, 110(6):434–439, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985796.

Lenzen:1965:CCP

[Len65] Victor F. Lenzen. The contributions of Charles S. Peirce to metrology. Proceedings of the American Philosophical Society, 109(1):29– 46, February 1965. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985776.

Lenher:1966:LPU

[Len66] Samuel Lenher. … for promoting useful knowledge. Proceedings of the American Philosophical Society, 110(5):326–331, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986021.

Leontief:1953:DPF

[Leo53] Wassily Leontief. Domestic production and foreign trade; the American capital position re-examined. Proceedings of the American Philosophical Society, 97(4):332–349, September 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3149288.

Leopold:1994:RMA

[Leo94] Luna B. Leopold. River morphology as an analog to Darwin’s theory of natural selection. Proceedings of the American Philosophical Society, 138(1):31–47, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986704.

Lerner:1959:CNS

[Ler59] I. Michael Lerner. The concept of natural selection: a centennial view. Proceedings of the American Philosophical Society, 103(2):

173–182, April 1959. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985147.

Leranbaum:1977:MOE

[Ler77] Miriam Leranbaum. “Mistresses of orthodoxy”: Education in the lives and writings of Late Eighteenth-Century English women writers. Proceedings of the American Philosophical Society, 121(4):281– 301, August 1977. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986417.

Levin:1962:ST

[Lev62] Harry Levin. Shakespeare today. Proceedings of the American Philosophical Society, 106(5):422–426, October 1962. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985486.

Levy:1988:TNL

[Lev88] Kenneth Levy. Is tonality a natural law? Proceedings of the American Philosophical Society, 132(3):299–303, September 1988. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143856.

Lewis:1954:OCR

[Lew54] Naphtali Lewis. On official corruption in Roman Egypt: The Edict of Vergilius Capito. Proceedings of the American Philosophical Society, 98(2):153–158, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143647.

Lewis:1956:FCN

[Lew56a] Oscar Lewis. The festival cycle in a North Indian Jāt. village. Proceedings of the American Philosophical Society, 100(3):168–196, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143771.

Lewis:1956:AMF

[Lew56b] Wilmarth S. Lewis. Acceptance of Mr. Franklin on behalf of Yale University. Proceedings of the American Philosophical Society, 100 (4):352–353, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143764.

Lewis:1963:EPC

[Lew63] Wilmarth Sheldon Lewis. Editing private correspondence. Proceedings of the American Philosophical Society, 107(4):289–293, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985671.

Lewis:1968:WTS

[Lew68] W. Arthur Lewis. World trade since the war. Proceedings of the American Philosophical Society, 112(6):362–366, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985935.

Lewis:1980:TA

[Lew80] Bernard Lewis. Translation from Arabic. Proceedings of the American Philosophical Society, 124(1):41–47, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986656.

Lewalski:1994:ORC

[Lew94] Barbara Kiefer Lewalski. Old Renaissance canons, new women’s texts: Some Jacobean examples. Proceedings of the American Philosophical Society, 138(3):397–406, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986746.

Lewis:1997:CCF

[Lew97] Bernard Lewis. Claude Cahen (26 February 1909–18 November 1991). Proceedings of the American Philosophical Society, 141(2): 218–220, June 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987303.

Lewis:1999:DH

[Lew99] Bernard Lewis. In defense of history. Proceedings of the American Philosophical Society, 143(4):573–587, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181989.

Lund:1952:HLN

[LF52] Henrik Lund and Svend Frederiksen. Henrik Lund, a national poet of Greenland. Proceedings of the American Philosophical Society, 96(6):653–659, December 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143634.

Logan:1978:FDP

[LG78] Jonothan L. Logan and Samuel A. Goudsmit. The first digit phenomenon. Proceedings of the American Philosophical Society, 122 (4):193–197, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986530.

Leeds-Hurwitz:1985:CRN

[LH85] Wendy Leeds-Hurwitz. The Committee on Research in Native American Languages. Proceedings of the American Philosophical Society, 129(2):129–160, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986986.

Lhermitte:1993:JSA

[Lhe93] François Lhermitte. Jean Stoetzel (23 April 1910–21 February 1987). Proceedings of the American Philosophical Society, 137 (1):174–177, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986954.

Li:1955:LCP

[Li55] Hui-Lin Li. Luigi Castiglioni as a pioneer in plant geography and plant introduction. Proceedings of the American Philosophical Society, 99(2):51–56, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143815.

Li:1972:HA

[Li72] Choh Hao Li. Hormones of the adenohypophysis. Proceedings of the American Philosophical Society, 116(5):365–382, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986068.

Likens:1999:SNN

[Lik99] Gene E. Likens. The science of nature, the nature of science: Long-term ecological studies at Hubbard Brook. Proceedings of the American Philosophical Society, 143(4):558–572, December 1999.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181988.

Lilly:1962:VBB

[Lil62] John C. Lilly. Vocal behavior of the bottlenose dolphin. Proceedings of the American Philosophical Society, 106(6):520–529, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985258.

Lingelbach:1950:FML

[Lin50] William E. Lingelbach. Franklin Medal Lecture: The story of “Philosophical Hall”. Proceedings of the American Philosophical Society, 94(3):185–213, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143553.

Lingelbach:1951:NLP

[Lin51] William E. Lingelbach. Notable letters and papers: Relating to (1) Early American diplomacy and winning the peace, (2) Darwinism and the great revolution in science, (3) a new Franklin letter to Jane Mecom and (4) Dr. Rush to Patrick Henry. Proceedings of the American Philosophical Society, 95(3):209–219, June 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143056.

Lingelbach:1952:FAI

[Lin52a] William E. Lingelbach. Franklin’s “American instructor”. Early Americanism in the art of writing. Proceedings of the American Philosophical Society, 96(4):367–387, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143795.

Lingelbach:1952:NPA

[Lin52b] William E. Lingelbach. Note on the papers on American Indian languages and archaeology. Proceedings of the American Philosophical Society, 96(4):??, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143794.

Lingelbach:1953:APS

[Lin53] William E. Lingelbach. The American Philosophical Society Library from 1942 to 1952 with a survey of its historical background. Proceedings of the American Philosophical Society, 97 (5):471–492, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3149265.

Lingelbach:1955:BFP

[Lin55a] William E. Lingelbach. Benjamin Franklin’s papers and the American Philosophical Society. Proceedings of the American Philosophical Society, 99(6):359–380, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143918.

Linton:1955:CFO

[Lin55b] M. Albert Linton. Controversial facets of old-age security. Proceedings of the American Philosophical Society, 99(4):205–218, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143699.

Lingelbach:1956:BFA

[Lin56] William E. Lingelbach. Benjamin Franklin and the American Philosophical Society in 1956. Proceedings of the American Philosophical Society, 100(4):354–368, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143765.

Link:1968:WWS

[Lin68] Arthur S. Link. Woodrow Wilson and the study of administration. Proceedings of the American Philosophical Society, 112(6):431–433, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985941.

Link:1987:R

[Lin87] Arthur S. Link. Remarks. Proceedings of the American Philosophical Society, 131(3):287–288, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987023.

Lindauer:1994:WRJ

[Lin94a] Martin Lindauer. Werner Reichardt (30 January 1924–18 September 1992). Proceedings of the American Philosophical Society, 138 (3):458–461, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986753.

Link:1994:DGP

[Lin94b] Arthur S. Link. Dr. Grayson’s predicament. Proceedings of the American Philosophical Society, 138(4):487–494, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986848.

Lipsey:1991:SFA

[Lip91] Robert E. Lipsey. Solomon Fabricant (August 15, 1906–September 13, 1989). Proceedings of the American Philosophical Society, 135 (4):606–611, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986821.

Litz:1996:CBO

[Lit96] A. Walton Litz. Cleanth Brooks (16 October 1906–10 May 1994). Proceedings of the American Philosophical Society, 140(1):88–91, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987278.

Levi:1974:MMR

[LL74] Annalina Levi and Mario Levi. The medieval map of Rome in the Ambrosian Library’s manuscript of Solinus (c 246 inf.). Proceedings of the American Philosophical Society, 118(6):567–594, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986406.

Lloyd:1950:DCJ

[Llo50] H. Alan Lloyd. Description of a clock by John Shelton, owned by the Royal Society and used by Maskelyne on his visit to St. Helena in 1761, and probably by Mason & Dixon in Pennsylvania. Proceedings of the American Philosophical Society, 94(3):268–271, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143561.

Lloyd:1989:WSC

[Llo89] Paul M. Lloyd. Where Spanish came from: The birth of Castile. Proceedings of the American Philosophical Society, 133(3):360–369, September 1989. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987105.

Locke:1997:VBW

[Loc97] Michael Locke. V. B. Wigglesworth (17 April 1899–12 February 1994). Proceedings of the American Philosophical Society, 141(1): 108–114, March 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987256.

Loftus:1998:IM

[Lof98] Elizabeth F. Loftus. Illusions of memory. Proceedings of the American Philosophical Society, 142(1):60–73, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3152264.

Lokken:1978:CCA

[Lok78] Roy N. Lokken. Cadwallader Colden’s attempt to advance natural philosophy beyond the Eighteenth-Century mechanistic paradigm. Proceedings of the American Philosophical Society, 122(6):365–376, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986452.

Lombardo:1997:SI

[Lom97] Agostino Lombardo. Shakespeare in Italy. Proceedings of the American Philosophical Society, 141(4):454–462, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987221.

Long:1957:DVH

[Lon57] Esmond R. Long. Development of the voluntary health movement in America as illustrated in the pioneer National Tuberculosis Association. Proceedings of the American Philosophical Society, 101 (2):142–148, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985505.

Long:1958:NAP

[Lon58] Herbert S. Long. Notes on Aeschylus’ “Prometheus Bound”. Proceedings of the American Philosophical Society, 102(3):229–280, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985576.

Longhurst:1959:LS

[Lon59] John E. Longhurst. Luther in Spain: 1520–1540. Proceedings of the American Philosophical Society, 103(1):66–93, February 1959. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985382.

Lopez:1964:BFL

[Lop64] Claude A. Lopez. Benjamin Franklin, Lafayette, and the “Lafayette”. Proceedings of the American Philosophical Society, 108(3):181–223, June 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985679.

Lopez:1977:OCS

[Lop77] Claude-Anne Lopez. One Christmas season in the life of Benjamin Franklin. Proceedings of the American Philosophical Society, 121 (5):373–376, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986341.

Lopez:1985:BFW

[Lop85] Claude-Anne Lopez. Benjamin Franklin and William Dodd: a new look at an old cause célèbre. Proceedings of the American Philosophical Society, 129(3):260–267, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987009.

Lowdin:1991:OTS

[Löw91] Per-Olov Löwdin. On the origin of the tone system and the theory of harmony. Proceedings of the American Philosophical Society, 135 (4):533–556, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986815.

Leighton:1955:CSS

[LS55] Alexander H. Leighton and Robert J. Smith. A comparative study of social and cultural change. Proceedings of the American Philosophical Society, 99(2):79–88, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143818.

Lubenow:1978:IGD

[Lub78] William C. Lubenow. Ireland, the great depression, and the railway rates: Political issues and backbench opinion in the house of commons of 1886. Proceedings of the American Philosophical Society, 122(4):204–213, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986532.

Luco:1970:LLE

[Luc70] J. V. Luco. Long-lasting effects of neuron activity. Proceedings of the American Philosophical Society, 114(6):432–434, December 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986145.

Ludmerer:1990:MIH

[Lud90] Kenneth M. Ludmerer. Methodological issues in the history of medicine: Achievements and challenges. Proceedings of the American Philosophical Society, 134(4):367–386, December 1990. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986894.

Lukacs:1981:DPN

[Luk81] John R. Lukacs. Dental pathology and nutritional patterns of South Asian megalith-builders: The evidence from Iron Age Mahurjhari. Proceedings of the American Philosophical Society, 125(3):220–237, June 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986483.

Lumiansky:1978:UUC

[Lum78] R. M. Lumiansky. The US–USSR Commission on the Humanities and the Social Sciences. Proceedings of the American Philosophical Society, 122(2):115–118, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986626.

Luria:1972:PML

[Lur72] S. E. Luria. Penrose Memorial Lecture. Slippery when wet being an essay on science, technology, and responsibility. Proceedings of the American Philosophical Society, 116(5):351–356, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986065.

Leopold:1998:VCM

[LVF98] Luna B. Leopold and Claudio Vita-Finzi. Valley changes in the Mediterranean and America and their effects on humans. Proceedings of the American Philosophical Society, 142(1):1–17, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152261.

Maass:1981:PWT

[Maa81] Eleanor A. Maass. A public watchdog: Thomas Pym Cope and the Philadelphia waterworks. Proceedings of the American Philosophical Society, 125(2):134–154, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986642.

MacInnes:1953:EFC

[Mac53] Duncan A. MacInnes. The electromotive force centrifuge, the development of a tool for research. Proceedings of the American Philosophical Society, 97(1):51–55, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143730.

MacIver:1957:ROS

[Mac57] Robert M. MacIver. The rights and obligations of the scholar. Proceedings of the American Philosophical Society, 101(5):455–458, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985200.

MacIver:1961:SSP

[Mac61] Robert M. MacIver. Science as a social phenomenon. Proceedings of the American Philosophical Society, 105(5):500–505, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985549.

MacDonald:1965:DHE

[Mac65a] Gordon J. F. MacDonald. Dynamical history of the Earth–Moon system. Proceedings of the American Philosophical Society, 109 (5):282–287, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986081.

Machlup:1965:FGC

[Mac65b] Fritz Machlup. The fascination of gold: Comment on the article “Gold” by Zay Jeffries in the Proceedings, volume 108, pages 437–442. Proceedings of the American Philosophical Society, 109 (2):105–107, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/985786. See [Jef64] and comments [Jef65].

Machlup:1965:WED

[Mac65c] Fritz Machlup. Why economists disagree. Proceedings of the American Philosophical Society, 109(1):1–7, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985773.

Mackey:1973:GTS

[Mac73] George W. Mackey. Group theory and its significance for mathematics and physics. Proceedings of the American Philosophical Society, 117(5):374–380, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986606.

MacLane:1984:PE

[Mac84] Saunders MacLane. Proof is eternal. Proceedings of the American Philosophical Society, 128(1):44–47, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986493.

MacLaury:1997:RVR

[Mac97] Bruce K. MacLaury. Robert V. Roosa (21 June 1918–23 December 1993). Proceedings of the American Philosophical Society, 141(2): 226–229, June 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987305.

Madden:1962:FWL

[Mad62] Edward H. Madden. Francis Wayland and the limits of moral responsibility. Proceedings of the American Philosophical Society, 106(4):348–359, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985268.

Maehl:1977:BFA

[Mae77] William H. Maehl. Bebel’s fight against the “Schlachtflotte”, nemesis to the primacy of foreign policy. Proceedings of the American Philosophical Society, 121(3):209–226, June 1977. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986411.

Maestro:1980:DPV

[Mae80] Marcello Maestro. The death penalty viewed as an act of selfdefense by two Italian jurists in the Eighteenth Century. Proceedings of the American Philosophical Society, 124(1):52–54, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986658.

Maguire:1956:DEE

[Mag56] Bassett Maguire. Distribution, endemicity, and evolution patterns among compositae of the Guayana Highland of Venezuela. Proceedings of the American Philosophical Society, 100(5):467–475, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143680.

Magill:1958:FFT

[Mag58] Roswell Magill. Fundamentals of Federal taxation. Proceedings of the American Philosophical Society, 102(2):160–163, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985280.

Maggs:1985:FTI

[Mag85] Barbara W. Maggs. The Franklin–Tolstoy influence controversy. Proceedings of the American Philosophical Society, 129(3):268–277, September 1985. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987010.

Magnusson:1999:IDT

[Mag99] David Magnusson. Individual development: Toward a developmental science. Proceedings of the American Philosophical Society, 143(1):86–96, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181976.

Mahoney:1991:MRP

[Mah91] Margaret E. Mahoney. Medical rights and the public welfare. Proceedings of the American Philosophical Society, 135(1):22–29, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987146.

Mahoney:1994:HBC

[Mah94] Margaret E. Mahoney. Healing broken communities: Building for the future. Proceedings of the American Philosophical Society, 138 (2):189–194, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987121.

Malone:1950:PWC

[Mal50] Kemp Malone. A poet at work: Chaucer revising his verses. Proceedings of the American Philosophical Society, 94(4):317–320, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143514.

Malik:1953:REW

[Mal53] Charles Malik. The relations of east and west. Proceedings of the American Philosophical Society, 97(1):1–7, February 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143724.

Malzberg:1955:TMD

[Mal55] Benjamin Malzberg. Trends of mental disease in New York State, 1920–1950. Proceedings of the American Philosophical Society, 99 (3):174–183, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143782.

Malone:1958:DHB

[Mal58] Dumas Malone. Discussion: Hamilton on balance. Proceedings of the American Philosophical Society, 102(2):129–135, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985276.

Malkiel:1998:TCE

[Mal98] Burton G. Malkiel. Two common errors in empirical financial research. Proceedings of the American Philosophical Society, 142(3): 432–441, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152249.

Mangelsdorf:1958:RAC

[Man58] Paul C. Mangelsdorf. Reconstructing the ancestor of corn. Proceedings of the American Philosophical Society, 102(5):454–463, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985591.

Mangelsdorf:1965:GAW

[Man65] Paul C. Mangelsdorf. Genetics, agriculture, and the world food problem. Proceedings of the American Philosophical Society, 109 (4):242–248, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985883.

Mangelsdorf:1983:MCN

[Man83] Paul C. Mangelsdorf. The mystery of corn: New perspectives. Proceedings of the American Philosophical Society, 127(4):215–247, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986272.

Marvin:1951:BVP

[Mar51] Mary Vaughan Marvin. The Benjamin Vaughan Papers: Discovery and provenance. Proceedings of the American Philosophical Society, 95(3):246–249, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143061.

Marceau:1953:PAT

[Mar53] Henri Marceau. Photographic aids and their uses in problems of authenticity in the field of paintings. Proceedings of the American Philosophical Society, 97(6):686–712, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3149284.

Margenau:1955:PSN

[Mar55] Henry Margenau. Present status and needs of the philosophy of science. Proceedings of the American Philosophical Society, 99(5): 334–337, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/3143932. See comments [Din55].

Margenau:1961:BMP

[Mar61] Henry Margenau. Bacon and modern physics: a confrontation. Proceedings of the American Philosophical Society, 105(5):487–492, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985547.

Martin:1965:MPL

[Mar65] Philip L. Martin. Madison’s precedent of legislative ratification for Constitutional amendments. Proceedings of the American Philosophical Society, 109(1):47–52, February 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985777.

Marrocco:1972:SQA

[Mar72] W. Thomas Marrocco. The string quartet attributed to Benjamin Franklin. Proceedings of the American Philosophical Society, 116 (6):477–485, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985761.

Martin:1977:VDH

[Mar77] John Rupert Martin. Van Dyck as history painter: “the mocking of Christ”. Proceedings of the American Philosophical Society, 121(3):227–234, June 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986412.

Martin:1986:PDR

[Mar86] John Rupert Martin. Portraits of doctors by Rembrandt and Rubens. Proceedings of the American Philosophical Society, 130(1): 7–20, March 1986. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987089.

Marshak:1988:PHB

[Mar88] Robert E. Marshak. The pragmatic humanism of Bohr, Einstein and Sakharov. Proceedings of the American Philosophical Society, 132(3):268–275, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143854.

Martin:1993:PE

[Mar93] Judith Martin. A philosophy of etiquette. Proceedings of the American Philosophical Society, 137(3):350–356, September 1993. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986996.

Marty:1998:FWF

[Mar98] Martin E. Marty. The future of world fundamentalisms. Proceedings of the American Philosophical Society, 142(3):367–377, September 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152242.

Massey:1992:PUK

[Mas92] Walter E. Massey. Promoting useful knowledge. Proceedings of the American Philosophical Society, 136(3):347–352, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986906.

Matthewson:1996:JNH

[Mat96] Tim Matthewson. Jefferson and the nonrecognition of Haiti. Proceedings of the American Philosophical Society, 140(1):22–48, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987274.

Matlock:1999:CCC

[Mat99] Jack F. Matlock, Jr. Can civilizations clash? Proceedings of the American Philosophical Society, 143(3):428–439, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181955.

Mauersberger:1999:ACN

[Mau99] Konrad Mauersberger. Alfred O. C. Nier (28 May 1911–16 May 1994). Proceedings of the American Philosophical Society, 143(4): 686–691, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181997.

Mayr:1959:IEF

[May59] Ernst Mayr. Isolation as an evolutionary factor. Proceedings of the American Philosophical Society, 103(2):221–230, April 1959. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985150.

Mayor:1960:SEE

[May60] John R. Mayor. Science education in the elementary and secondary schools. Proceedings of the American Philosophical Society, 104 (6):558–561, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985530.

Mayerson:1963:DSP

[May63a] Philip Mayerson. The desert of Southern Palestine according to Byzantine sources. Proceedings of the American Philosophical Society, 107(2):160–172, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985446.

Maynard:1963:ROI

[May63b] Douglas Maynard. Reform and the origin of the international organization movement. Proceedings of the American Philosophical Society, 107(3):220–231, June 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985581.

Mayer:1978:IVI

[May78] Hans Eberhard Mayer. Ibelin versus Ibelin: The struggle for the Regency of Jerusalem 1253–1258. Proceedings of the American Philosophical Society, 122(1):25–57, February 1978. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986261.

Mayer:1981:AC

[May81] Hans Eberhard Mayer. America and the Crusades. Proceedings of the American Philosophical Society, 125(1):38–45, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986185.

Mayer:1984:JJH

[May84] Hans Eberhard Mayer. John of Jaffa, his opponents, and his fiefs. Proceedings of the American Philosophical Society, 128(2):134–163, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986227.

Mayer:1989:AVN

[May89] Hans Eberhard Mayer. Angevins versus Normans: The new men of King Fulk of Jerusalem. Proceedings of the American Philosophical Society, 133(1):1–25, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987158.

Mayr:1991:IRD

[May91] Ernst Mayr. The ideological resistance to Darwin’s theory of natural selection. Proceedings of the American Philosophical Society, 135(2):123–139, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987030.

Mayr:1995:DIM

[May95] Ernst Mayr. Darwin’s impact on modern thought. Proceedings of the American Philosophical Society, 139(4):317–325, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987231.

Mayer:1997:KMS

[May97] Hans Eberhard Mayer. Kenneth Meyer Setton (17 June 1914–18 February 1995). Proceedings of the American Philosophical Society, 141(2):240–249, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987307.

Mazzeo:1955:DPL

[Maz55] Joseph Anthony Mazzeo. Dante, the poet of love. Dante and the Phaedrus tradition of poetic inspiration. Proceedings of the American Philosophical Society, 99(3):133–145, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143779.

McBride:1956:BFV

[McB56] John McBride. Benjamin Franklin as viewed in France during the Bourbon Restoration (1814–1830). Proceedings of the American Philosophical Society, 100(2):114–128, April 1956. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143711.

McCormmach:1966:OMM

[McC66] Russell McCormmach. Ormsby MacKnight Mitchel’s “Sidereal Messenger”, 1846–1848. Proceedings of the American Philosophical Society, 110(1):35–47, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986000.

McCarty:1984:UFM

[McC84] Maclyn McCarty. On the unexpected fruits of mission-oriented research. Proceedings of the American Philosophical Society, 128(1): 20–26, March 1984. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986488.

McCarthy:1985:SSA

[McC85] Kathleen D. McCarthy. The short and simple annals of the poor: Foundation funding for the humanities, 1900–1983. Proceedings of the American Philosophical Society, 129(1):3–8, March 1985. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986974.

McCorison:1991:JLH

[McC91] Marcus A. McCorison. The Jayne Lecture. Humanists and bytesized bibliography. Or, how to digest expanding sources of information. Proceedings of the American Philosophical Society, 135(1): 61–72, March 1991. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987150.

McCarty:1993:MHA

[McC93] Maclyn McCarty. Michael Heidelberger (29 April 1888–25 June 1991). Proceedings of the American Philosophical Society, 137(3): 432–437, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987002.

McCormmach:1998:MCW

[McC98] Russell McCormmach. Mr. Cavendish weighs the world. Proceedings of the American Philosophical Society, 142(3):355–366, September 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152241.

McDermott:1959:WJG

[McD59] John Francis McDermott. The Western journals of George Hunter, 1796–1805. Proceedings of the American Philosophical Society, 103 (6):770–773, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985390.

McEvoy:1984:JPS

[McE84] John G. McEvoy. Joseph Priestley, scientist, philosopher and divine. Proceedings of the American Philosophical Society, 128(3): 193–199, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986880.

McGaugh:1967:AMT

[McG67] James L. McGaugh. Analysis of memory transfer and enhancement. Proceedings of the American Philosophical Society, 111(6): 347–351, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985731.

McGucken:1979:SRS

[McG79] William McGucken. The social relations of science: The British Association for the Advancement of Science, 1931–1946. Proceedings of the American Philosophical Society, 123(4):236–264, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986583.

McGill:1988:EFI

[McG88] Dan M. McGill. Economic and financial implications of the aging phenomenon. Proceedings of the American Philosophical Society, 132(2):154–171, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986861.

McGhee:1995:HPC

[McG95] George C. McGhee. Humankind and the purpose of the cosmos. Proceedings of the American Philosophical Society, 139(2):163–170, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987263.

McGhee:1996:SUW

[McG96] George C. McGhee. The scourge of unemployment and what might be done about it. Proceedings of the American Philosophical Society, 140(4):459–468, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987288.

McGhee:1997:MWP

[McG97] George C. McGhee. The mid-weekend: a proposal for alternative days of rest. Proceedings of the American Philosophical Society, 141(2):213–215, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987302.

McGhee:1999:GWT

[McG99] George C. McGhee. George Washington’s trusteeship over the estate of the Reverend Dr. Charles Green, 1765–1772, which included the present area of Farmer’s Delight Farm, Loudoun County, Virginia. Proceedings of the American Philosophical Society, 143(4): 639–647, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181992.

McIntosh:1975:SAC

[McI75] Marjorie Keniston McIntosh. Sir Anthony Cooke: Tudor humanist, educator, and religious reformer. Proceedings of the American Philosophical Society, 119(3):233–250, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986673.

McKinsey:1973:JRP

[McK73] Elizabeth R. McKinsey. James Ralph: The professional writer comes of age. Proceedings of the American Philosophical Society, 117(1):59–78, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985948.

McKusick:1994:DGS

[McK94] Vincent L. McKusick. Discretionary gatekeeping: The Supreme Court’s management of its original jurisdiction docket since 1961 (through 25 April 1993). Proceedings of the American Philosophical Society, 138(2):195–250, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987122.

McKusick:1997:MHG

[McK97] Victor A. McKusick. Mapping the human genome: Retrospective, perspective and prospective. Proceedings of the American Philosophical Society, 141(4):417–424, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987218.

McNeill:1979:WHE

[McN79] William H. McNeill. What’s happening to European history in the United States? Proceedings of the American Philosophical Society, 123(6):341–343, December 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986614.

McPherson:1995:WFS

[McP95] James M. McPherson. Who freed the slaves? Proceedings of the American Philosophical Society, 139(1):1–10, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986716.

McPherson:1998:CBB

[McP98a] Mary Patterson McPherson. A century of building at Bryn Mawr. Proceedings of the American Philosophical Society, 142(3):399–415, September 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152246.

McPherson:1998:JMO

[McP98b] Mary Patterson McPherson. Jane Marion Oppenheimer (19 September 1911–19 March 1996). Proceedings of the American Philosophical Society, 142(2):288–291, June 1998. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3152306.

McPherson:1999:WBT

[McP99] James M. McPherson. Was blood thicker than water? Ethnic and civic nationalism in the American Civil War. Proceedings of the American Philosophical Society, 143(1):102–108, March 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181978.

Murphy:1964:EDG

[MD64] C. T. Murphy and R. H. Dicke. The effects of a decreasing gravitational constant in the interior of the Earth. Proceedings of the American Philosophical Society, 108(3):224–246, June 1964. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985680.

Mearns:1954:CSM

[Mea54] David C. Mearns. Comments on the Symposium on the Manuscript Sources of American History: Problems of their control, use, and publication. Proceedings of the American Philosophical Society, 98(3):185–188, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143844.

Mears:1978:SSH

[Mea78] James A. Mears. Some sources of the herbarium of Henry Muhlenberg (1753–1815). Proceedings of the American Philosophical Society, 122(3):155–174, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986550.

Meiss:1966:SVA

[Mei66] Millard Meiss. Sleep in Venice. Ancient myths and Renaissance proclivities. Proceedings of the American Philosophical Society, 110 (5):348–382, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986024.

Melville:1957:RCC

[Mel57] Ward Melville. The re-creation of a Colonial village. Proceedings of the American Philosophical Society, 101(3):290–295, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985370.

Mellink:1966:AON

[Mel66] Machteld J. Mellink. Anatolia: Old and new perspectives. Proceedings of the American Philosophical Society, 110(2):111–129, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985744.

Mellon:1975:NGA

[Mel75] Paul Mellon. The National Gallery of Art: an American experiment. Proceedings of the American Philosophical Society, 119 (5):349–352, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986677.

Mela:1979:DNM

[Mel79] Leena Mela. Development of the newborn to meet its own energy requirements. Proceedings of the American Philosophical Society, 123(5):293–297, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986591.

Melchers:1997:NKJ

[Mel97] Fritz Melchers. Niels Kaj Jerne (23 December 1911–7 October 1994). Proceedings of the American Philosophical Society, 141(1): 90–96, March 1997. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987253.

Mellink:1998:EPA

[Mel98] Machteld J. Mellink. Edith Porada (22 August 1912–24 March 1994). Proceedings of the American Philosophical Society, 142(4): 688–692, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152293.

Menzel:1951:CGP

[Men51] Margaret Young Menzel. The cytotaxonomy and genetics of physalis. Proceedings of the American Philosophical Society, 95 (2):132–183, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143331.

Mendelsohn:1961:CSH

[Men61] Everett Mendelsohn. The controversy over the site of heat production in the body. Proceedings of the American Philosophical Society, 105(4):412–420, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985498.

Menzel:1969:MAL

[Men69a] Donald H. Menzel. The Moon as an abode of life? Proceedings of the American Philosophical Society, 113(2):102–126, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985960.

Menzel:1969:VPD

[Men69b] Donald H. Menzel. Venus past, and the distance of the Sun. Proceedings of the American Philosophical Society, 113(3):197–202, June 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985735.

Menard:1989:WTM

[Men89] Russell R. Menard. Was there a “middle colonies demographic regime”? Proceedings of the American Philosophical Society, 133 (2):215–218, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987051.

Merrill:1953:SLC

[Mer53] Paul W. Merrill. Side lights on cosmic chemistry. Proceedings of the American Philosophical Society, 97(4):356–358, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3149290.

Merton:1961:SMS

[Mer61] Robert K. Merton. Singletons and multiples in scientific discovery: a chapter in the sociology of science. Proceedings of the American Philosophical Society, 105(5):470–486, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985546.

Meritt:1971:CPW

[Mer71] Benjamin D. Meritt. The chronology of the peloponnesian war. Proceedings of the American Philosophical Society, 115(2):97–124, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985851.

Meritt:1975:PAM

[Mer75] Benjamin D. Meritt. Perikles, the Athenian Mint, and the Hephaisteion. Proceedings of the American Philosophical Society, 119 (4):267–274, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986288.

Meritt:1977:ABA

[Mer77] Benjamin D. Meritt. The alliance between Athens and Egesta. Proceedings of the American Philosophical Society, 121(6):437–447, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986392.

Meritt:1978:CPW

[Mer78] Benjamin Dean Meritt. The chronology of the peloponnesian war-II. Proceedings of the American Philosophical Society, 122(5):287– 293, October 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986684.

Merrell:1983:RAE

[Mer83] James H. Merrell. Reading “an almost erased page”: a reassessment of Frank G. Speck’s Catawba studies. Proceedings of the American Philosophical Society, 127(4):248–262, 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986273.

Meritt:1984:SRA

[Mer84] Benjamin D. Meritt. The Samian Revolt from Athens in 440–439 B.C. Proceedings of the American Philosophical Society, 128(2): 123–133, June 1984. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986226.

Meritt:1988:PPB

[Mer88] Benjamin D. Meritt. Philokles and the Panathenaia of 406 B.C. Proceedings of the American Philosophical Society, 132(1):128– 135, March 1988. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143828.

Metzger:1987:HEG

[Met87] Bruce M. Metzger. History of editing the Greek New Testament. Proceedings of the American Philosophical Society, 131(2):148–158, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986788.

Metzger:1991:NRS

[Met91] Bruce M. Metzger. The new revised standard version of the Bible its making and character. Proceedings of the American Philosophical Society, 135(3):368–381, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986774.

Metcalf:1993:MTF

[Met93] William E. Metcalf. Margaret Thompson (22 February 1911–29 February 1992). Proceedings of the American Philosophical Society, 137(3):458–463, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987006.

Meyerson:1998:ESN

[Mey98] Martin Meyerson. Eliot Stellar (1 November 1919–12 October 1993). Proceedings of the American Philosophical Society, 142(4): 694–700, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152294.

Middleton:1985:IRP

[Mid85] Charles R. Middleton. Irish representative peerage elections and the Conservative Party, 1832–1841. Proceedings of the American Philosophical Society, 129(1):90–111, March 1985. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986981.

Mikol:1958:IST

[Mik58] Vera Mikol. The influence of Sigismund Thalberg on American musical taste, 1830–1872. Proceedings of the American Philosophical Society, 102(5):464–468, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985592.

Miller:1955:FTS

[Mil55] C. William Miller. Franklin’s type: Its study past and present. Proceedings of the American Philosophical Society, 99(6):418–432, December 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143924.

Miller:1959:CLC

[Mil59] Perry Miller. The Common Law and codification in Jacksonian America. Proceedings of the American Philosophical Society, 103 (3):463–468, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985477.

Miller:1964:ITP

[Mil64a] Maynard M. Miller. Inventory of terminal position changes in Alaskan coastal glaciers since the 1750’s. Proceedings of the American Philosophical Society, 108(3):257–273, June 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985682.

Miller:1964:MCP

[Mil64b] Maynard M. Miller. Morphogenetic classification of Pleistocene glaciations in the Alaska–Canada Boundary Range. Proceedings of the American Philosophical Society, 108(3):247–256, June 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985681.

Miller:1967:LLR

[Mil67] Neal E. Miller. Laws of learning relevant to its biological basis. Proceedings of the American Philosophical Society, 111(6):315–325, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985728.

Miller:1972:LM

[Mil72] George A. Miller. Lexical memory. Proceedings of the American Philosophical Society, 116(2):140–144, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986168.

Miller:1976:ASI

[Mil76] Sally M. Miller. Americans and the Second International. Proceedings of the American Philosophical Society, 120(5):372–387, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986268.

Mild:1978:SHL

[Mil78] Warren Mild. Susanna Highmore’s literary reputation. Proceedings of the American Philosophical Society, 122(6):377–384, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986453.

Miller:1982:AF

[Mil82a] J. Irwin Miller. An anatomy of freedom. Proceedings of the American Philosophical Society, 126(1):1–5, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986474.

Miller:1982:JOY

[Mil82b] Robert Ryal Miller. James Orton: a Yankee naturalist in South America, 1867–1877. Proceedings of the American Philosophical Society, 126(1):11–25, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986476.

Miller:1987:BCJ

[Mil87] S. J. Miller. A belated conversion from jurisdictionalism to Jansenism, Josse Le Plat, 1732–1810. Proceedings of the American Philosophical Society, 131(4):397–424, December 1987. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987139.

Miller:1995:AHN

[Mil95] Frank C. Miller. E. Adamson Hoebel (6 November 1906–23 July 1993). Proceedings of the American Philosophical Society, 139(1): 106–108, March 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986725.

Milsom:1998:SET

[Mil98] S. F. C. Milsom. Samuel Edmund Thorne (14 October 1907–7 April 1994). Proceedings of the American Philosophical Society, 142(1):162–168, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152274.

Miner:1973:TMA

[Min73] Horace M. Miner. Traditional mobility among the weavers of Fez. Proceedings of the American Philosophical Society, 117(1):17–36, February 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985945.

Mitchell:1958:AHF

[Mit58] Broadus Mitchell. Alexander Hamilton as Finance Minister. Proceedings of the American Philosophical Society, 102(2):117–123, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985274.

Mitchell:1962:TPT

[Mit62] Robert B. Mitchell. Transportation problems and their solution. Proceedings of the American Philosophical Society, 106(3):170–176, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985172.

Mizushima:1979:CSB

[Miz79] Sanichiro Mizushima. Cultural and social background of the rapid modernization of Japan. Proceedings of the American Philosophical Society, 123(3):164–167, June 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986219.

Munk:1972:VH

[MM72] Judith Munk and Walter Munk. Venice hologram. Proceedings of the American Philosophical Society, 116(5):415–442, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986073.

McDavid:1986:ILA

[MMK^{+}86] Raven I. McDavid, Jr., Virginia G. McDavid, William A. Kretzschmar, Jr., Theodore K. Lerud, and Martha Ratliff. Inside a linguistic atlas. Proceedings of the American Philosophical Society, 130(4):390–405, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986926.

Mayer-Oakes:1959:SER

[MO59] William J. Mayer-Oakes. A stratigraphic excavation at El Risco, Mexico. Proceedings of the American Philosophical Society, 103(3): 332–373, June 1959. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985473.

Mochon:1968:SMC

[Moc68] Marion Johnson Mochon. Stockbridge–Munsee cultural adaptations: “assimilated Indians”. Proceedings of the American Philosophical Society, 112(3):182–219, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986163.

Moe:1958:VPP

[Moe58] Henry Allen Moe. “The vision of Piers the plowman” and the law of foundations. Proceedings of the American Philosophical Society, 102(4):371–375, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985221.

Moe:1961:NOP

[Moe61] Henry Allen Moe. Notes on the origin of philanthropy in Christendom. Proceedings of the American Philosophical Society, 105 (2):141–144, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985624.

Moment:1950:EPE

[Mom50] Gairdner B. Moment. On the electrical potentials of an earthworm, Eisenia foetida, considered in relation to growth. Proceedings of the American Philosophical Society, 94(1):24–30, February 1950. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143248.

Moore:1960:CWS

[Moo60] Wilbert E. Moore. The corporation as a welfare state. Proceedings of the American Philosophical Society, 104(6):677–680, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985540.

Moody:1987:AHA

[Moo87] Michael E. Moody. The apostasy of Henry Ainsworth: a case study in early separatist historiography. Proceedings of the American Philosophical Society, 131(1):15–31, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987114.

Mooney:1999:PJK

[Moo99] Harold A. Mooney. Paul Jackson Kramer (8 May 1904–24 May 1995). Proceedings of the American Philosophical Society, 143(2): 340–343, June 1999. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3181944.

Morris:1958:WHG

[Mor58] Richard B. Morris. Washington and Hamilton: a great collaboration. Proceedings of the American Philosophical Society, 102 (2):107–116, April 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985273.

Morris:1960:CEI

[Mor60] Morris David Morris. Caste and the evolution of the industrial workforce in India. Proceedings of the American Philosophical Society, 104(2):124–133, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985655.

Morse:1964:CMR

[Mor64a] Philip M. Morse. Computing machines as research assistants. Proceedings of the American Philosophical Society, 108(4):291–297, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985909.

Moruzzi:1964:HDD

[Mor64b] Giuseppe Moruzzi. The historical development of the deafferentation hypothesis of sleep. Proceedings of the American Philosophical Society, 108(1):19–28, February 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985993.

Morgan:1966:PHT

[Mor66] Edmund S. Morgan. Puritan hostility to the theatre. Proceedings of the American Philosophical Society, 110(5):340–347, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986023.

Morgan:1969:ODC

[Mor69] William G. Morgan. The origin and development of the Congressional Nominating Caucus. Proceedings of the American Philosophical Society, 113(2):184–196, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985965.

Moriarty:1974:IAT

[Mor74a] Thomas F. Moriarty. The Irish Absentee Tax Controversy of 1773: a study in Anglo–Irish politics on the eve of the American Revolution. Proceedings of the American Philosophical Society, 118(4): 370–408, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986282.

Morris:1974:JQA

[Mor74b] Walter J. Morris. John Quincy Adams’s German library, with a catalog of his German books. Proceedings of the American Philosophical Society, 118(4):321–333, September 1974. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986279.

Morton:1977:RRC

[Mor77] Martin L. Morton. Relationship of reproductive cycle to environmental conditions and energy stores in white-crowned sparrows at high altitude. Proceedings of the American Philosophical Society, 121(5):377–382, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986342.

Morrison:1980:FFM

[Mor80a] Karl F. Morrison. “From form into form:” mimesis and personality in Augustine’s historical thought. Proceedings of the American Philosophical Society, 124(4):276–294, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986430.

Morrison:1980:WTL

[Mor80b] Philip Morrison. Warfare today: Limits to growth. Proceedings of the American Philosophical Society, 124(4):257–260, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986426.

Morgan:1983:WWP

[Mor83] Edmund S. Morgan. The world and William Penn. Proceedings of the American Philosophical Society, 127(5):291–315, October 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986499.

Morris:1996:SLP

[Mor96] Norval Morris. Street-level purposes of punishment. Proceedings of the American Philosophical Society, 140(2):197–203, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987325.

Mosely:1950:SRS

[Mos50a] Philip E. Mosely. Soviet research in the social field. Proceedings of the American Philosophical Society, 94(2):105–110, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143208.

Moshinsky:1950:IPP

[Mos50b] Marcos Moshinsky. Interference phenomena for particles obeying Bose or Fermi statistics. Proceedings of the American Philosophical Society, 94(1):53–58, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143251.

Mosely:1955:CIR

[Mos55a] Philip E. Mosely. Comments on ideology and reality in the Soviet system. Proceedings of the American Philosophical Society, 99(1):31–33, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143691.

Mosely:1955:RIC

[Mos55b] Philip E. Mosely. The Russian Institute of Columbia University. Proceedings of the American Philosophical Society, 99(1):36– 38, January 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143693.

Mosteller:1958:SML

[Mos58] Frederick Mosteller. Stochastic models for the learning process. Proceedings of the American Philosophical Society, 102(1):53–59, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985304.

Mosteller:1987:SSW

[Mos87] Frederick Mosteller. A statistical study of the writing styles of the authors of “the federalist” papers. Proceedings of the American Philosophical Society, 131(2):132–140, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986786.

Moscati:1997:W

[Mos97] Sabatino Moscati. Welcome. Proceedings of the American Philosophical Society, 141(4):392–393, December 1997. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987214.

Mote:1975:ARM

[Mot75] Victor L. Mote. Air resource management in Soviet Russia. Proceedings of the American Philosophical Society, 119(3):201–213, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986670.

Mould:1959:AGR

[Mou59] Richard A. Mould. An axiomatization of general relativity. Proceedings of the American Philosophical Society, 103(3):485–529, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985480.

Moulton:1983:SJL

[Mou83] Gary E. Moulton. The specialized journals of Lewis and Clark. Proceedings of the American Philosophical Society, 127(3):194–201, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986386.

Mountcastle:1991:I

[Mou91] Vernon B. Mountcastle. Introduction. Proceedings of the American Philosophical Society, 135(4):510–512, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986812.

Movius:1953:PMS

[Mov53] Hallam L. Movius, Jr. Palaeolithic and Mesolithic sites in Soviet Central Asia. Proceedings of the American Philosophical Society, 97(4):383–421, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3149292.

Moynihan:1992:ABU

[Moy92] Daniel Patrick Moynihan. Appendix b: The underclass: Toward a post-industrial social policy. Proceedings of the American Philosophical Society, 136(3):390–398, September 1992. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986912.

Moynihan:1997:JCM

[Moy97] Daniel Patrick Moynihan. James S. Coleman (12 May 1926–25 March 1995). Proceedings of the American Philosophical Society, 141(2):222–224, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987304.

Mohr:1967:SBC

[MS67] Albert Mohr and L. L. Sample. The sacred bundle complex among the Chumash. Proceedings of the American Philosophical Society, 111(1):38–45, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985752.

McPherson:1998:PFS

[MS98] Michael S. McPherson and Morton Owen Schapiro. Priorities for Federal student aid policy: Looking beyond Pell. Proceedings of the American Philosophical Society, 142(2):171–190, June 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152297.

Multhauf:1962:UBC

[Mul62] Robert P. Multhauf. On the use of the balance in chemistry. Proceedings of the American Philosophical Society, 106(3):210–218, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985176.

Muller:1972:ILM

[Mul72] Paul M. Muller. Implications of the lunar Mascon discovery. Proceedings of the American Philosophical Society, 116(5):362–364, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986067.

Mumford:1954:ASC

[Mum54] Lewis Mumford. Anticipations and social consequences of atomic energy. Proceedings of the American Philosophical Society, 98(2): 149–152, April 1954. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143646.

Mumford:1961:ST

[Mum61] Lewis Mumford. Science as technology. Proceedings of the American Philosophical Society, 105(5):506–511, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985550.

Muntzing:1959:DVV

[Mün59] Arne Müntzing. Darwin’s views on variation under domestication in the light of present-day knowledge. Proceedings of the American Philosophical Society, 103(2):190–220, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985149.

Murphy:1951:IMU

[Mur51] Robert Cushman Murphy. The impact of man upon nature in New Zealand. Proceedings of the American Philosophical Society, 95(6): 569–582, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143506.

Murrill:1955:RBP

[Mur55] Rupert Ivan Murrill. Racial blood pressure studies: a critique of methodology. With especial reference to the effect of age, nutrition, climate, and race on blood pressure in Puerto Rico. Proceedings of the American Philosophical Society, 99(4):277–324, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143707.

Murdock:1956:CEL

[Mur56] Kenneth B. Murdock. The Colonial experience in the literature of the United States. Proceedings of the American Philosophical Society, 100(2):129–132, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143712.

Murphy:1957:STR

[Mur57] Robert Cushman Murphy. The sketches of Titian Ramsay Peale (1799–1885). Proceedings of the American Philosophical Society, 101(6):523–531, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985520.

Murphy:1959:RFF

[Mur59] Robert Cushman Murphy. Robert Ferryman, forgotten naturalist. Proceedings of the American Philosophical Society, 103(6):774–777, December 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985391.

Murphey:1961:ISU

[Mur61a] Murray G. Murphey. The influence of science upon modern culture: The program. Proceedings of the American Philosophical Society, 105(5):461–463, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985544.

Murphree:1961:EAP

[Mur61b] Idus L. Murphree. The evolutionary anthropologists: The progress of mankind: The concepts of progress and culture in the thought of John Lubbock, Edward B. Taylor, and Lewis H. Morgan. Proceedings of the American Philosophical Society, 105(3):265–300, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985456.

Murphy:1962:MBP

[Mur62] Robert Cushman Murphy. Mitchill’s “Birds of Plandome,” Long Island. Proceedings of the American Philosophical Society, 106 (1):48–52, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985210.

Murty:1974:LPC

[Mur74] M. L. Krishna Murty. A Late Pleistocene cave site in Southern India. Proceedings of the American Philosophical Society, 118(2): 196–230, April 1974. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986666.

Musto:1969:YJQ

[Mus69] David F. Musto. The youth of John Quincy Adams. Proceedings of the American Philosophical Society, 113(4):269–282, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986152.

Musgrave:1972:FP

[Mus72] Richard A. Musgrave. Fiscal policy. Proceedings of the American Philosophical Society, 116(3):197–202, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986114.

Marquez-Villanueva:1994:DAO

[MV94] Francisco Márquez-Villanueva. Dámaso Alonso (22 October 1898– 25 January 1990). Proceedings of the American Philosophical Society, 138(1):146–148, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986708.

Mylonas:1955:CEM

[Myl55] George E. Mylonas. The cemeteries of Eleusis and Mycenae. Proceedings of the American Philosophical Society, 99(2):57–67, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143816.

Myrdal:1951:PTD

[Myr51] Alva Myrdal. Population trends in densely populated areas. Proceedings of the American Philosophical Society, 95(1):1–7, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143570.

Nash:1968:CPP

[Nas68] Gary B. Nash. City planning and political tension in the Seventeenth Century: The case of Philadelphia. Proceedings of the American Philosophical Society, 112(1):54–73, February 1968. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986101.

Nathan:1991:WBC

[Nat91] David G. Nathan. William Bosworth Castle (October 21, 1897– August 9, 1990). Proceedings of the American Philosophical Society, 135(4):592–597, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986819.

Nemec:1968:CED

[Nem68a] Ludvik Nemec. The communist ecclesiology during the church– state relationship in Czechoslovakia, 1945–1967. Proceedings of the American Philosophical Society, 112(4):245–276, August 1968. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985872.

Nemec:1968:CJA

[Nem68b] Ludvik Nemec. The Czech Jednota, the avant-garde of modern clerical progressivism and unionism. Proceedings of the American Philosophical Society, 112(1):74–100, February 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986102.

Nemec:1980:CRM

[Nem80] Ludvik Nemec. The Czech reform movement: “devotio moderna” in the Fourteenth and Fifteenth Centuries. Proceedings of the American Philosophical Society, 124(5):386–397, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986575.

Neugebauer:1951:BMC

[Neu51] O. Neugebauer. The Babylonian method for the computation of the last visibilities of Mercury. Proceedings of the American Philosophical Society, 95(2):110–116, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143328.

Neugebauer:1954:BPT

[Neu54] O. Neugebauer. Babylonian planetary theory. Proceedings of the American Philosophical Society, 98(1):60–89, February 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143671.

Neugebauer:1962:APO

[Neu62a] O. Neugebauer. Astronomical papyri and Ostraca: Bibliographical notes. Proceedings of the American Philosophical Society, 106 (4):383–391, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985270.

Neugebauer:1962:TBQ

[Neu62b] O. Neugebauer. Thãbit ben Qurra “on the solar year” and “on the motion of the eighth sphere”. Proceedings of the American Philosophical Society, 106(3):264–299, June 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985180.

Neugebauer:1963:SBM

[Neu63] O. Neugebauer. The survival of Babylonian methods in the exact sciences of Antiquity and Middle Ages. Proceedings of the American Philosophical Society, 107(6):528–535, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986109.

Neugebauer:1972:SAE

[Neu72] O. Neugebauer. On some aspects of Early Greek astronomy. Proceedings of the American Philosophical Society, 116(3):243–251, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986118.

Neugebauer:1986:IAS

[Neu86] Gerry Neugebauer. The infrared astronomical satellite. Proceedings of the American Philosophical Society, 130(2):155–174, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987179.

Neugebauer:1989:ARA

[Neu89] O. Neugebauer. From Assyriology to Renaissance art. Proceedings of the American Philosophical Society, 133(3):391–403, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987107.

Neustadt:1998:DKP

[Neu98] Richard E. Neustadt. Don K. Price (23 January 1910–9 July 1995). Proceedings of the American Philosophical Society, 142 (1):154–160, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152273.

Newell:1959:NFR

[New59] Norman D. Newell. The nature of the fossil record. Proceedings of the American Philosophical Society, 103(2):264–285, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985153.

Newsom:1961:SFS

[New61] Carroll V. Newsom. Science as a form of systematic knowledge: Some contributions from mathematics. Proceedings of the American Philosophical Society, 105(5):493–499, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985548.

Newman:1971:FMM

[New71] Eric P. Newman. Franklin making money more plentiful. Proceedings of the American Philosophical Society, 115(5):341–349, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985767.

Newell:1973:SCO

[New73] Norman D. Newell. Special creation and organic evolution. Proceedings of the American Philosophical Society, 117(4):323–331, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986698.

Newcomb:1990:CBE

[New90] Sally Newcomb. Contributions of British experimentalists to the discipline of geology: 1780–1820. Proceedings of the American Philosophical Society, 134(2):161–225, June 1990. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986843.

Nicholas:1950:DC

[Nic50] John S. Nicholas. Development of contractility. Proceedings of the American Philosophical Society, 94(2):175–183, April 1950. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143217.

Nichols:1956:FMT

[Nic56] Roy F. Nichols. Franklin’s message in the Twentieth Century. Proceedings of the American Philosophical Society, 100(4):346– 351, August 1956. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143763.

Nichols:1961:ISU

[Nic61a] Roy F. Nichols. The influence of science upon modern culture: Introduction. Proceedings of the American Philosophical Society, 105(5):459–460, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985543.

Nichols:1961:MS

[Nic61b] Roy F. Nichols. The magic square. Proceedings of the American Philosophical Society, 105(3):237–243, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985453.

Nicolaevsky:1961:WDH

[Nic61c] B. I. Nicolaevsky. Who is distorting history? (“Voprosy Istorii” and Karl Marx in 1848–1849). Proceedings of the American Philosophical Society, 105(2):209–236, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985632.

Nichols:1962:PCW

[Nic62] Roy F. Nichols. The problem of Civil War historiography: a discussion of the points of view of Professors Brogan, Duroselle, and Geyl. Proceedings of the American Philosophical Society, 106(1):36–40, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985208.

Nichols:1964:AVI

[Nic64] Roy F. Nichols. Adaptation versus invention as elements in historical analysis. Proceedings of the American Philosophical Society, 108(5):404–410, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985813.

Nichols:1967:DSL

[Nic67] Jeannette P. Nichols. Dollar strength as a liability in United States diplomacy. Proceedings of the American Philosophical Society, 111 (1):46–56, February 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985753.

Nierenberg:1986:CP

[Nie86] W. A. Nierenberg. Cyrano, physicist. Proceedings of the American Philosophical Society, 130(3):354–361, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986834.

Nierenberg:1992:RRD

[Nie92] William A. Nierenberg. Roger Randall Dougan Revelle (7 March 1909–15 July 1991). Proceedings of the American Philosophical Society, 136(4):596–600, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986765.

Nisbet:1975:TCP

[Nis75] Robert Nisbet. Turgot and the contexts of progress. Proceedings of the American Philosophical Society, 119(3):214–222, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986671.

Nisbet:1982:GM

[Nis82] Robert A. Nisbet. Genius & milieu. Proceedings of the American Philosophical Society, 126(6):441–451, December 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986250.

Noble:1968:STG

[Nob68] Joseph Veach Noble. Some trick Greek vases. Proceedings of the American Philosophical Society, 112(6):371–378, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985937.

Nordenfalk:1973:OIY

[Nor73] Carl Nordenfalk. Outdoors–indoors: a 2000-year-old space problem in Western Art. Proceedings of the American Philosophical Society, 117(4):233–258, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986694.

North:1993:EE

[Nor93] Helen F. North. Emblems of eloquence. Proceedings of the American Philosophical Society, 137(3):406–430, September 1993. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987001.

Notestein:1954:SDA

[Not54] Frank W. Notestein. Some demographic aspects of aging. Proceedings of the American Philosophical Society, 98(1):38–45, February 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143668.

Nourse:1950:EAP

[Nou50] Edwin G. Nourse. Economic analysis and political synthesis. Proceedings of the American Philosophical Society, 94(4):311–316, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143513.

Nowell:1995:CRP

[Now95] Peter C. Nowell. Cancer research: Promise and paradox. Proceedings of the American Philosophical Society, 139(1):32–43, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986718.

Noyes:1951:PTS

[Noy51] W. Albert Noyes, Jr. Photochemistry as a tool for the study of reaction mechanism. Proceedings of the American Philosophical Society, 95(5):542–543, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143239.

Neugebauer:1967:ATM

[NP67] O. Neugebauer and David Pingree. The astronomical tables of Mahādeva. Proceedings of the American Philosophical Society, 111 (2):69–92, April 1967. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986035.

Neugebauer:1981:DLE

[NPZ81] O. Neugebauer, Richard A. Parker, and Karl-Th. Zauzich. A demotic lunar eclipse text of the First Century, B.C. Proceedings of the American Philosophical Society, 125(4):312–327, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986333.

Nelson:1981:FVH

[NRF81] Clifford M. Nelson, Mary C. Rabbitt, and Fritiof M. Fryxell. Ferdinand Vandeveer Hayden: The U.S. Geological Survey years, 1879– 1886. Proceedings of the American Philosophical Society, 125(3): 238–243, June 1981. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986484.

Neugebauer:1964:APO

[NV64] O. Neugebauer and H. B. Van Hoesen. Astrological papyri and Ostraca: Bibliographical notes. Proceedings of the American Philosophical Society, 108(2):57–72, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985887.

Nygren:1970:RPE

[Nyg70] Edward John Nygren. Rubens Peale’s experiments with mesmerism. Proceedings of the American Philosophical Society, 114 (2):100–108, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986028.

Oaks:1977:PMO

[Oak77] Robert F. Oaks. Philadelphia merchants and the origins of American Independence. Proceedings of the American Philosophical Society, 121(6):407–436, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986391.

Oakes:1995:HSN

[Oak95] John B. Oakes. Harrison E. Salisbury (14 November 1908–5 July 1993). Proceedings of the American Philosophical Society, 139(2): 192–195, June 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987267.

Oakes:1998:JBR

[Oak98] John B. Oakes. James B. Reston (3 November 1909–6 December 1995). Proceedings of the American Philosophical Society, 142(2): 302–305, June 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152308.

ofBessborough:1990:WEE

[oB90] The Earl of Bessborough. Whither Europe: East and West. Proceedings of the American Philosophical Society, 134(3):321–327, September 1990. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986970.

Oberdorfer:1979:NPR

[Obe79] Don Oberdorfer. News and the perception of reality. Proceedings of the American Philosophical Society, 123(2):129–130, April 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986235.

Ogilvie:1981:NMS

[OC81] Marilyn Bailey Ogilvie and Clifford J. Choquette. Nettie Maria Stevens (1861–1912): Her life and contributions to cytogenetics. Proceedings of the American Philosophical Society, 125(4):292–311, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986332.

OConnor:1995:WBS

[O’C95] Sandra Day O’Connor. Warren E. Burger (17 September 1907– 25 June 1995). Proceedings of the American Philosophical Society, 139(3):266–269, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987192.

Oehler:1974:ASK

[Oeh74] Klaus Oehler. Aristotle on self-knowledge. Proceedings of the American Philosophical Society, 118(6):493–506, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986399.

OGorman:1978:MFR

[O’G78] Richard O’Gorman. The Middle French redaction of Robert de Boron’s “Joseph d’Arimathie”. Proceedings of the American Philosophical Society, 122(4):261–285, August 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986536.

Ohel:1978:CDP

[Ohe78] Milla Y. Ohel. Clactonian dispersals: a problem. Proceedings of the American Philosophical Society, 122(3):175–181, June 1978. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986551.

Ohlin:1974:CSC

[Ohl74] Lloyd E. Ohlin. Correctional strategies in conflict. Proceedings of the American Philosophical Society, 118(3):248–253, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986442.

Oldson:1994:RAS

[Old94] William O. Oldson. Rationalizing anti-Semitism: The Romanian gambit. Proceedings of the American Philosophical Society, 138(1):

25–30, March 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986703.

Olivier:1950:SRD

[Oli50] Charles P. Olivier. Some results from data secured by the American Meteor Society. Proceedings of the American Philosophical Society, 94(4):327–335, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143516.

Olivier:1957:LEM

[Oli57] Charles P. Olivier. Long enduring meteor trains and fireball orbits. Third paper. Proceedings of the American Philosophical Society, 101(3):296–315, June 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985371.

Olivier:1969:LEM

[Oli69] Charles P. Olivier. Long enduring meteor trains. Fourth paper. Proceedings of the American Philosophical Society, 113(2):127–139, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985961.

Olmsted:1960:VJR

[Olm60] J. W. Olmsted. The voyage of Jean Richer to Acadia in 1670: a study in the relations of science and navigation under Colbert. Proceedings of the American Philosophical Society, 104(6):612–634, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985537.

OMeara:1977:PWS

[O’M77] Dominic J. O’Meara. The philosophical writings, sources, and thought of Athanasius Rhetor (ca. 1571–1663). Proceedings of the American Philosophical Society, 121(6):483–499, December 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986395.

Ong:1976:MLE

[Ong76] Walter J. Ong. Milton’s logical epic and evolving consciousness. Proceedings of the American Philosophical Society, 120(4):295–305, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986322.

Oppenheimer:1982:EHH

[Opp82] Jane M. Oppenheimer. Ernst Heinrich Haeckel as an intermediary in the transmutation of an idea. Proceedings of the American Philosophical Society, 126(5):347–355, October 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986211.

Oppenheimer:1990:SNC

[Opp90] Jane M. Oppenheimer. Science and nationality: The case of Karl Ernst von Baer (1792–1876). Proceedings of the American Philosophical Society, 134(2):75–82, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986838.

Orlans:1971:SPH

[Orl71] Harold Orlans. Science and the polity, or how much knowledge does a nation need? Proceedings of the American Philosophical Society, 115(1):4–9, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985917.

Osborn:1953:NTH

[Osb53] Frederick Osborn. New trends in human evolution. Proceedings of the American Philosophical Society, 97(2):143–146, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143886.

Ostwald:1995:PEW

[Ost95] Martin Ostwald. Public expense: Whose obligation? Athens 600– 454 B.C.E. Proceedings of the American Philosophical Society, 139 (4):368–379, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987236.

Osterbrock:1998:SCO

[Ost98] Donald E. Osterbrock. Subrahmanyan Chandrasekhar (19 October 1910–21 August 1995). Proceedings of the American Philosophical Society, 142(4):658–665, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3152289.

Ostriker:1999:MSM

[Ost99] Jeremiah P. Ostriker. Martin Schwarzschild (31 May 1912–10 April 1997). Proceedings of the American Philosophical Society, 143(3): 486–489, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181962.

Oswalt:1964:TST

[Osw64] Wendell H. Oswalt. Traditional storyknife tales of Yuk girls. Proceedings of the American Philosophical Society, 108(4):310–336, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985911.

Otto:1986:ORC

[Ott86] John Solomon Otto. Open-range cattle-ranching in the Florida pinewoods: a problem in comparative agricultural history. Proceedings of the American Philosophical Society, 130(3):312–324, September 1986. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986829.

Otto:1989:FFS

[Ott89] John S. Otto. Forest fallowing in the Southern Appalachian Mountains: a problem in comparative agricultural history. Proceedings of the American Philosophical Society, 133(1):51–63, March 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987161.

Pace:1950:BFI

[Pac50] Antonio Pace. Benjamin Franklin and Italy since the Eighteenth Century. Proceedings of the American Philosophical Society, 94(3): 242–250, June 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143557.

Pace:1952:MGB

[Pac52] Antonio Pace. The manuscripts of Giambatista Beccaria, correspondent of Benjamin Franklin. Proceedings of the American Philosophical Society, 96(4):406–416, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143797.

Pace:1955:FVC

[Pac55] Antonio Pace. The Franklin–Volta correspondence: Legend or fact? Proceedings of the American Philosophical Society, 99(6): 436–439, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143926.

Pack:1982:MMP

[Pac82] Roger A. Pack. Milo: De Mundi Philosophia. Proceedings of the American Philosophical Society, 126(2):155a–182, April 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986358.

Paddayya:1977:ACH

[Pad77] K. Paddayya. The Acheulian culture of the Hunsgi Valley (Shorapur Doab), Peninsular India. Proceedings of the American Philosophical Society, 121(5):383–406, October 1977. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986343.

Painter:1954:RCE

[Pai54] Theophilus S. Painter. Regional cooperation in education. Proceedings of the American Philosophical Society, 98(4):266–269, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143879.

Palmer:1961:KCS

[Pal61] Robert R. Palmer. The Kingdom of Corsica and the science of history. Proceedings of the American Philosophical Society, 105 (4):354–360, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985494.

Palmer:1978:SSC

[Pal78] Robert E. A. Palmer. Silvanus, Sylvester, and the Chair of St. Peter. Proceedings of the American Philosophical Society, 122(4):222– 247, August 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986534.

Palmer:1981:TSH

[Pal81] Robert E. A. Palmer. The topography and social history of Rome’s Trastevere (southern sector). Proceedings of the American Philosophical Society, 125(5):368–397, October 1981. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986199.

Palmer:1987:TTT

[Pal87] R. R. Palmer. The two Tocquevilles and two kinds of history. Proceedings of the American Philosophical Society, 131(4):341–347, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987135.

Panofsky:1963:IAR

[Pan63] Erwin Panofsky. The ideological antecedents of the Rolls–Royce radiator. Proceedings of the American Philosophical Society, 107 (4):273–288, August 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985670.

Pandey:1972:AZ

[Pan72] Triloki Nath Pandey. Anthropologists at Zuni. Proceedings of the American Philosophical Society, 116(4):321–337, August 1972. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985902.

Panofsky:1989:TLE

[Pan89] Wolfgang K. H. Panofsky. Is there a limit to the energy attainable by high energy colliders for particle physics? Proceedings of the American Philosophical Society, 133(3):416–426, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987109.

Parran:1957:DNV

[Par57] Thomas Parran. Discussion of national, voluntary health agencies. Proceedings of the American Philosophical Society, 101(2):156–158, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985508.

Parsons:1963:CPP

[Par63] Talcott Parsons. On the concept of political power. Proceedings of the American Philosophical Society, 107(3):232–262, June 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985582.

Parsons:1972:SWS

[Par72] Coleman O. Parsons. Sir Walter Scott: Yesterday and today. Proceedings of the American Philosophical Society, 116(6):450–457, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985759.

Pardee:1976:CCN

[Par76] Arthur B. Pardee. Cancer cells and normal cells. Proceedings of the American Philosophical Society, 120(2):87–92, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986514.

Parry:1977:JTH

[Par77] J. H. Parry. Juan de Tovar and the history of the Indians. Proceedings of the American Philosophical Society, 121(4):316–319, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986420.

Parkerson:1983:ICM

[Par83] Phillip T. Parkerson. The Inca coca monopoly: Fact or legal fiction? Proceedings of the American Philosophical Society, 127(2): 107–123, April 1983. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986194.

Paret:1992:GH

[Par92] Peter Paret. God’s hammer. Proceedings of the American Philosophical Society, 136(2):226–246, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987170.

Paret:1993:JBE

[Par93] Peter Paret. Jefferson and the birth of European liberalism. Proceedings of the American Philosophical Society, 137(4):488–497, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987068.

Paret:1996:KBM

[Par96] Peter Paret. Kenneth Bourne (17 March 1930–13 December 1992). Proceedings of the American Philosophical Society, 140(4):564–570, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987292.

Patterson:1970:LCA

[Pat70a] Clair Patterson. Lead contamination of the atmosphere and the Earth’s surface. Proceedings of the American Philosophical Society, 114(1):9, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985719.

Patterson:1970:ACQ

[Pat70b] David S. Patterson. Andrew Carnegie’s quest for world peace. Proceedings of the American Philosophical Society, 114(5):371–383, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985802.

Patterson:1974:CMS

[Pat74] Elizabeth C. Patterson. The case of Mary Somerville: an aspect of Nineteenth-Century science. Proceedings of the American Philosophical Society, 118(3):269–275, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986445.

Patrick:1976:FMB

[Pat76a] Ruth Patrick. The formation and maintenance of benthic diatom communities. Proceedings of the American Philosophical Society, 120(6):475–484, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986600.

Patrides:1976:MAC

[Pat76b] C. A. Patrides. Milton and the Arian controversy. Or, some reflexions on contextual settings and the experience of deuteroscopy. Proceedings of the American Philosophical Society, 120(4):245–252, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986320.

Patrick:1984:STC

[Pat84] Ruth Patrick. Some thoughts concerning the importance of pattern in diverse systems. Proceedings of the American Philosophical Society, 128(1):48–78, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986494.

Patrick:1990:RSP

[Pat90] Ruth Patrick. Reevaluation of solutions to pollution problems. Proceedings of the American Philosophical Society, 134(2):90–94, June 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986840.

Patrick:1994:GEH

[Pat94a] Ruth Patrick. George Evelyn Hutchinson (30 January 1903–17 May 1991). Proceedings of the American Philosophical Society, 138(4): 530–535, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986851.

Patrick:1994:WON

[Pat94b] Ruth Patrick. Is water our next crisis? Proceedings of the American Philosophical Society, 138(3):371–376, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986742.

Patterson:1997:AUP

[Pat97] James T. Patterson. America’s “underclasses,” past and present: a historical perspective. Proceedings of the American Philosophical Society, 141(1):13–29, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987247.

Pauling:1952:HMH

[Pau52] Linus Pauling. The hemoglobin molecule in health and disease. Proceedings of the American Philosophical Society, 96(5):556–565, October 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143656.

Paul:1970:JPR

[Pau70] Charles B. Paul. Jean-Philippe Rameau (1683–1764), the musician as philosophe. Proceedings of the American Philosophical Society, 114(2):140–154, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986031.

Payne:1978:MAR

[Pay78] Harry C. Payne. Modernizing the ancients: The reconstruction of ritual drama 1870–1920. Proceedings of the American Philosophical Society, 122(3):182–192, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986552.

Payne:1981:MS

[Pay81] Harry C. Payne. Malinowski’s style. Proceedings of the American Philosophical Society, 125(6):416–440, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986298.

Payton:1988:VC

[Pay88] Robert L. Payton. Virtue and its consequences. Proceedings of the American Philosophical Society, 132(3):252–257, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143852.

Pleiner:1974:AIA

[PB74] Radomír Pleiner and Judith K. Bjorkman. The Assyrian Iron Age: The history of iron in the Assyrian civilization. Proceedings of the American Philosophical Society, 118(3):283–313, June 1974. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986447.

Pelikan:1961:CCS

[Pel61] Jaroslav Pelikan. Cosmos and creation: Science and theology in reformation thought. Proceedings of the American Philosophical Society, 105(5):464–469, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985545.

Pelikan:1993:HP

[Pel93] Jaroslav Pelikan. The historian as polyglot. Proceedings of the American Philosophical Society, 137(4):659–668, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987080.

Pelikan:1995:BGA

[Pel95] Jaroslav Pelikan. A. Bartlett Giamatti (4 April 1938–1 September 1989). Proceedings of the American Philosophical Society, 139(3): 272–274, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987193.

Pennell:1950:HBC

[Pen50] Francis W. Pennell. Historic botanical collections of the American Philosophical Society and the Academy of Natural Sciences of Philadelphia. Proceedings of the American Philosophical Society, 94(2):137–151, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143214.

Penfield:1954:SOF

[Pen54] Wilder Penfield. Some observations on the functional organization of the human brain. Proceedings of the American Philosophical Society, 98(5):293–297, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143906.

Penfield:1957:APC

[Pen57] Wilder Penfield. The Asclepiad physicians of Cnidus and Cos with a note on the probable site of the Triopion Temple of Apollo. Proceedings of the American Philosophical Society, 101(5):393–400, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985193.

Pendergast:1994:KKI

[Pen94] James F. Pendergast. The Kakouagoga or Kahkwas: an Iroquoian nation destroyed in the Niagara Region. Proceedings of the American Philosophical Society, 138(1):96–144, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986707.

Pepper:1950:BM

[Pep50] O. H. Perry Pepper. Babel in medicine. Proceedings of the American Philosophical Society, 94(4):364–368, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143522.

Perkins:1953:ASP

[Per53] Merle L. Perkins. The Abbé De Saint-Pierre and the Seventeenth-Century intellectual background. Proceedings of the American Philosophical Society, 97(1):69–76, February 1953. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143733.

Perkins:1958:LSC

[Per58] Merle L. Perkins. Late Seventeenth-Century scientific circles and the Abbé de Saint-Pierre. Proceedings of the American Philosophical Society, 102(4):404–412, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985225.

Perkins:1962:UF

[Per62] G. Holmes Perkins. Urban form. Proceedings of the American Philosophical Society, 106(3):190–194, June 1962. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985174.

Perutz:1969:PML

[Per69] M. F. Perutz. Penrose Memorial Lecture. The structure of proteins as revealed by X-ray analysis: a contribution of physics to biology. Proceedings of the American Philosophical Society, 113(4):247–264, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986150.

Perutz:1996:WRD

[Per96] M. F. Perutz. By what right do we invoke human rights? Proceedings of the American Philosophical Society, 140(2):135–147, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987320.

Perrin:1999:RAN

[Per99] Robert G. Perrin. Robert Alexander Nisbet (30 September 1913–9 September 1996). Proceedings of the American Philosophical Society, 143(4):694–710, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3181998.

Peterson:1951:LHH

[Pet51] Charles E. Peterson. Library Hall: Home of the Library Company of Philadelphia 1790–1880. Proceedings of the American Philosophical Society, 95(3):266–285, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143065.

Peterson:1961:TJN

[Pet61] Merrill D. Peterson. Thomas Jefferson and the national purpose. Proceedings of the American Philosophical Society, 105(6):517–520, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985160.

Pettit:1963:EYC

[Pet63] Henry Pettit. Edward Young and the case of “Lee” vs. “D’Aranda”. Proceedings of the American Philosophical Society, 107(2):145–159, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985445.

Petersen:1969:RA

[Pet69] Howard C. Petersen. The revolution in aspirations. Proceedings of the American Philosophical Society, 113(4):265–268, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986151.

Peterson:1973:WNP

[Pet73] Willard J. Peterson. Western natural philosophy published in Late Ming China. Proceedings of the American Philosophical Society, 117(4):295–322, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986697.

Peters:1999:RSC

[Pet99] Ellen Ash Peters. The role of state constitutions in our federal system. Proceedings of the American Philosophical Society, 143(3): 418–427, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181954.

Peyre:1958:ACA

[Pey58] Henri Peyre. Albert Camus, an anti-Christian moralist. Proceedings of the American Philosophical Society, 102(5):477–482, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985594.

Peyre:1970:PCR

[Pey70] Henri Peyre. Paul Claudel, a Romantic in spite of himself. Proceedings of the American Philosophical Society, 114(3):179–186, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985694.

Payne-Gaposchkin:1952:WC

[PG52] Cecilia Payne-Gaposchkin. “Worlds in collision”. Proceedings of the American Philosophical Society, 96(5):519–525, October 1952. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143652.

Philbrick:1953:NEE

[Phi53] Francis S. Philbrick. Notes on early editions and editors of Franklin. Proceedings of the American Philosophical Society, 97(5):525–564, October 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3149269.

Pick:1954:EHP

[Pic54] Joseph Pick. The evolution of homeostasis: The phylogenetic development of the regulation of bodily and mental activities by the autonomic nervous system. Proceedings of the American Philosophical Society, 98(5):298–303, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143907.

Pickering:1966:NMS

[Pic66] W. H. Pickering. Navigating The Mariner Spacecraft to Mars. Proceedings of the American Philosophical Society, 110(5):332–339, October 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986022.

Piel:1965:TC

[Pie65] Gerard Piel. The treason of the clerks. Proceedings of the American Philosophical Society, 109(5):259–266, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986077.

Piel:1968:IRS

[Pie68] Gerard Piel. Introductory remarks to symposium on gondwanaland revisited: New evidence for continental drift. Proceedings of the American Philosophical Society, 112(5):307–308, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986050.

Piel:1979:PUK

[Pie79] Gerard Piel. On promoting useful knowledge. Proceedings of the American Philosophical Society, 123(6):337–340, December 1979.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986613.

Piel:1980:SWW

[Pie80] Gerard Piel. Symposium: Weapons and war in the decade ahead: Introductory remarks. Proceedings of the American Philosophical Society, 124(4):256, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986425.

Piel:1995:I

[Pie95] Gerard Piel. Introduction. Proceedings of the American Philosophical Society, 139(3):231–232, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987187.

Piel:1999:I

[Pie99] Gerard Piel. Introduction. Proceedings of the American Philosophical Society, 143(2):159–160, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181929.

Pingree:1978:IA

[Pin78] David Pingree. Indian astronomy. Proceedings of the American Philosophical Society, 122(6):361–364, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986451.

Pingree:1999:AP

[Pin99] David Pingree. An astronomer’s progress. Proceedings of the American Philosophical Society, 143(1):73–85, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181975.

Pleasants:1967:TJE

[Ple67] Samuel A. Pleasants. Thomas Jefferson — educational philosopher. Proceedings of the American Philosophical Society, 111(1): 1–4, February 1967. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985750.

Plotkin:1978:AVN

[Plo78] Howard Plotkin. Astronomers versus the Navy: The revolt of American astronomers over the management of the United States Naval Observatory, 1877–1902. Proceedings of the American Philosophical Society, 122(6):385–399, December 1978. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986454.

Plummer:1977:FAB

[Plu77] Gayther L. Plummer. “Franklinia Alatamaha” Bartram ex Marshall: The Lost “Gordonia” (Theaceae). Proceedings of the American Philosophical Society, 121(6):475–482, December 1977. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986394.

Poesch:1956:MPF

[Poe56] Jessie J. Poesch. Mr. Peale’s “farm persevere”: Some documentary views. Proceedings of the American Philosophical Society, 100(6): 545–556, December 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143788.

Pollard:1977:AUZ

[Pol77] Helen Perlstein Pollard. An analysis of urban zoning and planning at prehispanic Tzintzuntzan. Proceedings of the American Philosophical Society, 121(1):46–69, February 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986566.

Pollak:1985:MLH

[Pol85] Louis H. Pollak. The making of law: Hierarchies of courts and hierarchies of values. Proceedings of the American Philosophical Society, 129(1):82–89, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986980.

Pomerantz:1954:GWI

[Pom54] Sidney I. Pomerantz. George Washington and the inception of aeronautics in the Young Republic. Proceedings of the American Philosophical Society, 98(2):131–138, April 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143643.

Paul:1997:G

[Pop97] Pope John Paul, II. Greeting. Proceedings of the American Philosophical Society, 141(4):389, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987212.

Porter:1974:WBR

[Por74] William W. Porter, II. William Bartram’s “Road to Xanadu” on modern Florida maps. Proceedings of the American Philosophical Society, 118(6):514–518, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986401.

Porter:1985:LTR

[Por85] Charlotte M. Porter. The lifework of Titian Ramsay Peale. Proceedings of the American Philosophical Society, 129(3):300–312, September 1985. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987013.

Porter:1992:CSP

[Por92] Lord Porter. Can science policy be left to the scientists? Proceedings of the American Philosophical Society, 136(4):521–525, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986760.

Porter:1996:PPR

[Por96] Barbara Nevling Porter. Politics and public relations campaigns in Ancient Assyria: King Esarhaddon and Babylonia. Proceedings of the American Philosophical Society, 140(2):164–174, June 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987322.

Pottle:1962:NIP

[Pot62] Frederick A. Pottle. Notes on the importance of private legal documents for the writing of biography and literary history. Proceedings of the American Philosophical Society, 106(4):327–334, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985266.

Pottle:1972:WPD

[Pot72] Frederick A. Pottle. Wordsworth in the present day. Proceedings of the American Philosophical Society, 116(6):443–449, December 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985758.

Poteet:1975:ULS

[Pot75] James M. Poteet. Unrest in the “land of steady habits”: The Hartford Riot of 1722. Proceedings of the American Philosophical Society, 119(3):223–232, June 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986672.

Pounds:1997:IPX

[Pou97] K. A. Pounds. International programmes in X-ray astronomy. Proceedings of the American Philosophical Society, 141(1):67–76, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987250.

Power:1995:EBP

[Pow95] Philip H. Power. Eugene Barnum Power (4 June 1905–6 December 1993). Proceedings of the American Philosophical Society, 139(3): 300–304, September 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987197.

Prager:1968:MTQ

[Pra68] Frank D. Prager. A manuscript of Taccola, quoting Brunelleschi, on problems of inventors and builders. Proceedings of the American Philosophical Society, 112(3):131–149, June 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986159.

Price:1962:SE

[Pri62] Don K. Price. The scientific establishment. Proceedings of the American Philosophical Society, 106(3):235–245, June 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985178.

Price:1998:CSA

[Pri98] Hugh B. Price. Cities: The soul of America. Proceedings of the American Philosophical Society, 142(3):388–395, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152244.

Proskouriakoff:1984:IAM

[Pro84] Tatiana Proskouriakoff. Incidents of Ancient Maya history. Proceedings of the American Philosophical Society, 128(2):164–166, June 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986228.

Perkins:1955:LSP

[PSP55] Merle L. Perkins and Abbé Saint-Pierre. The “Leviathan” and Saint-Pierre’s “projet de paix perpétuelle”. Proceedings of the American Philosophical Society, 99(4):259–267, August 1955. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143705.

Puknat:1957:CGT

[Puk57] Siegfried B. Puknat. Channing and German thought. Proceedings of the American Philosophical Society, 101(2):195–203, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985512.

Puknat:1958:WNE

[Puk58] Siegfried B. Puknat. De Wette in New England. Proceedings of the American Philosophical Society, 102(4):376–395, August 1958. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985222.

Pursell:1965:AFS

[Pur65] Carroll W. Pursell, Jr. The anatomy of a failure: The Science Advisory Board, 1933–1935. Proceedings of the American Philosophical Society, 109(6):342–351, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986136.

Putnam:1984:PE

[Put84] Hilary Putnam. Proof and experience. Proceedings of the American Philosophical Society, 128(1):31–34, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986490.

Pye:1978:ALC

[Pye78a] Lucian W. Pye. Aesopian language in Chinese politics. Proceedings of the American Philosophical Society, 122(5):336–339, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986688.

Pyenson:1978:ITE

[Pye78b] Lewis Pyenson. The incomplete transmission of a European image: Physics at Greater Buenos Aires and Montréal, 1890–1920. Proceedings of the American Philosophical Society, 122(2):92–114, April 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986625.

Quinlan:1953:MWC

[QC53] Maurice J. Quinlan and William Cowper. Memoir of William Cowper. Proceedings of the American Philosophical Society, 97(4): 359–382, September 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3149291.

Quine:1960:VE

[Qui60] Willard V. Quine. Variables explained away. Proceedings of the American Philosophical Society, 104(3):343–347, June 1960. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985250.

Radda:1993:MHD

[Rad93] George K. Radda. From molecules to human disease: a new window through magnetism. Proceedings of the American Philosophical Society, 137(4):587–601, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987075.

Rainey:1966:NTA

[Rai66] Froelich Rainey. New techniques in archaeology. Proceedings of the American Philosophical Society, 110(2):146–152, April 1966. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985747.

Raimi:1985:FDP

[Rai85] Ralph A. Raimi. The first digit phenomenon again. Proceedings of the American Philosophical Society, 129(2):211–219, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986989.

Rainger:1986:JBS

[Rai86] Ronald Rainger. Just before Simpson: William Diller Matthew’s understanding of evolution. Proceedings of the American Philosophical Society, 130(4):453–474, December 1986. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986929.

Rainger:1992:RDS

[Rai92] Ronald Rainger. The rise and decline of a science: Vertebrate paleontology at Philadelphia’s Academy of Natural Sciences, 1820– 1900. Proceedings of the American Philosophical Society, 136(1):1– 32, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986795.

Ramsey:1969:NAO

[Ram69] Elizabeth M. Ramsey. New appraisal of an old organ: The placenta. Proceedings of the American Philosophical Society, 113(4):296–304, August 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986154.

Ramo:1977:SAP

[Ram77] Simon Ramo. On a scientific approach to prediction. Proceedings of the American Philosophical Society, 121(4):302–308, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986418.

Ramsey:1999:EMP

[Ram99] Norman F. Ramsey. Edward Mills Purcell (30 August 1912–7 March 1997). Proceedings of the American Philosophical Society, 143(3):480–483, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181961.

Randle:1953:WCI

[Ran53] Martha Champion Randle. The Waugh Collection of Iroquois folktales. Proceedings of the American Philosophical Society, 97(5): 611–633, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3149274.

Randel:1970:HA

[Ran70] William Peirce Randel. Huxley in America. Proceedings of the American Philosophical Society, 114(2):73–99, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986027.

Raper:1960:LCI

[Rap60] Kenneth B. Raper. Levels of cellular interaction in amoeboid populations. Proceedings of the American Philosophical Society, 104 (6):579–604, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985535.

Ratcliffe:1968:CZE

[Rat68] Herbert L. Ratcliffe. Contribution of a zoo to an ecology of disease. Proceedings of the American Philosophical Society, 112(4):235–244, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985871.

Ratliff:1971:CC

[Rat71] Floyd Ratliff. Contour and contrast. Proceedings of the American Philosophical Society, 115(2):150–163, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985854.

Ratliff:1976:PBU

[Rat76] Floyd Ratliff. On the psychophysiological bases of universal color terms. Proceedings of the American Philosophical Society, 120(5): 311–330, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986265.

Rauschenberg:1973:JJB

[Rau73] Roy A. Rauschenberg. The journals of Joseph Banks’s voyage up Great Britain’s West Coast to Iceland and to the Orkney Isles, July to October, 1772. Proceedings of the American Philosophical Society, 117(3):186–226, June 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986542.

Ravitch:1985:PSC

[Rav85] Diane Ravitch. Politicization and the schools: The case of bilingual education. Proceedings of the American Philosophical Society, 129(2):121–128, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986985.

Ravitch:1991:LCO

[Rav91] Diane Ravitch. Lawrence A. Cremin (October 31, 1925–September 4, 1990). Proceedings of the American Philosophical Society, 135 (4):600–603, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986820.

Rawson:1978:MSC

[Raw78] Don C. Rawson. Mendeleev and the scientific claims of spiritualism. Proceedings of the American Philosophical Society, 122(1):1–8, February 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986258.

Rayback:1980:MRE

[Ray80] Joseph G. Rayback. A myth re-examined: Martin van Buren’s role in the Presidential Election of 1816. Proceedings of the American Philosophical Society, 124(2):106–118, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986206.

Roberts:1996:LAD

[RB96] John D. Roberts and Harold Brown. Lee Alvin Dubridge (21 September 1901–23 January 1993). Proceedings of the American Philosophical Society, 140(2):230–238, June 1996. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987328.

Rudwick:1967:CUG

[RC67] M. J. S. Rudwick and W. D. Conybeare. A critique of uniformitarian geology: a letter from W. D. Conybeare to Charles Lyell, 1841. Proceedings of the American Philosophical Society, 111(5):272–287, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986045.

Roberts:1999:OLP

[RC99a] John D. Roberts and Francis H. Clauser. An outsider’s look at a principal paradigm of modern cosmology. Proceedings of the American Philosophical Society, 143(1):12–27, March 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181970.

Rovee-Collier:1999:RMM

[RC99b] Carolyn Rovee-Collier. The roots of multiple memory systems. Proceedings of the American Philosophical Society, 143(2):266–279, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181937.

Rhoads:1973:WGG

[RD73] Jonathan E. Rhoads and Stanley J. Dudrick. Weight gain, growth, and development induced in man by the intravenous administration of pure chemicals. Proceedings of the American Philosophical Society, 117(3):152–161, June 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986540.

Redfield:1952:PWV

[Red52] Robert Redfield. The primitive world view. Proceedings of the American Philosophical Society, 96(1):30–36, February 1952. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143740.

Reder:1978:CDU

[Red78] Melvin W. Reder. The coming demise of unemployment. Proceedings of the American Philosophical Society, 122(3):139–144, June 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986548.

Reedy:1986:ASS

[Ree86] W. Jay Reedy. Art for society’s sake: Louis de Bonald’s sociology of aesthetics and the theocratic ideology. Proceedings of the American Philosophical Society, 130(1):101–129, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987093.

Rees:1993:BBC

[Ree93] Martin J. Rees. From the “big bang” to a complex cosmos: Galaxies, black holes, and dark matter. Proceedings of the American Philosophical Society, 137(4):545–557, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987073.

Rehn:1954:JEL

[Reh54] James A. G. Rehn. The John Eatton Leconte Collection of paintings of insects, arachnids, and myriopods. Proceedings of the American Philosophical Society, 98(6):442–448, December 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143866.

Reichard:1950:ECP

[Rei50] Gladys A. Reichard. The Elsie Clews Parsons Collection. Proceedings of the American Philosophical Society, 94(3):308–309, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143567.

Reinhold:1968:OCL

[Rei68] Meyer Reinhold. Opponents of classical learning in America during the Revolutionary Period. Proceedings of the American Philosophical Society, 112(4):221–234, August 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985870.

Reinhold:1970:GGA

[Rei70] Meyer Reinhold. The generation gap in antiquity. Proceedings of the American Philosophical Society, 114(5):347–365, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985800.

Reinhold:1975:QUK

[Rei75] Meyer Reinhold. The quest for “useful knowledge” in Eighteenth-Century America. Proceedings of the American Philosophical Society, 119(2):108–132, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986630.

Reiner:1986:WDY

[Rei86] Erica Reiner. Why do you cuss me? Proceedings of the American Philosophical Society, 130(1):1–6, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987088.

Reingold:1990:TA

[Rei90] Nathan Reingold. Tales from the archives. Proceedings of the American Philosophical Society, 134(4):340–348, December 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986892.

Reimers:1992:HRI

[Rei92] David H. Reimers. History of recent immigration regulation. Proceedings of the American Philosophical Society, 136(2):176–187, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987167.

Reuter:1997:WCW

[Reu97] Peter Reuter. Why can’t we make prohibition work better? Some consequences of ignoring the unattractive. Proceedings of the American Philosophical Society, 141(3):262–275, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987202.

Ronen:1972:UAK

[RGSS72] A. Ronen, D. Gilead, E. Shachnai, and A. Saull. Upper Acheulean in the Kissufim Region. Proceedings of the American Philosophical Society, 116(1):68–96, February 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985707.

Robbins:1958:AMN

[RH58] William J. Robbins and Mary Christine Howson. André Michaux’s New Jersey garden and Pierre Paul Saunier, journeyman gardener. Proceedings of the American Philosophical Society, 102(4):351–370, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985220.

Rhoads:1976:PRS

[Rho76] Jonathan E. Rhoads. Preliminary remarks for the symposium on new approaches in the study of neoplasia. Proceedings of the American Philosophical Society, 120(2):67–68, April 1976. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986512.

Rhoads:1988:MR

[Rho88] Jonathan E. Rhoads. Moderator’s remarks. Proceedings of the American Philosophical Society, 132(2):137–139, June 1988. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986858.

Rhodes:1998:ALP

[Rho98] Frank H. T. Rhodes. The advancement of learning: Prospects in a cynical age. Proceedings of the American Philosophical Society, 142(2):218–243, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152300.

Richards:1950:PMS

[Ric50a] Horace G. Richards. Postglacial marine submergence of Arctic North America with special reference to the Mackenzie Delta. Proceedings of the American Philosophical Society, 94(1):31–37, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143249.

Rickett:1950:JBE

[Ric50b] H. W. Rickett. John Bradbury’s explorations in Missouri Territory. Proceedings of the American Philosophical Society, 94(1):59– 89, February 1950. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143252.

Rice:1951:JGF

[Ric51a] Howard C. Rice, Jr. Jefferson’s gift of fossils to the Museum of Natural History in Paris. Proceedings of the American Philosophical Society, 95(6):597–627, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143509.

Richter:1951:WMR

[Ric51b] Gisela M. A. Richter. Who made the Roman portrait statues — Greeks or Romans? Proceedings of the American Philosophical Society, 95(2):184–191 + 193–208, April 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143332.

Rice:1953:WEC

[Ric53] Madeleine Hooke Rice. William Ellery Channing: The making of a social gospel. Proceedings of the American Philosophical Society, 97(1):31–43, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143728.

Richter:1968:ERR

[Ric68] Curt P. Richter. Experiences of a reluctant rat-catcher the common Norway rat — friend or enemy? Proceedings of the American Philosophical Society, 112(6):403–415, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985939.

Richards:1970:CSD

[Ric70] Horace G. Richards. Changes in shoreline during the past million years. Proceedings of the American Philosophical Society, 114(3): 198–204, June 1970. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985697.

Richards:1980:CWP

[Ric80] Robert J. Richards. Christian Wolff’s prolegomena to empirical and rational psychology: Translation and commentary. Proceedings of the American Philosophical Society, 124(3):227–239, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986371.

Riddick:1952:DUM

[Rid52] Thomas M. Riddick. Dowsing: an unorthodox method of locating underground water supplies or an interesting facet of the human mind. Proceedings of the American Philosophical Society, 96 (5):526–534, October 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143653.

Rieber:1984:BBC

[Rie84] Alfred J. Rieber. Businessmen and business culture in Imperial Russia. Proceedings of the American Philosophical Society, 128(3): 238–243, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986883.

Riley:1951:DFC

[Ril51] Edward M. Riley. The Deborah Franklin correspondence. Proceedings of the American Philosophical Society, 95(3):239–245, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143060.

Riley:1988:STT

[Ril88a] Mark Riley. Science and tradition in the “Tetrabiblos”. Proceedings of the American Philosophical Society, 132(1):67–84, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143825.

Riley:1988:ASP

[Ril88b] Matilda White Riley. The aging society: Problems and prospects. Proceedings of the American Philosophical Society, 132(2):148–153, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986860.

Ringe:1993:RPG

[Rin93] Donald A. Ringe, Jr. A reply to Professor Greenberg. Proceedings of the American Philosophical Society, 137(1):91–109, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986947.

Rivers:1954:SRP

[Riv54] Thomas M. Rivers. The story of research on poliomyelitis. Proceedings of the American Philosophical Society, 98(4):250–254, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143877.

Rivas:1958:OED

[Riv58] Luis Rene Rivas. The origin, evolution, dispersal, and geographical distribution of the Cuban poeciliid fishes of the Tribe Girardinini. Proceedings of the American Philosophical Society, 102(3):281–320, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985577.

Rose:1982:CNF

[RK82] Kenneth D. Rose and David W. Krause. Cyriacotheriidae, a new family of early tertiary pantodonts from Western North America. Proceedings of the American Philosophical Society, 126(1):26–50, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986477.

Roach:1966:TGT

[Roa66] William Roach. Transformations of the grail theme in the first two continuations of the Old French “Perceval”. Proceedings of the American Philosophical Society, 110(3):160–164, June 1966. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985984.

Roach:1970:FMP

[Roa70] William Roach. Francisque Michel: a pioneer in medieval studies. Proceedings of the American Philosophical Society, 114(3):168–178, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985693.

Roach:1998:BBD

[Roa98] Eleanor Roach. Bernhard Bischoff (20 December 1906–17 September 1991). Proceedings of the American Philosophical Society, 142 (2):266–273, June 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152303.

Robinson:1950:JDB

[Rob50] H. W. Robinson. Jeremiah Dixon (1733–1779): a biographical note. Proceedings of the American Philosophical Society, 94(3):272–274, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143562.

Robbins:1953:TBH

[Rob53] Caroline Robbins. Thomas Brand Hollis (1719–1804), English admirer of Franklin and intimate of John Adams. Proceedings of the American Philosophical Society, 97(3):239–247, June 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143746.

Robinson:1955:SVU

[Rob55] Geroid Tanquary Robinson. Stalin’s vision of utopia: The future communist society. Proceedings of the American Philosophical Society, 99(1):11–21, January 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143688.

Robbins:1957:FBF

[Rob57a] William J. Robbins. French botanists and the flora of the Northeastern United States: J. G. Milbert and Elias Durand. Proceedings of the American Philosophical Society, 101(4):362–368, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985402.

Robbins:1957:RDW

[Rob57b] William J. Robbins. The retirement of Dr. William E. Lingelbach. Proceedings of the American Philosophical Society, 101(6):506–507, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985518.

Robbins:1960:DHH

[Rob60] Christine Chapman Robbins. David Hosack’s herbarium and its Linnaean specimens. Proceedings of the American Philosophical Society, 104(3):293–313, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985245.

Robbins:1962:HHJ

[Rob62] Caroline Robbins. Honest heretic: Joseph Priestley in America, 1794–1804. Proceedings of the American Philosophical Society, 106 (1):60–76, February 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985212.

Robbins:1964:TPS

[Rob64] William J. Robbins. Topophysis, a problem in somatic inheritance. Proceedings of the American Philosophical Society, 108(5):395–403, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985812.

Robel:1967:SBG

[Rob67a] Robert J. Robel. Significance of booming grounds of greater prairie chickens. Proceedings of the American Philosophical Society, 111 (2):109–114, April 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986037.

Roberts:1967:MLS

[Rob67b] Richard Brooke Roberts. Memory and learning from the standpoint of computer model building. Proceedings of the American Philosophical Society, 111(6):352–358, December 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985732.

Roberts:1972:APM

[Rob72] William I. Roberts III. American potash manufacture before the American Revolution. Proceedings of the American Philosophical Society, 116(5):383–395, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986069.

Roberts:1976:CCQ

[Rob76] Walter Orr Roberts. Climate change and the quality of life for the Earth’s new millions. Proceedings of the American Philosophical Society, 120(3):230–232, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986562.

Roberts:1979:PML

[Rob79] Walter Orr Roberts. Penrose Memorial Lecture. Variations in the Sun and their effects on weather and climate. Proceedings of the American Philosophical Society, 123(3):151–159, June 1979. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986217.

Robbins:1983:LET

[Rob83a] Caroline Robbins. The lifelong education of Thomas Paine (1737– 1809): Some reflections upon his acquaintance among books. Proceedings of the American Philosophical Society, 127(3):135–142, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986383.

Robinson:1983:BMR

[Rob83b] Denis M. Robinson. British microwave radar 1939–41. Proceedings of the American Philosophical Society, 127(1):26–31, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986362.

Robertson:1986:PPB

[Rob86] Rutherford Robertson. Purple pigment, bacteria, sheep’s eyes and serendipity. Proceedings of the American Philosophical Society, 130 (4):374–381, December 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986924.

Roberts:1989:UKA

[Rob89] John D. Roberts. Useful knowledge about the application of nuclear magnetic resonance to medicine. Proceedings of the American Philosophical Society, 133(4):489–502, December 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986873.

Robbins:1991:JFE

[Rob91a] Frederick C. Robbins. John Franklin Enders (February 10, 1897– September 8, 1985). Proceedings of the American Philosophical Society, 135(3):452–457, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986779.

Robbins:1991:JHN

[Rob91b] Frederick C. Robbins. John Howard Northrop (July 5, 1891–May 27, 1987). Proceedings of the American Philosophical Society, 135 (2):314–320, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987038.

Roberts:1998:RWS

[Rob98] Paul Craig Roberts. A reconsideration of the welfare state. Proceedings of the American Philosophical Society, 142(3):396–398, September 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152245.

Rockefeller:1963:CRB

[Roc63] David Rockefeller. The changing role of business in our society. Proceedings of the American Philosophical Society, 107(2):112–116, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985440.

Rocher:1993:LBO

[Roc93a] Ludo Rocher. Law books in an oral culture: The Indian “Dharmaśāstras”. Proceedings of the American Philosophical Society, 137(2):254–267, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986732.

Rockefeller:1993:HJH

[Roc93b] John D. Rockefeller. Henry John Heinz III (23 October 1938–4 April 1991). Proceedings of the American Philosophical Society, 137(3):440–444, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987003.

Rochberg:1997:PMS

[Roc97] George Rochberg. Polarity in music: Symmetry and asymmetry and their consequences. Proceedings of the American Philosophical Society, 141(2):169–176, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987300.

Rodgers:1992:PCS

[Rod92] John Rodgers. Preston Cloud (September 26, 1912–January 16, 1991). Proceedings of the American Philosophical Society, 136(2): 280–287, June 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987173.

Rogers:1951:AL

[Rog51] Lindsay Rogers. Authority and liberty, 1950. Proceedings of the American Philosophical Society, 95(5):504–511, October 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143234.

Rolph:1956:PFM

[Rol56] S. Wyman Rolph. Presentation of the Franklin Medal to the American Philosophical Society. Proceedings of the American Philosophical Society, 100(4):340–341, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143761.

Romer:1956:EEL

[Rom56] Alfred Sherwood Romer. The early evolution of land vertebrates. Proceedings of the American Philosophical Society, 100(3):157–167, June 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143770.

Romer:1968:FG

[Rom68] Alfred Sherwood Romer. Fossils and Gondwanaland. Proceedings of the American Philosophical Society, 112(5):335–343, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986054.

Roosa:1972:NSS

[Roo72] Robert V. Roosa. The new services state: The changing role of the United States in the international economic system. Proceedings of the American Philosophical Society, 116(3):203–215, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986115.

Roosa:1990:DBP

[Roo90] Robert V. Roosa. The discipline of the balance of payments. Proceedings of the American Philosophical Society, 134(1):30– 35, March 1990. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143903.

Root:1994:SPC

[Roo94] Terry L. Root. Scientific/philosophical challenges of global change research: a case study of climatic changes on birds. Proceedings of the American Philosophical Society, 138(3):377–384, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986743.

Rosen:1954:DGC

[Ros54] Edward Rosen. Did Galileo claim he invented the telescope? Proceedings of the American Philosophical Society, 98(5):304–312, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143908.

Rosen:1957:MAT

[Ros57] Edward Rosen. Maurolico’s attitude toward Copernicus. Proceedings of the American Philosophical Society, 101(2):177–194, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985511.

Rosen:1960:CWP

[Ros60] Edward Rosen. Copernicus was not a priest. Proceedings of the American Philosophical Society, 104(6):635–661, December 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985538.

Rosenthal:1964:GBM

[Ros64] Franz Rosenthal. Gifts and bribes: The Muslim view. Proceedings of the American Philosophical Society, 108(2):135–144, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985891.

Rosenthal:1966:TGA

[Ros66] Franz Rosenthal. Two Graeco–Arabic works on music. Proceedings of the American Philosophical Society, 110(4):261–268, August 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985688.

Rossi:1975:SLH

[Ros75] John P. Rossi. The selection of Lord Hartington as Liberal leader in the House of Commons, February, 1875. Proceedings of the American Philosophical Society, 119(4):307–314, August 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986292.

Rose:1981:EPL

[Ros81] David J. Rose. Energy prospects for the long term. Proceedings of the American Philosophical Society, 125(4):269–273, August 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986328.

Rosenwaike:1989:DCJ

[Ros89] Ira Rosenwaike. Demographic characteristics of the Jewish population of the United States and Philadelphia in 1830. Proceedings of the American Philosophical Society, 133(2):333–338, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987059.

Rosenwald:1990:ALB

[Ros90] Julius Rosenwald, II. An amateur looks at birds. Proceedings of the American Philosophical Society, 134(1):19–23, March 1990. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143901.

Rosand:1998:MSS

[Ros98] David Rosand. Meyer Schapiro (23 September 1904–3 March 1996). Proceedings of the American Philosophical Society, 142(3):494–498, September 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152255.

Rotz:1977:LUD

[Rot77] Rhiman A. Rotz. The Lubeck Uprising of 1408 and the decline of the Hanseatic League. Proceedings of the American Philosophical Society, 121(1):1–45, February 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986565.

Rowe:1991:IJR

[Row91] G. S. Rowe. Infanticide, its judicial resolution, and criminal code revision in Early Pennsylvania. Proceedings of the American Philosophical Society, 135(2):200–232, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987032.

Roy:1996:HGC

[Roy96] David T. Roy. Herrlee Glessner Creel (19 January 1905–1 June 1994). Proceedings of the American Philosophical Society, 140(1): 94–97, March 1996. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987279.

Rostow:1992:LMA

[RR92a] W. W. Rostow and Elspeth D. Rostow. Letter to the Mayor of Austin. Proceedings of the American Philosophical Society, 136(3): 355–357, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986907.

Rostow:1992:RHB

[RR92b] W. W. Rostow and Elspeth D. Rostow. Reflections on how to bring people in the disadvantaged communities into the mainstream. Proceedings of the American Philosophical Society, 136(3): 359–377, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986909.

Roberts:1999:MSS

[RT99] Priscilla H. Roberts and James N. Tull. Moroccan Sultan Sidi Muhammad Ibn Abdallah’s diplomatic initiatives toward the United States, 1777–1786. Proceedings of the American Philosophical Society, 143(2):233–265, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3181936.

Rubinstein:1963:SIU

[Rub63] Alvin Z. Rubinstein. The Soviet image of the United Nations. Proceedings of the American Philosophical Society, 107(2):132–137, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985443.

Rubin:1988:DMUa

[Rub88a] Vera C. Rubin. Dark matter in the universe. Proceedings of the American Philosophical Society, 132(3):258–267, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143853.

Rubin:1988:DMUb

[Rub88b] Vera C. Rubin. Dark matter in the universe. Proceedings of the American Philosophical Society, 132(4):434–443, December 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986963.

Russell:1957:VPE

[Rus57] John M. Russell. The volunteer, the professional executive, and the scientist in the national volunteer health agencies. Proceedings of the American Philosophical Society, 101(2):153–155, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985507.

Russell:1962:MMA

[Rus62] J. C. Russell. The medieval Monedatge of Aragon and Valencia. Proceedings of the American Philosophical Society, 106(6):483–504, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985255.

Russell:1963:GHR

[Rus63] W. L. Russell. Genetic hazards of radiation. Proceedings of the American Philosophical Society, 107(1):11–17, February 1963. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985465.

Salinger:1950:WFO

[Sal50] Gerard Salinger. Was the futūwa an oriental form of chivalry? Proceedings of the American Philosophical Society, 94(5):481–493, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143614.

Salzman:1967:POA

[Sal67] Philip C. Salzman. Political organization among nomadic peoples. Proceedings of the American Philosophical Society, 111(2):115–131, April 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986038.

Salvini:1992:EAS

[Sal92] Giorgio Salvini. Edoardo Amaldi (5 September 1908–5 December 1989). Proceedings of the American Philosophical Society, 136(3): 424–429, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986916.

Salvini:1997:W

[Sal97] Giorgio Salvini. Welcome. Proceedings of the American Philosophical Society, 141(4):390–391, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987213.

Salpeter:1998:WFA

[Sal98] E. E. Salpeter. William A. Fowler (9 August 1911–14 March 1995). Proceedings of the American Philosophical Society, 142 (1):122–126, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152268.

Samuelsson:1999:NDR

[Sam99] Bengt Samuelsson. New drugs: The role of basic science. Proceedings of the American Philosophical Society, 143(1):50–51, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181973.

Sarma:1974:HPS

[Sar74] Akkaraju V. N. Sarma. Holocene paleoecology of South Coastal Ecuador. Proceedings of the American Philosophical Society, 118 (1):93–134, February 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986437.

Sarma:1978:PCT

[Sar78] Akkaraju Sarma. The paleoecology of Coastal Tamilnadu, South India: Chronology of raised beaches. Proceedings of the American Philosophical Society, 122(6):411–426, December 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986456.

Sasso:1980:SCP

[Sas80] James Sasso. The stages of the creative process. Proceedings of the American Philosophical Society, 124(2):119–132, April 1980. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986207.

Satin:1982:ELH

[Sat82] David George Satin. Erich Lindemann: The humanist and the era of community mental health. Proceedings of the American Philosophical Society, 126(4):327–346, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986178.

Sauer:1962:SPH

[Sau62] Carl O. Sauer. Seashore-primitive home of man? Proceedings of the American Philosophical Society, 106(1):41–47, February 1962. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985209.

Sawyer:1989:HFF

[Saw89] John E. Sawyer. How firm are foundations? Proceedings of the American Philosophical Society, 133(3):339–346, September 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987103.

Sawhill:1992:AUO

[Saw92] Isabel V. Sawhill. Appendix a: The underclass: an overview. Proceedings of the American Philosophical Society, 136(3):380–389, September 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986911.

Saxon:1996:JBW

[Sax96] David S. Saxon. Jerome Bert Wiesner (30 May 1915–21 October 1994). Proceedings of the American Philosophical Society, 140(2): 254–262, June 1996. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987331.

Sayles:1971:KRI

[Say71] George Osborne Sayles. King Richard II of England: a fresh look. Proceedings of the American Philosophical Society, 115(1):28–31, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985921.

Stapleton:1984:IBC

[SC84] Darwin H. Stapleton and Edward C. Carter, II. “I have the itch of botany, of chemistry, of mathematics … strong upon me”: The science of Benjamin Henry Latrobe. Proceedings of the American Philosophical Society, 128(3):173–192, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986879.

Simler:1989:BPM

[SC89] Lucy Simler and Paul G. E. Clemens. The “best poor man’s country” in 1783: The population structure of rural society in Late-Eighteenth-Century Southeastern Pennsylvania. Proceedings of the American Philosophical Society, 133(2):234–261, June 1989. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987054.

Scaglia:1976:MBW

[Sca76] Gustina Scaglia. A miscellany of bronze works and texts in the “Zibaldone” of Buonaccorso Ghiberti. Proceedings of the American Philosophical Society, 120(6):485–513, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986601.

Scarman:1987:DRF

[Sca87] Lord Scarman. The Dr. Richard A. F. Penrose Memorial Lecture: Remarks on the Constitutional celebration. Proceedings of the American Philosophical Society, 131(3):225–236, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987017.

Schultz:1950:PDM

[Sch50] Adolph H. Schultz. The physical distinctions of man. Proceedings of the American Philosophical Society, 94(5):428–449, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143610.

Schuyler:1953:BIT

[Sch53] Robert Livingston Schuyler. British Imperial theory and American territorial policy. A suggested relationship. Proceedings of the American Philosophical Society, 97(4):317–331, September 1953.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3149287.

Schlesinger:1955:PMA

[Sch55] Arthur Meier Schlesinger. Political mobs and the American Revolution, 1765–1776. Proceedings of the American Philosophical Society, 99(4):244–250, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143703.

Schmitt:1956:MIP

[Sch56a] Francis O. Schmitt. Macromolecular interaction patterns in biological systems. Proceedings of the American Philosophical Society, 100 (5):476–486, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143681.

Schramm:1956:MFA

[Sch56b] J. R. Schramm. The memorial to François André Michaux at the Morris Arboretum, University of Pennsylvania. Proceedings of the American Philosophical Society, 100(2):145–149, April 1956. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143714.

Schmidt:1957:HI

[Sch57a] Albert J. Schmidt. A household inventory, 1581. Proceedings of the American Philosophical Society, 101(5):459–480, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985201.

Schneirla:1957:TCC

[Sch57b] T. C. Schneirla. Theoretical consideration of cyclic processes in Doryline ants. Proceedings of the American Philosophical Society, 101(1):106–133, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985143.

Schramm:1957:IPP

[Sch57c] J. R. Schramm. Influence — past and present — of François-André Michaux on forestry and forest research in America. Proceedings of the American Philosophical Society, 101(4):336–343, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985400.

Schramm:1958:MFH

[Sch58] J. R. Schramm. The mechanism of frost heaving of tree seedlings. Proceedings of the American Philosophical Society, 102(4):333–350, August 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985219.

Schmitt:1959:FWW

[Sch59] Bernadotte E. Schmitt. The First World War, 1914–1918. Proceedings of the American Philosophical Society, 103(3):321–331, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985472.

Schmitt:1960:PT

[Sch60] Bernadotte E. Schmitt. The peace treaties of 1919–1920. Proceedings of the American Philosophical Society, 104(1):101–110, February 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985606.

Schneider:1963:FSW

[Sch63] Howard A. Schneider. A fourth strategy in the war on infectious diseases. Proceedings of the American Philosophical Society, 107 (5):443–445, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985649.

Schopf:1969:GDD

[Sch69] Thomas J. M. Schopf. Geographic and depth distribution of the phylum Ectoprocta from 200 to 6,000 meters. Proceedings of the American Philosophical Society, 113(6):464–474, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985829.

Schmidt:1970:WHS

[Sch70] Albert J. Schmidt. William Hastie, Scottish planner of Russian cities. Proceedings of the American Philosophical Society, 114(3): 226–243, June 1970. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985700.

Schultz:1971:MRR

[Sch71] Stanley K. Schultz. The making of a reformer: The Reverend Samuel Hopkins as an Eighteenth-Century abolitionist. Proceedings of the American Philosophical Society, 115(5):350–365, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985768.

Schweitzer:1973:NSI

[Sch73] Pierre-Paul Schweitzer. National sovereignty and international monetary cooperation. Proceedings of the American Philosophical Society, 117(4):227–232, August 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986693.

Schrieffer:1974:NSE

[Sch74] J. Robert Schrieffer. A new schema for entropy. Proceedings of the American Philosophical Society, 118(6):538–539, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986403.

Schmitt:1994:HDA

[Sch94] Francis O. Schmitt. Hallowell Davis (31 August 1896–22 August 1992). Proceedings of the American Philosophical Society, 138(3): 414–419, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986748.

Schmidt:1995:WCE

[Sch95a] Albert J. Schmidt. Westernization as consumption: Estate building in the Moscow region during the Eighteenth Century. Proceedings of the American Philosophical Society, 139(4):380–419, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987237.

Schmitt:1995:WJH

[Sch95b] Francis O. Schmitt. Walle J. H. Nauta (8 June 1916–24 March 1994). Proceedings of the American Philosophical Society, 139(4): 444–449, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987242.

Schuh:1995:WFP

[Sch95c] G. Edward Schuh. The world food problem and population growth. Proceedings of the American Philosophical Society, 139(3):240–246, September 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987189.

Scheide:1997:TJS

[Sch97] William H. Scheide. Thoughts on Johann Sebastian Bach (1685– 1750), history, and society. Proceedings of the American Philosophical Society, 141(2):160–168, June 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987299.

Schuyler:1998:MMM

[Sch98] David Schuyler. Myth, memory, and the meaning of community: Sunnyside (Lancaster, Pennsylvania) and the limits of planning. Proceedings of the American Philosophical Society, 142(4):614–638, December 1998. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3152285.

Schmidt:1999:HHA

[Sch99a] Albert J. Schmidt. Hastie and his albums: Publicizing and imposing architectural style in Catherinian and Alexandrian Russia. Proceedings of the American Philosophical Society, 143(2):208–232, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181935.

Schopf:1999:LMT

[Sch99b] J. William Schopf. Life on Mars: Tempest in a teapot? A firsthand account. Proceedings of the American Philosophical Society, 143(3):359–378, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181949.

Schorske:1999:LKA

[Sch99c] Carl E. Schorske. Leonard Krieger (28 August 1918–12 October 1990). Proceedings of the American Philosophical Society, 143(3): 456–458, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181958.

Scribner:1980:PDS

[Scr80] Charles Scribner, Jr. Publishing the “dictionary of scientific biography”. Proceedings of the American Philosophical Society, 124 (5):320–322, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986571.

Scribner:1990:SIP

[Scr90] Charles Scribner, Jr. Scientific imagery in Proust. Proceedings of the American Philosophical Society, 134(3):243–308, September 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986968.

Scribner:1991:TBL

[Scr91] Charles Scribner III. Transfigurations: Bernini’s last works. Proceedings of the American Philosophical Society, 135(4):490–509, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986811.

Seitz:1998:FUC

[SE98] Frederick Seitz and Norman G. Einspruch. First use of crystal rectifiers in wireless. Proceedings of the American Philosophical Society, 142(4):639–642, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152286.

Seaborg:1966:SHF

[Sea66] Glenn T. Seaborg. Science, the humanities, and the federal government-partners in progress. Proceedings of the American Philosophical Society, 110(5):318–325, October 1966. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986020.

Seaborg:1993:EMM

[Sea93] Glenn T. Seaborg. Edwin Mattison McMillan (18 September 1907– 7 September 1991). Proceedings of the American Philosophical Society, 137(2):286–291, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986736.

Secher:1965:CNG

[Sec65] H. P. Secher. Controlling the new German military elite: The political role of the Parliamentary Defense Commissioner in the Federal Republic. Proceedings of the American Philosophical Society, 109(2):63–84, April 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985782.

See:1951:HCS

[See51] T. J. J. See. The hyperbolic curves for the stresses of gravitation in space identical with St. Venant’s torsional stresses in solids. Proceedings of the American Philosophical Society, 95(5):544–545, October 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143240.

Seeley:1961:SPF

[See61] Stuart William Seeley. Shoran — a precision five hundred mile yardstick. Proceedings of the American Philosophical Society, 105 (4):447–451, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985500.

Seidel:1990:CCR

[Sei90] Robert W. Seidel. Clio and the complex: Recent historiography of science and national security. Proceedings of the American Philosophical Society, 134(4):420–441, December 1990. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986897.

Seitz:1996:TPA

[Sei96] Frederick Seitz. The tangled prelude to the age of silicon electronics. Proceedings of the American Philosophical Society, 140(3): 289–337, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987311.

Seitz:1998:PS

[Sei98] Frederick Seitz. Politics and science. Proceedings of the American Philosophical Society, 142(4):517–524, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3152278.

Sellers:1950:PPB

[Sel50] Charles Coleman Sellers. The Peale portraits of Benjamin Franklin. Proceedings of the American Philosophical Society, 94(3):251–257, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143558.

Sellers:1951:SPP

[Sel51] Charles Coleman Sellers. Sellers papers in the Peale–Sellers Collection. Proceedings of the American Philosophical Society, 95(3):

262–265, June 1951. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143064.

Sellin:1953:MCG

[Sel53] Thorsten Sellin. The measurement of criminality in geographic areas. Proceedings of the American Philosophical Society, 97(2): 163–167, April 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143890.

Sellers:1955:JML

[Sel55] Charles Coleman Sellers. Jane Mecom’s little picture. Proceedings of the American Philosophical Society, 99(6):433–435, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143925.

Sellers:1956:CSE

[Sel56] Charles Coleman Sellers. Catalogue of the Society’s exhibition of portraits of Benjamin Franklin: January 17–April 20, 1956. Proceedings of the American Philosophical Society, 100(4):369– 384, August 1956. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143766.

Sellin:1965:PSO

[Sel65] Thorsten Sellin. Penal servitude: Origin and survival. Proceedings of the American Philosophical Society, 109(5):277–281, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986080.

Sellers:1980:PMN

[Sel80] Charles Coleman Sellers. Peale’s Museum and “the new museum idea”. Proceedings of the American Philosophical Society, 124 (1):25–34, February 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986654.

Sela:1996:CBA

[Sel96] Michael Sela. Christian Boehmer Anfinsen (26 March 1916—14 May 1995). Proceedings of the American Philosophical Society, 140 (4):558–562, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987291.

Sera:1991:TPB

[Ser91] Luigi Sera. Trans-pontem: a bridging between Europe and America by Lorenzo Da Ponte. Proceedings of the American Philosophical Society, 135(3):388–400, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986776.

Schopf:1965:FIB

[SESB65] J. M. Schopf, E. G. Ehlers, D. V. Stiles, and J. D. Birle. Fossil iron bacteria preserved in pyrite. Proceedings of the American Philosophical Society, 109(5):288–308, October 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986082.

Setton:1956:BBI

[Set56] Kenneth M. Setton. The Byzantine background to the Italian Renaissance. Proceedings of the American Philosophical Society, 100 (1):1–76, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143807.

Setton:1960:MML

[Set60] Kenneth M. Setton. From medieval to modern library. Proceedings of the American Philosophical Society, 104(4):371–390, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985616.

Setton:1969:PML

[Set69] Kenneth M. Setton. Penrose Memorial Lecture. Pope Leo X and the Turkish Peril. Proceedings of the American Philosophical Society, 113(6):367–424, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985823.

Setton:1992:GKD

[Set92] Kenneth M. Setton. Gladys Krieble Delmas (July 3, 1913– November 20, 1991). Proceedings of the American Philosophical Society, 136(2):290–293, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987174.

Seymour:1957:WWP

[Sey57] Charles Seymour. Woodrow Wilson: a political balance sheet. Proceedings of the American Philosophical Society, 101(2):135–141, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985504.

Shapiro:1957:OIM

[SH57] Herbert Shapiro and Thomas S. Harvey. Ocular implantation of the mammalian ovum. Proceedings of the American Philosophical Society, 101(2):164–176, April 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985510.

Shackelford:1958:WSD

[Sha58] George Green Shackelford. William Short: Diplomat in Revolutionary France, 1785–1793. Proceedings of the American Philosophical Society, 102(6):596–612, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985569.

Shackleton:1970:EIP

[Sha70] Robert Shackleton. The “encyclopédie” as an international phenomenon. Proceedings of the American Philosophical Society, 114 (5):389–394, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985804.

Shapiro:1976:PPL

[Sha76] Stanley Shapiro. The passage of power: Labor and the new social order. Proceedings of the American Philosophical Society, 120(6): 464–474, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986599.

Sheppe:1961:SER

[She61] Walter Sheppe, Jr. Systematic and ecological relations of “Peromyscus Oreas” and “P. Maniculatus”. Proceedings of the American Philosophical Society, 105(4):421–446, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985499.

Sheriff:1983:AGF

[She83] Mary D. Sheriff. “Au génie de Franklin”: an allegory by J.-H. Fragonard. Proceedings of the American Philosophical Society, 127(3): 180–193, June 1983. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986385.

Shewmaker:1985:FGC

[She85] Kenneth E. Shewmaker. Forging the “Great Chain”: Daniel Webster and the origins of American foreign policy toward East Asia and the Pacific, 1841–1852. Proceedings of the American Philosophical Society, 129(3):225–259, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987008.

Shih:1951:HUD

[Shi51] Hu Shih. How to understand a decade of rapidly deteriorated Sino-American relations. Proceedings of the American Philosophical Society, 95(4):457–459, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143286.

Shingleton:1970:RPR

[Shi70] Royce Shingleton. The Republic of Porkdom revisited: a note, comment on the article “Hog Meat and Cornpone: Food Habits in the Ante-Bellum South” by Sam Hilliard in the Proceedings, Volume 113, Pages 1–13. Proceedings of the American Philosophical Society, 114(5):407–408, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985807. See reply [Hil69, Hil70].

Shklar:1993:PF

[Shk93] Judith Nisse Shklar. Politics and friendship. Proceedings of the American Philosophical Society, 137(2):207–212, June 1993. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986730.

Shortliffe:1993:DPC

[Sho93] Edward H. Shortliffe. Doctors, patients, and computers: Will information technology dehumanize health-care delivery? Proceedings of the American Philosophical Society, 137(3):390–398, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986999.

Shryock:1955:NCI

[Shr55] Richard H. Shryock. The nature of the conference and its implications. Proceedings of the American Philosophical Society, 99 (5):327–331, October 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143930.

Shryock:1960:LAP

[Shr60a] Richard H. Shryock. The Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 104 (4):398–403, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985618.

Shryock:1960:PFO

[Shr60b] Richard H. Shryock. The planning and formal opening of library hall. Proceedings of the American Philosophical Society, 104(4): 349–356, August 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985614.

Shryock:1961:HSA

[Shr61] Richard H. Shryock. The history of science in American universities. Proceedings of the American Philosophical Society, 105(5):512, October 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985551.

Shryock:1964:MAA

[Shr64] Richard H. Shryock. The Middle Atlantic area in American history. Proceedings of the American Philosophical Society, 108(2):147–155, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985893.

Shuster:1953:AOG

[Shu53] George N. Shuster. The American occupation and German education. Proceedings of the American Philosophical Society, 97(2): 159–162, April 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143889.

Shulim:1958:USV

[Shu58] Joseph I. Shulim. The United States views Russia in the Napoleonic Age. Proceedings of the American Philosophical Society, 102(2):

148–159, April 1958. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985279.

Shumway:1971:PSH

[Shu71] Norman E. Shumway. Present status of heart transplantation in man. Proceedings of the American Philosophical Society, 115 (4):267–270, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986087.

Shulman:1998:JNH

[Shu98] Marshall D. Shulman. John Newbold Hazard (5 January 1909–7 April 1995). Proceedings of the American Philosophical Society, 142(1):140–144, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152271.

Shy:1987:FWN

[Shy87] John Shy. Franklin, Washington, and a new nation. Proceedings of the American Philosophical Society, 131(3):308–324, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987026.

Sibley:1958:HSC

[Sib58] Charles G. Sibley. Hybridization in some Colombian tanagers, avian genus “Ramphocelus”. Proceedings of the American Philosophical Society, 102(5):448–453, October 1958. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985590.

Sidman:1964:A

[Sid64] Murray Sidman. Anxiety. Proceedings of the American Philosophical Society, 108(6):478–481, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985866.

Sider:1999:CRT

[Sid99] Robert D. Sider. Craig Ringwalt Thompson (25 June 1911–4 October 1996). Proceedings of the American Philosophical Society, 143(3):512–521, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3181965.

Silk:1995:TME

[Sil95] Leonard Silk. The transition to a market economy: How much freedom does capitalism require? Proceedings of the American Philosophical Society, 139(2):149–162, June 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987262.

Simpson:1955:CCR

[Sim55] George Eaton Simpson. Culture change and reintegration found in the cults of West Kingston, Jamaica. Proceedings of the American Philosophical Society, 99(2):89–92, April 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143819.

Simmel:1958:COP

[Sim58] Marianne L. Simmel. The conditions of occurrence of phantom limbs. Proceedings of the American Philosophical Society, 102(5): 492–500, October 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985596.

Simpson:1959:AMC

[Sim59] George Gaylord Simpson. Anatomy and morphology: Classification and evolution: 1859 and 1959. Proceedings of the American Philosophical Society, 103(2):286–306, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985154.

Simon:1962:AC

[Sim62] Herbert A. Simon. The architecture of complexity. Proceedings of the American Philosophical Society, 106(6):467–482, December 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985254.

Simpson:1977:TML

[Sim77] George Gaylord Simpson. Too many lines; the limits of the Oriental and Australian zoogeographic regions. Proceedings of the American Philosophical Society, 121(2):107–120, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986523.

Simpson:1978:EMS

[Sim78] George Gaylord Simpson. Early mammals in South America: Fact, controversy, and mystery. Proceedings of the American Philosophical Society, 122(5):318–328, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986686.

Simpson:1984:MC

[Sim84] George Gaylord Simpson. Mammals and cryptozoology. Proceedings of the American Philosophical Society, 128(1):1–19, March 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986487.

Simpson:1985:E

[Sim85] George Gaylord Simpson. Extinction. Proceedings of the American Philosophical Society, 129(4):407–416, December 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986936.

Simpson:1987:HNP

[Sim87] J. A. Simpson. The hierarchy of nuclear particle accelerators in astrophysics. Proceedings of the American Philosophical Society, 131(1):1–14, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987113.

Simon:1993:HMS

[Sim93] Herbert A. Simon. The human mind: The symbolic level. Proceedings of the American Philosophical Society, 137(4):638–647, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987078.

Singer:1994:GJH

[Sin94] Maxine F. Singer. From genomic junk to human disease. Proceedings of the American Philosophical Society, 138(1):11–24, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986702.

Sinfelt:1999:COC

[Sin99] John H. Sinfelt. Catalysis: an old but continuing theme in chemistry. Proceedings of the American Philosophical Society, 143(3): 388–399, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181951.

Sioussat:1951:DWS

[Sio51] George L. Sioussat. Dr. William Smith, David Rittenhouse, and the Canal Plate September 7, 1777. Proceedings of the American Philosophical Society, 95(3):223–231, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143058.

Sjogren:1972:ELG

[Sjo72] William L. Sjogren. The evolution of lunar gravity measurements. Proceedings of the American Philosophical Society, 116(5):357–361, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986066.

Sjoberg:1994:SNK

[Sjö94] Åke W. Sjöberg. Samuel Noah Kramer (28 September 1897–26 November 1990). Proceedings of the American Philosophical Society, 138(1):170–173, March 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986712.

Sjoberg:1999:TJJ

[Sjö99] Åke W. Sjöberg. Thorkild Jacobsen (7 June 1904–2 May 1993). Proceedings of the American Philosophical Society, 143(2):334–337, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3181943.

Skendi:1975:LFN

[Ske75] Stavro Skendi. Language as a factor of national identity in the Balkans of the Nineteenth Century. Proceedings of the American Philosophical Society, 119(2):186–189, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986634.

Skinner:1964:M

[Ski64] B. F. Skinner. Man. Proceedings of the American Philosophical Society, 108(6):482–485, December 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985867.

Slichter:1998:GAN

[Sli98] Charles P. Slichter. The golden anniversary of nuclear magnetic resonance NMR: Fifty years of surprises. Proceedings of the American Philosophical Society, 142(4):533–556, December 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152280.

Smelser:1993:RBF

[Sme93] Neil J. Smelser. Reinhard Bendix (25 February 1916–28 February 1991). Proceedings of the American Philosophical Society, 137 (1):118–123, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986949.

Smethurst:1995:OSF

[Sme95] Richard G. Smethurst. Oliver Shewell Franks (16 February 1905–15 October 1992). Proceedings of the American Philosophical Society, 139(1):82–87, March 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986723.

Smith:1960:CKS

[Smi60] Allan H. Smith. The culture of Kabira, Southern Ryūkyū Islands. Proceedings of the American Philosophical Society, 104(2):134–171, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985656.

Smith:1962:HMB

[Smi62] C. Earle Smith, Jr. Henry Muhlenberg — botanical pioneer. Proceedings of the American Philosophical Society, 106(5):443–460, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985489.

See erratum [Ano62m].

Smith:1971:BRB

[Smi71] Robert C. Smith. Baroque and Rococo Braga: Documenting Eighteenth-Century architecture and sculpture in Northern Portugal. Proceedings of the American Philosophical Society, 115(3): 214–220, June 1971. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985978.

Smith:1972:PML

[Smi72] Cyril Stanley Smith. Penrose Memorial Lecture. Metallurgical footnotes to the history of art. Proceedings of the American Philosophical Society, 116(2):97–135, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986166. See erratum [Ano72m].

Smith:1979:EGD

[Smi79a] Emil L. Smith. The evolution of glutamate dehydrogenases and a hypothesis for the insertion or deletion of multiple residues in the interior of polypeptide chains. Proceedings of the American Philosophical Society, 123(1):73–84, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986471.

Smith:1979:ILL

[Smi79b] Raoul N. Smith. The interest in language and languages in Colonial and Federal America. Proceedings of the American Philosophical Society, 123(1):29–46, February 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986469.

Smith:1983:PSD

[Smi83] Murphy D. Smith. Peter Stephen Du Ponceau and his study of languages: a historical account. Proceedings of the American Philosophical Society, 127(3):143–179, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986384.

Smith:1984:PCA

[Smi84] Murphy D. Smith. The public clock of the American Philosophical Society. Proceedings of the American Philosophical Society, 128(3): 244–247, September 1984. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986884.

Smith:1986:SCJ

[Smi86] Cornelia D. Smith. A Seventeenth-Century judge views the law: Oliver St. John’s introduction to his charge at the Thetford Assizes in 1658. Proceedings of the American Philosophical Society, 130(1):37–78, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987091.

Smith:1988:PEM

[Smi88] Billy G. Smith. Poverty and economic marginality in Eighteenth-Century America. Proceedings of the American Philosophical Society, 132(1):85–118, March 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143826.

Smith:1989:FLL

[Smi89] Billy G. Smith. The family lives of laboring Philadelphians during the late Eighteenth Century. Proceedings of the American Philosophical Society, 133(2):328–332, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987058.

Snow:1958:GTD

[Sno58] Vernon F. Snow. The grand tour diary of Robert C. Johnson, 1792–1793. Proceedings of the American Philosophical Society, 102 (1):60–105, February 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985305.

Snyderman:1953:PSA

[Sny53] George S. Snyderman. A preliminary survey of American Indian manuscripts in repositories of the Philadelphia area. Proceedings of the American Philosophical Society, 97(5):596–610, October 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3149273.

Snyderman:1954:FW

[Sny54] George S. Snyderman. The functions of wampum. Proceedings of the American Philosophical Society, 98(6):469–494, December 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143870.

Snyderman:1957:HJJ

[Sny57] George S. Snyderman. Halliday Jackson’s journal of a visit paid to the Indians of New York (1806). Proceedings of the American Philosophical Society, 101(6):565–588, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985525.

Snyderman:1958:MCP

[Sny58] George S. Snyderman. The Manuscript Collections of the Philadelphia Yearly Meeting of Friends pertaining to the American Indian. Proceedings of the American Philosophical Society, 102(6):613–620, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985570.

Snyderman:1961:FWI

[Sny61] George S. Snyderman. The function of wampum in Iroquois religion. Proceedings of the American Philosophical Society, 105(6): 571–608, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985167.

Snyderman:1982:EDA

[Sny82] George S. Snyderman. An ethnological discussion of Allegany Seneca wampum folklore. Proceedings of the American Philosophical Society, 126(4):316–326, August 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986177.

Snyderman:1983:WWA

[Sny83] George S. Snyderman. Witches, witchcraft, and Allegany Seneca medicine. Proceedings of the American Philosophical Society, 127 (4):263–277, 1983. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986274.

Soderlund:1989:BIM

[Sod89] Jean R. Soderlund. Black importation and migration into Southeastern Pennsylvania, 1682–1810. Proceedings of the American Philosophical Society, 133(2):144–153, June 1989. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987045.

Sokal:1972:PVC

[Sok72] Michael M. Sokal. Psychology at Victorian Cambridge: The unofficial laboratory of 1887–1888. Proceedings of the American Philosophical Society, 116(2):145–147, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986169.

Solmsen:1968:GIH

[Sol68] Friedrich Solmsen. Greek ideas of the hereafter in Virgil’s Roman epic. Proceedings of the American Philosophical Society, 112 (1):8–14, February 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986099.

Soltow:1971:ESS

[Sol71] James H. Soltow. Entrepreneurial strategy in small industry: Belgian metal fabricators. Proceedings of the American Philosophical Society, 115(1):32–64, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985922.

Soman:1976:PPC

[Som76] Alfred Soman. Press, pulpit, and censorship in France before Richelieu. Proceedings of the American Philosophical Society, 120(6): 439–463, December 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986598.

Sontag:1954:DDS

[Son54] Raymond J. Sontag. The democracies and the dictators since 1933. Proceedings of the American Philosophical Society, 98(5): 313–317, October 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143909.

Sonneborn:1965:GMV

[Son65] T. M. Sonneborn. Genetics and Man’s vision. Proceedings of the American Philosophical Society, 109(4):237–241, August 1965. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985882.

Soudek:1952:ATE

[Sou52] Josef Soudek. Aristotle’s theory of exchange: an inquiry into the origin of economic analysis. Proceedings of the American Philosophical Society, 96(1):45–75, February 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143742.

Sozanski:1992:HMJ

[Soz92] Edward J. Sozanski. Henry Moore (July 30, 1898–August 30, 1986). Proceedings of the American Philosophical Society, 136 (1):138–140, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986805.

Speiser:1951:ANE

[Spe51a] Ephraim A. Speiser. The ancient near east and modern philosophies of history. Proceedings of the American Philosophical Society, 95(6):583–588, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143507.

Spengler:1951:EFD

[Spe51b] Joseph J. Spengler. Economic factors in the development of densely populated areas. Proceedings of the American Philosophical Society, 95(1):20–53, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143572.

Spencer:1956:EKE

[Spe56] Robert F. Spencer. Exhortation and the Klamath ethos. Proceedings of the American Philosophical Society, 100(1):77–86, February 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143808.

Speiser:1963:CLH

[Spe63] E. A. Speiser. Cuneiform law and the history of civilization. Proceedings of the American Philosophical Society, 107(6):536–541, December 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986110.

Spengler:1969:KPL

[Spe69] Joseph J. Spengler. Kaut.ilya, Plato, Lord Shang: Comparative political economy. Proceedings of the American Philosophical Society, 113(6):450–457, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985827.

Spiller:1956:FAB

[Spi56] Robert E. Spiller. Franklin on the art of being human. Proceedings of the American Philosophical Society, 100(4):304–315, August 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143756.

Spiro:1969:RSS

[Spi69] Melford E. Spiro. Religious symbolism and social behavior. Proceedings of the American Philosophical Society, 113(5):341–349, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985971.

Spilhaus:1971:NIR

[Spi71] Athelstan Spilhaus. The next industrial revolution. Proceedings of the American Philosophical Society, 115(4):324–327, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986093.

Spiegelman:1976:SVH

[Spi76] Sol Spiegelman. The search for viruses in human cancer. Proceedings of the American Philosophical Society, 120(2):69–86, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986513.

Spilhaus:1983:WOM

[Spi83] Athelstan Spilhaus. World ocean maps: The proper places to interrupt. Proceedings of the American Philosophical Society, 127 (1):50–60, February 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986364.

Spilhaus:1984:PTG

[Spi84] Athelstan Spilhaus. Plate tectonics in geoforms and jigsaws. Proceedings of the American Philosophical Society, 128(3):257–269, September 1984. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986887.

Spilhaus:1993:NEA

[Spi93] Athelstan Spilhaus. New equal area world map projections to better display polar regions. Proceedings of the American Philosophical Society, 137(2):179–193, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986728.

Spoehr:1951:CSF

[Spo51] H. A. Spoehr. Chlorella as a source of food. Proceedings of the American Philosophical Society, 95(1):62–67, February 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143574.

Spoehr:1981:LHM

[Spo81] Alexander Spoehr. Lewis Henry Morgan and his Pacific collaborators: a Nineteenth-Century chapter in the history of anthropological research. Proceedings of the American Philosophical Society, 125(6):449–459, December 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986300.

Sprague:1998:LBF

[Spr98] James M. Sprague. Louis Barkhouse Flexner (7 January 1902–29 March 1996). Proceedings of the American Philosophical Society, 142(3):468–477, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152252.

Stokes:1957:RPS

[SS57] Joseph Stokes, Jr. and Joseph Smolens. Repeated plasmapheresis in the same person — a rationale for modern bloodletting. Proceedings of the American Philosophical Society, 101(4):330–335, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985399.

Simpson:1954:RPM

[ST54] George Gaylord Simpson and H. Tobien. The rediscovery of Peale’s mastodon. Proceedings of the American Philosophical Society, 98 (4):279–281, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143882.

Stadie:1950:CCI

[Sta50a] William C. Stadie. The chemical combination of insulin with muscle and its hormonal regulation. Proceedings of the American Philosophical Society, 94(2):171–174, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143216.

Stakman:1950:SNA

[Sta50b] Elvin C. Stakman. Science and national authoritarianism. Proceedings of the American Philosophical Society, 94(2):114–120, April 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143210.

Stassen:1950:AFP

[Sta50c] Harold E. Stassen. American foreign policy. Proceedings of the American Philosophical Society, 94(5):450–452, October 1950. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143611.

Stamm:1951:PNS

[Sta51] Alfred J. Stamm. Production of nutritive substances from inedible carbohydrates. Proceedings of the American Philosophical Society, 95(1):68–76, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143575.

Stankiewicz:1955:HDI

[Sta55] W. J. Stankiewicz. The Huguenot downfall: The influence of Richelieu’s policy and doctrine. Proceedings of the American Philosophical Society, 99(3):146–168, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143780.

Stanley:1957:PML

[Sta57] Wendell M. Stanley. Penrose Memorial Lecture. On the nature of viruses, cancer, genes, and life — a declaration of dependence. Proceedings of the American Philosophical Society, 101(4):317–324, August 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985397.

Stauffer:1960:ELM

[Sta60] Robert Clinton Stauffer. Ecology in the long manuscript version of Darwin’s “Origin of Species” and Linnaeus’ “Oeconomy of Nature”. Proceedings of the American Philosophical Society, 104(2):235–241, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985662.

Starzl:1971:TWH

[Sta71] Thomas E. Starzl. Transplantation of whole human organs. Proceedings of the American Philosophical Society, 115(4):258–266, August 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986086.

Stapleton:1991:TEI

[Sta91] Darwin H. Stapleton. The transfer of early industrial technologies to America with especial reference to the role of the American Philosophical Society. Proceedings of the American Philosophical Society, 135(2):286–298, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/987035.

Stanley:1998:DTC

[Sta98] Julian C. Stanley. Donald Thomas Campbell (20 November 1916– 6 May 1996). Proceedings of the American Philosophical Society, 142(1):114–120, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152267.

Stern:1952:CCC

[Ste52] Theodore Stern. Chickahominy: The changing culture of a Virginia Indian community. Proceedings of the American Philosophical Society, 96(2):157–225, April 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143721.

Stebbins:1959:RHE

[Ste59a] G. Ledyard Stebbins. The role of hybridization in evolution. Proceedings of the American Philosophical Society, 103(2):231–251, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985151.

Stern:1959:VHT

[Ste59b] Curt Stern. Variation and hereditary transmission. Proceedings of the American Philosophical Society, 103(2):183–189, April 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985148.

Stein:1963:IMB

[Ste63a] Robert Carrington Stein. Isolating mechanisms between populations of Traill’s flycatchers. Proceedings of the American Philosophical Society, 107(1):21–50, February 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985467.

Stewart:1963:PSD

[Ste63b] Duncan Stewart. Petrography of some dredgings collected by operation deep freeze IV. Proceedings of the American Philosophical Society, 107(5):431–442, October 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985648.

Stern:1965:MHG

[Ste65] Curt Stern. Mendel and human genetics. Proceedings of the American Philosophical Society, 109(4):216–226, August 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985880.

Stephens:1969:DTM

[Ste69] Olin J. Stephens, II. Designing a twelve meter yacht. Proceedings of the American Philosophical Society, 113(6):428–432, December 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985825.

Stellar:1974:BMH

[Ste74] Eliot Stellar. Brain mechanisms in hunger and other hedonic experiences. Proceedings of the American Philosophical Society, 118(3): 276–282, June 1974. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986446.

Steadman:1976:ISH

[Ste76] John M. Steadman. The idea of Satan as the hero of “Paradise Lost”. Proceedings of the American Philosophical Society, 120 (4):253–294, August 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986321.

Stewart:1977:NSR

[Ste77] T. D. Stewart. The Neanderthal skeletal remains from Shanidar Cave, Iraq: a summary of findings to date. Proceedings of the American Philosophical Society, 121(2):121–165, April 1977. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986524.

Stent:1986:HAC

[Ste86] Gunther S. Stent. Hermeneutics and the analysis of complex biological systems. Proceedings of the American Philosophical Society, 130(3):336–342, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986832.

Steiner:1987:BFB

[Ste87] Prudence L. Steiner. Benjamin Franklin’s biblical hoaxes. Proceedings of the American Philosophical Society, 131(2):183–196, June 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986791.

Stein:1989:CSO

[Ste89] Stephen J. Stein. “A candid statement of our principles”: Early Shaker theology in the West. Proceedings of the American Philosophical Society, 133(4):503–519, December 1989. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986874.

Stellar:1992:BNA

[Ste92] Eliot Stellar. The biological nature of appetite. Proceedings of the American Philosophical Society, 136(4):507–520, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986759.

Stellar:1993:CR

[Ste93a] Eliot Stellar. Closing remarks. Proceedings of the American Philosophical Society, 137(4):708, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987085.

Stellar:1993:OR

[Ste93b] Eliot Stellar. Opening remarks. Proceedings of the American Philosophical Society, 137(4):465–466, December 1993. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987064.

Stellar:1993:SAP

[Ste93c] Eliot Stellar. The state of the American Philosophical Society. Proceedings of the American Philosophical Society, 137(4):467–471, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987065.

Stent:1997:AML

[Ste97] Gunther S. Stent. André Michel Lwoff (8 May 1902–30 September 1994). Proceedings of the American Philosophical Society, 141 (1):98–101, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987254.

Stent:1998:EDM

[Ste98] Gunther S. Stent. Epistemic dualism of mind and body. Proceedings of the American Philosophical Society, 142(4):578–588, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152282.

Stebbins:1999:KEA

[Ste99a] George Ledyard Stebbins. Katherine Esau (3 April 1898–4 June 1997). Proceedings of the American Philosophical Society, 143(4): 664–672, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181994.

Stern:1999:GKO

[Ste99b] Fritz Stern. Grayson Kirk (12 October 1903–21 November 1997). Proceedings of the American Philosophical Society, 143(4):674–678, December 1999. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3181995.

Stigler:1959:BSS

[Sti59] George J. Stigler. Bernard Shaw, Sidney Webb, and the theory of Fabian socialism. Proceedings of the American Philosophical Society, 103(3):469–475, June 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985478.

Stillwell:1961:KT

[Sti61] Richard Stillwell. Kourion: The theater. Proceedings of the American Philosophical Society, 105(1):37–78, February 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985354.

Stinchcombe:1977:NMT

[Sti77] William Stinchcombe. A neglected memoir by Talleyrand on French–American relations, 1793–1797. Proceedings of the American Philosophical Society, 121(3):195–208, June 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986410.

Stigler:1992:AMS

[Sti92] Stephen M. Stigler. Apollo Mathematicus: a story of resistance to quantification in the Seventeenth Century. Proceedings of the American Philosophical Society, 136(1):93–126, March 1992. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986802.

Stouffer:1955:NRT

[Sto55] Samuel A. Stouffer. Needed research on the tolerance of nonconformity. Proceedings of the American Philosophical Society, 99 (4):239–243, August 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143702.

Stoker:1959:NSF

[Sto59] J. J. Stoker. Numerical solution of flood wave and river regulation problems. Proceedings of the American Philosophical Society, 103 (4):548–553, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985557.

Stone:1962:PP

[Sto62] George Winchester Stone, Jr. The poet and the players. Proceedings of the American Philosophical Society, 106(5):412–421, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985485.

Stokes:1985:PCA

[Sto85] Donald E. Stokes. The paradox of campaign appeals and election mandates. Proceedings of the American Philosophical Society, 129(1):20–25, March 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986977.

Stone:1987:RFA

[Sto87] Geoffrey R. Stone. Reflections on the First Amendment: The evolution of the American jurisprudence of free expression. Proceedings of the American Philosophical Society, 131(3):251–260, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987020.

Stow:1999:SSR

[Sto99a] George B. Stow. Stubbs, Steel, and Richard II as insane: The origin and evolution of an English historiographical myth. Proceedings of the American Philosophical Society, 143(4):601–638, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181991.

Stowe:1999:KVT

[Sto99b] Bruce B. Stowe. Kenneth V. Thimann (5 August 1904–15 January 1997). Proceedings of the American Philosophical Society, 143(3): 502–509, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181964.

Strayer:1961:PFC

[Str61] Joseph R. Strayer. The promise of the Fourteenth Century. Proceedings of the American Philosophical Society, 105(6):609–611, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985168.

Stratton:1964:PML

[Str64] Julius A. Stratton. Penrose Memorial Lecture. Learning and action. Proceedings of the American Philosophical Society, 108(5): 387–394, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985811.

Strong:1988:MGS

[Str88] Maurice Strong. Managing for global survival. Proceedings of the American Philosophical Society, 132(2):196–203, June 1988. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986865.

Stratton:1991:HEE

[Str91] Julius A. Stratton. Harold Eugene Edgerton (April 6, 1903– January 4, 1990). Proceedings of the American Philosophical Society, 135(3):444–450, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986778.

Strong:1995:ASS

[Str95] Maurice F. Strong. Agenda 21: Sustaining the spirit of rio. Proceedings of the American Philosophical Society, 139(3):233–239, September 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987188.

Strangway:1997:JTW

[Str97] David W. Strangway. J. Tuzo Wilson (24 October 1908–15 April 1993). Proceedings of the American Philosophical Society, 141(3): 384–386, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987208.

Sturtevant:1965:EM

[Stu65] Alfred H. Sturtevant. The early Mendelians. Proceedings of the American Philosophical Society, 109(4):199–204, August 1965. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985878.

Sturgis:1970:ISA

[Stu70] Katharine Boucot Sturgis. Introduction to the symposium on atmospheric pollution: Its long-term implications. Proceedings of the American Philosophical Society, 114(1):1–2, February 1970. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985717.

Suits:1953:DC

[Sui53] C. G. Suits. The defects in crystals. Proceedings of the American Philosophical Society, 97(6):681–685, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3149283.

Suits:1964:MMD

[Sui64] C. G. Suits. Man-made diamonds — a progress report. Proceedings of the American Philosophical Society, 108(5):443–449, October 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985819.

Sullivan:1968:NTD

[Sul68] Leon H. Sullivan. New trends in the development of self-help in the inner city. Proceedings of the American Philosophical Society, 112 (6):358–361, December 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985934.

Summers:1960:ECS

[Sum60] Roger Summers. Environment and culture in Southern Rhodesia: a study in the “personality” of a land-locked country. Proceedings of the American Philosophical Society, 104(3):266–292, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985244.

Shane:1950:DEN

[SW50] C. D. Shane and C. A. Wirtanen. Distribution of extragalactic nebulae in three selected areas. Proceedings of the American Philosophical Society, 94(1):13–17, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143246.

Swanson:1977:HEC

[SW77] Richard A. Swanson and Stanley Witkowski. Hopi ethnoanatomy: a comparative treatment. Proceedings of the American Philosophical Society, 121(4):320–337, August 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986421.

Swadesh:1952:LSD

[Swa52] Morris Swadesh. Lexico–statistic dating of prehistoric ethnic contacts: With special reference to North American Indians and Eskimos. Proceedings of the American Philosophical Society, 96(4):452– 463, August 1952. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143802.

Swerdlow:1973:DFD

[Swe73] Noel M. Swerdlow. The derivation and first draft of Copernicus’s planetary theory: a translation of the Commentariolus with commentary. Proceedings of the American Philosophical Society, 117 (6):423–512, December 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986461.

Swerdlow:1993:ONM

[Swe93] Noel M. Swerdlow. Otto E. Neugebauer (26 May 1899–19 February 1990). Proceedings of the American Philosophical Society, 137 (1):138–165, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986952.

Swett:1996:MW

[Swe96] Hannah Mickley Swett. Making waves. Proceedings of the American Philosophical Society, 140(4):546–556, December 1996. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987290.

Swings:1969:VTS

[Swi69] Pol Swings. Venus through a spectroscope. Proceedings of the American Philosophical Society, 113(3):229–246, June 1969. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985738.

Swiggers:1998:ALO

[Swi98] Pierre Swiggers. Americanist linguistics and the origin of linguistic typology: Peter Stephen Du Ponceau’s “comparative science of language”. Proceedings of the American Philosophical Society, 142(1):18–46, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152262.

Syme:1960:BRA

[Sym60] Ronald Syme. Bastards in the Roman aristocracy. Proceedings of the American Philosophical Society, 104(3):323–327, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985248.

Syme:1972:LGC

[Sym72] Ronald Syme. Lawyers in government: The case of Ulpian. Proceedings of the American Philosophical Society, 116(5):406–409, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986071.

Szabo:1972:SVL

[Sza72] László Szabó. Stress and vowel length in malecite. Proceedings of the American Philosophical Society, 116(4):338–342, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985903.

Taft:1962:LHL

[Taf62] Philip Taft. Labor history and the labor issues of today. Proceedings of the American Philosophical Society, 106(4):304–309, August 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985264.

Talalay:1999:WAC

[Tal99] Paul Talalay. The war against cancer: New hope. Proceedings of the American Philosophical Society, 143(1):52–72, March 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181974.

Tatel:1953:SEC

[TAT53] H. E. Tatel, L. H. Adams, and M. A. Tuve. Studies of the Earth’s crust using waves from explosions. Proceedings of the American Philosophical Society, 97(6):658–669, December 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3149281.

Tate:1954:POS

[Tat54] Vernon D. Tate. Photographic and other scientific aids to the editor and curator. Proceedings of the American Philosophical Society, 98(3):179–184, June 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143843.

Taussig:1976:DBT

[Tau76] Helen B. Taussig. The development of the Blalock–Taussig operation and its results twenty years later. Proceedings of the American Philosophical Society, 120(1):13–20, February 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986348.

Taylor:1950:ARC

[Tay50] Lily Ross Taylor. Annals of the Roman consulship on the Arch of Augustus. Proceedings of the American Philosophical Society, 94 (6):511–516, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143592.

Taylor:1953:RD

[Tay53] Archer Taylor. Riddles in dialogue. Proceedings of the American Philosophical Society, 97(1):61–68, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143732.

Taylor:1954:MIR

[Tay54] Francis Henry Taylor. The moral and intellectual responsibilities of expanding collections. Proceedings of the American Philosophical Society, 98(4):245–249, August 1954. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143876.

Taylor:1957:RAG

[Tay57] Harold Taylor. The role of the administrator as guardian of academic freedom. Proceedings of the American Philosophical Society, 101(5):450–454, October 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985199.

Taylor:1963:CCL

[Tay63] Allan R. Taylor. The classification of the Caddoan languages. Proceedings of the American Philosophical Society, 107(1):51–59, February 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985468.

Taylor:1968:PPI

[Tay68] Telford Taylor. Press and prejudice: The impact of news on justice. Proceedings of the American Philosophical Society, 112(2):121–125, April 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985929.

Taylor:1972:CPL

[Tay72] George W. Taylor. Competition in product and labor markets. Proceedings of the American Philosophical Society, 116(3):216–224, June 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986116.

Taylor:1993:TGB

[Tay93] J. H. Taylor. Testing gravity with binary pulsars. Proceedings of the American Philosophical Society, 137(3):357–363, September 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986997.

Taylor:1999:GLD

[Tay99] Lance Taylor. Globalization, liberalization, distribution, and growth: Developing and transition economies. Proceedings of the American Philosophical Society, 143(2):187–193, June 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181933.

Tedeschi:1987:QMW

[Ted87] John Tedeschi. The question of magic and witchcraft in two unpublished inquisitorial manuals of the Seventeenth Century. Proceedings of the American Philosophical Society, 131(1):92–111, March 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987117.

Teitelbaum:1964:A

[Tei64] Philip Teitelbaum. Appetite. Proceedings of the American Philosophical Society, 108(6):464–472, December 1964. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985864.

Teitelbaum:1988:DCT

[Tei88] Michael S. Teitelbaum. Demographic change through the lenses of science and politics. Proceedings of the American Philosophical Society, 132(2):173–184, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986863.

Teitelbaum:1992:AAA

[Tei92] Michael S. Teitelbaum. Advocacy, ambivalence, ambiguity: Immigration policies and prospects in the United States. Proceedings of the American Philosophical Society, 136(2):208–225, June 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987169.

Thimann:1991:PCM

[TG91] Kenneth V. Thimann and Walton C. Galinat. Paul Christoph Mangelsdorf (July 20, 1899–July 22, 1989). Proceedings of the American Philosophical Society, 135(3):468–472, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986781.

Thier:1993:RAE

[Thi93] Samuel O. Thier. Responding to the AIDS epidemic: Science vs. politics. Proceedings of the American Philosophical Society, 137 (1):110–116, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986948.

Thomas:1950:SSA

[Tho50a] Dorothy Swaine Thomas. Some social aspects of Japanese– American demography. Proceedings of the American Philosophical Society, 94(5):459–480, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143613.

Thorndike:1950:PCA

[Tho50b] Lynn Thorndike. Pre-Copernican astronomical activity. Proceedings of the American Philosophical Society, 94(4):321–326, August 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143515.

Thompson:1953:AHP

[Tho53] Homer A. Thompson. Athens and the Hellenistic princes. Proceedings of the American Philosophical Society, 97(3):254–261, June 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143748.

Thorndike:1960:SMR

[Tho60] Lynn Thorndike. Some Medieval and Renaissance manuscripts on physics. Proceedings of the American Philosophical Society, 104 (2):188–201, April 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985658.

Thompson:1966:CL

[Tho66] Homer A. Thompson. Classical lands. Proceedings of the American Philosophical Society, 110(2):100–104, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985742.

Thomas:1967:PNS

[Tho67] Dorothy S. Thomas. Prefatory note to symposium on population problems. Proceedings of the American Philosophical Society, 111(3):133–135, June 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985710.

Thomson:1971:RCW

[Tho71] Robert Polk Thomson. The reform of the College of William and Mary, 1763–1780. Proceedings of the American Philosophical Society, 115(3):187–213, June 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985977.

Thorpe:1990:PHA

[Tho90] James Thorpe. In Philosophical Hall: The American Philosophical Society Held at Philadelphia for Promoting Useful Knowledge. Proceedings of the American Philosophical Society, 134(3):227–229, September 1990. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986966.

Thompson:1991:BDM

[Tho91] Homer A. Thompson. Benjamin Dean Meritt (March 31, 1899– July 7, 1989). Proceedings of the American Philosophical Society, 135(1):110–115, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987155.

Thomas:1998:RRC

[Tho98] John Meurig Thomas. Rumford’s remarkable creation. Proceedings of the American Philosophical Society, 142(4):597–613, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152284.

Thurstone:1950:SPA

[Thu50] L. L. Thurstone. Some primary abilities in visual thinking. Proceedings of the American Philosophical Society, 94(6):517–521, December 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143593.

Tibbitts:1954:SVA

[Tib54] Clark Tibbitts. A sociological view of aging. Proceedings of the American Philosophical Society, 98(2):144–148, April 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143645.

Tierney:1998:SGK

[Tie98] Brian Tierney. Stephan George Kuttner (24 March 1907–12 August 1996). Proceedings of the American Philosophical Society, 142(3): 486–491, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152254.

Tinkcom:1970:SRC

[Tin70] Margaret B. Tinkcom. Southwark, a river community: Its shape and substance. Proceedings of the American Philosophical Society, 114(4):327–342, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985956.

True:1971:TWS

[TNN71] D. L. True, Lautaro Núñez, and Patricio Núñez. Tarapaca 10: a workshop site in Northern Chile. Proceedings of the American Philosophical Society, 115(5):398–421, October 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985770.

Tobin:1991:WWH

[Tob91] James Tobin. Walter W. Heller (August 27, 1915–June 15, 1987). Proceedings of the American Philosophical Society, 135(1):100–107, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987154.

Tobias:1996:PDS

[Tob96] Phillip V. Tobias. Premature discoveries in science with especial reference to “Australopithecus” and “Homo Habilis”. Proceedings of the American Philosophical Society, 140(1):49–64, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987275.

Tobin:1997:JTA

[Tob97] James Tobin. Jan Tinbergen (12 April 1903–9 June 1994). Proceedings of the American Philosophical Society, 141(4):510–514, December 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987227.

Tobin:1999:FG

[Tob99] James Tobin. Financial globalization. Proceedings of the American Philosophical Society, 143(2):161–167, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181930.

Todd:1992:PLM

[Tod92] Bruce Todd. Preface. Letter from Mayor Bruce Todd. Proceedings of the American Philosophical Society, 136(3):358, September 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986908.

Tolles:1951:GLA

[Tol51a] Frederick B. Tolles. George Logan and the agricultural revolution. Proceedings of the American Philosophical Society, 95(6): 589–596, December 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143508.

Tolley:1951:FHW

[Tol51b] Howard R. Tolley. Farmers in a hungry world. Proceedings of the American Philosophical Society, 95(1):54–61, February 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143573.

Tolles:1963:NCQ

[Tol63] Frederick B. Tolles. Nonviolent contact: The Quakers and the Indians. Proceedings of the American Philosophical Society, 107(2): 93–101, April 1963. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985436.

Tooker:1989:DHL

[Too89] Elisabeth Tooker. On the development of the handsome lake religion. Proceedings of the American Philosophical Society, 133(1): 35–50, March 1989. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987160.

Topping:1976:VLI

[Top76] Peter Topping. Venice’s last imperial venture. Proceedings of the American Philosophical Society, 120(3):159–165, June 1976. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986555.

Toynbee:1955:OEW

[Toy55] Arnold Joseph Toynbee. The Ottoman Empire in world history. Proceedings of the American Philosophical Society, 99(3):119–126, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143777.

Traut:1998:MEJ

[Tra98] Thomas W. Traut. Mary Ellen Jones (25 December 1922–23 August 1996). Proceedings of the American Philosophical Society, 142 (3):480–484, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152253.

Trever:1953:SPD

[Tre53] John C. Trever. Studies in the problem of dating the dead sea scrolls. Proceedings of the American Philosophical Society, 97(2): 184–193, April 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143893.

Trenn:1974:CDF

[Tre74] Thaddeus J. Trenn. Charles Darwin, fossil Cirripedes, and Robert Fitch: Presenting sixteen hitherto unpublished Darwin letters of 1849 to 1851. Proceedings of the American Philosophical Society, 118(5):471–491, October 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986311.

Trimmer:1980:PSQ

[Tri80] John D. Trimmer. The present situation in quantum mechanics: a translation of Schrödinger’s “Cat Paradox” paper. Proceedings of the American Philosophical Society, 124(5):323–338, October 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986572.

Truman:1959:PCL

[Tru59] David B. Truman. The Presidency and Congressional leadership: Some notes on our changing Constitution. Proceedings of the American Philosophical Society, 103(5):687–692, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985425.

Tuden:1969:TPA

[Tud69] Arthur Tuden. Trends in political anthropology. Proceedings of the American Philosophical Society, 113(5):336–340, October 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985970.

Turner:1971:WFC

[Tur71] Ralph V. Turner. William De Forz, Count of Aumale: an Early Thirteenth-Century English Baron. Proceedings of the American Philosophical Society, 115(3):221–249, June 1971. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985979.

Tomlinson:1966:VBP

[TZ66] P. B. Tomlinson and M. H. Zimmermann. Vascular bundles in palm stems — their bibliographic evolution. Proceedings of the American Philosophical Society, 110(3):174–181, June 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985986.

Urey:1971:RSM

[Ure71] Harold C. Urey. A review of the structure of the Moon. Proceedings of the American Philosophical Society, 115(2):67–73, April 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985847.

Urton:1981:AAQ

[Urt81] Gary Urton. Animals and astronomy in the Quechua universe. Proceedings of the American Philosophical Society, 125(2):110–127, April 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986640.

VanHeyningen:1970:CT

[Van70] W. E. Van Heyningen. Cholera toxin. Proceedings of the American Philosophical Society, 114(5):395–397, October 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985805.

Vanderhoof:1953:HTA

[VD53] Mary B. Vanderhoof and Jean François Ducis. Hamlet: a tragedy adapted from Shakespeare (1770) by Jean François ducis. A critical edition. Proceedings of the American Philosophical Society, 97 (1):88–142, February 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143735.

vanDoren:1977:WTI

[vD77] Llewain Scott van Doren. War taxation, institutional change, and social conflict in provincial France — the royal “taille” in Dauphiné, 1494–1559. Proceedings of the American Philosophical Society, 121 (1):70–96, February 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986567.

vanDoren:1986:MAI

[vD86] Llewain Scott van Doren. Military administration and intercommunal relations in Dauphiné, 1494–1559. Proceedings of the American Philosophical Society, 130(1):79–100, March 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987092.

Vernon:1994:ESM

[VD94] Raymond Vernon and John T. Dunlop. Edward Sagendorph Mason (22 February 1899–29 February 1992). Proceedings of the American Philosophical Society, 138(2):342–345, June 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987130.

vandeHulst:1994:JHO

[vdH94] H. C. van de Hulst. Jan Hendrik Oort (28 April 1900–5 November 1992). Proceedings of the American Philosophical Society, 138(3): 448–455, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986752.

Vendler:1999:SHO

[Ven99] Helen Vendler. Seamus Heaney and the “Oresteia”: “Mycenae Lookout” and the usefulness of tradition. Proceedings of the American Philosophical Society, 143(1):116–129, March 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3181980.

Vermeule:1958:ASA

[Ver58] Cornelius C. Vermeule. Aspects of scientific archaeology in the Seventeenth Century: Marble reliefs, Greek vases, manuscripts, and minor objects in the Dal Pozzo-Albani drawings of classical antiquities. Proceedings of the American Philosophical Society, 102(2): 193–214, April 1958. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985284.

Vermeule:1964:GRP

[Ver64] Cornelius C. Vermeule. Greek and Roman portraits in North American collections open to the public: a survey of important monumental likenesses in marble and bronze which have not been published extensively. Proceedings of the American Philosophical Society, 108(2):99–134, April 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985890.

Vermeule:1965:GTS

[Ver65] Cornelius C. Vermeule III. A Greek theme and its survivals: The Ruler’s shield (Tondo image) in tomb and temple. Proceedings of the American Philosophical Society, 109(6):361–397, December 1965. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986138.

Vermeule:1993:JH

[Ver93] Emily Townsend Vermeule. Jefferson and Homer. Proceedings of the American Philosophical Society, 137(4):689–703, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987083.

vonFrisch:1956:LOB

[vF56] Karl von Frisch. The “language” and orientation of the bees. Proceedings of the American Philosophical Society, 100(5):515– 519, October 1956. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143684.

Voegelin:1951:MDI

[VH51] C. F. Voegelin and Zellig S. Harris. Methods for determining intelligibility among dialects of natural languages. Proceedings of the American Philosophical Society, 95(3):322–329, June 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143070.

Voegelin:1953:SNA

[VH53] C. F. Voegelin and D. H. Hymes. A sample of North American Indian dictionaries with reference to acculturation. Proceedings of the American Philosophical Society, 97(5):634–644, October 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3149275.

vanHorne:1981:JSO

[vH81] John C. van Horne. Joseph Solomon Ottolenghe (ca. 1711–1775): Catechist to the Negroes, superintendent of the silk culture, and public servant in Colonial Georgia. Proceedings of the American Philosophical Society, 125(5):398–409, October 1981. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986200.

vanHorne:1993:EWD

[vH93] John C. van Horne. Edwin Wolf 2nd (6 December 1911–20 February 1991). Proceedings of the American Philosophical Society, 137 (2):300–304, June 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986738.

vanHelden:1996:GSS

[vH96] Albert van Helden. Galileo and Scheiner on sunspots: a case study in the visual language of astronomy. Proceedings of the American Philosophical Society, 140(3):358–396, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987314.

Vine:1968:ESG

[Vin68] Frederick J. Vine. Evidence from submarine geology. Proceedings of the American Philosophical Society, 112(5):325–334, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986053.

Visscher:1972:NA

[Vis72] Maurice B. Visscher. The newer antivivisectionists. Proceedings of the American Philosophical Society, 116(2):157–162, April 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986171.

Vivian:1974:CBA

[Viv74] James F. Vivian. The Commercial Bureau of American Republics, 1894–1902: The advertising policy, the State Department, and the governance of the International Union. Proceedings of the American Philosophical Society, 118(6):555–566, December 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986405.

Vassar:1970:WKD

[VK70] Rena Vassar and William Knox. William Knox’s defense of slavery (1768). Proceedings of the American Philosophical Society, 114 (4):310–326, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985955.

vonMuralt:1980:PBP

[vM80] Alexander von Muralt. The physical basis of perception. Proceedings of the American Philosophical Society, 124(2):100–105, April 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986205.

Voegelin:1950:MRA

[Voe50] C. F. Voegelin. Magnetic recording of American Indian languages and the relationship of this to other kinds of memory. Proceedings of the American Philosophical Society, 94(3):295–300, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143565.

Voegelin:1952:BPP

[Voe52] C. F. Voegelin. The Boas Plan for the presentation of American Indian languages. Proceedings of the American Philosophical Society, 96(4):439–451, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143801.

Vogt:1973:LSH

[Vog73] John L. Vogt. The Lisbon Slave House and African trade, 1486– 1521. Proceedings of the American Philosophical Society, 117(1): 1–16, February 1973. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
985944.

Vryonis:1999:MVA

[Vry99] Speros Vryonis, Jr. Milton V. Anastos (10 July 1909–10 April 1997). Proceedings of the American Philosophical Society, 143(3): 444–450, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181956.

Vermeule:1978:NAR

[VW78] Emily Vermeule and Florence Wolsky. New Aegean relations with Cyprus: The Minoan and Mycenaean pottery from Toumba Tou Skourou, Morphou. Proceedings of the American Philosophical Society, 122(5):294–317, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986685.

Waara:1974:FIF

[Waa74] Elizabeth R. Waara. Franklin’s “ingenious friend” and scientific heir. Proceedings of the American Philosophical Society, 118(4): 315–320, September 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986278.

Wace:1953:EM

[Wac53] Alan J. B. Wace. Excavations at Mycenae 1952. Proceedings of the American Philosophical Society, 97(3):248–253, June 1953. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143747.

Wade:1957:MVC

[Wad57] Ira O. Wade. A manuscript of Voltaire’s “Candide”. Proceedings of the American Philosophical Society, 101(1):93–105, February 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985142.

Walford:1951:IEA

[Wal51a] Lionel A. Walford. On increasing the exploitation of aquatic resources for world food needs. Proceedings of the American Philosophical Society, 95(1):77–83, February 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143576.

Wallace:1951:FGS

[Wal51b] Anthony F. C. Wallace. The Frank G. Speck Collection. Proceedings of the American Philosophical Society, 95(3):286–289, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143066.

Wallace:1951:TKI

[Wal51c] Paul A. W. Wallace. They knew the Indian: The men who wrote the Moravian Records. Proceedings of the American Philosophical Society, 95(3):290–295, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143067.

Wallace:1952:JHI

[Wal52] Paul A. W. Wallace. John Heckewelder’s Indians and the Fenimore Cooper tradition. Proceedings of the American Philosophical Society, 96(4):496–504, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143804.

Wallace:1956:CMF

[Wal56] Henry A. Wallace. Corn and the midwestern farmer. Proceedings of the American Philosophical Society, 100(5):455–466, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143679.

Walton:1970:RS

[Wal70] Craig Walton. Ramus and Socrates. Proceedings of the American Philosophical Society, 114(2):119–139, April 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986030.

Walker:1980:UAF

[Wal80] Paul F. Walker. The U.S. Armed Forces: Prospect and proposal. Proceedings of the American Philosophical Society, 124(4):271–275, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986429.

Wallace:1983:PRN

[Wal83] Anthony F. C. Wallace. The perception of risk in Nineteenth Century anthracite mining operations. Proceedings of the American Philosophical Society, 127(2):99–106, April 1983. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986193.

Walther:1993:LMD

[Wal93] Herbert Walther. Light — magic doorway to new technology. Proceedings of the American Philosophical Society, 137(4):558–584, December 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/987074.

Walther:1999:KZL

[Wal99] Fritz R. Walther. Konrad Z. Lorenz (7 November 1903–27 February 1989). Proceedings of the American Philosophical Society, 143(3): 460–469, September 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181959.

Warren:1963:ERP

[War63] Shields Warren. Exposure rates and protective measures against radiation. Proceedings of the American Philosophical Society, 107 (1):18–20, February 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985466.

Warren:1966:CDF

[War66] Shields Warren. Civil defense — fact or fraud. Proceedings of the American Philosophical Society, 110(3):153–155, June 1966. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985982.

Ware:1977:CPP

[War77a] Willis H. Ware. Computers and personal privacy. Proceedings of the American Philosophical Society, 121(5):355–359, October 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986339.

Warren:1977:HNT

[War77b] Shields Warren. Hiroshima and Nagasaki thirty years after. Proceedings of the American Philosophical Society, 121(2):97–99, April 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986521.

Warner:1985:OPD

[War85] Deborah Jean Warner. Optics in Philadelphia during the Nineteenth Century. Proceedings of the American Philosophical Society, 129(3):291–299, September 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987012.

Wasson:1958:DMP

[Was58] R. Gordon Wasson. The divine mushroom: Primitive religion and hallucinatory agents. Proceedings of the American Philosophical Society, 102(3):221–223, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985574.

Washburn:1963:DSO

[Was63] Wilcomb E. Washburn. Discussion of the symposium on our contacts with American Indians, Polynesians, and Africans. Proceedings of the American Philosophical Society, 107(2):110–111, April 1963. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985439.

Waserman:1973:DZW

[Was73] Manfred J. Waserman. Dr. Zuriel Waterman: an itinerant surgeon in the Revolutionary Era. Proceedings of the American Philosophical Society, 117(5):388–403, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986608.

Watson:1953:NSR

[Wat53] David Meredith Seares Watson. On the nature of scientific research. Proceedings of the American Philosophical Society, 97(2):168–172, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143891.

Watts:1961:GFR

[Wat61] George B. Watts. The Geneva Folio reprinting of the “Encyclopédie”. Proceedings of the American Philosophical Society, 105 (4):361–367, August 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985495.

Watkins:1978:LUN

[Wat78] Calvert Watkins. Let us now praise famous grains. Proceedings of the American Philosophical Society, 122(1):9–17, February 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986259.

Watson:1995:FOH

[Wat95] Bernard C. Watson. Fritz Otto Haas (9 March 1915–2 January 1994). Proceedings of the American Philosophical Society, 139(4): 434–438, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987240.

Watson:1999:SLA

[Wat99] James D. Watson. Salvador E. Luria (13 August 1912–6 February 1991). Proceedings of the American Philosophical Society, 143(4): 680–683, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3181996.

Woronzoff-Dashkoff:1996:PRD

[WD96] A. Woronzoff-Dashkoff. Princess E. R. Dashkova: First woman member of the American Philosophical Society. Proceedings of the American Philosophical Society, 140(3):406–417, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987316.

West-Eberhard:1979:SSS

[WE79] Mary Jane West-Eberhard. Sexual selection, social competition, and evolution. Proceedings of the American Philosophical Society, 123(4):222–234, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986582.

Witkop:1998:THF

[WE98] Bernhard Witkop and Manfred Eigen. Theodor Hermann Felix Wieland (5 June 1913–24 November 1995). Proceedings of the American Philosophical Society, 142(2):316–319, June 1998. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3152310.

Weaver:1954:MML

[Wea54] Warren Weaver. The mathematical manuscripts of Lewis Carroll. Proceedings of the American Philosophical Society, 98(5):377– 381, October 1954. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143914.

Weaver:1960:IS

[Wea60] Warren Weaver. The imperfections of science. Proceedings of the American Philosophical Society, 104(5):419–428, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985228.

Webster:1976:LMF

[Web76] David Webster. Lowland Maya fortifications. Proceedings of the American Philosophical Society, 120(5):361–371, October 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986267.

Weems:1954:CN

[Wee54] Philip Van Horn Weems. Celestial navigation. Proceedings of the American Philosophical Society, 98(4):270–272, August 1954. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143880.

Weinreich:1959:HYL

[Wei59] Max Weinreich. History of the Yiddish language: The problems and their implications. Proceedings of the American Philosophical Society, 103(4):563–570, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985559.

Weiss:1960:KGP

[Wei60] Paul Weiss. Knowledge: a growth process. Proceedings of the American Philosophical Society, 104(2):242–247, April 1960. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985663.

Weitzmann:1966:MSC

[Wei66] Kurt Weitzmann. The mosaic in St. Catherine’s Monastery on Mount Sinai. Proceedings of the American Philosophical Society, 110(6):392–405, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985793.

Weiss:1969:PRF

[Wei69] Paul A. Weiss. “Panta’ rhei” — and so flow our nerves. Proceedings of the American Philosophical Society, 113(2):140–148, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985962.

Weinberg:1980:NEN

[Wei80] Alvin M. Weinberg. Is nuclear energy necessary? Proceedings of the American Philosophical Society, 124(6):399–403, December 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986241.

Weinberg:1981:EST

[Wei81] Steven Weinberg. Einstein and spacetime: Then and now. Proceedings of the American Philosophical Society, 125(1):20–24, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986183.

Weinberg:1986:BSN

[Wei86] Alvin M. Weinberg. Are breeders still necessary? Proceedings of the American Philosophical Society, 130(3):343–353, September 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986833.

Weintraub:1993:IBE

[Wei93] Sidney B. Weintraub. The interplay between economic and political opening in Mexico. Proceedings of the American Philosophical Society, 137(1):64–78, March 1993. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986945.

Weinberg:1997:EPW

[Wei97] Alvin M. Weinberg. Eugene Paul Wigner (17 November 1902–1 January 1995). Proceedings of the American Philosophical Society, 141(3):376–381, September 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987207.

Weinberg:1999:SM

[Wei99] Alvin M. Weinberg. Scientific millenarianism. Proceedings of the American Philosophical Society, 143(4):531–539, December 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3181986.

Wells:1974:WTE

[Wel74] Lloyd Allan Wells. “Why not try the experiment?” the scientific education of Edward Jenner. Proceedings of the American Philosophical Society, 118(2):135–145, April 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986662.

Wellek:1975:CE

[Wel75] René Wellek. Criticism as evaluation. Proceedings of the American Philosophical Society, 119(5):397–400, October 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986680.

Wells:1989:DRH

[Wel89] Robert V. Wells. The demography of a region: Historical reality or historian’s creation? Proceedings of the American Philosophical Society, 133(2):219–222, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/987052.

Wertenbaker:1953:ACW

[Wer53] Thomas J. Wertenbaker. The archaeology of Colonial Williamsburg. Proceedings of the American Philosophical Society, 97(1):44– 50, February 1953. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143729.

Wertenbaker:1957:J

[Wer57] Thomas J. Wertenbaker. Jamestown, 1607–1957. Proceedings of the American Philosophical Society, 101(4):369–374, August 1957. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985403.

Werner:1981:TJM

[Wer81] George S. Werner. Traveling journeymen in Metternichian South Germany. Proceedings of the American Philosophical Society, 125 (3):190–219, June 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986482.

Weston:1956:BCT

[Wes56] Corinne Comstock Weston. Beginnings of the classical theory of the English Constitution. Proceedings of the American Philosophical Society, 100(2):133–144, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143713.

Westheimer:1978:PLM

[Wes78] F. H. Westheimer. Photoaffinity labeling: Marking the receptors for biologically active molecules. Proceedings of the American Philosophical Society, 122(5):355–360, October 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986690.

Westheimer:1998:PDB

[Wes98a] F. H. Westheimer. Paul Doughty Bartlett (14 August 1907–11 October 1997). Proceedings of the American Philosophical Society, 142(3):446–456, September 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3152250.

Weston:1998:LSC

[Wes98b] Corinne Comstock Weston. Lord Salisbury: Conservative political leader and preeminent politician in Victorian England. Proceedings of the American Philosophical Society, 142(1):74–80, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3152265.

Wheeler:1981:LBH

[Whe81] John Archibald Wheeler. The lesson of the black hole. Proceedings of the American Philosophical Society, 125(1):25–37, February 1981. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986184.

Whitaker:1954:OWH

[Whi54] Arthur P. Whitaker. The origin of the Western Hemisphere idea. Proceedings of the American Philosophical Society, 98(5):323–326, October 1954. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143911.

Whipple:1958:OTA

[Whi58a] Fred L. Whipple. The optical tracking of artificial Earth satellites. Proceedings of the American Philosophical Society, 102(3):215–220, June 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985573.

Whitaker:1958:ESA

[Whi58b] Arthur P. Whitaker. The Enlightenment in Spanish America. Proceedings of the American Philosophical Society, 102(6):555–559, December 1958. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985566.

Whitaker:1960:AHS

[Whi60] Arthur P. Whitaker. Alexander von Humboldt and Spanish America. Proceedings of the American Philosophical Society, 104(3):317– 322, June 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985247.

Whitaker:1970:CUI

[Whi70a] Arthur P. Whitaker. Changing and unchanging interpretations of the enlightenment in Spanish America. Proceedings of the American Philosophical Society, 114(4):256–271, August 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985952.

White:1970:OC

[Whi70b] Lynn White, Jr. The origins of the coach. Proceedings of the American Philosophical Society, 114(6):423–431, December 1970. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986144.

Whipple:1976:CKR

[Whi76] Fred L. Whipple. Comet Kohoutek in retrospect. Proceedings of the American Philosophical Society, 120(1):1–6, February 1976. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986346.

White:1994:CED

[Whi94] Gilbert F. White. Coping with environmental degradation. Proceedings of the American Philosophical Society, 138(3):394–396, September 1994. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986745.

Wiener:1967:MOP

[Wie67] Frederick Bernays Wiener. The military occupation of Philadelphia in 1777–1778. Proceedings of the American Philosophical Society, 111(5):310–313, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986047.

Wiesner:1971:CUC

[Wie71] Jerome B. Wiesner. Can the university continue to be an important source of scientific research? Proceedings of the American Philosophical Society, 115(1):22–27, February 1971. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985920.

Wigner:1950:LS

[Wig50] Eugene P. Wigner. The limits of science. Proceedings of the American Philosophical Society, 94(5):422–427, October 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143609.

Wigner:1969:WM

[Wig69] Eugene P. Wigner. Are we machines? Proceedings of the American Philosophical Society, 113(2):95–101, April 1969. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985959.

Wilks:1951:ADM

[Wil51a] Samuel Stanley Wilks. The age distribution and mortality of the resident members of the American Philosophical Society. Proceedings of the American Philosophical Society, 95(4):393–400, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143281.

Williams:1951:MWM

[Wil51b] Alfred H. Williams. Money will not manage itself. Proceedings of the American Philosophical Society, 95(4):449–452, August 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143284.

Wilks:1955:SAD

[Wil55] Samuel S. Wilks. Statistical aspects of the design of experiments. Proceedings of the American Philosophical Society, 99(3):169–173, June 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143781.

Wilson:1956:RMA

[Wil56] John A. Wilson. The royal myth in Ancient Egypt. Proceedings of the American Philosophical Society, 100(5):439–442, October 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143677.

Wilkins:1957:AA

[Wil57] Ernest Hatch Wilkins. Arcadia in America. Proceedings of the American Philosophical Society, 101(1):4–30, February 1957. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985138.

Willits:1961:EMU

[Wil61a] Joseph H. Willits. Excellence and mediocrity in a university. Proceedings of the American Philosophical Society, 105(3):244–258, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985454.

Wilson:1961:LUN

[Wil61b] Edwin B. Wilson. The last unpublished notes of J. Willard Gibbs. Proceedings of the American Philosophical Society, 105(6):545–558, December 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985165.

Wilson:1961:NER

[Wil61c] Robert E. Wilson. The need for the early resumption of underground nuclear weapons tests. Proceedings of the American Philosophical Society, 105(2):206–208, April 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985631.

Willey:1966:NWA

[Wil66] Gordon R. Willey. New World archaeology in 1965. Proceedings of the American Philosophical Society, 110(2):140–145, April 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985746.

Wilson:1967:NCE

[Wil67] John A. Wilson. The Nubian Campaign: an exercise in international archaeology. Proceedings of the American Philosophical Society, 111(5):268–271, October 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986044.

Wilson:1968:SME

[Wil68] J. Tuzo Wilson. Static or mobile Earth: The current scientific revolution. Proceedings of the American Philosophical Society, 112 (5):309–320, October 1968. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986051.

Williams:1970:SOS

[Wil70] Robin M. Williams, Jr. Social order and social conflict. Proceedings of the American Philosophical Society, 114(3):217–225, June 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985699.

Williams:1972:WPO

[Wil72] John Hoyt Williams. Woodbine Parish and the “opening” of Paraguay. Proceedings of the American Philosophical Society, 116 (4):343–349, August 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985904.

Wilkins:1974:DC

[Wil74] Leslie T. Wilkins. Directions for corrections. Proceedings of the American Philosophical Society, 118(3):235–247, June 1974. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986441.

Williams:1975:OFL

[Wil75a] Harry F. Williams. Old French lives of Saint Barbara. Proceedings of the American Philosophical Society, 119(2):156–185, April 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986633.

Williams:1975:BGS

[Wil75b] John Hoyt Williams. From the barrel of a gun: Some notes on Dr. Francia and Paraguayan militarism. Proceedings of the American Philosophical Society, 119(1):73–86, February 1975. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986650.

Williams:1976:CSP

[Wil76] Stephen E. Williams. Comparative sensory physiology of the Droseraceae — the evolution of a plant sensory system. Proceedings of the American Philosophical Society, 120(3):187–204, June 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986558.

Wilson:1979:ECS

[Wil79] Edward O. Wilson. The evolution of caste systems in social insects. Proceedings of the American Philosophical Society, 123(4):204–210, August 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986579.

Wilson:1980:GEC

[Wil80] Leonard G. Wilson. Geology on the eve of Charles Lyell’s first visit to America, 1841. Proceedings of the American Philosophical Society, 124(3):168–202, June 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986369.

Wilkie:1983:FSH

[Wil83] Everett C. Wilkie, Jr. Franklin and “the sale of the Hessians”: The growth of a myth. Proceedings of the American Philosophical Society, 127(3):202–212, June 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986387.

Wills:1987:MPU

[Wil87] Garry Wills. A more perfect union. Proceedings of the American Philosophical Society, 131(3):289–297, September 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987024.

Willey:1990:AMP

[Wil90] Gordon R. Willey. Ancient Maya politics. Proceedings of the American Philosophical Society, 134(1):1–9, March 1990. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143899.

Wilson:1992:PIC

[Wil92] William Julius Wilson. The plight of the inner-city black male. Proceedings of the American Philosophical Society, 136(3):320–325, September 1992. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
986903.

Willey:1994:EWH

[Wil94a] Gordon R. Willey. Emil Walter Haury (2 May 1904–5 December 1992). Proceedings of the American Philosophical Society, 138(3): 426–430, September 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986750.

Williams:1994:CAJ

[Wil94b] John Williams. Carl Adam Johan Nordenfalk (13 December 1907– 13 June 1992). Proceedings of the American Philosophical Society, 138(4):554–557, December 1994. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/986854.

Wilson:1996:MI

[Wil96a] James Q. Wilson. Moral intuitions. Proceedings of the American Philosophical Society, 140(1):65–76, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/987276.

Wilson:1996:HSS

[Wil96b] R. R. Wilson. Hiroshima: The scientists’ social and political reaction. Proceedings of the American Philosophical Society, 140(3): 350–357, September 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987313.

Willey:1998:GAK

[Wil98] Gordon R. Willey. George Alexander Kubler (26 July 1912–3 October 1996). Proceedings of the American Philosophical Society, 142 (4):672–675, December 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152291.

Winternitz:1967:ATI

[Win67] Emanuel Winternitz. Anatomy the teacher: On the impact of Leonardo’s anatomical research on his musical and other machines. Proceedings of the American Philosophical Society, 111(4):234–247, August 1967. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986010.

Winks:1979:DSR

[Win79] Robin W. Winks. The decapitated society recovers its head. Proceedings of the American Philosophical Society, 123(5):265–269, October 1979. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986586.

Wingate:1983:PHM

[Win83] P. J. Wingate. The philosophy of H. L. Mencken. Proceedings of the American Philosophical Society, 127(2):95–98, April 1983. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986192.

Witthoft:1950:HPS

[Wit50] John Witthoft. The history and present status of Pennsylvania archeology. Proceedings of the American Philosophical Society, 94(3):301–307, June 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143566.

Witthoft:1951:IAM

[Wit51] John Witthoft. Iroquois archeology at the Mid-Century. Proceedings of the American Philosophical Society, 95(3):311–321, June 1951. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143069.

Witthoft:1952:PIS

[Wit52] John Witthoft. A Paleo-Indian site in Eastern Pennsylvania: an early hunting culture. Proceedings of the American Philosophical Society, 96(4):464–495, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143803.

Witkop:1999:PEH

[Wit99] Bernhard Witkop. Paul Ehrlich and his magic bullets, revisited. Proceedings of the American Philosophical Society, 143(4):540–557, December 1999. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3181987.

Weinstein:1970:IAP

[WM70] Leonard H. Weinstein and Delbert C. McCune. Implications of air pollution for plant life. Proceedings of the American Philosophical Society, 114(1):18–21, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/985721.

Wokeck:1989:GII

[Wok89] Marianne S. Wokeck. German and Irish immigration to Colonial Philadelphia. Proceedings of the American Philosophical Society, 133(2):128–143, June 1989. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987044.

Wolf:1952:CCP

[Wol52] Edwin Wolf. The 1706 charter of the City of Philadelphia. Proceedings of the American Philosophical Society, 96(4):388–405, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143796.

Wolf:1955:BFS

[Wol55] Edwin Wolf, 2nd. Benjamin Franklin’s Stamp Act cartoon. Proceedings of the American Philosophical Society, 99(6):388–396, December 1955. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143920.

Wolfson:1959:PIP

[Wol59] Harry A. Wolfson. Philosophical implications of the pelagian controversy. Proceedings of the American Philosophical Society, 103 (4):554–562, August 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985558.

Wolfgang:1973:CBC

[Wol73] Marvin E. Wolfgang. Crime in a birth cohort. Proceedings of the American Philosophical Society, 117(5):404–411, October 1973. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986609.

Wolfe:1982:AGS

[Wol82] Robert J. Wolfe. Alaska’s great sickness, 1900: an epidemic of measles and influenza in a virgin soil population. Proceedings of the American Philosophical Society, 126(2):91–121, April 1982. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986355.

Wolfart:1988:BAL

[Wol88] H. Christoph Wolfart. The beginnings of Algonquian lexicography. Proceedings of the American Philosophical Society, 132(1): 119–127, March 1988. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
3143827.

Wolfgang:1991:CAO

[Wol91] Lenora D. Wolfgang. Caxton’s “Aesop”: The origin and evolution of a fable: Or, do not believe everything you hear. Proceedings of the American Philosophical Society, 135(1):73–83, March 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/987151.

Wolfe:1994:LJG

[Wol94] Nancy Travis Wolfe. Lay judges in German criminal courts: The modification of an institution. Proceedings of the American Philosophical Society, 138(4):495–515, December 1994. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/986849.

Wolfgang:1995:WJR

[Wol95a] Lenora D. Wolfgang. William J. Roach (20 June 1907–30 July 1993). Proceedings of the American Philosophical Society, 139(2): 184–189, June 1995. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
987266.

Wolfram:1995:GBL

[Wol95b] Herwig Wolfram. Gerhart B. Ladner (3 December 1905–21 September 1993). Proceedings of the American Philosophical Society, 139 (4):440–442, December 1995. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987241.

Wolfgang:1996:TSO

[Wol96] Marvin E. Wolfgang. Thorsten Sellin (26 October 1896–17 September 1994). Proceedings of the American Philosophical Society, 140 (4):580–586, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987294.

Woodward:1953:PDH

[Woo53] Llewellyn Woodward. The place of diplomatic history in the study of international relations during the twentieth century. Proceedings of the American Philosophical Society, 97(2):151–158, April 1953. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143888.

Woodring:1966:PLB

[Woo66a] W. P. Woodring. The Panama Land Bridge as a sea barrier. Proceedings of the American Philosophical Society, 110(6):425–433, December 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985795.

Woodward:1966:SFR

[Woo66b] C. Vann Woodward. Seeds of failure in radical race policy. Proceedings of the American Philosophical Society, 110(1):1–9, February 1966. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985998.

Woodward:1992:RPW

[Woo92] C. Vann Woodward. Robert Penn Warren (April 24, 1905– September 15, 1989). Proceedings of the American Philosophical Society, 136(1):150–155, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/986807.

Woods:1998:JWF

[Woo98] Randall Bennett Woods. J. William Fulbright (9 April 1905–9 February 1995). Proceedings of the American Philosophical Society, 142(1):128–132, March 1998. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3152269.

Wood:1999:RTA

[Woo99] Gordon S. Wood. The real treason of aaron burr. Proceedings of the American Philosophical Society, 143(2):280–295, June 1999. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3181938.

Wright:1950:HIO

[Wri50a] Louis B. Wright. Historical implications of optimism in expanding America. Proceedings of the American Philosophical Society, 94 (1):18–23, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143247.

Wright:1950:PUE

[Wri50b] William Hammond Wright. On a proposal to use the extragalactic nebulae in measuring the proper motions of stars, and in evaluating the precessional constant. Proceedings of the American Philosophical Society, 94(1):1–12, February 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143245.

Wright:1951:LIR

[Wri51] Quincy Wright. Law and international relations. Proceedings of the American Philosophical Society, 95(5):490–503, October 1951. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143233.

Wriston:1956:FSR

[Wri56] Henry M. Wriston. The foreign service and representation abroad. Proceedings of the American Philosophical Society, 100(2):105–113, April 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143710.

Wright:1962:SE

[Wri62] Louis B. Wright. Shakespeare for everyman. Proceedings of the American Philosophical Society, 106(5):393–400, October 1962. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985483.

Wright:1972:LRI

[Wri72] Louis B. Wright. Life in a relatively innocent era. Proceedings of the American Philosophical Society, 116(5):396–405, October 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/986070.

Wright:1997:JBD

[Wri97] Gordon Wright. Jean-Baptiste Duroselle (17 November 1917–12 September 1994). Proceedings of the American Philosophical Society, 141(1):78–81, March 1997. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/987251.

Whitford:1970:TDP

[WW70] Kathryn Whitford and Philip Whitford. Timothy Dwight’s place in Eighteenth-Century American science. Proceedings of the American Philosophical Society, 114(1):60–71, February 1970. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985724.

Wyllie:1959:SDB

[Wyl59] Irvin G. Wyllie. Social Darwinism and the businessman. Proceedings of the American Philosophical Society, 103(5):629–635, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985421.

Yang:1996:SP

[Yan96] Chen Ning Yang. Symmetry and physics. Proceedings of the American Philosophical Society, 140(3):267–288, September 1996. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/987310.

Yngve:1960:MHL

[Yng60] Victor H. Yngve. A model and an hypothesis for language structure. Proceedings of the American Philosophical Society, 104(5): 444–466, October 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985230.

Yngve:1964:IMT

[Yng64] Victor H. Yngve. Implications of mechanical translation research. Proceedings of the American Philosophical Society, 108(4):275–281, August 1964. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985907.

Yochelson:1992:MPH

[Yoc92] Ellis L. Yochelson. Mr. Peale and his Mammoth Museum. Proceedings of the American Philosophical Society, 136(4):487–506, December 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986758.

Yochelson:1996:DCD

[Yoc96] Ellis L. Yochelson. Discovery, collection, and description of the Middle Cambrian Burgess Shale biota by Charles Doolittle Walcott. Proceedings of the American Philosophical Society, 140(4): 469–545, December 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987289.

Yoder:1982:SMC

[Yod82] Hatten S. Yoder, Jr. Strategic minerals: a critical research need and opportunity. Proceedings of the American Philosophical Society, 126(3):229–241, June 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986509.

Young:1963:GRR

[You63] Rodney S. Young. Gordion on the Royal Road. Proceedings of the American Philosophical Society, 107(4):348–364, August 1963. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985675.

Zakai:1991:OEN

[Zak91] Avihu Zakai. Orthodoxy in England and New England: Puritans and the issue of religious toleration, 1640–1650. Proceedings of the American Philosophical Society, 135(3):401–441, September 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986777.

Zaller:1991:HPR

[Zal91] Robert Zaller. Henry Parker and the regiment of true government. Proceedings of the American Philosophical Society, 135(2):255–285, June 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/987034.

Zangwill:1992:MFA

[Zan92] Oliver Zangwill. Meyer Fortes (April 25, 1906–January 27, 1983). Proceedings of the American Philosophical Society, 136(1):128–131, March 1992. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/986803.

Zeidenstein:1988:VET

[Zei88] George Zeidenstein. A view of the early twenty-first century. Proceedings of the American Philosophical Society, 132(2):204–220, June 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986866.

Zieger:1982:UCC

[Zie82] Robert H. Zieger. The union comes to Covington: Virginia paperworkers organize, 1933–1952. Proceedings of the American Philosophical Society, 126(1):51–89, February 1982. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/986478.

Zigrosser:1957:MSA

[Zig57] Carl Zigrosser. The medallic sketches of Augustin Dupré in American collections. Proceedings of the American Philosophical Society, 101(6):535–550, December 1957. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/985522.

Zimring:1974:TPI

[Zim74] Franklin E. Zimring. Threat of punishment as an instrument of crime control. Proceedings of the American Philosophical Society, 118(3):231–234, June 1974. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/986440.

Ziolkowski:1987:SMR

[Zio87] Theodore Ziolkowski. Schinkel’s Museum: The Romantic temple of art. Proceedings of the American Philosophical Society, 131(4): 367–377, December 1987. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/987137.

Zirkle:1959:SBD

[Zir59] Conway Zirkle. Species before Darwin. Proceedings of the American Philosophical Society, 103(5):636–644, October 1959. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985422.

Zubrod:1976:TSC

[Zub76] C. Gordon Zubrod. Treatment strategy for the cure of cancer. Proceedings of the American Philosophical Society, 120(2):100–102, April 1976. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986516.

Zuckerman:1980:SAS

[Zuc80] Lord Zuckerman. Science advisers and scientific advisers. Proceedings of the American Philosophical Society, 124(4):241–255, August 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/986424.

Zworykin:1961:MEP

[Zwo61] Vladimir K. Zworykin. Medical electronics-promise and challenge. Proceedings of the American Philosophical Society, 105(3):340–347, June 1961. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985460.